A Complete Bibliography of Publications in Mathematics and Computers in Simulation: 1980–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

22 August 2025 Version 1.00

Title word cross-reference

#9 [1116].

$10.00 [208, 212]. $102.00/Dfl. [428]. $109.75 [352]. $113.50 [1095]. $12.50 [129]. $131.50 [1063]. $14.95 [220, 317, 311]. $145.00 [518]. $15.95 [315]. $16.60 [316]. $17.00 [323]. $17.50 [321]. $184.25 [1062]. $19.50 [128, 405]. $20.00 [210, 280]. $223.75 [1059]. $24.95 [354]. $25.00 [211]. $28.00 [471]. $28.50 [313]. $29.50 [153, 219]. $29.95 [820]. $30.00 [206, 207]. $32.00 [357, 472]. $34.50 [446, 447]. $35.00 [407]. $36 [745]. $36.00 [209, 470]. $37.25 [409]. $39.00/Dfl.110.00 [778]. $39.75 [876]. $42.00 [1094]. $43.00 [403]. $44.50/Dfl.120.00 [758]. $44.90 [1092]. $45 [1090]. $45.00 [274]. $47.50 [449]. $48.00 [517]. $49.00/Dfl.120.00 [777]. $49.50 [1091]. $49.80 [270]. $49.95 [1088, 1098]. $50.00 [812]. $51.75 [814]. $54.20 [878]. $55.00 [351]. $55.50/Dfl.150.00 [757]. $58.25 [406]. $59.00/Dfl. [937]. $64.75/Dfl.175.00 [761]. $65.00/Dfl. [935]. $65/Dfl.210.00 [760]. $66.00 [401]. $69.00/Dfl. [938]. $69.50 [740]. $71.40 [448]. $74.00/Dfl.200.00 [765]. $79.50 [1054]. $79.50/Dfl.275.00 [767]. $81.50 [452]. $89.00/UK [888]. $93.00 [451]. $94.75 [1095]. $95.50 [743]. $99.00/Dfl. [933]. ^{1} [1079, 974]. ^{2} [136]. _{2} [442]. _{∞} [144]. A [442]. C [1079]. C^{1} [975]. H [1050]. k [947]. N [65, 948]. ph(t)/ph(t)/1/c [1002]. R [203, 1109]. θ [203]. u_{t} + cu_{x} = 0 [975]. ε [512]. ∥ [1137]. x [821]. Z [940, 942].

-body [65]. -collocation [974, 975]. -coupled [512]. -cyclic [416, 947]. -estimates [136]. -function [1050]. -stable [442]. -step [947]. -webs [1025].

/£ [1025].

0 [872, 983, 995, 979, 984, 991, 987, 41, 101, 80, 105, 42, 43, 73, 79, 96, 97, 99, 106, 107, 46, 77, 98, 102, 103, 48, 74, 81, 78, 37, 47, 104, 95, 75, 100, 108, 109, 213, 31, 122, 123, 125, 128, 153, 217, 218, 157, 158, 151, 196, 126, 214, 150, 154, 155, 156, 193, 129, 149, 148, 124, 185, 186, 194, 216, 215, 317, 357, 243, 250, 297, 314, 354, 359, 240, 294, 356, 358, 272, 277, 323, 353, 320, 322, 278, 316]. 0 [352, 298, 312, 351, 321, 273, 276, 295, 318, 355, 279, 402, 410, 452, 456, 380, 387, 404, 428, 429, 448, 449, 466, 478, 431, 434, 457, 458, 382, 385, 407, 455, 433, 468, 450, 475, 495, 405, 408, 469, 384, 390, 451, 502, 435, 383, 430, 474, 409, 476, 386, 403, 406, 467, 477, 497, 536, 559, 560, 593, 553, 558, 533, 557, 537, 555, 590, 596, 594, 595, 519, 520, 535, 552, 554, 588, 589, 591, 620, 626, 627, 653, 654, 621]. 0 [658, 657, 622, 624, 655, 707, 758, 779, 716, 717, 718, 719, 761, 706, 710, 770, 715, 729, 708, 713, 767, 743, 744, 745, 709, 739, 765, 768, 711, 731, 780, 760, 740, 757, 762, 769, 712, 855, 859, 882, 818, 820, 822, 847, 887, 889, 857, 873, 851, 874, 815, 814, 812, 813, 819, 825, 811, 821, 878, 884, 885, 891, 856, 858, 881, 860, 823, 853, 879, 826, 876, 816, 890, 968, 1030, 928, 935, 971, 925, 923, 1026, 929, 953, 1010, 1023]. 0 [930, 934, 941, 958, 1009, 927, 956, 1008, 962, 966, 1011, 1104, 1142, 1069, 1099, 1116, 1146, 1147, 1064, 1110, 1111, 1145, 1088, 1067, 1101, 1055, 1060, 1065, 1102, 1091, 1105, 1053, 1095, 1098, 1109, 1114, 1144, 989, 993, 742, 607]. 0-07-021910-9 [156]. 0-08-027624-5 [457]. 0-12-012120-4 [357]. 0-12-012127-1 [1104]. 0-12-044370-8 [554]. 0-12-049260-1 [1098]. 0-12-085555-0 [979]. 0-12-122920-3 [380]. 0-12-122952-1 [1099]. 0-12-164780-3 [430]. 0-12-170620-6 [740]. 0-12-237459-2 [1115]. 0-12-237460-6 [1115]. 0-12-249050-9 [1146]. 0-12-249930-1 [1055]. 0-12-269990-4 [1142]. 0-12-293475-X [1060]. 0-12-346350-5 [73]. 0-12-363480-6 [558]. 0-12-378650-9 [75]. 0-12-378660-6 [78]. 0-12-417635-6 [853]. 0-12-434100-4 [987]. 0-12-468662-1 [384]. 0-12-49850-7 [95]. 0-12-545075-3 [1064]. 0-12-550355-5 [1145]. 0-12-592520-4 [502]. 0-12-632620-7 [385]. 0-12-637475-9 [1065]. 0-12-659050-8 [359]. 0-12-665250-3 [77]. 0-12-697545-0 [1067]. 0-12-700850-0 [405]. 0-12-706960-7 [536]. 0-12-728450-8 [1144]. 0-12-742689-2 [1140]. 0-12-742690-6 [1140]. 0-12-747252-5 [74]. 0-12-774861-X [1069]. 0-12-784390-6 [1088]. 0-13-201229-4 [710]. 0-13-317107-8 [654]. 0-19-853190-7 [815]. 0-19-853364-0 [884]. 0-19-853612-7 [882]. 0-19-853614-3 [879]. 0-19-853912-6 [929]. 0-19-858181-5 [971]. 0-201-00018-0 [555]. 0-201-02988-X [155]. 0-201-04144-8 [594]. 0-201-04297-5 [353]. 0-201-06414-6 [475]. 0-201-06672-6 [560]. 0-201-07890-2 [596]. 0-201-08725-1 [356]. 0-201-09286-7 [984, 1009]. 0-201-10102-5 [557]. 0-201-10271-4 [589]. 0-201-10806-2 [874]. 0-201-11148-9 [872]. 0-201-14400-X [467]. 0-201-14473-5 [590]. 0-262-23091-7 [157]. 0-262-62052-9 [1011]. 0-262-7110-X [1008]. 0-275-92229-4 [981]. 0-275-92630-3 [981]. 0-306-40788-4 [448]. 0-306-40847-3 [449]. 0-387-15283-0 [1054]. 0-387-18140-7 [1096]. 0-387-18141-5 [1096]. 0-387-18143-1 [1096]. 0-387-96146-1 [729]. 0-408-01153-X [466]. 0-412-27960-6 [930]. 0-442-22646-2 [816]. 0-442-27344-4 [821]. 0-442-80325-7 [354]. 0-444-00049-6 [129]. 0-444-00238-3 [148]. 0-444-00243-X [150]. 0-444-00256-1 [151]. 0-444-00322-3 [272].

0-444-00350-9 [213]. 0-444-00366-5 [295]. 0-444-00381-9 [320]. 0-444-00421-1 [314]. 0-444-00581-1 [351]. 0-444-01173-0 [991]. 0-444-01181-1 [935]. 0-444-01250-8 [995]. 0-444-41722-2 [149]. 0-444-70003-X [860]. 0-444-70011-0 [812]. 0-444-70022-6 [881]. 0-444-70028-5 [851]. 0-444-70051-X [890]. 0-444-70053-6 [889]. 0-444-70056-0 [887]. 0-444-70057-9 [873]. 0-444-70065-X [857]. 0-444-70067-6 [891]. 0-444-70074-9 [814]. 0-444-70126-5 [858]. 0-444-70140-0 [885]. 0-444-70209-1 [1059]. 0-444-70225-3 [941]. 0-444-70288-1 [962]. 0-444-70305-5 [958]. 0-444-70349-7 [988, 957]. 0-444-70350-0 [988, 957]. 0-444-70351-9 [988, 957]. 0-444-70400-0 [1062]. 0-444-70430-2 [1063]. 0-444-80095-1 [965]. 0-444-85028-7 [98]. 0-444-85038-4 [97, 122]. 0-444-85096-1 [102]. 0-444-85105-4 [96, 123]. 0-444-85106-2 [99, 125]. 0-444-85107-0 [126]. 0-444-85143-7 [100]. 0-444-85181-X [103]. 0-444-85186-0 [37, 124]. 0-444-85188-7 [31]. 0-444-85288-3 [104]. 0-444-85299-9 [106]. 0-444-85332-4 [47]. 0-444-85340-5 [81]. 0-444-85413-4 [298]. 0-444-85447-9 [109]. 0-444-85449-5 [216]. 0-444-85471-1 [107]. 0-444-86008-8 [217]. 0-444-86012-6 [294]. 0-444-86053-3 [218]. 0-444-86056-8 [214]. 0-444-86109-2 [240]. 0-444-86123-8 [215]. 0-444-86129-7 [196]. 0-444-86133-5 [273]. 0-444-86134-3 [408]. 0-444-86142-4 [250, 297]. 0-444-86199-8 [352]. 0-444-86216-1 [318]. 0-444-86223-4 [312]. 0-444-86224-2 [322]. 0-444-86279-X [429]. 0-444-86284-6 [382]. 0-444-86307-9 [387, 428]. 0-444-86330-3 [383]. 0-444-86338-9 [456]. 0-444-86356-7 [451]. 0-444-86358-3 [452]. 0-444-86359-1 [407]. 0-444-86372-9 [386]. 0-444-86374-5 [458]. 0-444-86377-X [409]. 0-444-86380-X [435]. 0-444-86392-3 [390]. 0-444-86399-0 [410]. 0-444-86406-7 [455]. 0-444-86429-6 [431]. 0-444-864329-X [406].

0-444-86440-7 [433]. 0-444-86453-9 [620]. 0-444-86459-8 [450]. 0-444-86462-8 [468]. 0-444-86474-1 [434]. 0-444-86479-2 [607]. 0-444-86517-9 [519]. 0-444-86550-0 [497]. 0-444-86669-8 [591]. 0-444-86729-5 [621]. 0-444-86770-8 [653]. 0-444-86795-3 [588]. 0-444-86797-X [622]. 0-444-86852-6 [655]. 0-444-86892-5 [657]. 0-444-87695-2 [770]. 0-444-87696-0 [765]. 0-444-87743-6 [758]. 0-444-87773-8 [767]. 0-444-87798-3 [757]. 0-444-87826-2 [761]. 0-444-87881-5 [760]. 0-444-87900-5 [768]. 0-444-87955-2 [847]. 0-444-99767-9 [193]. 0-444-99780-6 [108]. 0-470-20088-X [1101]. 0-470-21406-6 [1114]. 0-471-01751-5 [185]. 0-471-0259-4 [154]. 0-471-05179-9 [186]. 0-471-06336-3 [316]. 0-471-08684-3 [323]. 0-471-09725-X [595]. 0-471-09988-0 [278]. 0-471-10027-7 [358]. 0-471-10345-4 [520]. 0-471-10419-1 [478]. 0-471-80963-2 [706]. 0-471-81929-8 [825]. 0-471-83966-3 [989]. 0-471-88814-1 [731]. 0-471-89067-7 [708]. 0-471-90170-9 [819]. 0-471-90322-1 [713]. 0-471-90490-2 [717]. 0-471-90558-5 [715]. 0-471-90849-5 [744]. 0-471-90854-1 [826]. 0-471-90934-3 [742]. 0-471-91046-5 [811]. 0-471-91130-5 [820]. 0-471-91201-8 [855]. 0-471-91547-5 [927]. 0-471-91661-7 [1103]. 0-471-91952-7 [1086]. 0-471-91958-6 [1110]. 0-471-92079-7 [1085]. 0-519-02101-3 [714]. 0-521-22767-4 [153]. 0-521-30245-5 [718]. 0-521-30674-4 [1091]. 0-521-30811-9 [934]. 0-521-33058-0 [923]. 0-521-33107-2 [993]. 0-521-36929-0 [1147]. 0-566-00558-1 [477]. 0-669-0214-9 [158]. 0-7201-0734-2 [79]. 0-7204-0406-1 [42]. 0-7204-0499-1 [101]. 0-7204-0500-9 [41]. 0-7204-0757-5 [43]. 0-7273-0203-5 [553]. 0-7273-0501-8 [46]. 0-7273-1303-7 [48]. 0-7458-0100-5 [818]. 0-7458-0101-3 [878]. 0-7458-0139-0 [1102]. 0-7458-0288-5 [928]. 0-7458-0581-7 [1082]. 0-7458-0635-X [1114]. 0-7923-0023-8 [1111].

0-7923-0072-6 [1113]. 0-7923-0073-4 [1113]. 0-8039-2956-0 [983]. 0-8176-3413-4 [1092]. 0-8176-3414-2 [1090]. 0-8176-3417-7 [1141]. 0-8218-0114-7 [1010]. 0-8247-1387-7 [355]. 0-8247-1421-0 [404]. 0-8247-1543-8 [469]. 0-8247-1634-5 [537]. 0-8247-6548-6 [402]. 0-8247-7322-5 [876]. 0-8247-7570-8 [925]. 0-8247-8019-1 [1095]. 0-8247-8025-6 [1084]. 0-8357-1327 [474]. 0-8357-1364 [476]. 0-85226-434-8 [711]. 0-85296-242-8 [401]. 0-85312-164-8 [321]. 0-85312-491-4 [535]. 0-85312-495-7 [822]. 0-85312-563-5 [559, 593]. 0-85312-565-1 [626]. 0-85312-597-X [658]. 0-85312-670-4 [712]. 0-85312-703-4 [627]. 0-85312-736-0 [719]. 0-85312-772-7 [707]. 0-85312-807-3 [716]. 0-85312-848-0 [709]. 0-85312-887-1 [739]. 0-85312-946-0 [1082]. 0-85501-452-0 [128]. 0-86103-036-2 [194]. 0-86103-051-6 [279]. 0-86380-033-5 [744]. 0-86380-033-5/0-471-90849-5 [744]. 0-88385-059-1 [1116]. 0-89116-461-829f83-540-16308-5 [743]. 0-89838-230-0 [956]. 0-89871-207-6 [813]. 0-905451-15-5 [533]. 0-905451-16-3 [552]. 0-905451-49-X [1112]. 0-905451-67-8 [1096]. 0-905451-74-0 [1100]. 0-905451-80-5 [1097]. 0-905451-94-5 [1096]. 0-905451-95-3 [1096]. 0-906048-15 [105]. 0-906048-15-X [80]. 0-906048-44-3 [276]. 0-906048-48-6 [277]. 0-906048-63-X [403]. 0-90648-15-X [243]. 0-931215-70-6 [1097]. 0-946536-00-7 [592]. 0-946536-05-8 [592]. 0-9614088-0-4 [745]. 0. [127]. 0.-444-85217-4 [127]. 0.12.047650.9 [446]. 0.12.175520.7 [447]. 0.12.747254.1 [436]. 0.444.86386.9 [453]. 0.471.10449.3 [454]. 008176-2217-9 [1094]. 0149 [471]. 0149-7111 [471]. 0167 [389, 391]. 0167-50036 [389]. 0167-5036 [391]. 0276 [470]. 0276-6574 [470]. 1 [994, 101, 42, 107, 102, 209, 151, 244, 314, 356, 248, 351, 303, 318, 478, 434, 407, 384, 388, 477, 626, 729, 713, 739, 731, 854, 824, 875, 846, 826, 1029, 935, 960, 936, 962, 1104, 1070, 1099, 1116, 1055, 1095, 1098, 678, 104, 1105]. 1-55608-000-X [970]. 1-55608-010-7 [970]. 1-D [678]. 105.00/US$42.50/£ [956]. 10th [59, 191, 198]. 11 [243]. 110.00/US$54.00/£ [954]. 112.50 [970]. 118 [1083]. 11th [985, 891]. 12.50/$26.50 [477]. 120.00/£ [969]. 12th [832]. 13 [164, 163, 302, 311, 315, 300, 239, 246, 247, 249, 301, 303, 501, 500]. 13-056607-7 [164]. 13-165183-8 [302]. 13-283796-X [501]. 13-308254-7 [311]. 13-383950-8 [315]. 13-461566-2 [300]. 13-539908-4 [239]. 13-605154-5 [246]. 13-605162-6 [500]. 13-661983-5 [163, 249, 301]. 13-709873-1 [303]. 13-731729-8 [247]. 130.00/US$64.00/£ [994]. 1327 [474]. 135.00/£ [964]. 1364 [476]. 140.00/US$51.75 [873, 890]. 140.00/US$56.00/£ [982]. 140.00/US$69.00/£ [986]. 145.00/£ [937]. 148.00/US$49.00/£ [953]. 15 [105]. 150.00/US$69.00/£ [940]. 150.00/US$69.00/UK [875]. 16.00.ISBN [299]. 160.00/£ [938]. 160.00/US$59.25 [887, 889]. 168.00/US$68.00/£ [959]. 17th [53]. 185.00/£ [924]. 185.00/US$79.00/£ [990]. 185.00/US$97.25 [957]. 185.00US$68.50 [891]. 18th [381]. 19.50 [592]. 190.00/US$89.00/£ [980]. 195 [1096]. 195.00/US$78.00/£ [966, 967]. 1975 [74]. 1976 [79, 102]. 1977 [99, 126]. 1978 [181, 38, 39, 51, 52, 53, 54, 55, 56, 50]. 1979 [56, 58, 191, 109, 217, 218, 216, 215, 298]. 198 [401]. 1980 [85, 19, 2, 62, 107, 213, 206, 210, 214, 211, 219, 245, 250, 297]. 1981 [221, 199, 132, 113, 166]. 1982 [252, 325, 226, 282, 306, 319, 470]. 1983 [393, 438, 481, 363, 413, 460]. 1984 [523, 565, 599, 506, 541, 581]. 1985 [630, 663, 612, 672]. 1986 [722, 747, 783, 698, 733, 772]. 1987 [803, 794, 862, 1090]. 1988 [832, 1014, 945, 946, 997, 973, 1150]. 1989 [1041, 1072]. 19th [52]. 1C [55]. 1HQ [401, 403]. 1SA [271]. 1st [988, 46, 245, 472, 966, 967].

2 [79, 99, 125, 161, 207, 149, 194, 299, 240, 300, 322, 275, 381, 596, 623, 717, 761, 706, 770, 778, 777, 764, 847, 874, 849, 883, 848, 877, 816, 893, 959, 1032, 1004, 924, 1025, 1117, 1143, 993, 678, 607, 986, 498, 561]. 2-04-015748-4 [1004]. 2-7261-0232-8 [152]. 2-87080-000-2 [381]. 2-D [678]. 215.00/US [888]. 21st [206, 207]. 235.00/£ [932]. 24.95 [1086]. 245.00/£ [933]. 24th [381]. 25.00 [556]. 250.00/US$80.00 [881]. 25th [860]. 27 [1104]. 270.00/US$129.00/£ [960]. 273.60/US$144.00 [957]. 28.00/DM [479]. 29.00 [473]. 29.95 [878, 1082]. 29.95/U.S. [878]. 295.00/US$118.00/£ [961]. 295.45/US$155.50 [965]. 2AA [533, 552]. 2JQ [159]. 2nd [54, 56, 50, 60, 458, 708, 855, 873, 964].

3 [982, 985, 44, 29, 77, 34, 104, 160, 161, 218, 162, 148, 211, 195, 269, 248, 272, 323, 270, 316, 275, 276, 452, 380, 449, 495, 408, 390, 383, 430, 411, 561, 517, 521, 552, 588, 659, 623, 624, 759, 779, 716, 730, 741, 757, 762, 764, 766, 769, 763, 880, 817, 824, 849, 852, 883, 810, 878, 856, 848, 823, 877, 879, 968, 955, 1028, 1006, 893, 1026, 926, 936, 941, 1007, 963, 1012, 1024, 1070, 1057, 1107, 1064, 1056, 1108, 1106, 1061, 1053]. 3 [1068, 1109, 989, 992, 742, 293]. 3-211-81566-X [269]. 3-211-81606-2 [275]. 3-211-81800-6 [659]. 3-211-81848-0 [856]. 3-258-03076-6 [521]. 3-258-03175-4 [561]. 3-258-03417-6 [759]. 3-258-03423-0 [769]. 3-258-03537-7 [852]. 3-258-03596-2 [848]. 3-519-02150-1 [1103]. 3-519-02448-9 [880]. 3-528-03340-1 [248]. 3-540-08442-8 [29]. 3-540-08680-3 [34]. 3-540-08775-3 [44]. 3-540-09822-4 [270]. 3-540-09823-2 [161]. 3-540-09853-4 [160]. 3-540-10159-4 [162]. 3-540-11590-0 [495]. 3-540-12761-5 [741]. 3-540-12772-0 [624]. 3-540-12883-2 [623]. 3-540-13371-2 [764]. 3-540-13373-9 [763]. 3-540-13623-1 [824]. 3-540-13625-8 [810]. 3-540-15283-0 [1053]. 3-540-15485-X [817]. 3-540-15851-0 [779]. 3-540-16031-0 [762]. 3-540-16854-0 [823]. 3-540-17145-2 [877]. 3-540-17505-9 [955]. 3-540-17767-1 [936]. 3-540-17936-4 [1112]. 3-540-18023-0 [968]. 3-540-18115-6 [1107]. 3-540-18300-0 [1026]. 3-540-18372-8 [1024]. 3-540-18395-7 [1006]. 3-540-18468-6 [1028]. 3-540-18546-1 [1070]. 3-540-18584-4 [1100]. 3-540-18726-X [1108]. 3-540-18861-4 [1061]. 3-540-18869-X [1056]. 3-540-18916-5 [1068]. 3-540-18958-0 [1109]. 3-540-90482-0 [625]. 3-540-90540-5 [195]. 3-540-90910-9 [766]. 3-540-96302-2 [849]. 3-540-96516-5 [963]. 3-540-96598-X [1012]. 3-540-96608-0 [1087]. 3-7643-1733-7 [730]. 3-7643-1761-2 [883]. 3-7643-1855-4 [926]. 3-7643-1940-2 [893]. 3-7643-1970-4 [985]. 3-7643-2217 [1094]. 3-7643-2220-9 [1106]. 3-7643-2221-7 [1057]. 3-7643-2242-X [1007]. 3-7643-2255-1 [1083]. 3-7643-3183-6 [992]. 3-7643-3413-4 [1092]. 3-7643-3414-2 [1090]. 3-7643-3417-7 [1141]. 3-88538-800-6 [411]. 3-D [293]. 310.00/US$124.00 [319]. 34.00 [479]. 35 [1085]. 358 [131]. 35th [129]. 365.00/US$149.00/£ [970]. 3C [56]. 3rd [991, 41, 30, 208, 241, 887].

4 [990, 985, 987, 96, 97, 47, 160, 122, 123, 153, 162, 154, 127, 357, 250, 297, 239, 270, 298, 312, 448, 502, 561, 534, 520, 535, 589, 627, 710, 745, 712, 821, 926, 931, 1004, 1033, 1142, 1061, 1091, 608, 384]. 41C/CV [411]. 48.00/DM [759]. 49.00/DM [848]. 4C [54]. 4th [38, 51, 102, 209, 304, 244].

5 [43, 73, 74, 81, 206, 163, 185, 216, 195, 245, 353, 246, 249, 301, 273, 295, 496, 457, 458, 559, 593, 553, 533, 557, 537, 590, 562, 621, 657, 718, 715, 743, 744, 741, 768, 760, 818, 820, 851, 811, 858, 876, 928, 971, 958, 927, 937, 963, 1145, 1068]. 5.5 [159]. 5/0 [744]. 50036 [389]. 5036 [391]. 52.00 [856]. 55.00/DM [852]. 5536E [129]. 58H [279]. 5th [40].

6 [986, 78, 108, 129, 294, 277, 279, 402, 431, 382, 475, 500, 411, 560, 558, 518, 521, 659, 655, 758, 759, 740, 888, 886, 889, 813, 891, 881, 853, 938, 1028, 961, 929, 954, 930, 939, 969, 1107, 1110, 1093, 1088, 1114, 992]. 6574 [470]. 67 [248]. 69 [1087]. 690.- [269]. 6th [953, 959, 1036].

7 [984, 98, 48, 95, 100, 164, 31, 157, 196, 354, 311, 358, 355, 499, 385, 455, 433, 451, 498, 536, 707, 730, 708, 882, 822, 815, 852, 964, 1006, 1010, 933, 1009, 1057]. 7111 [471]. 7273 [313, 496, 388, 498]. 7273-0203-5 [496]. 7273-1404-1 [388]. 7273-1610-9 [313]. 7273-1611-7 [498]. ’78 [39, 53]. 78CH1289 [45]. 78CH1289-8C [45]. 78CH1321 [57]. 78CH1321-9C [57]. 78CH1338 [56]. 78CH1338-3C [56]. 78CH1387 [50]. 78CH1387-0 [50]. 78CH1388 [53]. 78CH1388-8C [53]. 78CH1397 [52]. 78CH1397-9C [52]. 78CH1400 [55]. 78CH1400-1C [55]. 78CH1412 [51]. 78CH1412-6 [51]. 78CH1413 [54]. 78CH1413-4C [54]. 78CH1415 [59]. 78CH1415-9SMC [59]. ’79 [215]. 7th [564].

8 [995, 46, 29, 217, 214, 215, 302, 359, 315, 352, 247, 321, 468, 450, 469, 594, 554, 591, 653, 654, 767, 711, 855, 850, 825, 810, 925, 1031, 1024, 1111, 1144]. 80-81351 [241]. 80CH1498 [206]. 80CH1498-5 [206]. 80CH1511 [245]. 80CH1511-5 [245]. 80CH1569 [211]. 80CH1569-3 [211]. 80CH1571 [210]. 80CH1571-9 [210]. 80CH1598 [207]. 80CH1598-2 [207]. 80CH1607 [209]. 80CH1607-1 [209]. 80CH170 [244]. 80CH170-1 [244]. ’81 [279, 312, 382, 383]. 829f83 [743]. ’83 [621, 653]. 8359 [317, 499]. 8359-0208-0 [317]. 8359-9874-7 [499]. ’84 [757, 767]. ’85 [891, 881]. ’86 [857]. 8C [53, 45]. 8th [470].

9 [872, 980, 41, 106, 75, 109, 213, 158, 210, 156, 193, 186, 320, 313, 456, 387, 428, 386, 519, 620, 766, 763, 880, 873, 814, 819, 955, 1027, 1005, 934, 967, 1011, 940, 942, 1146, 1106, 1065]. 90 [990, 982, 980, 994, 986, 562, 517, 518, 534, 778, 777, 780, 854, 859, 888, 850, 875, 846, 938, 1029, 1030, 960, 964, 961, 1027, 953, 954, 959, 1005, 1023, 1031, 1032, 931, 933, 937, 924, 939, 966, 967, 969, 940, 942, 1025, 1033, 1089, 1093, 1143, 1105, 608]. 90-247-3408-8 [850]. 90-247-3423-1 [875]. 90-247-3581-5 [932]. 90-247-3582-3 [932]. 90-247-3656-0 [1023]. 90-247-3751-6 [1093]. 90-277-1497-5 [562]. 90-277-1590-4 [534]. 90-277-1865-2 [778]. 90-277-1917-9 [940, 942]. 90-277-1928-4 [990]. 90-277-2077-0 [859]. 90-277-2186-6 [961]. 90-277-2188-2 [777]. 90-277-2211-0 [780]. 90-277-2253-6 [888]. 90-277-2289-7 [933]. 90-277-2328-1 [854]. 90-277-2336-2 [1032]. 90-277-2376-1 [846]. 90-277-2379-6 [938]. 90-277-2383-4 [1033]. 90-277-2437-7 [964]. 90-277-2443-1 [1029]. 90-277-2506-3 [982]. 90-277-2513-6 [939, 969]. 90-277-2514-4 [931]. 90-277-2515-2 [924]. 90-277-2550-0 [966]. 90-277-2551-9 [967]. 90-277-2553-5 [937]. 90-277-2580-2 [959]. 90-277-2581-0 [953]. 90-277-2592-6 [954]. 90-277-2605-1 [960]. 90-277-2629-9 [980, 1005].

90-277-2639-6 [986]. 90-277-2641-8 [1031]. 90-277-2653-1 [994]. 90-277-2699-X [1089]. 90-277-2726-0 [1030]. 90-277-2730-9 [1027]. 90-277-2756-2 [1025]. 90-277-2788-0 [1105]. 90-277-2837-2 [1143]. 90-6191-225-3 [517]. 90-6191-246-6 [518]. 90-90-00599-4 [608]. 98.- [469]. 9971 [886]. 9971-5-0090-6 [886]. 9C [52, 57]. 9SMC [59]. 9th [211, 215, 471, 875].

A.D.I. [203]. Aarts [939]. abbruch [763]. Ablaufsteuerung [759]. abstract [559, 593, 629, 1060]. abstracts [82]. Academic [979, 987, 73, 77, 74, 78, 95, 75, 219, 357, 359, 446, 380, 385, 405, 447, 436, 384, 502, 430, 536, 558, 592, 554, 740, 888, 875, 853, 1030, 970, 1032, 956, 1025, 1033, 1104, 1140, 1142, 1069, 1099, 1146, 1064, 1089, 1111, 1117, 1145, 1093, 1088, 1067, 1055, 1060, 1065, 1105, 1113, 1098, 1115, 1144]. academy [929]. accelerated [236]. acceleration [137, 292, 416, 145]. accelerator [615]. accuracy [28, 542, 1022, 1017]. accurate [601]. ACM [52, 44]. acoustic [420]. acoustics [988, 957]. Acquisition [210, 317]. active [344]. activity [1131, 530]. Aczél [846]. Adaptation [216, 976]. adaptive [667, 8, 900, 367, 605, 1131, 734, 147, 402]. Added [86]. Addison [872, 984, 155, 274, 356, 353, 475, 467, 560, 557, 590, 596, 594, 589, 874, 1009]. Addison-Wesley [872, 984, 155, 274, 356, 353, 475, 467, 560, 557, 590, 596, 594, 589, 874, 1009]. addressable [161]. Adey [46, 553, 533, 552]. adjoint [999]. Administrative [264]. Adomian [554]. Advanced [1093, 371, 1140, 1142]. Advances [1057, 1094, 606, 41, 304, 357, 660, 1104]. Advantages [642]. advisor [1134]. advisory [1134, 1133]. AE [608]. aeration [921]. aeronautical [73, 502]. Afdeling [608]. AG [355, 402, 404, 469]. Agarwal [1083]. aggregation [925]. aggregative [487]. agign [350]. AGNES [1123]. Aho [555]. AI [1135]. aid [31]. aided [41, 106, 45, 149, 359, 497, 1112, 1146, 287, 404]. air [754, 642, 896]. Akademische [597]. Akaike [888]. Akin [446]. al [96, 35, 123, 181, 190, 475, 812, 816, 1011]. Albert [34]. Alder [477, 1098]. Alder-shot [477]. Alefeld [269]. algebra [932, 1082, 559, 593]. algebraic [728, 1012, 1117, 1029]. algebriques [705]. algorithm [26, 605, 284, 536]. Algorithm-structured [536]. Algorithmes [242]. Algorithmic [990]. Algorithmische [160]. Algorithms [988, 474, 680, 343, 314, 678, 549, 694, 900, 737, 547, 162, 707, 706, 713, 884, 956, 410, 560, 855, 882]. Also [932]. alternating [9, 670]. alternating-direction [9]. Alternative [570, 443]. alternatives [39]. Alto [40]. Ambrosetti [937]. America [191, 1116]. American [1010]. amplitude [170]. Amsterdam [988, 995, 991, 41, 101, 42, 43, 96, 97, 99, 106, 107, 98, 102, 103, 81, 37, 47, 49, 104, 100, 108, 109, 31, 122, 123, 125, 151, 196, 126, 150, 193, 149, 184, 148, 124, 127, 272, 273, 410, 452, 456, 387, 389, 428, 429, 431, 434, 458, 382, 407, 455, 433, 468, 450, 475, 453, 408, 390, 451, 391, 435, 383, 409, 386, 406, 467, 497, 519, 588, 591, 620, 653, 621, 657, 622, 655, 758, 761, 770, 767, 765, 768, 760, 757]. Amsterdam [847, 887, 889, 857, 873, 851, 814, 812, 885, 891, 858, 881, 860, 890, 935, 957, 965, 941, 958, 962, 1062, 1063, 1059, 607]. analog [235, 723, 317]. analogue [90]. analyses [485]. Analysis [780, 885, 684, 668, 665, 807, 1020, 1018, 995, 984, 95, 152, 148, 354, 269, 352, 499, 455, 406, 432, 537, 589, 713, 731, 762, 817, 874, 849, 883, 811, 853, 961, 954, 1031, 1009, 924, 1036, 1096, 1111, 1068, 237, 785, 756, 511, 582, 118, 444, 1047, 1135, 1052, 349, 202, 735, 975, 679, 1021, 773, 639, 371, 117, 418, 607, 1016, 1048, 478, 477, 75, 240, 716, 719, 765, 854, 815, 879, 1029, 955, 932, 930, 936, 659, 859]. Analytic [774, 274, 396, 685, 400]. Anatomy [1123]. Anderson [872, 153, 311]. Andreotti [1092]. Andreotti-Grauert [1092]. Andrews [126]. animal [919]. Ann [52]. Annals [318]. annealing [939, 969]. anniversary [860]. Annotated [717, 970]. Announcement [223, 198, 828]. Annual [40, 213, 191, 389, 391, 875]. anomalies [910]. Anwendungsaspekte [479]. any [1132]. Application [960, 12, 369, 23, 441, 1019, 103, 74, 321, 495, 762, 20, 683, 444, 775, 896, 1000, 257, 288, 418, 615, 345, 142, 774, 316, 730, 446]. Applications [60, 208, 209, 218, 961, 1057, 1066, 688, 988, 980, 994, 987, 986, 42, 54, 56, 154, 181, 194, 195, 244, 457, 436, 564, 758, 712, 818, 819, 938, 1029, 1032, 933, 958, 924, 969, 940, 942, 1025, 1033, 1096, 1094, 1089, 1117, 1114, 677, 583, 1016, 641, 208, 780, 151, 187, 274, 429, 535, 761, 741, 874, 891, 1031, 1007, 939, 940, 942, 1096, 1057, 1062, 1110, 1111, 1085, 1082, 1106, 452]. Applied [217, 469, 914, 952, 984, 408, 745, 878, 955, 1023, 1009, 1095, 176, 284, 488, 559, 593, 707, 854]. Approach [475, 402, 589, 702, 785, 333, 91, 647, 668, 1135, 425, 375, 734, 773, 443, 527, 1016]. approaches [966, 967, 334]. Approximate [567, 432, 796]. Approximating [1002]. approximation [32, 537, 626, 730, 740, 813, 883, 810, 885, 926, 1035, 1085, 838, 993, 445, 231, 177, 462, 136, 89, 665, 144, 616, 777, 882]. approximations [705, 9, 840, 893]. April [44, 214, 216, 471, 381, 533, 828, 252, 393, 523, 630, 722, 803, 40]. aquatic [999, 204, 309]. aquifer [259]. Arato [47]. Arbor [52]. Architecture [471, 693, 40, 50, 1003, 158]. Architectures [536, 448]. area [655]. arithmetic [51, 880, 1037, 228, 229, 1022]. Arlington [743]. Armand [381]. Arnold [556]. arrays [536]. Arrillaga [556]. arrivals [753]. Art [812, 75, 164, 879]. arterial [750]. artificial [964, 119, 263, 758]. Ashurst [533, 552]. ASI [1023, 937, 1105]. aspects [654, 1096, 1114, 342]. Ass [129]. assess [268]. Assessing [398]. assessment [642]. Assisting [1135]. associated [1109]. Association [1116]. Asymptotic [231, 1047, 976, 936]. asymptotics [1109, 968]. Asynchronous [476, 737, 169]. Athens [788]. Atluri [847]. atmosphere [604, 572, 738]. atmospheric [911]. Auerbach [213]. Auflage [561]. August [988, 106, 50, 57, 198, 218, 126, 409, 857, 438, 564, 565, 663, 747]. Austin [471]. Australia [451, 910]. Austria [214, 953, 959]. Author [111, 224, 361, 504, 610, 696, 792, 1039, 6]. Authority [852]. auto [347, 348]. auto-regression [348]. auto-regressive [347]. autocorrelated [234]. automata [155, 285, 718]. Automated [101, 456]. automaten [275]. automatic [1126, 756, 116, 1123, 1121, 633]. automotive [202]. AUTOSIM [633]. Autowave [938]. availability [78]. available [932, 1133]. Aven [924]. average [713]. averaged [1086]. averaging [276].

B [872, 73, 96, 33, 81, 123, 153, 152, 192, 185, 186, 274, 296, 353, 295, 410, 408, 390, 451, 381, 467, 520, 655, 758, 767, 768, 822, 880, 875, 883, 878, 858, 823, 876, 923, 934, 1033, 1086, 1103, 1110, 1095, 1098, 1109, 992, 953]. B-4000 [381]. Baber [386]. background [310]. backward [399]. Bad [953, 959, 625, 714, 965]. badly [8]. Balboo [314]. Balkema [517, 518]. Ball [739]. banded [143]. Bangladesh [703]. Barnard [510]. Barnard-Birnbaum [510]. Barnett [1068]. Barnhill [757]. Barrekette [35]. Barrier [916]. Basch [76]. base [101]. based [1138, 966, 967, 1127, 683, 605, 1132, 796, 1131, 347, 257, 94, 1067, 1105]. Basel [985, 355, 402, 404, 469, 537, 730, 883, 893, 926, 1007, 1092, 1057, 1094, 1141, 1106, 1083, 1090, 992]. Bases [468, 107, 434]. basic [1097, 1115, 597, 496, 553, 561]. BASIC-Software [597]. basis [751, 378]. Batam [634]. Bauer [160]. Bay [513]. Bd [242]. be [660, 656]. Beach [59, 210]. beacon [908]. beam [604]. Beauty [779]. bed [492, 493]. beef [922]. behaved [8]. behavior [231]. behaviour [345, 866]. Belgium [381]. Belitsky [1106]. Bellaire [57]. Bellman [777, 886]. Beltrami [979]. Belytschko [607]. benefits [398]. Bensoussan [452, 406, 1004]. Berg [936]. Berigir [848]. Berlin [44, 33, 29, 34, 160, 161, 162, 179, 181, 188, 192, 183, 216, 195, 495, 411, 625, 623, 624, 779, 762, 764, 763, 817, 824, 810, 823, 877, 1028, 1026, 1024, 1070, 1096, 1100, 1107, 1112, 1056, 1108, 1061, 1053, 1068, 1087, 1109]. Bern [479, 561, 597, 521, 759, 769, 852, 848]. Berne [31]. Bernold [873]. Best [213, 231, 637]. Besturingsystemen [608]. Bethlehem [304]. between [510, 409, 576]. Beyond [725]. BFr [381]. bibliographies [717]. Bibliography [110, 130, 165, 197, 222, 251, 281, 305, 324, 360, 392, 412, 437, 459, 480, 503, 538, 539, 540, 563, 580, 598, 661, 61, 83, 522, 609, 628]. Biebel [758]. Biendl [759]. bifurcation [995, 798]. Bifurcations [1048]. bilinear [291, 1045]. biological [93, 327, 999, 329, 80, 105, 243]. biologischer [764]. Biomathematics [1030]. biomedical [828, 454]. biotechnological [787]. bipolar [149, 748]. Birkhäuser [985, 730, 883, 893, 926, 1007, 1092, 1057, 1094, 1141, 1106, 1083, 1090, 992]. Birkhäuser-Verlag [1007, 1057, 1106]. Birnbaum [510]. birth [526]. bivariate [1050]. Blahut [557]. block [684, 138]. block-pulse [684]. blocks [647].

Blutkonserven [479]. Board

[3, 114, 227, 364, 507, 613, 699, 795, 894, 1042]. Bockmann [858]. body [917, 65, 551]. Boehm [96, 123, 757]. boiler [284]. Boisvert [766]. Boite [387, 428]. Bollobás [923]. Bologna [45]. Bolyai [1059]. bond [608]. Book [872, 990, 988, 983, 995, 982, 979, 980, 984, 991, 994, 985, 987, 986, 981, 41, 101, 80, 105, 30, 38, 39, 40, 42, 43, 51, 52, 53, 54, 55, 56, 58, 73, 79, 96, 97, 99, 106, 44, 107, 33, 46, 29, 77, 98, 36, 45, 102, 103, 48, 50, 74, 81, 78, 32, 35, 34, 37, 47, 49, 57, 104, 95, 59, 60, 75, 100, 108, 109, 76, 160, 164, 213, 31, 122, 123, 125, 128, 153, 161, 206, 208, 209, 217, 218, 220, 152, 157]. Book [158, 207, 210, 151, 196, 212, 126, 162, 214, 150, 154, 155, 156, 159, 187, 189, 193, 129, 149, 178, 179, 181, 184, 188, 190, 191, 148, 192, 124, 163, 183, 185, 186, 194, 211, 216, 219, 221, 127, 180, 215, 195, 182, 242, 304, 317, 357, 243, 244, 245, 250, 274, 297, 299, 302, 314, 319, 354, 359, 240, 294, 296, 269, 311, 356, 358, 315, 248, 272, 277, 300, 323, 353, 239, 246, 270, 271, 320, 322, 278, 316, 352, 298, 312]. Book [351, 247, 249, 275, 301, 303, 313, 321, 273, 276, 280, 295, 318, 355, 241, 279, 402, 410, 446, 452, 456, 454, 496, 380, 387, 389, 404, 428, 429, 448, 449, 466, 470, 471, 472, 478, 499, 431, 434, 457, 458, 479, 382, 385, 407, 455, 501, 433, 468, 473, 450, 475, 495, 405, 447, 453, 408, 436, 469, 500, 384, 388, 390, 451, 502, 391, 435, 381, 383, 430, 474, 498, 409, 476, 386, 401, 403, 406, 411, 467, 477, 497, 432, 579]. Book [536, 559, 560, 593, 553, 561, 597, 556, 558, 533, 557, 537, 555, 590, 596, 594, 595, 562, 592, 517, 518, 534, 519, 520, 535, 521, 552, 554, 588, 589, 591, 659, 660, 620, 625, 626, 627, 653, 654, 621, 623, 658, 657, 656, 622, 624, 655, 707, 758, 759, 779, 716, 717, 718, 719, 761, 706, 710, 730, 770, 715, 729, 708, 713, 714, 778, 767, 743, 744, 745, 709, 739, 741, 765, 768, 777, 711, 731, 780, 760, 740, 757, 762, 764, 766].

Book [769, 712, 763, 854, 855, 859, 882, 888, 886, 818, 820, 822, 847, 880, 887, 889, 817, 824, 857, 873, 851, 874, 849, 815, 850, 875, 814, 812, 813, 819, 825, 852, 883, 810, 811, 821, 878, 884, 885, 891, 856, 858, 881, 848, 860, 823, 846, 853, 877, 879, 826, 876, 816, 890, 938, 968, 1029, 1030, 928, 935, 955, 957, 960, 971, 1028, 964, 970, 932, 961, 1027, 965, 925, 1006, 893, 923, 1026, 929, 953, 954, 959, 1005, 1010, 1023, 1031, 1032]. Book [926, 930, 931, 933, 934, 936, 941, 958, 1009, 1007, 927, 937, 963, 1004, 956, 1008, 1012, 924, 939, 962, 966, 967, 969, 1011, 940, 942, 1025, 1033, 1024, 1086, 1092, 1104, 1140, 1142, 1070, 1096, 1097, 1100, 1057, 1062, 1063, 1069, 1094, 1099, 1107, 1112, 1116, 1146, 1147, 1064, 1089, 1103, 1110, 1111, 1117, 1145, 1141, 1056, 1084, 1085, 1093, 1088, 1067, 1143, 1082, 1108, 1101, 1106, 1055, 1059, 1060, 1061, 1065, 1083, 1102, 1058, 1091, 1105, 1113, 1053, 1054, 1066, 1068, 1087]. Book [1090, 1095, 1098, 1109, 1114, 1115, 1144, 989, 993, 992, 742, 13, 14, 15, 16, 17, 18, 607, 608, 1100]. boolean [751]. Bordas [1004]. Boston [979, 987, 1092, 1104, 1140, 1142, 1094, 1099, 1146, 1117, 1141, 1088, 1090, 1098, 1115, 1144]. Bottazzini [849]. Boulder [409]. bound [185, 186]. boundary [994, 313, 498, 1070, 1100, 1094, 618, 66, 24, 665, 837, 569, 796, 509, 921, 619, 1078, 1070, 1096, 1097]. boundary-layer [665]. boundary-layers [921]. boundary-value [796, 509]. bounded [951, 89]. bounds [664]. Bowker [474, 476]. Box [271, 280, 279, 401, 403]. Boyer [380, 1099]. Bradley [299, 875]. Braess [810]. brain [349]. braking [903]. Brans [312]. Brauner [1023]. Braunschweig [34]. breakdown [748]. Brebbia [313, 466, 553, 1100, 1112]. Brebbra [1096]. Breedveld [608]. breezes [788]. Brever [429]. Brownian [574]. Bruckmann [358]. Bruell [314]. Brunswick [660, 656]. bubble [270]. Bubnicki [193]. Bucharest [985]. Buchhandlung [597]. Bucy [534]. Budapest [407]. Buffer [237]. building [300]. Bullen [980, 1005]. Bultheel [893]. Burger [588]. Burnett [874]. bus [903, 797]. Butcher [811]. Butler [58]. Butrimenko [47]. Butterworths [466]. Bywater [315].

C [983, 985, 30, 38, 39, 40, 51, 52, 53, 55, 56, 73, 107, 36, 50, 81, 32, 60, 189, 194, 127, 357, 314, 359, 313, 410, 466, 405, 390, 411, 553, 561, 556, 592, 517, 654, 761, 743, 731, 882, 822, 815, 812, 883, 811, 821, 881, 1027, 1006, 1023, 930, 937, 1104, 1140, 1096, 1100, 1062, 1063, 1112, 1116, 1090, 608, 974]. CA [983, 984, 280]. cable [907]. cable-stayed [907]. CAD [322, 468]. CAD/CAM [322]. CAGD [757]. Calculateur [516]. calculation [170, 863, 586, 514]. calculus [849, 928]. Caldwell [388, 875]. Calendar [721, 830, 944, 1038, 782, 802, 1148]. calibration [899]. Call [84, 829, 832]. CAM [322]. Cambridge [153, 157, 718, 745, 923, 934, 1008, 1011, 1147, 1091, 993]. Canada [126, 298, 518, 876, 1095, 198]. CANDU [485]. can’t [182]. capabilities [371]. capturing [601]. car [238]. car-following [238]. Caracas [30]. carbon [804]. Cardiology [470]. care [54, 244]. Carey [654]. Carlo [825, 510, 790, 908, 422, 440, 839, 117]. Carlo-type [839]. Carrick [128]. Carter [981]. cascade [1078]. case [807, 713, 299]. Cashwell [100]. Casulli [984, 1009]. catchment [912]. cattle [922]. causality [686]. caused [617]. Cebedoc [381]. celebration [860]. cell [20]. Cellier [430]. cellular [350]. Center [39, 40, 51, 52, 53, 54, 55, 56, 50, 57, 59, 60]. centralized [808]. Centre [45, 206, 209, 207, 210, 211, 244, 245]. centro [143]. centro-symmetric [143]. century [860]. cereal [921]. certain [143]. CESTAC [1019, 1020]. CH [402, 404, 880]. CH-4010 [402, 404]. CH12849 [40]. CH13110 [38]. CH13284 [39]. CH4010 [355]. Chadeev [103]. chain [775]. Chalmer [876]. Chandi [122]. Chandy [97]. change [1105]. Changes [897, 868, 910]. channels [904]. chaos [995]. Chapman [930]. Chapter [44]. char [491]. characteristic [10]. characteristics [530, 96, 123]. characterization [890, 465]. Charles [1147]. Chatelin [740]. CHDL [429]. Chebyshev [140]. checkerboard [416]. checking [151]. chemical [30, 133, 605, 493]. Chen [1006]. Cheney [813]. Chernavskii [938]. Chesnay [152]. Chester [245]. Chicago [56, 60, 209]. Chichester [478, 473, 559, 556, 707, 716, 717, 719, 706, 713, 714, 709, 739, 712, 818, 820, 822, 819, 811, 878, 826, 927, 1086, 1085, 1082, 1102]. Chief [1066]. Chinese [971]. choice [146]. choices [168]. Chon [881]. Chow [1083]. Chrestenson [677]. chronology [915]. Chui [1006]. Ciarlet [98]. circadian [917]. circles [848]. circuit [285, 387, 428]. Circuits [245, 151, 149, 1036, 128, 245]. circular [421]. circulation [910]. Citadelle [221]. City [984]. Clapeyron [1077]. Clarendon [882, 815, 884, 971, 929]. Clark [447]. class [1073, 704, 345, 568]. classes [985, 195, 724]. Classical [960]. Classification [43]. Cliffs [180, 710]. climate [1105]. climatic [918]. clinical [347, 348]. clogged [205]. closed [314, 785]. Cloth [878, 716]. Cluster [719]. CML [533]. CND [621, 622]. CO [50, 990, 982, 980, 984, 991, 994, 986, 448, 449, 474, 476, 1029, 1027, 1031, 1143]. coal [634, 902]. coal-to-electricity [902]. coastline [703]. code [1001, 629, 485, 490, 604, 511, 82]. codes [151, 323, 557]. coding [693, 1117, 904, 501]. Coefficient [614]. coefficients [1049, 71]. Coffman [924]. Cohen [467]. cold [492]. Collatz [926]. Collected [1107]. collectors [573]. College [220, 451, 533, 38]. collision [587]. Collocation [949, 7, 1079, 12, 975, 974, 441]. Colloquia [1059]. Colophon [182]. Colorado [409]. Columbia [220]. Comb [368]. combination [871]. Combinatorial [717]. Combinatorics [818]. combined [774]. Combining [642]. combustion [1023, 490, 202]. Command [851]. comments [6]. committees [831]. Common [603, 1069]. communication [158, 322, 247, 770, 285, 808, 851]. Communications [53, 472, 881, 1117, 472, 622]. Comp [1100]. Company [41, 101, 30, 42, 43, 96, 97, 99, 106, 107, 98, 36, 102, 103, 32, 35, 37, 47, 49, 104, 100, 108, 109, 213, 31, 122, 123, 125, 217, 218, 151, 196, 126, 214, 150, 193, 149, 184, 148, 124, 216, 127, 215, 250, 297, 314, 319, 240, 294, 296, 272, 320, 322, 352, 298, 312, 351, 273, 295, 318, 410, 452, 456, 387, 389, 428, 429, 431, 434, 458, 382, 407, 455, 433, 450, 453, 408, 390, 451, 391, 435, 383, 409, 386]. Company [406, 477]. comparative [246, 585, 887]. Comparison [255, 670]. compartmental [335, 338, 331, 337]. Compatibility [212]. COMPCON [39, 207, 53]. compiler [1138, 104, 150]. compiler-based [1138]. Complementary [664]. Complex [467, 327, 188, 849, 1111, 527, 1130, 328]. complexes [516]. complexity [520, 855, 723, 555, 596, 1067]. compressible [439, 949]. compression [692, 683, 701]. COMPSAC [56, 60, 209]. computable [648]. Computation [198, 891, 1075, 29, 269, 409, 880, 723, 1078, 476, 155, 194, 711, 1012, 1091, 626, 718]. Computational [552, 847, 1097, 1053, 1054, 607, 988, 496, 553, 555, 654, 877, 1030, 963, 1096, 1100, 1099, 1112, 1098, 231, 604, 988, 73, 314, 466, 478, 558, 716, 822, 847, 955, 957, 1062, 1063, 607]. computations [592, 729]. computed [1021]. Computer [420, 40, 51, 52, 53, 54, 55, 56, 60, 213, 206, 208, 209, 220, 207, 210, 212, 149, 269, 280, 471, 475, 656, 875, 1067, 1055, 492, 582, 173, 755, 419, 922, 587, 652, 704, 41, 38, 97, 45, 37, 47, 31, 122, 196, 150, 148, 124, 304, 240, 356, 278, 298, 321, 279, 380, 433, 405, 453, 497, 536, 594, 564, 660, 620, 623, 657, 629, 767, 887, 857, 851, 881, 890, 924, 1037, 1146, 1102, 977, 177, 90, 749, 115, 176, 202, 736, 138, 773]. computer [92, 82, 171, 787, 120, 287, 39, 40, 42, 51, 52, 53, 54, 55, 56, 58, 79, 97, 106, 31, 122, 206, 208, 209, 244, 250, 297, 302, 319, 354, 359, 389, 404, 429, 448, 471, 499, 556, 627, 761, 820, 880, 889, 1112]. Computer-aided [149, 404]. Computer-oriented [269]. computerized [530]. Computers [131, 321, 472, 103, 182, 245, 389, 391, 1104, 1116, 635, 11, 357, 277, 386, 658, 962, 1098, 30, 128, 218, 470, 472, 1116]. Computertechniek [608]. Computing [217, 207, 210, 269, 1067, 1055, 1119, 1150, 962, 1035, 1099, 1128, 1125, 800, 454, 153, 217]. concept [427]. concepts [104, 126, 874, 1031, 1097, 371]. conceptual [625, 255, 375]. concerning [415]. concise [816, 971]. condition [66]. conditions [618]. conduction [618, 71, 619]. cones [1060]. Conf [456, 451]. Conference [41, 30, 52, 53, 56, 46, 29, 45, 37, 57, 59, 60, 208, 209, 207, 210, 214, 211, 245, 351, 470, 472, 381, 401, 533, 743, 887, 851, 1150, 985, 99, 106, 31, 125, 126, 124, 298, 279, 407, 409, 518, 857, 873, 875, 829, 1036, 59, 191, 965]. conflict [395]. conflicts [754]. Congress [198, 215, 891, 832]. Conley [1147]. connected [259]. connection [510, 998, 333]. consequent [897]. constant [674]. constrained [1019]. constraints [870, 602]. constructing [99, 125, 99, 125]. construction [104, 327, 805]. Content [161]. Content-addressable [161]. Contents [505, 611, 697, 793, 1040]. context [1129]. Continued [274, 1094].

Continuous

[636, 238, 64, 702, 838, 87, 133, 5, 841]. continuous-system [5]. continuous-time [841]. continuously [703]. continus [516]. continuum [986, 1044, 886]. Contraction [176]. contraintes [242]. contribution [1077]. Contributions [1057]. Control [210, 271, 1056, 1061, 105, 128, 157, 193, 190, 243, 457, 501, 435, 406, 557, 828, 1028, 961, 1026, 1024, 1070, 1066, 1068, 515, 549, 902, 754, 531, 308, 614, 808, 737, 144, 284, 674, 841, 94, 288, 527, 1022, 652, 147, 418, 512, 287, 441, 605, 80, 105, 103, 178, 402, 410, 452, 562, 519, 710, 770, 853, 1004, 449]. controlled [372]. controller [868, 293]. convection [365, 349, 388]. Convergence [137, 88, 687, 670]. conversion [317]. converting [680]. Convertisseurs [221]. convolution [162, 1132]. Cooperation [35]. Cooperative [852]. coordinate [1080]. coordinates [1075, 638]. coordination [516]. Corbett [872]. Corduneanu [743]. Cornelius [822]. correct [416, 1021]. correcting [323]. correctness [380]. correlated [640, 835]. cosine [692]. cost [25]. costs [398]. Coupled [546, 64, 121, 512, 797]. Coupling [1130, 1076]. course [846]. Courtois [657]. cover [762, 1061, 1115]. Cox [58, 882]. Coxeter [812]. creation [1124]. creative [396]. creativity [848]. criteria [1031]. criterion [888]. critical [1129]. Cross [48, 299]. Crouch [1024]. crown [916]. crude [529]. cryptography [742]. CSDT [134, 167]. CSSL [5]. CTAC [1062]. CTAC-87 [1062]. Cullum [814]. current [906, 374, 169, 935]. currents [464]. curriculum [38]. curved [1145, 665]. Curves [1146]. CV [411]. cycle [804, 376]. cyclic [416, 947]. cylinder [421]. cylinders [86].

D [990, 988, 982, 980, 984, 994, 985, 987, 986, 41, 101, 80, 105, 42, 48, 78, 37, 49, 164, 196, 150, 154, 155, 156, 187, 124, 304, 297, 294, 358, 248, 303, 410, 404, 450, 390, 381, 556, 555, 520, 744, 765, 764, 854, 859, 874, 883, 810, 846, 879, 938, 1029, 1027, 929, 1005, 1031, 1009, 1116, 1056, 1088, 1143, 1082, 1059, 1060, 1091, 1087, 1109, 678, 293]. D-6200 [248]. D.U [129]. d’activité [221]. Dalkowski [411]. Dallas [655]. damit [275]. dark [358]. Data [1020, 210, 468, 469, 693, 38, 107, 247, 434, 458, 954, 339, 640, 692, 684, 228, 683, 637, 1135, 808, 677, 347, 152, 294, 296]. data-communication [808]. data-flow [693]. Datenbanken [44]. David [620]. Davis [589, 1063]. DC [432, 207, 244]. DDA [25]. death [526]. Debnath [991]. debugging [1124]. decade [358, 674]. December [55, 59, 217, 1150, 85, 199, 325, 481, 599, 783, 1014]. Decentralized [527]. decision [184, 964, 1031, 394, 706]. decisions [131, 398, 1]. decomposed [693]. decomposition [976, 952, 690, 516, 691]. decoupling [284, 293]. decoupling-controller [293]. deep [917]. defined [642]. Deformations [1111]. deforming [703]. Deitel [590]. Dekker [355, 402, 404, 469, 537, 876, 925, 1084, 1095]. Delahaye [1053, 1054]. Deland [36]. delay [808, 730]. delays [93]. Delhi [711]. Delobel [107]. delta [578]. delta-modulated [578]. d’ensembles [221]. density [838, 948]. Department [206, 209, 207, 210, 211, 401, 403]. dependence [131, 1]. dependent [836, 71, 493, 147]. depletion [120]. deposition [915, 573]. Dept [244, 245, 660, 656]. DEQSOL [1124]. derivation [416]. description [429, 564, 761]. descriptions [126]. Design [45, 986, 164, 149, 192, 317, 240, 315, 239, 241, 594, 853, 1146, 1085, 1091, 1068, 906, 368, 677, 293, 92, 287, 778, 45, 164, 150, 456, 387, 404, 428, 499, 497, 589]. design-modeling [1091]. Designing [983, 710]. designs [343, 342, 1130]. desire [230, 725]. dessen [479]. desulfurization [489]. Desynchronization [584]. detecting [151]. Determination [798, 63, 1131]. Determining [483]. deterministic [276, 642, 340]. Deutsch [411]. Development [489, 11, 100, 899]. developments [1097, 421]. device [700]. devices [391, 1036, 866, 748, 659]. Dewilde [387, 428]. Dfi [965]. DFl [408, 990, 982, 980, 994, 986, 41, 101, 42, 43, 96, 97, 99, 106, 107, 98, 102, 103, 81, 37, 47, 49, 104, 100, 108, 109, 76, 31, 122, 123, 125, 126, 193, 149, 124, 127, 215, 250, 297, 314, 319, 294, 272, 320, 322, 298, 312, 273, 295, 318, 410, 456, 429, 431, 434, 458, 455, 433, 468, 450, 453, 435, 497, 562, 518, 588, 591, 620, 653, 621, 657, 622, 655, 888, 887, 889, 873, 875, 885, 891, 881, 890, 957, 960, 964]. Dfl [970, 961, 953, 954, 959, 941, 958, 956, 966, 967, 969, 940, 1089, 1111, 1093, 1143, 1105, 1113]. Dfl.100.00 [382, 519]. Dfl.125.00 [517]. Dfl.130.00/U.S.$69.00/£ [1027]. DFl.150.00 [383]. Dfl.183.00 [534]. Dfl.225.00 [607]. Dfl.240.00 [387]. Dfl.295.00/U.S.$149.00/£ [1032]. Dfl.295.00/U.S.$195.00/£ [1030]. Dfl.60.00 [390]. Dfl.80.00 [386]. Dfl120.00/U.S.$59.00/£ [1031]. Dfl185.00/U.S.$89.00/£ [1033]. Dfl195.00/U.S.$99.00/£ [1023]. Dfl260.00/U.S.$139.00 [1025]. Dfl320.00/U.S.$149.00/£ [1029]. diagnosis [1124]. diagnostics [966, 967]. Diaz [760]. Diego [1069, 1064, 1145, 1067, 1055, 1060, 1065]. difference [987, 712, 426, 72, 174, 461]. differencing [617]. different [237, 724, 978]. Differential [219, 991, 129, 304, 451, 660, 744, 712, 811, 885, 877, 1088, 1055, 1065, 8, 575, 1044, 426, 139, 290, 1122, 116, 203, 235, 1137, 1149, 616, 1121, 1003, 4, 369, 1000, 568, 839, 336, 415, 786, 1089]. diffusion [388, 345, 668, 365, 349, 840, 736, 572, 494]. digital [164, 315, 500, 90, 636, 809, 11, 685, 690, 141, 905, 157, 770, 1117]. digitally [372]. dimensional [485, 67, 600, 638, 20, 378, 379, 483, 1079, 784, 202, 439, 140, 690, 551, 738, 171, 615, 932]. dimensions [640, 866, 351]. dimpled [736]. d’integration [63]. direct [21, 725, 796, 616]. direct-executing [725]. directe [152, 705]. direction [9, 233, 670]. discharges [263]. discontinuities [116]. discontinuous [1089, 205]. discovery [1011]. Discrete [323, 676, 318, 977, 1051, 692, 998, 168, 978, 136, 89, 377, 1074, 1045, 840, 431, 824, 1064]. discrete-event [1074]. discrete-time [998, 136]. discretizations [700, 22, 200]. discussion [1129]. diseases [346]. diskreter [33]. disordered [347]. dispatching [261]. dispersion [897]. dispersive [426]. displacement [72, 145, 548]. Disposition [479]. dissection [719]. distorting [234]. Distributed [302, 94, 58, 107, 109, 181, 353, 886, 1116, 1128, 614, 142, 486, 288, 147, 512, 107, 158, 207, 210, 240, 434, 457, 458, 1028]. Distributed-parameter [94, 288, 512]. Distributed-processor [158]. distribution [68, 489, 308, 753, 896, 836, 1050, 234, 602, 307]. Distributor [191]. distributors [1114]. disturbances [265]. Division [129]. DM [44, 33, 29, 34, 160, 128, 161, 195, 248, 479, 495, 561, 659, 625, 623, 624, 759, 779, 741, 880, 817, 824, 849, 852, 810, 856, 848, 823, 968, 955, 1006, 936, 963, 1012, 1070, 1056, 1108, 1061, 1053, 1068, 1087]. DM.30.00 [411]. DM138.00 [766]. DM34.00 [769]. DM36.00 [1024]. DM39.00 [762]. DM39.50 [764, 763]. DM84.00 [1028, 1026]. Do [856, 182]. Dodes [148]. Dodrecht [1025]. domain [1056, 366]. domains [175]. d’optimisation [242]. Dordrecht [76, 562, 534, 778, 777, 780, 854, 859, 888, 850, 875, 846, 938, 1029, 1030, 960, 964, 970, 932, 961, 1027, 953, 954, 959, 1005, 1023, 1031, 1032, 931, 933, 937, 956, 924, 939, 966, 967, 969, 940, 942, 1033, 1089, 1111, 1093, 1143, 1105, 1113]. down [164]. Dreyfus [182]. drill [1069]. Drozzinov [1033]. drug [31]. Dual [377, 694]. Dubois [381]. Ducrot [31]. ducts [800]. Dudkin [925]. d’un [63]. Dungar [517]. durations [754]. durch [44]. during [921]. Dyadic [680]. Dyadic- [680]. Dynamic [68, 551, 979, 154, 187, 192, 186, 131, 435, 702, 571, 543, 1, 817]. dynamical [704, 95, 271, 829, 1142, 1110, 838]. Dynamics [475, 531, 1001, 956, 310, 231, 486, 735, 328, 73, 180, 502, 941, 1063, 1103, 1110, 1145].

E.J [126]. E.N.S.E.M [221]. each [280]. earliest [1077]. Early [230]. East [980, 1029, 961, 940, 942]. Eastern [711]. ECG [298, 683]. ecological [381]. Economic [852, 848, 1077, 135]. economics [5, 184]. ecosystem [259, 488]. ecosystems [266, 644]. Ed [770, 767, 765, 780, 760, 757, 764, 847, 887, 889, 857, 873, 860, 890, 923, 941, 1104, 1070, 1111, 1093, 1061, 1098, 855]. Ed. [708]. Edition [241, 991]. Editions [381]. Editor [41, 80, 46, 36, 102, 74, 34, 37, 49, 57, 60, 75, 124, 357, 297, 319, 269, 312, 351, 275, 313, 280, 318, 279, 389, 448, 431, 434, 457, 382, 385, 407, 455, 433, 447, 408, 469, 391, 435, 381, 383, 430, 533, 562, 518, 552, 591, 653, 1010, 1007, 1066, 1030]. Editor-in-Chief [1066]. Editorial [326, 253, 673, 3, 114, 227, 364, 507, 613, 699, 795, 894, 1042]. Editors [988, 985, 105, 30, 73, 97, 106, 44, 107, 77, 45, 48, 35, 47, 104, 109, 122, 191, 304, 299, 294, 322, 298, 410, 456, 380, 387, 428, 429, 458, 453, 384, 388, 517, 534, 588, 660, 656, 655, 730, 607, 1028, 1027, 1023]. Edmonton [518]. Eds [761, 882, 880, 851, 815, 875, 814, 812, 883, 884, 891, 858, 881, 823, 879, 935, 926, 931, 937, 1096, 1100, 1057, 1062, 1063, 1116, 1103, 1110, 1059, 1083, 1113, 1087]. Education [38, 30, 36, 450]. effect [618, 703, 263, 639]. effective [354, 462, 327]. effects [600, 649, 651, 268, 753, 463, 776]. Efficiency [603, 283, 520, 855]. Efficient [116, 175, 168, 978, 648, 547]. EH01420 [60]. EH0161 [241]. EH0161-0 [241]. EHV [265]. eigenelements [1017]. eigenfunction [687, 586]. eigenvalue [814, 735, 577]. Einführung [33]. Eisenstein [518]. Ekeland [937]. elasticity [543]. electric [235, 1076]. electric-analog [235]. Electrical [401, 556, 871, 866]. electrically [865]. electricity [902]. Electromagnetic [212, 863, 774, 1076]. electron [604]. electronic [149, 499, 715]. electronics [389, 391, 871]. Electrotechniek [608]. element [1001, 98, 313, 446, 498, 874, 1097, 1100, 1085, 664, 26, 483, 235, 9, 1079, 71, 543, 800, 514, 619, 949, 1078]. Elementary [478, 150]. elements [74, 436, 654, 1096, 7, 636, 774, 12]. Elisabethenstrasse [355]. Elliot [815]. Ellipic [1122]. elliptic [98, 766, 933, 8, 1122, 72, 203, 175, 235, 174, 461, 786, 385]. Ellis [707, 716, 719, 739, 712, 818, 822, 878, 1082, 1102, 1114, 1115]. ELLPACK [1137, 1149, 766]. Elsevier [995, 991, 108, 193, 962]. elsewhere [401, 403, 1095]. Elzas [81]. EMB [54]. Embodiment [901]. EMC [212, 212]. EMC-22 [212]. Encarnacao [382, 468]. enclosure [420]. enclosures [419]. Encyclopaedia [970, 970]. encyclopedia [816, 1066]. endocrine [330]. Energy [477, 200, 280, 397, 398, 525, 120]. Engineering [46, 217, 219, 271, 762, 1060, 987, 986, 105, 38, 185, 186, 243, 316, 273, 295, 381, 401, 552, 588, 711, 821, 878, 955, 1097, 1112, 1130, 208, 984, 76, 743, 873, 1009, 1093, 46, 533]. Engineers [401, 991, 496, 553, 1114]. engines [202]. England [279, 454, 478, 499, 501, 473, 500, 474, 476, 477, 559, 593, 556, 533, 595, 592, 552, 658, 271]. Englewood [180, 710]. English [898]. Enriques [1141]. Enschede [608]. entrained [490]. entrained-flow [490]. entropy [838, 948]. environment [914]. Environmental [477, 265, 902, 908, 264, 262, 254, 455]. environments [273]. environmetrics [914]. enzyme [817]. epidemic [345]. epitaxial [544]. Epping [474, 476]. equation [689, 445, 10, 290, 21, 238, 71, 365, 1123, 140, 1080, 587, 525, 1097]. Equations [219, 728, 991, 304, 660, 730, 877, 933, 1089, 8, 575, 1044, 567, 426, 139, 26, 1122, 116, 72, 203, 232, 235, 136, 1137, 1149, 1079, 484, 665, 950, 174, 461, 616, 700, 1121, 134, 1003, 4, 369, 1000, 1076, 494, 1081, 440, 839, 336, 22, 200, 201, 415, 167, 786, 576, 987, 129, 195, 451, 656, 744, 731, 762, 712, 811, 885, 846, 1088, 1055, 1065]. equilibrium [995, 1077, 648]. Equivalence [332, 947, 975]. ERA [221]. ergodic [1058, 1052]. Erratum [131]. error [1020, 1018, 323, 664, 365, 144, 1021, 1016, 151, 501, 557, 729]. error-correcting [323]. Error-free [729]. errors [951, 617, 913]. Eschenfelder [270]. Escher [812]. Essential [872]. establish [915]. estimate [261, 144]. estimates [1051, 136, 365, 341, 577]. Estimation [789, 570, 102, 32, 157, 181, 694, 342, 999, 616, 443, 27]. estimator [27]. Euler [849, 947, 421]. Euroconf [964]. EUROGRAPHICS [653, 382, 857]. European [980, 109, 857, 1029, 961, 1027, 966, 967, 940, 942]. evaluating [1138, 28, 260, 1017]. Evaluation [124, 685, 49, 1008, 735, 196, 354]. even [292]. event [1051, 978, 1074]. Events [721, 782, 830, 802, 944, 1148, 377, 1038]. Evolution [948, 267]. Evolutionary [407]. excitation [907]. executing [725]. exhaustive [789]. exhibition [857, 533]. exit [574]. expansions [686, 586]. Experiences [1120]. experiment [192, 571]. Experimental [778, 196]. experiments [1113, 343]. Expert [1139, 366, 1119, 1150, 1120, 1134, 1122, 1137, 1149, 1129, 873, 964]. experts [1135]. Explicit [796, 1081, 786, 118]. Exploiting [635]. exponentials [675]. extended [174]. extensions [70]. Exton [535]. extraction [347]. extrapolated [414, 203]. extremal [1109].

F [985, 73, 106, 100, 160, 152, 151, 221, 354, 277, 280, 468, 383, 430, 559, 593, 558, 596, 592, 534, 654, 743, 740, 764, 766, 820, 824, 935, 1028, 1031, 1008, 1096, 1107, 1089, 1067, 1108, 1109, 992]. F-54011 [221]. F-78150 [152]. Fachtagungen [44, 34]. facilities [752]. factorisation [143]. factorization [174]. Fairweather [931]. Fak [591]. Fall [53, 207]. Farber [183]. farms [922]. Farrar [129]. Fast [232, 751, 683, 21, 141, 162]. Fatunla [1055]. fault [966, 967, 311]. fault-tolerant [967]. FCT [29]. FCT-Conference [29]. Feature [347]. February [38, 39, 226, 363, 506, 612, 698, 794, 945, 1041]. Fed [968, 955, 1028, 1006, 1026, 963, 1012, 1024, 169]. Feedback [293, 667, 1056]. FEoLL [250, 297]. FEoLL-workshop [297]. Ferrante [466]. Ferrari [37, 196, 124]. Ferziger [316]. FF [152]. FFT [1132]. Fiacco [469]. fibers [391, 865]. field [258, 89]. fields [1008, 845, 431]. Fifth [433]. File [468]. Filippov [1089]. filter [146, 257, 690]. Filtering [67, 410, 147, 1006]. filters [368]. Finite [426, 9, 71, 1001, 98, 74, 446, 436, 664, 26, 483, 7, 235, 617, 1079, 774, 12, 543, 800, 94, 514, 949, 654, 874, 932, 1085]. Finite-dimensional [932]. Firmware [214, 214]. First [1150, 358, 10, 601, 441, 828, 1027]. first-order [10]. Fischbein [982]. fisheries [488, 77]. fit [637]. fitting [339, 462]. fixed [602, 1022]. fixed-point [1022]. Fl [38].

Flannery [934]. flat [776]. flat-plate [776]. Fletcher [927, 1062, 1063]. Floating [1018, 1021]. Floating-point [1018, 1021]. flood [641]. Florence [470]. flow [693, 1096, 1075, 600, 490, 119, 751, 65, 665, 121, 508, 463, 202, 12, 439, 257, 550, 421, 738, 171, 525, 514, 949, 1078, 558]. flows [790, 632, 23]. fluctuations [1113]. flue [489]. fluid [73, 502, 941, 1096, 1063, 600, 869, 231, 550, 776]. fluidized [492, 493]. fluidized-bed [493]. flux [439]. fluxes [266]. Follow [309]. Follow-up [309]. following [238]. Fontainebleau [152]. foolproof [664]. forage [922]. forces [863]. forecasting [844, 641]. Forest [533, 644, 645]. Foreword [1043, 1015]. Form [219, 90]. formal [901, 126]. formation [152, 911]. formulas [1073, 1092]. formulated [509]. formulation [762, 330]. FORSIM [87]. Fortran [33, 33]. Fortran-programmpaketes [33]. forward [399]. Fosdick [49]. fossil [262]. Foundation [851]. Foundations [52, 206, 714, 1044, 778]. Fourier [162, 28, 1049, 975, 679]. Fourth [217, 290, 542]. fourth-order [290]. Fowler [929]. fractal [1087]. Fractals [779]. fractions [274]. fracture [847]. framework [867]. France [217, 152, 221, 832, 1004]. Francisco [39, 73, 77, 74, 78, 75]. Fraser [156]. free [65, 729]. French [532]. frequency [515, 1056]. fresh [263]. Freudenberg [1056]. Friedr [248]. fronts [1145]. fruit [647]. fuels [262]. führungsstil [852]. function [1094, 682, 681, 836, 1050, 948, 307, 773, 228, 896]. Functional [780, 406, 716, 883, 846, 530, 537, 961]. functions [321, 535, 339, 838, 992, 689, 680, 684, 688, 677, 369, 685, 1000, 674, 627, 1033, 993]. Fundamental [356]. fundamentals [1097, 29, 269, 708, 713, 1084]. Fundy [513]. Furskov [1032]. Future [851, 427]. Fuzzy [473].

G [990, 99, 106, 98, 81, 57, 59, 109, 125, 158, 154, 187, 129, 188, 191, 183, 299, 314, 359, 294, 269, 358, 300, 320, 457, 453, 435, 432, 558, 537, 596, 595, 562, 592, 554, 657, 730, 770, 745, 882, 888, 880, 850, 878, 858, 877, 890, 938, 1006, 926, 931, 1012, 924, 1092, 1096, 1097, 1063, 1116, 1064, 1103, 1110, 1067, 1106, 1065, 1066, 337]. G.B [105, 277, 316, 276]. G.’F [100]. Gaithersburg [55]. Galerkin [177, 136, 866, 975]. Gallitano [296]. games [983]. GAMM [1037]. Garda [124]. Gardens [159]. Gardone [37]. Garrett [317]. gas [489, 508]. gasification [490]. Gaussian [11]. GB [164, 163]. Gelenbe [47]. Gellert [816]. general [807, 567, 648, 583, 910, 171, 666]. general-purpose [583]. Generalized [1052, 1033, 688, 582, 174, 685]. Generating [234]. generation [433, 1126, 664, 289, 1049, 978, 1125, 776]. generator [775, 172, 1123, 11, 633]. Geneva [351]. geomechanics [517, 518]. geometric [497, 1146]. geometry [932, 203, 574]. Geothermal [1093]. Ger [1096]. Gerasoulis [656]. Germain [242]. German [44, 852, 1114, 848]. Germany [297, 968, 955, 1028, 1006, 1026, 963, 1012, 1024, 1107, 1108, 1109]. Gertsbakh [1095]. Getreu [149]. GI [34, 44]. Gi-Fachtagungen [44]. Gibson [48, 299]. Gipson [1097]. Gladwell [850]. Global [910, 334, 358, 1102, 262, 583]. Glowinski [352]. Godunov [601]. Gohberg [1007, 1057]. Goldberg [31, 1025]. Goldstein [1090]. Golub [592]. Gower [477]. GPSS [33, 521]. GPSS-Fortran [33]. grain [921]. Gram [689]. graphic [120]. graphics [857, 1134, 749]. graphs [318, 608]. Grauert [1092]. Gray [1058]. Great [916]. Greece [109, 218, 966, 967]. Green [448]. Greenblat [983]. Greengard [1008]. greenhouse [545]. Greenleaf [106]. Greenspan [984, 1009, 1059]. grid [8, 784, 1131, 741]. Griffel [1082]. Groesen [1027]. Grone [935]. Groping [358]. Gross [355]. groundwater [258, 121, 257]. Group [888, 875, 617, 666, 97, 122, 200, 786]. groupe [221]. growth [1113, 919, 920, 20, 267]. Grujic [1026]. GU2 [279]. Guest [673]. guide [1133, 1146]. Guilford [279]. Gulf [464]. Guo [1060].

H

[988, 985, 30, 73, 77, 78, 59, 160, 31, 162, 191, 185, 186, 182, 317, 315, 270, 278, 316, 298, 273, 456, 454, 434, 479, 411, 559, 593, 590, 596, 594, 592, 535, 521, 620, 779, 859, 887, 812, 881, 860, 853, 929, 1010, 934, 937, 939, 1116, 1103, 1110, 1067, 1082, 1102, 1113, 1087, 989, 836]. H-function [836]. H.-O [779]. Haar [965, 369, 1000]. Hackbusch [741]. Hadamard [680, 141, 615]. Hadamard-ordered [680]. Hagen [653]. Haifa [106]. Hairer [877]. Haji [743]. Haji-Sheikh [743]. Halifax [298]. Hall [164, 163, 180, 317, 302, 311, 315, 300, 239, 246, 247, 249, 301, 303, 499, 501, 500, 654, 710, 930]. Hamiltonian [937, 1024]. Hampshire [533]. hand [1089]. handbook [128, 1099, 1112, 1088]. Handbooks [847]. handling [294]. Hansen [318]. Harbor [211]. Hard [592, 1115]. Hardback [932, 1113]. hardbound [981]. Hardware [214, 429, 564, 761, 315]. Hardware/software [315]. Harker [556]. Harper [182]. Harrogate [279]. Hartenstein [429]. Haslinger [1085]. HAUPT [597, 479, 561, 521, 759, 769, 852, 848]. Hawgood [407]. Hayashi [817]. Hazewinkel [1027]. head [587]. head-on [587]. health [36, 399]. heat [618, 638, 71, 140, 776, 619]. Heidelberg [33, 29, 34, 160, 161, 162, 741, 849, 968, 955, 1006, 936, 963, 1012]. held [206, 209, 218, 207, 210, 211, 351, 470, 471, 472, 381, 743, 890]. Helderman [411]. helicopter [531]. Helmholtz [21]. Helson [985]. Helton [1057]. Hemisphere [743]. Hemisphere/Springer [743]. Henkin [1092]. hepatic [347]. Herman [106]. Herts [271, 401, 403]. Hewitt [73]. Heyden [128, 159]. Hibbard [99, 125]. Hice [100]. Hierarchical [27]. hierarchies [131, 1]. hierarchy [396, 400]. high [604, 11, 748]. high-voltage [748]. Higher [167, 849, 26]. Highland [59]. Hilbert [170, 1045]. Hilfe [33]. Hilfinger [356]. Hill [156, 190]. Hillview [159]. history [849, 971]. Hitchin [80, 271]. Hoes [206, 209, 207, 210, 211, 244, 245]. Hoie [519]. Hoigne [31]. holistic [425]. Holland [988, 41, 101, 42, 43, 96, 97, 99, 106, 107, 98, 102, 103, 81, 37, 47, 49, 104, 100, 109, 213, 31, 122, 123, 125, 217, 218, 151, 196, 126, 214, 150, 149, 184, 148, 124, 216, 127, 215, 250, 297, 314, 319, 240, 294, 296, 272, 320, 322, 352, 298, 312, 351, 273, 295, 318, 410, 452, 456, 387, 389, 428, 429, 431, 434, 458, 382, 407, 455, 433, 468, 450, 453, 408, 390, 451, 391, 435, 383, 409, 386, 406, 497, 562]. Holland [534, 519, 588, 591, 620, 653, 621, 657, 622, 655, 758, 761, 770, 767, 765, 768, 760, 757, 847, 887, 889, 857, 873, 851, 814, 812, 885, 891, 858, 881, 860, 890, 935, 957, 965, 941, 958, 962, 1062, 1063, 1059, 607]. Holler [44, 45]. holography [755]. holomorphy [537]. Holst [79]. homogenizing [600]. Hopcroft [155, 555]. Hopkins [453]. Hopper [851]. Hopscotch [290, 292, 416]. Hordijk [657]. horizon [650]. Horwood [707, 716, 719, 739, 712, 818, 822, 878, 1082, 1102, 1114]. hourly [233]. House [159, 279, 401, 403]. Howard [194]. HP [411]. HP-41C [411]. HP-41C/CV [411]. Huang [743]. Huens [104]. Hughes [607]. Hull [59]. human [295, 851, 868, 767]. human-computer [851]. Hungary [407]. Hunke [273]. Huyalorn [558]. Huybrechts [104]. Hybrid [871, 177, 378, 379, 284, 509, 92, 169, 704, 513, 88, 176, 4, 138, 1080]. hybride [516]. hydraulic [485]. Hydraulics [466]. hydro [92]. hydro-mechanical [92]. hydrological [255]. hydromechanics [1032]. Hydrothermal [261]. hyperbolic [10, 232, 136, 542, 587, 22, 200, 201, 441]. hyperfunctions [1143]. hypergeometric [535].

I.E.E.A [221]. I/O [317]. Iazelooa [657]. ICPC’78 [124]. ICPCI [37]. ideas [328]. identifiability [332, 335, 333, 338, 337, 336, 334, 495]. identifiable [789]. Identification [291, 142, 449, 682, 951, 330, 342, 486, 257, 337, 192, 102, 103, 193]. IEE [105, 243, 271, 401]. IEEE [38, 39, 40, 51, 52, 53, 54, 55, 56, 45, 50, 57, 59, 60, 206, 208, 209, 207, 210, 211, 244, 245, 241, 470, 471, 472, 208, 212]. IFAC [41, 102, 828, 1150]. IFIP [41, 97, 99, 106, 31, 122, 125, 214, 298, 456, 407, 409, 591, 887, 851, 860, 564, 860]. IFIP-TC [887]. IFIP/IFAC [41]. II [74, 186, 957, 967, 1096, 1112, 864, 490, 379, 229]. III [100, 1096, 533, 654, 889]. ill [642]. ill-defined [642]. Illingworth [73]. IMACS [988, 109, 198, 215, 304, 660, 656, 891, 832, 1150, 662, 671, 695, 721, 782, 720, 732, 746, 771, 781, 791, 829, 802, 831, 801, 827, 861, 892, 1038, 1037, 1035, 1034, 943, 972, 996, 1013, 1148, 1071, 1118]. IMACS- [1034, 1071]. IMACS-GAMM [1037]. IMACS/IFAC [1150]. Image [692, 536, 694]. images [693, 1087]. imaginary [805]. Impact [240, 870, 262]. Imperial [533]. Implementation [583, 446, 378, 94, 547]. implementations [725]. implicit [232]. implicitness [514]. improved [289, 756]. Improvement [66]. inaccurately [897]. inclusion [700]. incompressible [1075, 119]. Index [79, 111, 112, 224, 225, 361, 362, 504, 610, 696, 792, 860, 1039, 841]. indicators [1131].

Indonesia [634]. induced [910, 576]. Induction [272]. industrial [103, 173, 300, 815, 875]. Industry [218, 115, 1027]. inequalities [980, 452, 1005]. inequations [352]. inference [768, 156]. infiltration [205]. informatics [152]. informatik [624]. Information [1070, 1056, 1061, 1068, 35, 456, 407, 888, 1028, 1026, 1024, 834, 536, 36, 596, 621, 623, 1067]. Information-based [1067]. initial [1055, 837]. injection [333]. Input [913, 449, 682]. INRIA [152]. INRIA-service [152]. instability [911]. installations [124, 37, 124]. Institute [852, 848, 1093]. Institution [401]. instruction [1116]. Instrument [191]. instrumentation [128]. Int [458]. intangibles [398]. integer [826, 159]. integral [195, 656, 1092, 567, 689, 23, 743]. Integrated [902, 1036, 578]. integration [870, 116, 524, 542, 805, 685, 931]. intelligence [758, 964]. Intelligent [1150]. intensity [420]. Interact [767]. interaction [767, 295]. interactions [423]. Interactive [45, 1124, 45]. Interactive/visual [1124]. interface [1125, 1129]. interfacing [128]. interlocking [143]. internal [202, 776]. International [41, 30, 53, 56, 46, 29, 45, 37, 57, 60, 164, 208, 209, 217, 218, 207, 163, 211, 317, 245, 311, 315, 300, 239, 246, 351, 247, 249, 301, 303, 499, 457, 501, 500, 381, 533, 885, 890, 1150, 985, 107, 124, 517, 518, 564, 873, 1037, 1035, 1036]. interpolation [1044]. Intersciences [323]. intertemporal [647]. intervals [229]. Introduction [986, 895, 566, 631, 833, 394, 164, 150, 277, 353, 590, 626, 744, 326, 253, 154, 155, 187, 189, 248, 323, 475, 555, 658, 745, 1143]. Intuition [982]. Invariance [482]. invariant [798, 443]. Inventory [734]. inverse [1139, 331, 796, 850]. Inversion [727]. inverter [424]. invertible [483]. investigation [333, 906]. involving [93, 450]. ionosphere [528]. IPC [194, 279]. Irons [878]. ISBN [872, 990, 988, 983, 995, 982, 979, 980, 984, 991, 994, 985, 987, 986, 981, 41, 101, 80, 105, 42, 43, 73, 79, 96, 97, 99, 106, 44, 107, 46, 29, 77, 98, 102, 103, 48, 74, 81, 78, 34, 37, 47, 104, 95, 75, 100, 108, 109, 160, 164, 213, 31, 122, 123, 125, 128, 153, 161, 217, 218, 152, 157, 158, 151, 196, 126, 162, 214, 150, 154, 155, 156, 193, 129, 149, 148, 124, 163, 185, 186, 194, 216, 127, 215, 195, 317, 357, 243]. ISBN [244, 245, 250, 297, 314, 354, 359, 240, 294, 269, 311, 356, 358, 315, 248, 272, 277, 323, 353, 239, 246, 270, 320, 322, 278, 316, 352, 298, 312, 351, 247, 249, 275, 313, 321, 273, 276, 295, 318, 355, 279, 402, 410, 446, 452, 456, 454, 496, 380, 387, 389, 404, 428, 429, 448, 449, 466, 478, 499, 431, 434, 457, 458, 382, 385, 407, 455, 501, 433, 468, 450, 475, 495, 405, 447, 453, 408, 436, 469, 500, 384, 388, 390, 451]. ISBN [502, 435, 381, 383, 430, 474, 498, 409, 476, 386, 401, 403, 406, 411, 467, 477, 497, 536, 559, 560, 593, 553, 561, 558, 533, 557, 537, 555, 590, 596, 594, 595, 562, 592, 517, 518, 534, 519, 520, 535, 521, 552, 554, 588, 589, 591, 659, 620, 625, 626, 627, 653, 654, 621, 623, 658, 657, 622, 624, 655, 707, 758, 759, 779, 716, 717, 718, 719, 761, 706, 710, 730, 770, 715, 729, 708, 713, 714, 778, 767, 743, 744, 745, 709]. ISBN [739, 741, 765, 768, 777, 711, 731, 780, 760, 740, 757, 762, 764, 766, 769, 712, 763, 854, 855, 859, 882, 888, 886, 818, 820, 822, 847, 880, 887, 889, 817, 824, 857, 873, 851, 874, 849, 815, 850, 875, 814, 812, 813, 819, 825, 852, 883, 810, 811, 821, 878, 884, 885, 891, 856, 858, 881, 848, 860, 823, 846, 853, 877, 879, 826, 876, 816, 890, 938, 968, 1029, 1030, 928, 935, 955, 957, 960, 971, 1028, 964, 970, 932, 961, 1027]. ISBN [965, 925, 1006, 893, 923, 1026, 929, 953, 954, 959, 1005, 1010, 1023, 1031, 1032, 926, 930, 931, 933, 934, 936, 941, 958, 1009, 1007, 927, 937, 963, 1004, 956, 1008, 1012, 924, 939, 962, 966, 967, 969, 1011, 940, 942, 1025, 1033, 1024, 1086, 1092, 1104, 1140, 1142, 1070, 1096, 1097, 1100, 1057, 1062, 1063, 1069, 1094, 1099, 1107, 1112, 1116, 1146, 1147, 1064, 1089, 1103, 1110, 1111, 1145, 1141, 1056, 1084, 1085, 1093, 1088, 1067, 1143, 1082, 1108, 1101, 1106, 1055, 1059]. ISBN [1060, 1061, 1065, 1083, 1102, 1091, 1105, 1113, 1053, 1054, 1068, 1087, 1090, 1095, 1098, 1109, 1114, 1115, 1144, 989, 993, 992, 742, 607, 608, 1117]. ISBN- [970]. ISBN-0-12-506995-2 [1117]. Iserles [879]. Ishiguro [888]. Island [634]. ISNM [730, 883, 926, 1083]. Isomorphic [461]. isoparametric [483]. Isotopic [349]. isotropic [43]. ISSN [470, 471, 391]. issue [208, 212]. Itô [576]. Italy [45, 37, 124, 215, 470, 890, 828]. iterations [824]. Iterative [24, 138, 1018, 414, 724, 947, 88, 461, 4, 925]. ITO [1044]. IV [304, 436]. IX [1096, 1084, 1010, 936, 1009, 1007, 1012].

J [872, 73, 106, 77, 48, 74, 81, 32, 57, 104, 59, 75, 158, 151, 162, 155, 184, 163, 185, 186, 194, 274, 359, 296, 358, 316, 352, 312, 301, 355, 279, 410, 446, 452, 456, 380, 466, 478, 382, 407, 468, 436, 388, 390, 451, 383, 409, 556, 555, 596, 517, 534, 588, 653, 654, 657, 761, 766, 882, 875, 814, 811, 884, 891, 823, 846, 853, 879, 876, 1010, 1032, 937, 939, 1024, 1086, 1092, 1070, 1100, 1057, 1062, 1099, 1112, 1116, 1111, 1056]. J [1085, 1067, 1101, 1083, 1102, 1091, 1053, 1054, 1098, 607]. Jacobs [78]. Jacques [930]. James [58, 178, 954]. Jan [297, 456]. János [1059]. Janssen [184]. January [250]. Japan [389, 391, 564]. Jarny [588]. Javor [431]. Jindalee [650]. John [872, 358, 323, 316, 321, 1116, 1103, 1110, 1101, 1114]. Johnson [928]. Jolla [280]. Jones [274]. Jr [924, 1115]. Judd [930]. July [218, 304, 832, 19, 132, 1072]. June [41, 37, 124, 304, 298, 458, 518, 282, 413, 541, 733, 946].

Kaikan [564]. Kalaba [449]. Kalman [146, 257, 1006]. Kandel [473]. Kappel [1028]. Kapur [1101]. Karlsruhe [44]. Karpinski [29]. Kaucher [880]. Keast [931]. Keidanren [564]. Keramidas [432]. Khalil [853]. Kiechl [852]. Kim [559, 593, 881]. kinetic [938]. kinetics [404, 20]. Kingdom [707, 716, 717, 719, 706, 713, 714, 709, 712, 878, 529]. Kinnmark [762]. Kirchgraber [1103, 1110]. Kishinev [829]. Kitagawa [319, 389, 888]. Kleinrechnern [44]. Klir [81]. Kluwer [888, 875, 1030, 970, 1032, 956, 1025, 1033, 1089, 1111, 1093, 1105, 1113]. Knodel [275]. Knowledge [1126, 1132, 873, 966, 967, 1127, 1131]. Knowledge-based [1132, 966, 967, 1127, 1131]. known [1049]. Kogan [924]. Kohonen [161]. Kokotovic [853]. Kolmogorov [510]. Komkov [986]. Koomen [761]. kooperativen [852]. Kopcke [763]. Korte [408]. Krall [854]. Kranakis [742]. Krause [468]. kreativität [848]. Krishnamurthy [729, 818]. Kronsjö [855]. Krylov [933]. Kuhn [1096]. Kulisch [880]. Kunisch [1028]. Kutta [167]. Kwong [476]. Kylstra [383].

L [988, 994, 73, 49, 59, 100, 160, 129, 184, 163, 182, 131, 300, 301, 452, 404, 478, 434, 382, 447, 384, 386, 536, 537, 855, 850, 884, 885, 1030, 925, 1026, 926, 1008, 939, 1057, 1116, 1064, 1068, 136, 144]. L.-S [1068]. Laarhoven [939]. Laboratory [432, 347]. Laguerre [688]. Laine [1023]. Lake [124, 351, 915]. Lakshmanan [477]. Lakshmikantham [987, 1060]. Lamé [1077]. Lamit [300]. land [788]. land-breezes [788]. Landau [402, 1010]. landuse [912]. Lane [206, 209, 207, 210, 211, 244, 245, 750].

Langley [1011]. language [549, 372, 5]. Languages [41, 851, 155, 429, 564, 761, 427, 901, 725, 303, 409, 880]. Laplace [177]. Laplace-modified [177]. Large [1046, 95, 109, 127, 93, 669, 338, 912, 308, 542, 645, 284, 736, 648, 288, 27, 814, 1026]. Large-scale [1046, 109, 288]. Larson [185, 186]. late [753]. Laurent [893]. law [686]. Lawrynowicz [1111]. layer [665]. layers [921]. LC [241]. Leadership [852]. learning [285]. Lecture [469, 762, 968, 1070, 1056, 1061, 1068, 1028, 1026, 1024, 1109]. Lee [988, 311]. leg [649]. Leinbach [1116]. Leiterer [1092]. leukemia [348]. level [768]. Lewi [104]. Lewin [594]. Lewis [450]. Lexington [158]. Liège [381]. Lientz [353]. life [739]. lightning [806]. LILLIPUT [532]. Limaye [883]. lime [489]. lime/limestone [489]. limestone [489]. limit [174]. limits [39, 542, 805]. linéaire [242]. Linear [578, 370, 985, 156, 740, 712, 826, 137, 998, 333, 669, 338, 232, 724, 584, 174, 1000, 975, 440, 443, 841, 334, 576, 595, 709, 1082]. lines [265, 10, 138]. Linkens [105]. Linkers [80]. Linz [214]. Lions [352, 452]. Lipovski [57]. Lipow [127]. liquid [65]. Lisbon [857]. Lisp [1069, 872]. literature [79]. Littlewood [923]. Litwin [107, 458]. live [856]. Livelocks [476]. LMFBR [511]. LNM [936]. load [515]. load- [515]. loading [908]. Loan [592]. local [655, 28, 405, 453]. location [1077]. Lock [73]. Lodge [533, 552]. Lodi [1115]. logic [1099, 677, 990, 42, 447, 594]. logical [101, 239]. London [58, 73, 79, 77, 74, 78, 75, 128, 159, 357, 243, 359, 496, 380, 466, 385, 405, 384, 388, 502, 430, 498, 536, 533, 554, 853, 930]. loosely [64]. Looze [1056]. lost [1108]. Loveland [101]. low [25]. low-cost [25]. LSI [5]. Ltd [80, 105, 159, 243, 358, 277, 323, 271, 316, 321, 276, 279, 454, 403, 477, 1082, 1102]. Lubinsky [1109]. Ludena [30].

Luenberger [154, 187]. Lunteren [964]. Lyapunov [1081]. Lyubich [1106].

M

[988, 980, 985, 97, 29, 103, 48, 81, 47, 104, 76, 31, 122, 158, 196, 191, 192, 127, 357, 299, 296, 356, 272, 320, 454, 429, 385, 384, 381, 590, 594, 534, 739, 760, 854, 882, 888, 815, 850, 812, 819, 878, 856, 879, 938, 1030, 928, 1027, 925, 1026, 1005, 1023, 1031, 1032, 939, 1092, 1104, 1094, 1107, 1084, 1083, 1058, 1105, 1066, 905]. M-QAM [905]. M. [812]. M.I.T [157]. MA [872, 157, 155, 274, 560, 557, 555, 596, 594, 589, 874, 1009, 1008, 1011]. MAA [1116]. Macfarlane [298]. Machine [41, 295, 404, 668, 169, 322]. Machines [221, 871]. Macht [852]. Maciejowski [179]. Mackerle [1100]. MACKSIM [87]. macroelements [774]. Madison [954]. MADP [670]. Madrid [885]. MAFELAP [74]. magnetic [1076, 270]. magnetostatic [286]. Maintenance [216]. making [184, 706]. Man [322]. Man-machine [322]. Management [218, 852, 848, 124, 240, 296, 381, 258, 896, 645, 546, 204, 309, 922, 455]. Managing [320]. Mangasarian [384]. Mangler [73]. Manipulation [287, 549, 1130]. Mansell [79]. manual [620]. manufacture [41]. many [417, 378, 379]. mapping [176]. Marcel [355, 402, 404, 469, 537, 876, 925, 1084, 1095]. March [39, 107, 31, 152, 743, 2, 113]. Marchuk [188]. marine [958, 310, 488]. Market [91]. Marketed [1116]. Markov [775]. Mars [276]. Martin [58]. Martinus [850, 932, 1023]. Martynyuk [1026]. Marz [34]. Mason [882]. mass [86, 638]. Masson [158, 242]. masts [907]. matching [422]. Mathematica [1059]. Mathematical [875, 1010, 937, 1116, 330, 1, 423, 299, 381, 859, 1096, 1111, 575, 1133, 992, 867, 268, 1150, 572, 530, 131, 970, 178, 469, 657, 819, 953, 959, 1023, 1032, 1101]. Mathematics [219, 131, 323, 469, 938, 968, 960, 961, 933, 924, 940, 942, 1060, 1053, 1054, 980, 984, 994, 987, 74, 76, 318, 408, 891, 877, 1029, 955, 971, 1027, 929, 1032, 1009, 1025, 1033, 1092, 1116, 1089, 1090, 1109, 1114, 982, 77, 626, 745, 816, 970, 1010, 1083, 979, 436, 739]. matrice [989]. matrices [727, 1114, 333, 947]. Matrinović [980]. Matrix [484, 729, 935, 1007, 796, 1017, 592, 1106]. matrix-valued [1007]. matrizen [1114]. Mattheij [1027]. MAXI [532]. MAXI-LILLIPUT [532]. maximum [524, 805]. May [99, 125, 214, 351, 279, 472, 518]. mbH [248]. McDonnell [73]. McGraw [156, 190]. McGraw-Hill [156, 190]. McKee [815]. McLaughlin [277]. McPherson [41]. Meadows [358]. mean [341, 841]. mean-square [841]. means [1094, 980, 1005]. Measurement [834, 128, 562]. measurements [756, 94]. measures [576]. Measuring [398]. Mech [1100]. mechanical [870, 92]. Mechanics [1097, 986, 552, 847, 1096, 1100, 1112, 607]. media [843]. medical [54, 244]. medicine [106]. Medium [918]. Medium-term [918]. meeting [109]. Meinardus [730, 926]. Melbourne [451]. membership [789]. Memorandum [432]. memorial [965, 1147]. memories [161]. Mende [856]. mesh [664, 201]. Meta [768]. Meta-level [768]. metabolic [330]. meteorological [617]. method [1019, 976, 952, 1020, 98, 520, 1044, 64, 445, 10, 292, 870, 442, 72, 143, 168, 236, 643, 1079, 999, 616, 586, 866, 145, 548, 670, 439, 4, 138, 975, 648, 800, 283, 440, 117, 23, 619, 1078]. methode [516, 63]. methodology [104, 60, 758, 373, 258, 100, 81]. Methods [217, 478, 607, 1018, 987, 73, 76, 304, 316, 502, 498, 558, 518, 660, 627, 658, 743, 739, 741, 711, 847, 821, 878, 853, 1031, 926, 927, 1004, 1037, 1055, 1102, 1144, 567, 66, 601, 669, 203, 232, 724, 947, 118, 9, 896, 524, 542, 805, 906, 461, 796, 1131, 376, 134, 138, 1021, 494, 465, 1017, 415, 167, 786, 577, 313, 446, 406, 822, 1059, 129, 777, 883]. metric [824]. Meyer [384]. MI [57, 211]. Miami [59, 210]. Micchelli [32, 883]. Michel [95]. Mickle [191]. micro [277, 632]. micro-simulation [632]. microclimate [545]. Microcomputers [50, 646, 50]. microprocessor [239, 683]. Microprocessors [562]. Microprogramming [214]. micros [450]. Middleton [31]. Miller [95]. Minc [989]. Mini [964, 277]. Mini-Euroconf [964]. Minimization [808]. minimizing [753]. minimum [784]. Minoux [819]. Mirkowska [990]. miscellany [923]. Mises [1075]. Mitrinović [1005]. mixing [790]. modal [402]. Modcir [485]. Model [681, 300, 513, 917, 751, 638, 920, 902, 489, 424, 637, 397, 632, 750, 268, 775, 133, 172, 493, 121, 645, 546, 735, 348, 204, 309, 868, 572, 738, 910, 491, 1080, 267, 532, 488, 530, 752, 330, 856]. Modeling [346, 367, 806, 350, 163, 131, 243, 249, 301, 355, 954, 1023, 1091, 65, 257, 135, 1, 394, 789, 797, 120, 979, 108, 149, 181, 190, 191]. modelled [897]. Modellesystem [479]. Modelling [875, 544, 515, 919, 266, 649, 651, 643, 865, 787, 905, 650, 80, 105, 81, 179, 359, 358, 388, 430, 403, 556, 828, 823, 829, 1105, 575, 600, 869, 330, 804, 900, 896, 605, 774, 1045, 913, 573, 97, 122, 299, 296, 381, 625, 820, 875, 958, 1101, 48, 188, 271, 765]. Models [477, 529, 704, 154, 187, 178, 185, 186, 295, 495, 536, 517, 706, 739, 69, 339, 898, 702, 335, 345, 464, 255, 442, 912, 911, 327, 749, 168, 642, 571, 341, 899, 648, 922, 734, 773, 844, 585, 334, 1048, 641, 212, 43, 156, 278, 435, 1113, 184]. Modern [408]. modified [177, 548, 670]. modular [867]. modulated [578]. moduli [1086]. Mohammed [884]. Mohapatra [454]. Mohler [1061]. moisture [921]. Moller [764]. moment [422]. moments [1049, 341, 1010]. Monds [277]. Monica [51]. monitoring [31, 31]. monoparametric [724]. montagefertigungen [759]. Monte [510, 790, 908, 422, 440, 839, 117, 825]. Montreal [198]. Moore [380, 1099]. Morph [462]. Morph-fitting [462]. Morris [354, 478]. MOS [424]. MOS-transistor [424]. Moscardini [388]. motion [574]. Moto [433, 761]. Moto-Oka [433, 761]. Mountain [50]. Moura [534]. movements [752]. MPSD [548]. multi [683, 750, 784, 171, 741]. multi-dimensional [784, 171]. Multi-grid [741]. multi-lane [750]. multi-microprocessor [683]. multiarea [515]. multibody [870]. multicodimensional [1025]. multicompartmental [342]. Multidimensional [585, 845, 378, 379]. multifluid [604]. Multigrid [376, 175]. Multiple [1074, 42, 239, 1031]. multiple-microprocessor [239]. multiple-valued [42]. multiprocessor [693, 1003, 797, 474]. Multiprocessors. [246]. multirate [684, 118, 500]. multivariable [1056, 1068]. Multivariate [813, 954]. Multsim [750]. Munich [860]. Murty [595]. muscle [865, 530]. Myint [991]. Myint-U [991].

N [30, 102, 103, 95, 59, 271, 454, 475, 453, 474, 517, 655, 847, 817, 878, 933, 1033, 1101, 1113, 1025]. N.S.W. [912, 918]. Naffah [453]. Nancy [221]. Naples [464]. NASECODE [1036]. Nash [403]. national [184, 887]. NATO [1023, 937, 1093, 1105]. nature [377]. Naval [432]. Navier [26]. navigational [25]. needs [78]. Neelamkavil [820]. neighborhood [748]. Neittaanmäki [1085]. Nelson [127]. Nemirovsky [520]. nerve [865]. Net [249, 163, 301]. Netherlands [517, 518, 888, 875, 938, 1029, 1030, 957, 960, 964, 970, 932, 961, 1027, 965, 953, 954, 959, 1005, 1023, 1031, 1032, 931, 933, 958, 937, 956, 924, 939, 962, 966, 967, 969, 940, 942, 1025, 1033, 608]. Network [329, 448, 405, 582, 808, 606, 551, 623]. networking [55]. Networks [53, 302, 58, 536, 889, 785, 285, 1136, 571, 314, 247, 453, 655, 825, 890]. Neuhold [126]. neural [1136]. New-York [29, 1054]. Newbury [983]. Newcastle [875]. Newman [851]. News [662, 671, 695, 720, 732, 746, 771, 781, 791, 801, 827, 861, 892, 943, 972, 996, 1013, 1118]. newspaper [753]. Newtonian [550]. next [1125]. Nicholson [271]. Nicoletti [837]. Nielsen [59]. Nijhoff [850, 932, 1023]. Nijkamp [477]. Nikiforov [992]. nitrogen [423]. NJ [206, 209, 207, 210, 211, 244, 245, 660, 656]. No [208, 212, 221, 318, 401, 432, 159, 321]. noise [186, 775, 576]. Noll [932]. nombres [1090]. non [43, 242, 137, 174, 550, 370, 789]. non-isotropic [43]. non-linear [137, 174, 370]. non-Newtonian [550]. non-scalar [137]. non-uniquely [789]. nonaliased [784]. nonconservative [1144]. Nondifferential [390]. nongaussian [570]. Nonlinear [842, 254, 807, 709, 731, 417, 838, 338, 24, 1047, 65, 367, 1079, 950, 1003, 4, 307, 787, 117, 334, 530, 809, 384, 534, 780, 810, 933, 1060, 1061]. Nonnegative [989]. nonpoint [260]. Nonstandard [936]. nonstationary [1002]. nonstiff [877]. norms [1106]. Norsett [877]. North [988, 41, 101, 42, 43, 96, 97, 99, 106, 107, 98, 102, 103, 81, 37, 47, 49, 104, 100, 109, 213, 31, 122, 123, 125, 217, 218, 151, 196, 126, 214, 150, 149, 184, 148, 124, 216, 127, 215, 250, 297, 314, 319, 240, 294, 296, 272, 320, 322, 352, 298, 312, 351, 273, 295, 318, 410, 452, 456, 387, 389, 428, 429, 431, 434, 458, 382, 407, 455, 433, 468, 450, 453, 408, 390, 451, 391, 435, 383, 409, 386, 406, 497, 519]. North [588, 591, 620, 653, 621, 657, 622, 655, 758, 761, 770, 767, 765, 768, 760, 757, 847, 887, 889, 857, 873, 851, 814, 812, 885, 891, 858, 881, 860, 890, 935, 957, 965, 941, 958, 962, 1062, 1063, 1059, 607, 910]. North-Holland [988, 41, 101, 42, 43, 96, 97, 99, 106, 107, 98, 102, 103, 81, 37, 47, 49, 104, 100, 109, 213, 31, 122, 123, 125, 217, 218, 151, 196, 126, 214, 150, 149, 184, 148, 124, 216, 127, 215, 250, 297, 314, 319, 240, 294, 296, 272, 320, 322, 352, 298, 312, 351, 273, 295, 318, 410, 452, 456, 387, 389, 428, 429, 431, 434, 458, 382, 407, 455, 433, 468, 450, 453, 408, 390, 451, 391, 435, 383, 409, 386, 406, 497, 519]. North-Holland [588, 591, 620, 653, 621, 657, 622, 655, 758, 761, 770, 767, 765, 768, 760, 757, 847, 887, 889, 857, 873, 851, 814, 812, 885, 891, 858, 881, 860, 890, 935, 957, 965, 941, 958, 962, 1062, 1063, 1059, 607]. Northeast [875]. northern [918]. Nostrand [354, 821, 816]. note [134]. notebook [1108]. Notes [469, 762, 968, 1070, 1116, 1056, 1060, 1061, 1068, 1028, 1026, 1024, 1109]. Nova [298]. November [54, 56, 60, 208, 212, 244, 997]. Novosibirsk [99, 35, 125]. Noye [451, 1062]. Nü [926]. nuclear [262, 288]. number [43, 289, 1132]. numbers [1074, 603]. Numerical [987, 170, 478, 885, 464, 1127, 863, 1119, 1150, 1022, 776, 49, 75, 129, 269, 409, 626, 658, 744, 822, 815, 811, 879, 1037, 1036, 1065, 1114, 544, 1126, 528, 66, 119, 870, 703, 442, 365, 1124, 1123, 545, 1121, 439, 494, 1129, 1130, 201, 415, 984, 78, 148, 278, 316, 352, 388, 451, 502, 517, 518, 656, 711, 731, 821, 878, 884, 885, 891, 926, 930, 931, 934, 941, 958, 1009, 1055, 1059, 1083]. Numérique [221]. Nürnberger [730]. Nussbaumer [162]. NW4 [159]. NY [206, 245].

O [185, 186, 317, 384, 388, 779, 924, 1103, 1110, 1055, 1087]. Observation [998, 570, 179, 146, 333]. observer [94]. observer-based [94]. observers [367]. obstructions [600, 869]. O’Carroll [48]. ocean [908, 1080]. October [51, 52, 109, 206, 209, 829, 306, 460, 581, 672, 772, 862]. odd [292]. odd-even [292]. ODE [1073]. Oden [654]. ODEs [1120]. off [1018, 1021, 1016]. offshore [807, 864]. Ogata [180]. oil [909]. Oka [433, 761]. Okandan [1093]. Okla [129]. Oleinick [474]. One [524, 542, 378, 379, 1079, 805, 439, 543, 440, 171, 415]. one- [171]. one-dimensional [1079, 439]. one-step [805, 415]. only [1049]. open [1080]. operating [590]. operation [237]. operational [435, 785, 312]. operations [408, 751, 809]. operator [985, 1057, 1094]. operators [985, 740]. Optical [35, 35, 391, 622]. Optimal [682, 342, 602, 512, 986, 1004, 1085, 1091, 343, 308, 737, 144, 92, 841, 32, 192]. Optimality [603]. optimisation [694, 403]. Optimization [69, 977, 339, 1019, 1020, 188, 278, 408, 825, 960, 1035, 417, 702, 87, 255, 799, 173, 546, 547, 449, 390, 520, 717, 823, 927, 1064, 1102, 298, 435]. optimization-simulation [546]. optimizer [583]. optimizing [346]. Optimum [778, 414, 63]. options [398]. oracle [715]. order [159, 414, 10, 26, 601, 290, 136, 21, 542, 738, 167, 441, 933]. ordered [660, 656, 680]. Ordering [351]. ordinary [811, 877, 1055, 1065, 575, 116, 839, 415]. ore [639]. O’Reilly [853]. Oren [81]. organization [487, 594]. organized [152]. oriented [269]. Orlando [38]. Orleans [456]. Ortez [885]. Orthogonal [7, 686, 968, 678, 687, 784, 12, 679, 674, 441, 1007]. orthogonalization [689]. orthonormalized [675]. oscillations [842, 442]. oscillators [93]. Oslo [891]. Ostrowsky [1113]. OT [1007, 1106]. other [985, 1108]. our [386]. outbreaks [916]. outcome [526]. output [371]. over-relaxation [786]. over-the-horizon [650]. overrelaxation [947, 70, 236]. overview [260, 329, 418, 585]. Oxford [457, 592, 882, 815, 884, 879, 971, 929, 1066]. oxidation [491].

P [980, 981, 79, 99, 98, 81, 31, 125, 152, 317, 354, 311, 356, 353, 278, 298, 312, 276, 318, 387, 428, 448, 458, 453, 474, 403, 477, 558, 653, 657, 759, 779, 780, 764, 848, 853, 877, 1005, 931, 934, 937, 939, 1011, 1024, 1070, 1056, 1085, 1059, 1083, 1091, 1053, 1054, 993, 608, 416]. P.D.E. [378, 379]. P.O [271, 280, 279, 401, 403]. Pacific [881]. package [620]. packages [87]. Padé [893]. Paderborn [250, 297]. Pages [85, 19, 2, 62, 199, 132, 113, 166, 252, 325, 226, 282, 306, 393, 438, 481, 363, 413, 460, 523, 565, 599, 506, 541, 581, 630, 663, 612, 672, 722, 747, 783, 698, 733, 772, 803, 794, 862, 1014, 945, 946, 997, 973, 1041, 1072, 872, 990, 988, 983, 995, 982, 979, 980, 984, 991, 994, 985, 987, 986, 981, 41, 101, 80, 105, 30, 38, 39, 40, 42, 43, 51, 52, 53, 54, 55, 73, 79, 96, 97, 99, 106, 46, 29, 77, 98, 36, 102, 50, 74, 81, 78]. pages [32, 35, 34, 37, 47, 49, 57, 104, 95, 59, 60, 75, 100, 76, 164, 213, 31, 122, 123, 125, 128, 153, 161, 209, 217, 218, 152, 157, 158, 207, 210, 151, 196, 126, 162, 214, 150, 154, 155, 156, 193, 149, 148, 124, 163, 185, 186, 194, 211, 216, 127, 215, 195, 317, 357, 243, 244, 245, 250, 274, 297, 299, 302, 314, 319, 354, 359, 240, 294, 269, 311, 356, 358, 315, 248, 272, 277, 300, 323, 353, 239, 246, 270, 271]. pages [320, 322, 278, 316, 352, 298, 312, 351, 247, 249, 275, 301, 303, 313, 321, 273, 276, 280, 295, 318, 355, 241, 279, 402, 410, 446, 452, 456, 454, 496, 380, 387, 389, 404, 428, 429, 448, 449, 466, 470, 471, 472, 478, 499, 431, 434, 457, 458, 479, 382, 385, 407, 455, 501, 433, 468, 473, 450, 495, 405, 447, 453, 408, 436, 469, 500, 384, 388, 390, 451, 502, 435, 381, 383, 430, 474, 498, 409, 476, 386, 401, 403, 406, 411]. pages [467, 477, 497, 432, 536, 559, 593, 553, 556, 558, 533, 537, 590, 596, 595, 562, 517, 518, 534, 519, 520, 535, 552, 554, 588, 591, 659, 660, 620, 625, 626, 627, 653, 654, 621, 623, 658, 657, 656, 622, 624, 655, 707, 758, 779, 716, 717, 718, 719, 761, 706, 710, 730, 770, 715, 729, 708, 713, 714, 778, 743, 744, 745, 709, 741, 765, 768, 777, 711, 731, 780, 760, 740, 757, 762, 764, 766, 769, 712, 763, 854, 855, 859, 882, 888]. pages [886, 818, 820, 822, 847, 880, 887, 889, 817, 824, 857, 873, 851, 874, 849, 815, 850, 875, 814, 812, 813, 819, 825, 852, 883, 810, 811, 821, 878, 884, 885, 891, 856, 858, 881, 848, 860, 823, 846, 853, 877, 879, 826, 876, 816, 890, 938, 968, 928, 935, 955, 957, 960, 971, 964, 970, 932, 961, 965, 925, 1006, 893, 923, 929, 953, 954, 959, 1005, 1010, 926, 930, 931, 933, 934, 936, 941, 958, 1009, 1007, 927, 937, 963, 1004, 956]. pages [1008, 1012, 924, 939, 962, 966, 967, 969, 1011, 940, 942, 1086, 1062, 1063, 1069, 1107, 1085, 1082, 1108, 1106, 1059, 1083, 1054, 1087, 1109, 989, 993, 992, 742, 607, 608, 220, 187, 189, 178, 179, 184, 188, 190, 192, 183, 219, 180, 182, 242, 304, 1084]. Pai [883]. Pain [396]. Palo [40]. Pande [517]. Pannonian [953, 959]. Papalambros [1091]. paper [716, 717, 711, 878, 1140, 1113, 1115, 115]. paperback [981, 932, 1069]. Papers [213, 885, 84, 401, 891, 829, 832, 1107, 1108]. parabolic [933, 290, 24, 614, 176, 174, 134, 4, 167]. Parallel [57, 211, 476, 143, 1073, 658, 64, 998, 1136, 1137, 1149, 796, 726, 1137, 1149, 57, 211, 474]. Parallele [275]. parallelization [1138]. Parameter [87, 336, 102, 109, 181, 457, 1028, 67, 702, 682, 951, 416, 482, 978, 89, 342, 616, 341, 94, 288, 512]. Parameters [999, 417, 292, 142, 735, 147, 789]. parametric [173]. parametrized [731]. Paris [107, 242, 832, 1004, 1090]. Park [983, 50]. Parker [1064]. parking [752].

Parma [890]. Part [212, 1105, 807, 864, 104, 70]. partial [129, 304, 451, 660, 885, 1065, 8, 1044, 426, 290, 1122, 203, 235, 1137, 1149, 616, 1121, 1003, 1000, 336, 786, 991]. Particle [701, 1008, 491, 573, 615]. particles [909, 572]. particular [713, 689]. Partioned [147]. partitioning [726]. Parts [280]. Pascal [725]. passage [268]. patch [9]. paths [978]. Patras [109, 218]. Pattern [473, 799, 1084, 1113]. patterns [911, 644]. Pau [184]. Paul [479, 561, 759, 769, 852, 848]. Pavel [1084]. Pavella [1026]. Payne [743]. PDE [177, 1128, 176, 1129]. PDEs [1126]. Pearcy [985]. Pearson [821]. pedestrian [632]. peers [981]. Pegase [604]. Pegden [189]. Peitgen [779, 1087]. penetration [91]. Pentech [46, 48, 299, 313, 496, 388, 498, 553]. perceived [754]. Perception [344]. percepts [536]. Peregrinus [80, 105, 243, 277, 271, 276, 403]. Performance [146, 904, 97, 122, 196, 124, 354, 657, 765, 1051, 841, 117, 379, 37, 47, 49, 383, 519]. Pergamon [457, 1066]. periodic [24, 937]. perspective [617]. perspectives [1099]. perturbation [853, 785, 1004]. Perturbations [469, 1047, 1026]. perturbed [1046, 461, 288, 512]. Peschel [856]. Peter [80, 105, 243, 277, 271, 321, 276, 403]. Peterson [163, 301]. Petkoff [758]. Petri [582, 163, 249, 301]. Petrov [866]. Petrushev [993]. Pey [1112]. Pezman [778]. pH [133, 605]. pH-control [605]. pH-process [133]. pharmacokinetic [343, 341]. phase [170, 869, 172, 171, 489]. phenomena [1144]. phenomenas [432]. Philadelphia [128, 813]. Phillips [822]. Phoenix [472]. phosphate [463]. Physical [120, 842, 937, 1111, 914, 575, 328, 608]. Physically [1105]. Physically-based [1105]. physico [133, 605]. physico-chemical [133, 605]. physics [1044, 859, 992]. physiologic [343]. piecewise [674]. Pilot [605]. Pinder [558]. Pingree [50]. Piscataway [206, 209, 207, 210, 211, 244, 245]. Pittsburg [191]. plain [898]. planning [1031, 397, 645, 399, 527]. plant [605, 204, 309, 532]. plantation [645]. plants [902, 193]. plate [1001, 776]. platforms [649]. Plato [929]. Plenum [30, 36, 32, 35, 448, 449]. Pless [323]. plus [860]. Plymouth [46, 48, 553]. PMICE [373]. point [1018, 963, 345, 1021, 1022]. points [798, 1132]. Poiseuille [550]. Poisson [1097]. policies [887]. policy [397, 398, 735, 399]. Polish [990]. Polishchuk [1094]. polluted [263]. pollution [178, 261, 260]. polygonal [419]. polynomial [729, 679]. polynomials [1007, 1109, 229, 968]. Polytechnics [875]. Popov [1086, 993]. Poppelbaum [276]. populations [268]. Port [245]. Porter [1116]. Posch [6]. Posnan [29]. possible [1021]. postage [152]. Postbus [608]. Postfach [248, 355, 402, 404]. potential [1008, 1096, 1075, 7, 286]. Potier [765]. pour [63]. Powell [879]. power [556, 485, 515, 604, 902, 871, 532, 27, 963]. powers [836]. pp [1029, 1030, 1028, 1027, 1026, 1023, 1031, 1032, 1025, 1033, 1024, 1092, 1104, 1140, 1142, 1070, 1096, 1097, 1100, 1094, 1099, 1112, 1116, 1146, 1147, 1064, 1089, 1103, 1110, 1111, 1117, 1145, 1141, 1056, 1093, 1088, 1067, 1143, 1101, 1055, 1060, 1061, 1065, 1102, 1091, 1105, 1113, 1053, 1068, 1090, 1095, 1098, 1114, 1115, 1144]. pp. [1066]. Practical [1017, 995, 128, 1146, 342, 194, 927]. practice [48, 216]. Prager [981]. precipitation [910]. precision [11]. preconditioned [669, 72, 145, 548]. Preconditioning [670]. predator [856].

Prediction [400, 348]. preferences [395]. Prekopa [823]. Preliminary [258]. Prentice [164, 163, 180, 317, 302, 311, 315, 300, 239, 246, 247, 249, 301, 303, 499, 501, 500, 654, 710]. Prentice-Hall [164, 163, 180, 317, 302, 311, 315, 300, 239, 246, 247, 249, 301, 303, 499, 501, 500, 654, 710]. preparation [296]. presented [885]. Presenting [898]. Press [979, 987, 73, 46, 77, 48, 74, 78, 95, 75, 153, 220, 157, 158, 194, 219, 357, 299, 359, 313, 279, 446, 496, 380, 457, 385, 405, 447, 436, 384, 388, 502, 430, 498, 536, 553, 558, 592, 554, 718, 744, 745, 740, 882, 815, 884, 853, 879, 971, 923, 929, 934, 1008, 1011, 1104, 1140, 1142, 1069, 1099, 1146, 1147, 1064, 1117, 1145, 1088, 1067, 1055, 1060, 1065, 1091, 1105, 1066, 1098, 1115, 1144, 993]. Press/Wiley [744]. pressurized [532]. prey [856]. Primality [742]. primitive [1080]. principle [805, 286]. principles [986, 960, 303, 1091]. Printing [129]. Priorities [395, 131, 1]. Pritsker [189]. prize [1034, 1071]. Probabilistic [577, 91, 185, 186]. probability [953, 959, 1095, 838, 948, 1058]. Problem [726, 994, 380, 385, 618, 333, 24, 668, 173, 331, 342, 12, 587, 520]. problems [1139, 1020, 73, 766, 850, 1032, 1094, 1055, 1060, 332, 25, 378, 379, 175, 837, 569, 737, 796, 376, 337, 509, 619, 98, 534, 552, 814, 877, 1030, 960]. Proc [964, 953, 954, 959, 966, 967, 988]. procedure [88]. procedures [290, 444]. Proceedings [41, 30, 38, 39, 40, 51, 53, 54, 55, 56, 99, 106, 107, 46, 29, 45, 102, 50, 35, 37, 57, 59, 60, 109, 31, 125, 209, 217, 218, 152, 207, 126, 214, 191, 124, 211, 216, 215, 304, 244, 245, 250, 297, 298, 351, 456, 470, 471, 472, 458, 407, 451, 381, 409, 533, 517, 518, 743, 887, 857, 873, 851, 875, 881, 860, 985, 279, 1150]. process [256, 489, 396, 668, 173, 142, 344, 526, 652, 400, 133]. processes [103, 536, 938, 1058, 67, 399, 394, 787, 527, 835, 185]. Processing [35, 57, 211, 302, 38, 157, 298, 562, 621, 1117, 1137, 1149, 636, 675, 677, 141, 58, 35, 57, 500, 1061]. processor [726, 158]. processors [796, 276]. production [103, 529]. products [836]. Prognostic [488]. program [714, 171, 633]. programmentwicklung [160]. programmes [242]. Programmiersprachen [34]. programmierung [411]. Programming [41, 469, 104, 126, 303, 384, 409, 386, 880, 1115, 373, 549, 647, 307, 370, 159, 272, 318, 447, 595, 709, 819, 826, 1140]. programmpaketes [33]. programs [499, 620]. progress [1092, 1090, 7, 313, 381, 430, 498]. project [127]. Propagation [22, 201, 1139, 1018, 988, 513, 604, 843, 800]. properties [1056, 333, 754, 22, 1058]. property [776]. proportional [639]. Proposal [790]. protection [264]. Protocol [760, 858]. protocols [448, 247]. prototypical [1134]. Providence [1010]. proving [101]. prozesse [275, 764]. Prozessrechnentechnik [183]. Prüfungsauggaben [624]. Przeworska [1029]. Przeworska-Rolewicz [1029]. PSD [72]. Pseudo [10]. Pseudo-characteristic [10]. Public [58]. Publication [206, 209, 244, 245, 470, 471, 472, 401]. Publications [152, 207, 210, 211, 533, 1097, 1096, 1100, 1112]. Published [152, 191, 304]. Publisher [1116]. Publishers [990, 981, 517, 518, 888, 875, 1030, 970, 1023, 1032, 956, 1025, 1033, 1089, 1111, 1093, 1105, 1113]. Publishing [990, 983, 982, 980, 984, 991, 994, 986, 41, 101, 30, 42, 43, 79, 96, 97, 99, 106, 107, 98, 36, 102, 103, 32, 35, 37, 47, 49, 104, 100, 108, 109, 213, 31, 122, 123, 125, 217, 218, 151, 196, 126, 214, 150, 193, 149, 184, 148, 124, 216, 127, 215, 250, 297, 314, 319, 240, 294, 296, 272, 320, 322, 352, 298, 312, 351, 273, 295, 318, 410, 452, 456, 387, 389, 428, 429, 448, 449, 431, 434, 458, 382, 407, 455, 433, 450]. Publishing [453, 408, 390, 451, 391, 435, 383, 474, 409, 476, 386, 401, 403, 406, 477, 1029, 1027, 1031, 1143]. Puig [1112]. Puig-Pey [1112]. pulp [115]. Pulsating [550]. pulse [684]. Purdue [1150]. Pure [469, 1095]. purpose [1098, 583, 171]. purposes [755]. PWN [990]. pyrolysis [909].

Q [416, 535]. q-hypergeometric [535]. QAM [905]. quadrant [143]. quadratic [483, 219]. Qualitative [95]. Quality [848, 99, 125, 642, 896, 546, 263, 96, 123, 190, 848]. quarter [860]. quasi [64]. quasi-parallel [64]. Queens [451]. queueing [1002, 708]. queuing [314, 825, 785, 571]. Quick [396]. Quinney [744]. quotients [836].

R [872, 988, 995, 73, 46, 48, 74, 95, 59, 60, 75, 31, 150, 304, 299, 315, 277, 352, 303, 321, 410, 454, 380, 387, 428, 429, 449, 458, 450, 436, 384, 390, 386, 477, 560, 553, 533, 557, 517, 534, 521, 552, 660, 656, 743, 777, 757, 766, 886, 814, 825, 852, 891, 928, 935, 1027, 941, 927, 940, 942, 1099, 1107, 1116, 1064, 1108, 1106, 1061, 1083, 1058, 1109, 607]. Rabinovich [925]. Rabinowitz [937]. radar [650]. radio [905]. Rainfall [641, 256, 912, 913]. rainfall-landuse-runoff [912]. rainfall-runoff [256, 913]. Rajbman [102, 103]. Ramamoorthy [60, 881]. Ramanujan [1108]. Ramm [195]. Randell [153]. Random [185, 186, 417, 603, 1058, 567, 845, 289, 1049, 843, 89, 1074, 837, 234, 283, 1113]. range [627]. rapid [1008]. Rapport [221]. Rardin [1064]. rate [463, 341]. rates [670]. Rational [993]. rationalized [1000]. Ratschek [1102]. Ravasio [453]. Raviv [106]. Rayna [1012]. RC [531]. reacting [508]. reaction [345, 736, 494]. reactive [790, 493]. reactor [544, 602, 288, 532]. reactors [485, 133, 493]. Reading [872, 155, 274, 560, 557, 555, 596, 594, 589, 874, 1009]. real [849, 993, 256]. real-time [256]. reality [127]. realization [690]. Rechnernetzen [44]. recipes [934]. Recognition [473, 1084]. recombination [700]. reconstruction [694]. Record [210]. Records [52]. recovery [317]. rectangular [420, 175]. recursion [272]. Recursive [321]. Redevelopment [647]. redistribution [869]. Reduce [1012]. Reduction [476, 444, 681, 374]. Redwood [984]. Reef [916]. reference [402, 1100]. refinement [8, 702, 201]. reflected [386]. Reformulation [951]. regenerative [903]. Regier [454]. region [647]. regional [1031, 546]. regions [524]. regression [756, 348, 578]. regressive [347]. regularly [693]. regulating [868]. Reid [409]. Reidel [990, 982, 980, 994, 986, 76, 562, 534, 778, 777, 780, 854, 859, 846, 938, 1029, 960, 964, 970, 961, 1027, 953, 954, 959, 1005, 1031, 931, 933, 937, 956, 924, 939, 966, 967, 969, 940, 942, 1143]. Reinhold [354, 821, 816]. Reiser [872, 97, 122]. Rektorys [76]. related [431, 1030, 1023, 937, 966, 967]. relations [23]. relationship [409]. Relationships [576]. relativistic [604]. relaxation [416, 737, 516, 786]. reliability [127, 710, 966, 967, 1095, 153, 320, 657]. relief [396]. remarks [5, 619]. Rep [968, 955, 1028, 1006, 1026, 963, 1012, 1024]. repair [668]. reply [6]. Report [58]. reported [1110, 1103]. Reports [1060]. representations [1126, 1051]. Requicha [857]. Research [432, 744, 103, 312, 435, 30, 408]. reserve [639, 120]. reservoir [1093]. reservoirs [121, 263]. residence [341]. residuals [445]. Resolution [705, 395]. resource [135]. resources [258, 254]. resp [957]. respective [268]. response [649, 908]. responses [868]. responsible [386]. Rest [621, 622]. Restructurable [214]. results [258, 444, 1021, 835]. retention [463]. Review [872, 990, 988, 983, 995, 982, 979, 980, 984, 991, 994, 985, 987, 986, 981, 41, 101, 80, 105, 30, 38, 39, 40, 42, 43, 51, 52, 53, 54, 55, 56, 58, 73, 79, 96, 97, 99, 106, 44, 107, 33, 46, 29, 77, 98, 36, 45, 102, 103, 48, 50, 74, 81, 78, 32, 35, 34, 37, 47, 49, 57, 104, 95, 59, 60, 75, 100, 108, 109, 76, 160, 164, 213, 31, 122, 123, 125, 128, 153, 161, 206, 208, 209, 217, 218, 220, 152, 157]. Review [158, 207, 210, 151, 196, 212, 126, 162, 214, 150, 154, 155, 156, 159, 187, 189, 193, 129, 149, 178, 179, 181, 184, 188, 190, 191, 148, 192, 124, 163, 183, 185, 186, 194, 211, 216, 219, 221, 127, 180, 215, 195, 182, 242, 304, 317, 357, 243, 244, 245, 250, 274, 297, 299, 302, 314, 319, 354, 359, 240, 294, 296, 269, 311, 356, 358, 315, 248, 272, 277, 300, 323, 353, 239, 246, 270, 271, 320, 322, 278, 316, 352, 298, 312]. Review [351, 247, 249, 275, 301, 303, 313, 321, 273, 276, 280, 295, 318, 355, 241, 279, 402, 410, 446, 452, 456, 454, 496, 380, 387, 389, 404, 428, 429, 448, 449, 466, 470, 471, 472, 478, 499, 431, 434, 457, 458, 479, 382, 385, 407, 455, 501, 433, 468, 473, 450, 475, 495, 405, 447, 453, 408, 436, 469, 500, 384, 388, 390, 451, 502, 391, 435, 381, 383, 430, 474, 498, 409, 476, 386, 401, 403, 406, 411, 467, 477, 497, 432, 536]. Review [559, 560, 593, 553, 561, 597, 556, 558, 533, 557, 537, 555, 590, 596, 594, 595, 562, 592, 517, 518, 534, 519, 520, 535, 521, 552, 554, 588, 589, 591, 659, 660, 620, 625, 626, 627, 653, 654, 621, 623, 658, 657, 656, 622, 624, 655, 707, 758, 759, 779, 716, 717, 718, 719, 761, 706, 710, 730, 770, 715, 729, 708, 713, 714, 778, 767, 743, 744, 745, 709, 739, 741, 765, 768, 777, 711, 731, 780, 760, 740, 757, 762, 764, 766, 769]. Review [712, 763, 854, 855, 859, 882, 888, 886, 818, 820, 822, 847, 880, 887, 889, 817, 824, 857, 873, 851, 874, 849, 815, 850, 875, 814, 812, 813, 819, 825, 852, 883, 810, 811, 821, 878, 884, 885, 891, 856, 858, 881, 848, 860, 823, 846, 853, 877, 879, 826, 876, 816, 890, 938, 968, 1029, 1030, 928, 935, 955, 957, 960, 971, 1028, 964, 970, 932, 961, 1027, 965, 925, 1006, 893, 923, 1026, 929, 953, 954, 959, 1005, 1010, 1023, 1031, 1032, 926]. Review [930, 931, 933, 934, 936, 941, 958, 1009, 1007, 927, 937, 963, 1004, 956, 1008, 1012, 924, 939, 962, 966, 967, 969, 1011, 940, 942, 1025, 1033, 1024, 1086, 1092, 1104, 1140, 1142, 1070, 1096, 1097, 1100, 1057, 1062, 1063, 1069, 1094, 1099, 1107, 1112, 1116, 1146, 1147, 1064, 1089, 1103, 1110, 1111, 1117, 1145, 1141, 1056, 1084, 1085, 1093, 1088, 1067, 1143, 1082, 1108, 1101, 1106, 1055, 1059, 1060, 1061, 1065, 1083, 1102, 1058, 1091, 1105, 1113, 1053, 1054, 1066, 1068, 1087, 1090]. Review [1095, 1098, 1109, 1114, 1115, 1144, 989, 993, 992, 742, 607, 608, 1016, 262, 494, 865, 579]. Reviews [13, 14, 15, 16, 17, 18]. revisited [643]. Rheinboldt [731]. Rheine [128]. Rhodes [966, 967]. rhythm [917]. Ribbens [1026]. Ribbens-Pavella [1026]. Ricciardi [1030]. Rice [766]. Richards [73]. Richardson [358]. Richter [779]. Riemann [994]. Riemannian [574]. Riesz [1107]. Riet [458]. right [1089]. rigid [551]. Rinaldi [190, 455]. Rine [42]. Ring [655]. rings [484]. Ritz [144]. river [190, 263]. riverland [647]. Riviera [37]. Rivlin [32]. roads [750]. Robert [824]. Roberts [475]. Robinson [384]. Robot [549, 956]. robots [287]. Rocky [50]. Rodin [994]. Rodman [1057]. Rokne [1102]. Role [335, 911, 749, 514]. Rolewicz [1029]. Romanovskii [938]. Rome [851]. root [920]. Rootmap [920]. Rosenberg [129]. Rossmannler [294]. Roth [354, 777, 886]. Rotterdam [517, 518]. round [1018, 1021, 1016]. round-off [1018, 1021, 1016]. Rouse [295]. Roush [559, 593]. route [444]. routing [285]. Royal [707]. Rózsa [1059]. Rubinstein [825]. rue [221, 381]. Runge [167]. runoff [256, 912, 913, 641]. Rüschlikon [873]. Rutgers [660, 656]. Ryan [404].

S [983, 980, 981, 43, 99, 102, 103, 48, 81, 35, 100, 109, 125, 157, 156, 159, 190, 304, 314, 294, 351, 355, 454, 380, 457, 455, 384, 435, 476, 558, 562, 534, 520, 589, 659, 660, 770, 777, 780, 886, 847, 889, 874, 815, 812, 877, 938, 1005, 934, 1033, 1097, 1099, 1056, 1108, 1055, 1083, 1068, 1109]. S.A [447]. S04 [533]. Saaty [131]. Sabelli [30]. Sackman [887]. Saddle [345]. safety [710]. Saff [1109]. Sage [983]. Saint [242]. Saint-Germain [242]. Sakamoto [888, 817]. Sakawa [1031]. Sales [206, 209, 207, 210, 211, 244, 245]. Salwicki [990]. SAM [216]. sample [510, 978, 837]. sampled [684]. sampled-data [684]. samples [422]. sampling [343, 342]. San [39, 73, 77, 74, 78, 75, 1069, 1064, 1145, 1067, 1055, 1060, 1065]. Santa [51]. SAOR [236]. Sarikaya [858]. SASSYS [511]. Sata [322]. satellite [904]. Saupe [1087]. Sawaragi [181]. scalar [1056, 137]. scale [95, 109, 814, 1026, 1046, 284, 288]. Schaft [1024]. Schappacher [1028]. schedule [343]. scheme [508]. schemes [414, 416]. Schendel [658, 1114]. Scheneider [456]. Schlesinger [1105]. Schmidt [33, 1023, 6, 689]. Schmidt-Laine [1023]. Schneider [434]. Schoemaker [889]. School [220]. Schrijver [826]. Schultz [988, 385]. Schuman [99, 125]. Schwefel [278]. Science [35, 206, 194, 219, 279, 660, 1093, 1067, 1055, 1060, 982, 984, 987, 38, 42, 52, 36, 185, 319, 294, 389, 743, 1009, 1087, 148, 356, 380, 812, 821, 878]. Sciences [217, 270, 656, 1070, 1056, 1061, 1068, 76, 1028, 1026, 1023, 937, 1024, 186, 588, 739]. Scientific [990, 108, 198, 193, 886, 891, 1067, 1055, 711, 880, 643, 1125, 962, 1011]. scientists [991, 1114]. score [773]. Scotia [298]. Scotland [41]. SE [208]. SE-6 [208]. sea [788, 910]. sea- [788]. search [417, 799, 283]. SEC’83 [591]. Second [381, 832, 933, 414, 601, 136, 738]. second-order [738]. Security [591]. Sedgewick [560]. sediment [915]. seismic [755]. Seiten [160, 561]. Selberherr [659]. Selected [813]. Selecting [637]. selection [891]. Self [900, 151, 1037, 1131]. Self-adaptive [900, 1131]. self-checking [151]. self-validating [1037]. semi [897, 796, 22, 200]. semi-direct [796]. semi-discretizations [22]. semi-variograms [897]. semiconductor [659, 1036, 700, 866]. semidiscretization [975]. Séminaire [1090]. Seminar [35, 218, 458]. Sendov [1086]. sensitivity [825, 339, 1051, 146, 1052, 773, 606]. Seo [1031]. Seoul [881]. Sept [470]. September [53, 46, 29, 45, 102, 207, 215, 407, 517, 62, 166, 973]. sequence [234, 1053, 1054]. sequences [137, 289]. Sequency [680]. Sequency- [680]. sequential [702, 343]. Serazzi [890]. Series [58, 323, 270, 271, 457, 469, 847, 960, 961, 1093, 1053, 1054, 980, 994, 105, 243, 302, 321, 318, 852, 848, 877, 938, 968, 1029, 893, 1023, 1032, 933, 937, 963, 940, 942, 1033, 1089, 1110, 1061, 1105, 1114, 687, 140, 679]. seriesed [347]. Service [39, 40, 51, 52, 53, 54, 55, 56, 45, 50, 57, 59, 60, 206, 209, 207, 210, 211, 244, 245, 152]. set [988, 957, 970, 1096, 789]. sets [179]. Sewell [1065]. Seydel [995]. SFR [479, 355, 402, 404, 469, 759, 852, 883, 848, 893, 926, 1007, 1106, 1083, 992]. SFr.156.00 [537]. SFr.28.00 [769]. SFr.28.00/DM34.00 [769]. SG1 [401, 403]. SG5 [271]. Shackel [767]. shale [909]. shallow [762]. shape [1085]. shaped [419]. Shapiro [355]. sharing [458]. Shaw [356]. Sheikh [743]. Shelah [43]. Shieh [1068]. ship [284]. shock [601]. shocks [23]. short [846]. shortest [444]. shot [477]. Shrive [878]. Shubert [185, 186]. SIAM [813]. sic [147]. sides [1089]. Signal [675, 157, 500, 562, 1117, 1061, 636, 141]. signals [186, 682, 683, 578]. Silver [768]. Similitude [108]. simmer [490]. simmer-II [490]. simple [442, 508, 204, 309]. Simplification [638]. Simplified [257, 532]. Simpson [63]. SIMUL [372]. SIMULA [248]. Simulated [205, 463, 649, 939, 969]. Simulating [915, 644, 526, 750]. Simulation [667, 650, 33, 59, 198, 221, 131, 280, 629, 875, 513, 93, 634, 640, 528, 285, 20, 909, 911, 372, 308, 906, 508, 840, 646, 736, 348, 788, 1076, 82, 903, 921, 797, 748, 420, 79, 48, 109, 189, 191, 250, 297, 431, 659, 770, 820, 889, 825, 1105, 366, 69, 977, 544, 515, 310, 635, 237, 64, 427, 139, 87, 845, 119, 68, 751, 464, 867, 492, 703, 346, 871, 378, 379, 790, 424, 235, 632, 749, 701, 168, 908, 367, 259, 425, 636, 173, 264, 377, 775, 133, 115, 487]. simulation [1074, 238, 755, 665, 121, 1124, 725, 809, 368, 950, 422, 545, 571, 546, 899, 419, 233, 329, 347, 439, 543, 254, 726, 5, 516, 865, 916, 551, 375, 773, 639, 423, 839, 465, 371, 787, 1130, 652, 905, 904, 633, 1048, 120, 441, 169, 835, 914, 475, 81, 359, 279, 430, 59, 109, 215, 280, 521, 588]. simulations [615, 983]. simulator [256, 582]. simulators [1126]. simultaneous [72, 145, 548]. sinc [689]. sinewaves [172]. Singapore [886, 1083]. Singh [780, 1066]. Single [491, 797]. singular [656, 1026, 841, 853]. Singularity [288]. singularly [1046, 461, 512]. Sinha [351]. sixth [21]. size [602]. skeletal [530]. skeleton [328]. SLAM [189]. slotted [66]. small [179, 482, 1047, 646]. Smalltalk [1125]. Smirnov [510]. Smith [1116]. smoothness [1086]. SO4 [552]. social [739, 398, 394]. socially [386]. Societatis [1059]. Society [53, 56, 207, 191, 280, 707, 1010]. Society/Ellis [707]. socio [135]. socio-economic-resource [135]. Soft [592, 762, 1061]. Software [46, 60, 208, 209, 216, 337, 56, 96, 99, 50, 78, 123, 125, 302, 320, 241, 533, 1012, 28, 1133, 343, 175, 1136, 1128, 1150, 375, 370, 46, 49, 125, 60, 127, 273, 386, 401, 597, 552, 315]. Sohn [248]. soils [463, 423]. Solid [270]. Solid-State [270]. Solution [728, 307, 1073, 170, 129, 656, 744, 1065, 567, 689, 10, 618, 177, 378, 379, 24, 143, 203, 232, 724, 365, 176, 174, 1121, 140, 4, 138, 369, 1000, 648, 1081, 509, 587, 949]. solutions [976, 451, 432, 937, 426, 1123, 837]. solver [21, 1129]. solvers [385]. Solving [139, 25, 669, 950, 461, 796, 134, 1003, 337, 440, 839, 167, 786, 766, 877]. Some [5, 619, 835, 195, 677, 494]. Son [128, 159]. Sons [358, 323, 316, 321, 1103, 1116]. SOR [947, 70]. Sorensen [410, 390]. Sorrento [215]. Sortimentsplanung [769]. sound [106, 419, 800]. Source [620, 260]. sources [205]. South [1100]. Southaampton [1100]. Southampton [46, 552, 1096, 1097, 1112]. Southgate [401, 403]. Soviet [994, 938, 960, 970, 1032, 933, 1033, 1089]. space [495, 845, 442, 89, 1045, 293, 147, 585, 334]. space-dependent [147]. spaces [932, 1094]. Spalti [769]. Spandoni [196]. sparse [1114, 174, 1114]. spatial [644, 94]. spatially [640, 142]. Spatio [1077]. Spatio-economic [1077]. Special [208, 212, 962, 985, 1098, 992]. Species [489]. Specification [467, 760, 858, 237]. spectral [694, 740]. Spectrum [159, 170, 804]. speed [233]. Spingarn [449]. splines [975]. Split [168]. Split-simulation [168]. splitting [439]. splittings [70]. sprache [160]. Spratt [655]. Spriet [359]. Spring [39]. Springer [44, 33, 29, 34, 160, 161, 162, 179, 181, 188, 192, 183, 195, 269, 270, 275, 495, 659, 625, 623, 624, 779, 729, 743, 741, 762, 764, 766, 763, 817, 824, 849, 810, 856, 823, 877, 968, 955, 1028, 1006, 1026, 936, 963, 1012, 1024, 1070, 1096, 1100, 1107, 1112, 1056, 1108, 1061, 1053, 1054, 1068, 1087, 1109]. Springer-Verlag [44, 33, 29, 34, 160, 161, 162, 179, 181, 188, 192, 183, 195, 269, 270, 275, 495, 1006, 1012, 1070, 1096, 1100, 1107, 1056, 1108, 1061, 1053, 1054, 1068, 1087, 1109]. Springs [211]. sprl [381]. square [841]. St [126]. Stability [704, 118, 568, 995, 1026, 482, 524, 542, 805, 134, 415]. stabilized [172]. Stable [4, 442]. Stanley [1113]. starfish [916]. STATCAT [620]. State [270, 570, 75, 495, 879, 146, 442, 486, 293, 443, 585, 488, 334, 374, 27]. state-space [585]. states [347, 348]. stationary [414, 724]. Statiques [221]. Statistical [707, 896, 117, 954, 1032, 642, 341, 355, 1095]. statistics [454, 707, 953, 959, 888, 876]. status [39]. stayed [907]. Steel [150, 296, 529]. Steele [77]. step [947, 524, 542, 805, 415]. Stephenson [279]. Stepleman [304, 660]. stepsize [1073]. Sternberg [988]. Stévard [381]. Stevenage [401, 403]. stiff [952, 139, 482, 442]. stimulated [865]. Stirling [41]. stirred [133]. Stochastic [1018, 1044, 340, 569, 573, 844, 807, 185, 452, 534, 69, 977, 575, 67, 838, 444, 1047, 1052, 1045, 1046, 571, 572, 574, 568, 948, 841, 839, 117, 1016, 1048, 1021, 276, 406, 554, 924]. Stockholm [887]. stocks [237]. Stocky [44, 45]. Stokes [26]. storage [924]. storm [703]. Strang [745]. Straszak [184]. Strategies [753, 1138]. strategy [998]. Strazicky [823]. Street [129]. stress [1096]. stringent [618]. Strong [1109]. structural [1026, 332, 333, 338, 1068]. structure [465]. structured [589, 1115, 536]. Structures [468, 807, 864, 356, 1111]. Structuring [912]. Studer [517]. Studies [744, 299, 318, 266, 490, 804, 168, 487, 840, 318]. study [127, 824, 258, 285, 604, 870, 379, 341, 916, 465, 606, 776, 904, 246]. studying [530]. Stuttgart [479, 880, 1103, 1110]. Subject [112, 225, 362]. subsurface [558]. subsystems [64]. successive [947, 70]. Suematsu [391]. sum [754]. summation [1021]. Supercomputing [962]. supermacros [427]. Supplementum [269]. supply [237]. support [964]. surface [266, 756, 65, 910]. surfaces [994, 1146, 1141, 665, 736, 497, 757]. surges [703]. Surrey [279]. switched [285]. switching [401]. Switzerland [355, 469, 561, 521, 769, 873, 1007, 1106, 597]. Sylvester [1081]. symbolic [194, 1130]. Symmetric [236, 143, 1017]. Symmetrical [691]. Symp [953, 954, 959, 988]. Symposium [40, 51, 52, 54, 50, 206, 217, 210, 471, 660, 656, 885, 881, 55, 107, 102, 152, 304, 244, 517, 564, 828, 1027, 1037, 1035]. synthesis [1132, 679, 841, 287]. Synthetische [411]. Syracuse [206]. System [198, 1051, 679, 97, 102, 122, 192, 296, 1120, 515, 310, 1134, 1133, 345, 333, 1122, 683, 372, 651, 978, 1137, 1149, 636, 614, 142, 487, 1131, 999, 735, 736, 868, 5, 606, 141, 797, 100, 180, 823, 966, 967]. Systemanalyse [764]. Systeme [33]. systemes [705, 516]. Systems [215, 457, 475, 1119, 260, 842, 952, 667, 1139, 704, 80, 105, 50, 81, 109, 153, 187, 179, 186, 243, 240, 456, 455, 758, 828, 873, 890, 829, 964, 1026, 958, 937, 1112, 1110, 366, 977, 417, 575, 682, 838, 1051, 998, 10, 490, 684, 330, 68, 549, 669, 870, 511, 338, 291, 143, 1127, 1136, 724, 1047, 308, 1150, 331, 642, 753, 774, 584, 1052, 1045, 1046, 950, 486, 646, 329, 254, 726, 1002, 293, 1076, 337, 568, 674, 551, 135, 948, 92, 1129, 443, 841, 288, 117, 147, 418, 512, 608]. systems [903, 905, 1048, 441, 328, 27, 302, 467, 47, 95, 154, 181, 188, 163, 315, 353, 239, 271, 249, 301, 280, 457, 458, 407, 433, 495, 401, 556, 590, 554, 817, 938, 1028, 964, 1008, 1024, 1142, 1056, 1068, 295, 403, 477, 589, 1066]. Szelezsan [823]. Szucs [108].

T [872, 991, 106, 48, 81, 32, 60, 153, 161, 127, 131, 319, 296, 311, 322, 276, 389, 433, 477, 519, 654, 761, 873, 1026, 934, 1069, 1068, 607]. tablelands [918]. Tagg [450]. Tanenbaum [302]. tank [133]. Tarnlund [447]. tasks [1135]. Tatzmannsdorf [953, 959]. Tauberian [1033]. Tbilisi [102]. TC [106, 31, 298, 407, 887]. TC-4 [106, 31, 298]. TC2 [409]. Teachers [220]. teaching [250, 297, 549]. Technical [831]. technique [528, 289, 462]. Techniques [473, 45, 354, 247, 496, 390, 553, 770, 765, 1062, 1098, 417, 255, 338, 444, 264, 681, 176, 174, 796, 551, 355, 241, 967]. technologies [319, 389, 405]. Technology [194, 279, 1066, 39, 36, 623, 655, 571, 294, 270]. telecommunications [401, 389, 391]. teleconnection [911]. temperature [917, 71, 868, 910]. Temperlmeier [521]. ten [653]. Tennent [303]. tension [756, 649]. tension-leg [649]. term [918, 700]. terminal [634, 602]. terms [608]. test [510, 284]. testing [760, 858, 605, 11, 334]. Teubner [880, 1103, 1110, 742]. Teukolsky [934]. Texas [743]. th [948]. Thayer [127]. their [980, 181, 321, 429, 564, 761, 855, 893, 1005, 1110, 1106, 1135, 268, 234, 327]. theorem [101]. theorems [1033, 415]. Theoretical [310, 378]. théorie [1090]. theories [739, 1117]. Theory [42, 187, 219, 818, 819, 939, 1066, 1139, 985, 986, 43, 154, 155, 163, 274, 323, 249, 301, 410, 387, 428, 457, 883, 953, 959, 969, 940, 942, 1092, 1142, 1057, 1094, 1085, 67, 235, 1077, 70, 418, 608, 789, 29, 32, 157, 321, 537, 813, 810, 935, 961, 925, 926, 1095, 987, 195, 712, 826, 1025]. therapiestudien [763]. therapy [31]. thermal [485, 142, 866, 921]. thermal-hydraulic [485]. thermodynamic [329]. thermometer [776]. thermoviscoelasticity [1079]. thin [1001]. Third [533, 542]. third-fourth [542]. thorns [916]. Three [70, 202, 485, 483, 172, 551, 615]. Three-dimensional [202, 485, 551, 615]. Three-part [70]. Thron [274]. THT [608].

Thurber [158]. tidal [513]. tightly [797]. time [131, 1061, 93, 67, 998, 256, 136, 493, 840, 341, 347, 1, 443, 841]. time-dependent [493]. time-invariant [443]. time-seriesed [347]. times [574]. tissue [349]. today [427]. Tokyo [564]. tolerance [311]. tolerant [967]. tomography [106]. Tool [220, 1138, 346]. Tools [41, 354, 456, 765]. top [164]. top-down [164]. Topics [962, 985, 813, 1142, 1097, 1023, 937]. topographic [911]. tori [798]. Tournantes [221]. traces [649]. traffic [751, 754, 750]. Transactions [208, 212, 891]. transfer [638]. Transform [162, 690, 693, 170, 940, 942, 28, 692, 683, 418, 615]. transform-based [683]. transformation [495, 845, 638, 341, 691, 350]. transformations [137, 678, 483, 1053, 1054]. transformer [906]. transforms [676]. transhipment [634]. Transient [259, 485, 618, 607, 619]. transistor [424, 149]. transit [903]. translated [1114]. translation [76, 970]. transmission [265]. transonic [421, 23]. transport [432, 445, 121]. Traub [596, 1067]. travail [221]. treatment [346]. treats [809]. Trefftz [144]. Tremolieres [352]. trends [935]. triangular [284]. Tricks [809]. Tridiagonal [974]. Triebel [859]. Trigiante [987]. Tropper [405]. truck [753]. Trzebiner [521]. Tsay [1068]. Tseng [743]. Tulsa [129]. tumor [20, 346]. tumorigenic [350]. Turbo [1115, 1140]. turbulent [266, 790, 907, 738]. Turner [100]. Turschmann [561]. Tutee [220]. Tutor [220]. tutorial [60, 241]. Tutorials [241]. Two [600, 869, 510, 351, 67, 640, 638, 20, 424, 483, 836, 866, 140, 690, 738, 171]. two-dimensional [638, 20, 690, 738]. Two-fluid [869]. two-phase [171]. two-sample [510]. TX [471]. Tyne [875]. type [1044, 70, 543, 839, 167]. types [228]. Tzafestas [109, 457, 435, 562, 770].

U [991, 658, 880, 849, 1103, 1110, 1114]. U.K [279, 457, 1089, 1093]. U.S. [878]. U.S.$24.50 [49]. U.S.A [356, 353, 472, 590, 596, 594, 589, 1010, 1009, 1008, 1011]. U.S.S.R. [829]. Uhlig [935]. Uhr [536]. UHV [265]. UK [194, 401, 403, 882, 886, 878, 884, 879, 1096, 1097, 1100, 1112, 1111, 1105, 1113]. UK£ [1143]. UKSC [279, 279]. Ullman [155, 555]. Ullrich [880]. Ultra [228, 229, 106]. Ultra-arithmetic [228, 229]. unbounded [843]. uncertainty [596, 643, 808, 734, 834]. unclogged [205]. unconstrained [1020]. undergraduate [1116]. Understanding [876]. Unified [694]. unilateral [870]. uniquely [789]. United [707, 716, 717, 719, 706, 713, 714, 709, 712, 878, 529]. Univ [988]. University [46, 109, 153, 218, 220, 451, 660, 656, 718, 743, 852, 848, 879, 923, 954, 934, 1147, 1091, 993, 1150, 250, 297]. unknown [951, 307, 147]. unknown-but-bounded [951]. unpublished [1108]. unsteady [26, 23]. unternehmung [769]. Unterstützung [44]. updated [970]. upon [875, 897]. upper [738]. upwind [508]. urban [915]. US$10.00 [44]. US$102.00 [620]. US$11.95 [923]. US$119 [1111]. US$12.00 [38]. US$13.50 [1011]. US$14.00 [60, 55]. US$146.25 [215]. US$149 [1143, 1105]. US$16.00 [50]. US$16.50 [150]. US$17.65 [77]. US$170.75 [102]. US$18.80 [195]. US$19.00 [33]. US$19.95 [983, 154, 1069]. US$20.00 [625, 39, 40]. US$20.75 [148]. US$20.95 [155]. US$21.50 [29, 1101]. US$21.95 [1114]. US$22.00 [151, 57, 51, 52]. US$22.50 [95, 157]. US$22.50. [535]. US$23.00 [158]. US$23.95 [1115]. US$24.00 [78]. US$24.95 [872, 185]. US$25.00 [96, 123, 1010, 53, 54, 241]. US$26.95 [186]. US$27.50 [979, 385]. US$29 [1113]. US$29.05 [239, 246]. US$29.50 [240, 380, 32]. US$29.95 [981, 1102]. US$30.40 [163, 249]. US$32.50 [1103]. US$33.70 [164]. US$33.75 [1008]. US$34.35 [73]. US$34.40 [623]. US$34.50 [156]. US$34.95 [1110, 1060]. US$35.00 [213]. US$35.00/Dfl.120.00 [768]. US$36.50 [196, 655]. US$37.50 [30, 35]. US$37.75 [100]. US$37.95 [1117]. US$39.00 [214]. US$39.50 [456, 450]. US$39.95 [1099, 1065]. US$39.95. [1142, 1146]. US$40.00 [127, 384]. US$40.45 [247]. US$40.50 [80, 105, 243]. US$41.00 [75]. US$41.50 [104]. US$42.50 [382, 1141, 36]. US$43.50 [149]. US$4350 [74]. US$44.50 [37, 124, 1055]. US$45.00/Dfl [860]. US$45.95. [1140]. US$46.10 [161]. US$46.25 [103, 653]. US$46.50 [458]. US$46.70 [160]. US$47.75 [876]. US$48.00 [59]. US$48.75 [101, 106, 107, 218, 250]. US$49.50 [846]. US$49.95 [987]. US$50.00 [245, 657]. US$51.50 [436]. US$52.00 [194]. US$53.75 [99, 31, 125]. US$55.00 [995]. US$58.50 [97, 109, 122, 216]. US$59.00 [562]. US$59.95. [1144]. US$61.00 [42]. US$61.60 [41]. US$63.00/Dfl.170.00 [770]. US$63.50 [98]. US$64.50 [1067]. US$65.00 [453]. US$65.75 [81, 108]. US$67.00 [76]. US$68.00 [985, 244]. US$68.25 [43, 217, 193]. US$68.50/Dfl [851]. US$69 [1104]. US$69.00 [877, 942]. US$69.50 [993]. US$69.50. [1147, 1145]. US$69.75 [383]. US$69.95 [780, 1064]. US$7.65 [1116]. US$70.00 [457]. US$70.75 [47]. US$72.25/Dfl [857]. US$73.25 [126]. US$75.00/Dfl [847]. US$76.00/Dfl. [858]. US$77.75 [885]. US$85.00 [622]. US$88.00 [607]. US$89 [1093]. US$89.00 [859, 1005, 931]. US$89.00/Dfl [924]. US$89.50 [854]. US$9.00 [34]. US$9.95 [981]. US$93.00 [455]. US$97.50/Dfl [932]. US$98.00 [621]. US$99 [1089]. US$99.50 [1057]. USA [129, 304, 351, 280, 456, 401, 403, 621, 622, 1095, 1150, 660, 656, 471]. usage [264]. use [1116, 8, 25, 1136, 327, 115, 1074, 174, 899, 619]. used [485]. user [620, 1129]. using [766, 1115, 513, 1075, 339, 285, 177, 143, 582, 235, 396, 372, 649, 636, 646, 736, 551, 94]. USSR [102, 35, 125]. Uvarov [992].

V [984, 987, 986, 30, 103, 60, 323, 469, 555, 591, 729, 760, 818, 825, 883, 858, 881, 938, 1032, 1009, 1025, 1033, 1086, 1099, 1060, 993, 992, 376, 660]. V-cycle [376]. Vahl [1063]. Vakgroep [608]. Vakhutinsky [925]. validating [1037, 749]. validation [335, 330]. validity [370]. value [1094, 1055, 24, 837, 569, 796, 509]. valued [42, 1007]. values [978]. Vansteenkiste [359]. Varga [70]. variable [1073, 977, 776]. variables [185]. Variance [444, 374, 897]. variant [1019]. variate [1049]. variates [836, 234]. variation [482, 918, 376]. Variational [286, 352, 452, 390, 669, 986, 76, 432, 1024, 1144]. variations [1070]. variograms [897]. varying [142]. Vasić [980, 1005]. Vasilescu [985]. Vasiliev [938]. vector [962, 635, 736, 551]. Vector-Network [551]. vectors [146]. vehicle [425]. vehicles [752]. vehicular [750]. velocity [617, 200, 666]. Venezuela [30]. Venice [828]. Vente [152]. Verification [490, 858, 549, 714, 760]. Verlag [44, 33, 29, 34, 160, 161, 162, 179, 181, 188, 192, 183, 195, 269, 270, 275, 402, 404, 479, 495, 411, 521, 759, 769, 1006, 1007, 1012, 1070, 1096, 1100, 1057, 1107, 1112, 1056, 1108, 1106, 1061, 1053, 1054, 1068, 1087, 1109]. Verlagsgesellschaft [248]. Versailles [217]. versus [340]. Vetterling [934]. VI [1107, 968, 1036, 858]. via [378, 379, 679, 789]. Vibration [807, 864, 875, 850]. Vich [940, 942]. Vichnevetsky [304, 660, 656, 891, 941, 1034, 1071]. view [963]. Vieweg [248]. Vignes [891]. VII [1109, 849, 955, 930]. VIII [1106, 814, 957, 963]. viscous [514]. Vishik [1032]. visual [1124]. visualization [1124]. Vladimorov [1033]. Vlaminck [104]. VNR [816]. Vogt [191]. Voiculescu [985]. Vol [208, 212, 186, 219, 469, 762, 968, 957, 970, 937, 962, 1092, 1104, 1096, 1097, 1057, 1094, 1112, 1103, 1110, 1093, 1082, 1060, 1083, 1090, 1095, 988, 987, 986, 185, 271, 877, 1028, 932, 1026, 953, 959, 1023, 926, 966, 967, 1024, 1059, 1054, 1109]. Vollert [479]. Vols [518]. voltage [748]. Volterra [856]. Volume [357, 457, 847, 852, 848, 1070, 1056, 1067, 1055, 1061, 1053, 1068, 1098, 607, 35, 224, 191, 225, 361, 323, 362, 504, 498, 505, 610, 611, 696, 654, 697, 792, 793, 1039, 960, 961, 1040, 111, 112, 1147]. Volumes [1112, 244, 245, 957, 965]. vorticity [119]. vorzeitiger [763].

W

[101, 73, 96, 44, 107, 45, 100, 123, 179, 191, 274, 356, 298, 275, 295, 458, 411, 559, 593, 561, 589, 653, 758, 741, 731, 757, 763, 813, 856, 816, 1028, 932, 1027, 934, 1096, 1057, 1067, 1115]. Wahl [30]. Wakerly [151]. Walkden [73]. Walker [320]. wall [66]. Walsh [884, 682, 680, 683, 681, 141, 418]. Walsh-Hadamard [141]. Walther [1103, 1110]. Wand [272]. Wanner [877]. Warman [322]. Washington [54, 207, 244, 432]. Wasilkowski [1067]. Wasilowski [596]. Wasserman [456]. water [178, 762, 266, 68, 308, 546, 868, 263, 532]. Watson [294]. Watt [280]. Wave [988, 1139, 762, 843]. waveforms [666]. wavelength [784]. waves [701]. way [27]. weak [4]. weakly [509, 835]. weapon [651]. Weber [561, 597, 521]. webs [1025]. Weierstrass [849]. weighted [445]. weights [1109]. Wellesley [745]. Wellesley-Cambridge [745]. Wendland [1096]. Wenger [1116]. Wesley [872, 984, 155, 274, 356, 353, 475, 467, 560, 557, 555, 590, 596, 594, 589, 874, 1009]. Westbury [279]. wet [489]. wetland [259]. Wets [410, 390]. WG [456, 851]. white [576]. Whiteman [74, 75, 436]. Wickes [411]. Wien [269, 275, 659, 856]. Wiesbaden [248]. Wilde [1091]. Wildenberg [297]. Wiley [154, 187, 189, 178, 191, 185, 186, 358, 323, 278, 316, 321, 454, 478, 473, 559, 593, 556, 595, 520, 535, 626, 627, 658, 717, 706, 715, 708, 713, 714, 744, 709, 711, 731, 855, 886, 820, 819, 825, 811, 826, 928, 927, 1086, 1116, 1103, 1110, 1085, 1101, 1114, 989, 742]. Wiley-distributors [1114]. Wiley-Teubner [742]. Wilkinson [48]. Willoughby [814]. Willsky [157]. Wilson [1083]. wind [649, 907, 775, 233]. Winkel [164]. Winograd [159]. Winter [59]. Wisconsin [351]. without [234]. Wolf [298]. woods [267]. Working [214, 456, 97, 99, 106, 31, 122, 125, 126, 298, 407, 409]. Workload [890]. Workshop [38, 216, 250, 890, 966, 967, 297]. World [198, 886, 891, 621, 622, 856, 832, 120]. worldwide [887]. Wossner [160]. Wozniakowski [596, 1067]. writing [396, 981]. Wulf [356].

X [970, 1086, 1112, 1082, 1114, 850, 80, 103, 150, 155, 243, 269, 429, 466, 501, 435, 409, 403, 406, 467, 595, 658, 622, 817, 857, 860, 890, 1007, 1008, 1012, 1069, 1089, 1056, 1108, 1101, 1060]. xi [854, 824, 825, 826, 893, 939, 993]. XII [1085, 855, 884, 823, 942]. XIII [1082, 1083, 817, 971, 931, 1063, 989]. XIV [1087, 812, 810, 879, 925, 1005, 927, 1004]. xix [1062]. xv [778, 777, 811, 1006]. xvi [882, 820]. xvii [815]. xviii [992]. xx [934]. xxi [929, 1054, 131]. xxii [886, 111, 112]. XXIII [224, 225]. XXIV [361, 362]. XXV [504, 505]. XXVI [610, 611]. XXVII [1108]. xxviii [859, 819]. xxxii [818]. xxxxxx [152].

Y.-S [476]. Yakhmo [938]. Yale [988]. Yeh [60]. York [981, 101, 30, 42, 43, 73, 96, 97, 99, 106, 44, 107, 33, 29, 77, 98, 36, 102, 103, 74, 78, 32, 35, 34, 104, 95, 75, 100, 160, 31, 122, 123, 125, 153, 161, 151, 162, 150, 154, 156, 187, 189, 149, 178, 190, 191, 148, 124, 185, 186, 182, 357, 354, 270, 278, 446, 380, 448, 449, 385, 405, 447, 436, 384, 520, 535, 659, 626, 627, 715, 729, 708, 731, 740, 766, 855, 825, 821, 876, 816, 928, 925, 1103, 1110, 1084]. York [1101, 1091, 1054, 1095, 989]. Yovits [357, 1104]. Yu [994, 938, 1033, 1106]. Yuasa [1069]. Yudin [520].

Z [193, 518]. Zapata [537]. Zarrop [192]. Zavialov [1033]. Zehnder [937]. Zeigler [81]. Zemanek [860]. Zolésio [1070]. zooplankton [268]. Zurich [517, 848, 852]. zusammenhängende [275]. Zwiechenauswertungen [763]. Zwillinger [1088].

References

Saaty:1979:MMD

[1] Thomas L. Saaty. Mathematical modeling of dynamic decisions; priorities and hierarchies with time dependence. Mathematics and Computers in Simulation, 21(4):352–358, December 1979. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                              /www.sciencedirect.com/science/
                                              article/pii/0378475479900041. See

erratum [131].

Anonymous:1980:PM

[2] Anonymous. Pages 1–80 (March 1980). Mathematics and Computers in Simulation, 22(1):??, March 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1980:EB

[3] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 22(1):iv, March 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900932.

Neundorf:1980:SIH

[4] W. Neundorf and Re. Schoenefeld. Stable iterative hybrid method for the solution of weak nonlinear parabolic differential equations. Mathematics and Computers in Simulation, 22(1): 1–6, March 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475480900944.

Posch:1980:SRC

[5] Bernd Posch and Günther Schmidt. Some remarks on a continuoussystem simulation language (CSSL) for LSI economics. Mathematics and Computers in Simulation, 22(1): 7–9, March 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475480900956.

Auslander:1980:ARC

[6] D. M. Auslander. Author’s reply to comments by Posch and Schmidt. Mathematics and Computers in Simulation, 22(1):10, March 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900968.

Finlayson:1980:OCF

[7] Bruce A. Finlayson. Orthogonal collocation on finite elements — progress and potential. Mathematics and Computers in Simulation, 22(1):11– 17, March 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548090097X.

Bank:1980:UAG

[8] Randolph E. Bank and Andrew H. Sherman. The use of adaptive grid refinement for badly behaved elliptic partial differential equations. Mathematics and Computers in Simulation, 22(1):18–24, March 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900981.

Hayes:1980:FEP

[9] Linda J. Hayes. Finite element patch approximations and alternatingdirection methods. Mathematics and Computers in Simulation, 22(1):25– 29, March 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475480900993.

Carver:1980:PCM

[10] M. B. Carver. Pseudo-characteristic method of lines solution of first-order hyperbolic equation systems. Mathematics and Computers in Simulation, 22(1):30–35, March 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480901007.

Maritsas:1980:DTH

[11] Dimitris G. Maritsas and Gerasimos G. Frangakis. Development and testing of a high precision digital Gaussian generator for Computers. Mathematics and Computers in Simulation, 22(1):36–48, March 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480901019.

Masliyah:1980:AOC

[12] Jacob H. Masliyah and Dileep Kumar. Application of orthogonal collocation on finite elements to a flow problem. Mathematics and Computers in Simulation, 22(1):49– 54, March 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475480901020.

VanRemoortere:1980:BRa

[13] P. Van Remoortere. Book reviews: 1977. Mathematics and Computers in Simulation, 22(1):55– 56, March 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475480901032.

VanRemoortere:1980:BRb

[14] P. Van Remoortere. Book reviews: 1978. Mathematics and Computers in Simulation, 22(1):56– 60, March 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475480901044.

VanRemoortere:1980:BRc

[15] P. Van Remoortere. Book reviews: 1979. Mathematics and Computers in Simulation, 22(1): 60, March 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475480901056.

VanRemoortere:1980:BRd

[16] P. Van Remoortere. Book reviews: 1977. Mathematics and Computers in Simulation, 22(1):61– 63, March 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475480901068.

VanRemoortere:1980:BRe

[17] P. Van Remoortere. Book reviews: 1978. Mathematics and Computers in Simulation, 22(1):63– 78, March 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847548090107X.

VanRemoortere:1980:BRf

[18] P. Van Remoortere. Book reviews: 1979. Mathematics and Computers in Simulation, 22(1):78– 79, March 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475480901081.

Anonymous:1980:PJ

[19] Anonymous. Pages 81–176 (July 1980). Mathematics and Computers in Simulation, 22(2):??, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Duchting:1980:STD

[20] W. Düchting and G. Dehl. Simulation of two-dimensional cell growth with application to tumor kinetics. Mathematics and Computers in Simulation, 22(2):81–90, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900014.

Houstis:1980:SOF

[21] E. N. Houstis, R. E. Lynch, and T. S. Papatheodorou. A sixth order fast direct Helmholtz equation solver. Mathematics and Computers in Simulation, 22(2):91–97, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900026.

Vichnevetsky:1980:PPS

[22] R. Vichnevetsky. Propagation properties of semi-discretizations of hyperbolic equations. Mathematics and Computers in Simulation, 22(2): 98–102, July 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475480900038.

Wadia:1980:AMI

[23] Aspi Rustom Wadia and Fred R. Payne. Application of the method of integral relations to unsteady transonic flows with shocks. Mathematics and Computers in Simulation, 22 (2):103–112, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548090004X.

Evans:1980:ISN

[24] D. J. Evans and A. Benson. Iterative solution of the nonlinear parabolic periodic boundary value problem. Mathematics and Computers in Simulation, 22(2):113–117, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475480900051.

Cennamo:1980:ULC

[25] F. Cennamo. The use of lowcost DDA for solving navigational problems. Mathematics and Computers in Simulation, 22(2):118– 126, July 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475480900063.

Cooke:1980:HOF

[26] Charlie H. Cooke and Doris K. Blanchard. A higher order finite element algorithm for the unsteady Navier–Stokes equations. Mathematics and Computers in Simulation, 22 (2):127–132, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900075.

vanCutsem:1980:HSE

[27] Th. van Cutsem, J. L. Horward, and M. Ribbens-Pavella. Hierarchical state estimator: a new way for large power systems estimation. Mathematics and Computers in Simulation, 22(2):133–140, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900087.

Bois:1980:SEL

[28] P. Bois and J. Vignes. A software for evaluating local accuracy in the Fourier transform. Mathematics and Computers in Simulation, 22(2):141–150, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900099.

Anonymous:1980:BRFa

[29] Anonymous. Book review: Fundamentals of computation theory: M. Karpinski: Proceedings of the 1977 International FCT-Conference, Posnan, 19–23 September 1977. Springer-Verlag, Berlin, Heidelberg, New-York, 542 pages, DM 43.00 US$21.50. ISBN 3-540-08442-8. Mathematics and Computers in Simulation, 22 (2):151, July 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475480900105.

Anonymous:1980:BRCa

[30] Anonymous. Book review: Computers in chemical education and research: E. V. Ludena, N. H. Sabelli and A. C. Wahl (Editors): Proceedings of the 3rd International Conference, Caracas, Venezuela, 1977. Plenum Publishing Company, New York, 1977, 488 pages, US$37.50. Mathematics and Computers in Simulation, 22(2):151, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900129.

Anonymous:1981:BRCa

[31] Anonymous. Book review: Computer aid to drug therapy and to drug monitoring. Proceedings of the IFIP TC-4 working conference on computer aid to drug therapy and to drug monitoring: H. Ducrot, M. Goldberg, R. Hoigne and P. Middleton: Berne, 5–10 March, 1978. 449 pages, Dfl. 110.00, US$53.75. North-Holland Publishing Company, Amsterdam, New York, 1978. ISBN 0-444-85188-7. Mathematics and Computers in Simulation, 23(1):74, March 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900112.

Anonymous:1980:BROa

[32] Anonymous. Book review: Optimal estimation in approximation theory: C. A. Micchelli and T. J. Rivlin: Plenum Publishing Company, New York, 1977, 310 pages, US$29.50. Mathematics and Computers in Simulation, 22(2):151, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900130.

Anonymous:1980:BREa

[33] Anonymous. Book review: Einführung in die Simulation diskreter Systeme mit Hilfe eines Fortran-programmpaketes: B. Schmidt: GPSS-Fortran. Springer-Verlag, Berlin, Heidelberg, New York, 1977, 298, DM 38.00 US$19.00. Mathematics and Computers in Simulation, 22(2):151, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900142.

Anonymous:1980:BRPa

[34] Anonymous. Book review: Programmiersprachen: K. Albert (Editor): Fachtagungen der GI Braunschweig 8– 9 Marz 1978. Springer-Verlag, Berlin, Heidelberg, New York, 1978, 179 pages, DM 18.00 US$9.00. ISBN 3-540-08680-3. Mathematics and Computers in Simulation, 22(2):151, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900154.

Anonymous:1980:BROb

[35] Anonymous. Book review: Optical information processing, volume 2: E. S. Barrekette et al. (Editors): Proceedings of the US–USSR Science Cooperation Seminar on Optical Information Processing, Novosibirsk, USSR. Plenum Publishing Company, New York, 1978, 415 pages, US$37.50. Mathematics and Computers in Simulation, 22(2):151, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475480900166.

Anonymous:1980:BRIa

[36] Anonymous. Book review: Information technology in health science education: E. C. Deland (Editor): Plenum Publishing Company, New York, 1978, 615 pages, US$42.50. Mathematics and Computers in Simulation, 22 (2):151–152, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475480900178.

Anonymous:1980:BRPb

[37] Anonymous. Book review: Performance of computer installations: D. Ferrari (Editor): Proceedings of the International Conference ICPCI 78, Gardone Riviera, Italy, 22–23 June 1978. North-Holland Publishing Company, Amsterdam, 1978, 352 pages, US$44.50, Dfl. 100.00. ISBN 0-444-85186-0. Mathematics and Computers in Simulation, 22 (2):152, July 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847548090018X.

Anonymous:1980:BRCb

[38] Anonymous. Book review: College curriculum in computer science engineering and data processing: IEEE C.: Proceedings of the 4th Education Workshop, Orlando, Fl, 2–3 February 1978. CH13110, 1978, 138 pages, US$12.00.

Mathematics and Computers in Simulation, 22(2):152, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900191.

Anonymous:1980:BRCc

[39] Anonymous. Book review: Computer technology: status, limits, alternatives: IEEE C.: Proceedings of the Spring COMPCON ’78, San Francisco, 28 February–3 March 1978. IEEE Service Center CH13284, 1978, 367 pages, US$20.00. Mathematics and Computers in Simulation, 22 (2):152, July 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475480900208.

Anonymous:1980:BRCd

[40] Anonymous. Book review: Computer architecture: IEEE C.: Proceedings of the 5th Annual Computer Architecture Symposium, Palo Alto, 3–5 April, IEEE Service Center, 1978. CH12849, 1978, 256 pages, US$20.00. Mathematics and Computers in Simulation, 22(2):152–153, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548090021X.

Anonymous:1980:BRAa

[41] Anonymous. Book review: Advances in computer aided manufacture: D. McPherson (Editor): Proceedings of the 3rd International IFIP/ IFAC Conference on Programming Languages for Machine Tools, Stirling, Scotland, 15–18 June, 1976. North-Holland Publishing Company, Amsterdam, 1977, 468 pages, Dfl 125.00 US$61.60. ISBN 0-7204-0500-9. Mathematics and Computers in Simulation, 22(2):153, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475480900221.

Anonymous:1980:BRCe

[42] Anonymous. Book review: Computer science and multiple-valued logic: Theory and applications: D. Rine: North-Holland Publishing Company, Amsterdam, New York, 1977, 548 pages, Dfl. 125.00 US$61.00. ISBN 0-7204-0406-1. Mathematics and Computers in Simulation, 22(2):153, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475480900233.

Anonymous:1980:BRCf

[43] Anonymous. Book review: Classification theory and the number of nonisotropic models: S. Shelah: North-Holland Publishing Company, Amsterdam, New York, 1978, 544 pages, Dfl. 140.00 US$68.25. ISBN 0-7204-0757-5. Mathematics and Computers in Simulation, 22(2):153, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475480900245.

Anonymous:1980:BRDa

[44] Anonymous. Book review: Datenbanken in Rechnernetzen mit Kleinrechnern: W. Stocky and E. Holler (Editors): Gi-Fachtagungen mit Unterstützung durch das German Chapter der ACM, Karlsruhe, 11–12 April 1978. Springer-Verlag Berlin, New York, 1978, DM 20.00, US$10.00. ISBN 3-540-08775-3. Mathematics and Computers in Simulation, 22 (2):153–154, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900257.

Anonymous:1980:BRIb

[45] Anonymous. Book review: Interactive techniques in computer aided design: W. Stocky and E. Holler (Editors): Proceedings of the International Conference on Interactive Techniques in Computer Aided Design, Bologna, Italy, 21–23 September 1978. IEEE Service Centre, 78CH1289-8C. Mathematics and Computers in Simulation, 22(2):154, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900269.

Anonymous:1980:BREb

[46] Anonymous. Book review: Engineering software: R. A. Adey (Editor): Proceedings of the 1st International Conference on Engineering Software, Southampton University, September 1979. Pentech Press, Plymouth, 701 pages. ISBN 0-7273-0501-8. Mathematics and Computers in Simulation, 22(2):154, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900270.

Anonymous:1980:BRPc

[47] Anonymous. Book review: Performance of computer systems: M. Arato, A. Butrimenko and E. Gelenbe (Editors): North-Holland Publishing Company, Amsterdam 1979, 566 pages, Dfl. 145.00, US$70.75. ISBN 0-444-85332-4. Mathematics and Computers in Simulation, 22(2):154, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475480900282.

Anonymous:1980:BRMa

[48] Anonymous. Book review: Modelling and simulation in practice: M. Cross, R. D. Gibson, M. J. O’Carroll and T. S. Wilkinson (Editors): Pentech Press, Plymouth, 1979. ISBN 0-7273-1303-7. Mathematics and Computers in Simulation, 22(2):154–155, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475480900294.

Anonymous:1980:BRPd

[49] Anonymous. Book review: Performance evaluation of numerical software: L. D. Fosdick (Editor): North-Holland Publishing Company, Amsterdam, 1979, 350 pages, Dfl. 50.00, U.S.$24.50. Mathematics and Computers in Simulation, 22 (2):155, July 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475480900300.

Anonymous:1980:BRMb

[50] Anonymous. Book review: Microcomputers: systems, software, architecture: IEEE C: Proceedings of the 2nd Rocky Mountain Symposium on Microcomputers, Pingree Park, CO, 27–30 August 1978. IEEE Service Center, 78CH1387-0 C, 1978, 418 pages, US$16.00. Mathematics and Computers in Simulation, 22(2):155, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900312.

Anonymous:1980:BRCg

[51] Anonymous. Book review: Computer arithmetic: IEEE C: Proceedings of the 4th Symposium on Computer Arithmetic, Santa Monica, 25–27 October 1978. IEEE Service Center, 78CH1412-6, 1979. 274 pages, US$22.00. Mathematics and Computers in Simulation, 22(2):155–156, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900348.

Anonymous:1980:BRCh

[52] Anonymous. Book review: Computer science: IEEE C and ACM: Conference Records of the 19th Symposium on Foundations of Computer Science, Ann Arbor, 16–17 October 1978. IEEE Service Center, 78CH1397-9C, 1979, 290 pages, US$22.00. Mathematics and Computers in Simulation, 22(2):155, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900336.

Anonymous:1980:BRCi

[53] Anonymous. Book review: Computer Society International Conference: IEEE C: Proceedings of the 17th IEEE (Compcon Fall ’78) Computer Communications Networks, 5–8 September 1978. IEEE Service Center, 78CH1388-8C, 1979, 434 pages, US$25.00. Mathematics and Computers in Simulation, 22(2):155, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475480900324.

Anonymous:1980:BRCj

[54] Anonymous. Book review: Computer applications in medical care: IEEE EMB: Proceedings of the 2nd Symposium on Computer Applications in Medical Care, Washington, 5–9 November 1978. IEEE Service Center, 78CH1413-4C, 1979, 667 pages, US$25.00. Mathematics and Computers in Simulation, 22 (2):156, July 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475480900373.

Anonymous:1980:BRCk

[55] Anonymous. Book review: Computer networking symposium: IEEE C: Proceedings of the Symposium on Computer Networking, Gaithersburg, 13 December 1978. IEEE Service Center, 78CH1400-1C, 1979, 137 pages, US$14.00. Mathematics and Computers in Simulation, 22 (2):156, July 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475480900361.

Anonymous:1980:BRCl

[56] Anonymous. Book review: Computer software and applications: IEEE C: Proceedings of the 2nd IEEE Computer Society International Computer Software and Applications Conference (COMPSAC 78), Chicago, 13– 16 November 1978. IEEE Service Center, 78CH1338-3C, 1979. Mathematics and Computers in Simulation, 22(2):156, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548090035X.

Anonymous:1980:BRPe

[57] Anonymous. Book review: Parallel processing: G. J. Lipovski (Editor): Proceedings of the 1978 International Conference on Parallel Processing, Bellaire, MI, 22–25 August 1978. IEEE Service Center, 78CH1321-9C, 1979, 269 pages US$22.00. Mathematics and Computers in Simulation, 22(2):156, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900385.

Anonymous:1980:BRCm

[58] Anonymous. Book review: Computer networks and distributed processing: James Martin Public Report Series: Butler Cox, London, 1979. Mathematics and Computers in Simulation, 22(2):156, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475480900397.

Anonymous:1980:BRSa

[59] Anonymous. Book review: Simulation conference: N. R. Nielsen, H. J. Highland and L. G. Hull: Proceedings of the 10th Winter Simulation Conference, Miami Beach, 4–6 December 1978. IEEE Service Center, 78CH1415-9SMC, 1979, 1051 pages, US$48.00. Mathematics and Computers in Simulation, 22(2):156, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475480900403.

Anonymous:1980:BRSb

[60] Anonymous. Book review: Software methodology, tutorial: C. V. Ramamoorthy and R. T. Yeh (Editor): Proceedings of the 2nd International Computer Software and Applications Conference (COMPSAC 78), Chicago, 13–16 November 1978. IEEE Service Center, EH01420, 1979, 832 pages, US$14.00. Mathematics and Computers in Simulation, 22 (2):156, July 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475480900415.

VanRemoortere:1980:Ba

[61] P. Van Remoortere. Bibliography. Mathematics and Computers in Simulation, 22(2):157–176, July 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/0378475480900427.

Anonymous:1980:PS

[62] Anonymous. Pages 177–290 (September 1980). Mathematics and Computers in Simulation, 22(3):??, September 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Vergnes:1980:DOI

[63] J. Vergnes. Determination d’un pas optimum d’integration pour la methode de Simpson. Mathematics and Computers in Simulation, 22(3):177– 188, September 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475480900440.

Birta:1980:QPM

[64] Louis G. Birta. A quasi-parallel method for the simulation of loosely coupled continuous subsystems. Mathematics and Computers in Simulation, 22(3):189–199, September 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475480900452.

Greenspan:1980:BMN

[65] Donald Greenspan. N-body modeling of nonlinear, free surface liquid flow. Mathematics and Computers in Simulation, 22(3):200–207, September 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475480900464.

Cooke:1980:ISW

[66] Charlie H. Cooke. Improvement in the slotted wall boundary condition by numerical methods. Mathematics and Computers in Simulation, 22(3): 208–212, September 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900476.

Bensoussan:1980:FTS

[67] A. Bensoussan, P. L. Chow, and J. L. Lions. Filtering theory for stochastic processes with two dimensional time parameter. Mathematics and Computers in Simulation, 22(3): 213–221, September 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900488.

Coulbeck:1980:DSW

[68] B. Coulbeck. Dynamic simulation of water distribution systems. Mathematics and Computers in Simulation, 22 (3):222–230, September 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548090049X.

Azadivar:1980:OSS

[69] F. Azadivar and J. Talavage. Optimization of stochastic simulation models. Mathematics and Computers in Simulation, 22(3):231–241, September 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475480900506.

Hadjidimos:1980:TPS

[70] A. Hadjidimos. Three-part splittings and Varga’s type extensions of the successive overrelaxation (SOR) theory. Mathematics and Computers in Simulation, 22(3):242–247, September 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475480900518.

Keramidas:1980:FEH

[71] George A. Keramidas. Finite element of the heat conduction equation with temperature dependent coefficients. Mathematics and Computers in Simulation, 22(3):248–255, September 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548090052X.

Evans:1980:PSD

[72] D. J. Evans and N. M. Missirlis. The preconditioned simultaneous displacement method (PSD method) for elliptic difference equations. Mathematics and Computers in Simulation, 22(3): 256–263, September 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900531.

Anonymous:1980:BRCn

[73] Anonymous. Book review: Computational methods and problems in aeronautical fluid dynamics: B. L. Hewitt, C. R. Illingworth, R. C. Lock, K. W. Mangler, J. H. McDonnell, C. Richards and F. Walkden(Editors), Academic Press, London, New York, San Francisco 1976, 526 pages, £16.60, US$34.35. ISBN 0-12-346350-5. Mathematics and Computers in Simulation, 22(3):264, September 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900543.

Anonymous:1980:BRMc

[74] Anonymous. Book review: The mathematics of finite elements and application II, MAFELAP 1975: J. R. Whiteman(Editor), Academic Press, London, New York, San Francisco, 1976, 573 pages, £21.00 US$4350. ISBN 0-12-747252-5. Mathematics and Computers in Simulation, 22(3): 264–265, September 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900555.

Anonymous:1980:BRSc

[75] Anonymous. Book review: The state of the art in numerical analysis: J. R. Whiteman(Editor), Academic Press, London, New York, San Francisco, 1977, 978 pages, £20.000, US$41.00. ISBN 0-12-378650-9. Mathematics and Computers in Simulation, 22(3):265, September 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900567.

Anonymous:1980:BRV

[76] Anonymous. Book review: Variational methods in mathematics, sciences and engineering (translation by M. Basch).: K. Rektorys, Reidel, Dordrecht, 1977, 516 pages, Dfl. 135.00, US$67.00. Mathematics and Computers in Simulation, 22(3):265– 266, September 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475480900579.

Anonymous:1980:BRFb

[77] Anonymous. Book review: Fisheries mathematics: J. H. Steele(Editors), Academic Press, London, New York, San Francisco, 1977, 198 pages, £8.50, US$17.65. ISBN 0-12-665250-3. Mathematics and Computers in Simulation, 22(3):266, September 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900580.

Anonymous:1980:BRN

[78] Anonymous. Book review: Numerical software — needs and availability: D. A. H. Jacobs, Academic Press, London, New York, San Francisco, 1978, 410 pages, £11.60, US$24.00. ISBN 0-12-378660-6. Mathematics and Computers in Simulation, 22(3): 266–267, September 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900592.

Anonymous:1980:BRCo

[79] Anonymous. Book review: Computer simulation 1951–1976: P. Holst, An Index to the literature. Mansell Publishing, London, 1979, 438 pages, £25.00. ISBN 0-7201-0734-2. Mathematics and Computers in Simulation, 22(3):267, September 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475480900609.

Anonymous:1980:BRBa

[80] Anonymous. Book review: Biological systems, modelling and control: D. A. Linkers(Editor), Peter Peregrinus Ltd., Hitchin, 350 pages, £15.00, US$40.50. ISBN 0-906048-15-X. Mathematics and Computers in Simulation, 22(3):267, September 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475480900610.

Anonymous:1980:BRMd

[81] Anonymous. Book review: Methodology in systems modelling and simulation: B. P. Zeigler, M. S. Elzas, G. J. Klir and T. C. Oren, North-Holland, Amsterdam, 1979, 536 pages, US$65.75, Dfl. 135.00. ISBN 0-444-85340-5. Mathematics and Computers in Simulation, 22(3):267–268, September 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475480900622.

Saphier:1980:SCC

[82] D. Saphier. Simulation computer code abstracts. Mathematics and Computers in Simulation, 22(3):269– 270, September 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475480900634.

VanRemoortere:1980:Bb

[83] P. Van Remoortere. Bibliography. Mathematics and Computers in Simulation, 22(3):271–287, September 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475480900646.

Anonymous:1980:CP

[84] Anonymous. Call for papers. Mathematics and Computers in Simulation, 22(3):288–289, September 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475480900658.

Anonymous:1980:PD

[85] Anonymous. Pages 291–401 (December 1980). Mathematics and Computers in Simulation, 22(4):??, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Birkhoff:1980:AMC

[86] Garrett Birkhoff and Robert E. Lynch. Added mass of cylinders. Mathematics and Computers in Simulation, 22 (4):291–297, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548090066X.

Carver:1980:POC

[87] M. B. Carver. Parameter optimization in the continuous simulation packages FORSIM and MACKSIM. Mathematics and Computers in Simulation, 22 (4):298–318, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900671.

Grandek:1980:CHI

[88] Karol Grandek. Convergence of a hybrid iterative procedure. Mathematics and Computers in Simulation, 22 (4):319–323, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900683.

Hryniewicz:1980:ARF

[89] Z. Hryniewicz. An approximation of random field with a bounded discrete parameter space. Mathematics and Computers in Simulation, 22(4): 324–329, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900695.

Forsythe:1980:ACD

[90] W. Forsythe. The analogue computer in digital form. Mathematics and Computers in Simulation, 22(4): 330–339, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900701.

Gerasoulis:1980:MPP

[91] A. Gerasoulis and A. S. Kydes. Market penetration — a probabilistic approach. Mathematics and Computers in Simulation, 22(4):340–352, December 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475480900713.

Sankar:1980:HCO

[92] S. Sankar and J. V. Svoboda. Hybrid computer in the optimal design of hydro-mechanical systems. Mathematics and Computers in Simulation, 22 (4):353–367, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900725.

Balasubramanian:1980:SBO

[93] R. Balasubramanian. Simulation of biological oscillators involving large time delays. Mathematics and Computers in Simulation, 22(4):368–372, December 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475480900737.

Stavroulakis:1980:DPO

[94] P. Stavroulakis and S. Tzafestas. Distributed-parameter observer-based control implementation using finite spatial measurements. Mathematics and Computers in Simulation, 22(4): 373–379, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900749.

Anonymous:1980:BRQ

[95] Anonymous. Book review: Qualitative analysis of large scale dynamical systems: A. N. Michel and R. K. Miller: Academic Press, New York, 1977, 289 pages, US$22.50. ISBN 0-12-49850-7. Mathematics and Computers in Simulation, 22 (4):380, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900750.

Anonymous:1980:BRCp

[96] Anonymous. Book review: Characteristics of software quality: B. W. Boehm et al.: North-Holland Publishing Company, Amsterdam, New York, 1978, 216 pages, Dfl. 65,00, US$25.00. ISBN 0-444-85105-4. Mathematics and Computers in Simulation, 22(4):380, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900762.

Anonymous:1980:BRCq

[97] Anonymous. Book review: Computer performance. IFIP working group 7.3 on computer system modelling: M. K. Chandy and M. Reiser (Editors): North-Holland Publishing Company, Amsterdam, New York, 1978, 546 pages, Dfl. 120.00 US$58.50 ISBN 0-444-85038-4. Mathematics and Computers in Simulation, 22 (4):380, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900774.

Anonymous:1980:BRFc

[98] Anonymous. Book review: The finite element method for elliptic problems: P. G. Ciarlet: North-Holland Publishing Company, Amsterdam, New York, 1978, 534 pages, Dfl. 130.00, US$63.50. ISBN 0-444-85028-7. Mathematics and Computers in Simulation, 22(4):380, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900786.

Anonymous:1980:BRCr

[99] Anonymous. Book review: Constructing quality software. Proceedings of the IFIP working conference on constructing quality software, Novosibirsk, 23– 27 May 1977: P. G. Hibbard and S. A. Schuman: North-Holland Publishing Company, Amsterdam, New York, 1978, 510 pages, Dfl. 110.00, US$53.75. ISBN 0-444-85106-2. Mathematics and Computers in Simulation, 22(4):380, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900798.

Anonymous:1980:BRSd

[100] Anonymous. Book review: System development methodology: G.’F. Hice, W. S. Turner III and L. F. Cashwell: North-Holland Publishing Company, Amsterdam, New York, 1978, 440 pages, Dfl. 85.00, US$37.75. ISBN 0-444-85143-7. Mathematics and Computers in Simulation, 22 (4):380, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475480900804.

Anonymous:1980:BRAb

[101] Anonymous. Book review: Automated theorem proving: a logical base: D. W. Loveland: North-Holland Publishing Company, Amsterdam, New York, 1978, 432 pages, Dfl. 100.00 US$48.75. ISBN 0-7204-0499-1. Mathematics and Computers in Simulation, 22(4):380, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475480900816.

Anonymous:1980:BRIc

[102] Anonymous. Book review: Identification and system parameter estimation. Proceedings of the 4th IFAC symposium in Tbilisi, USSR, 21–27 September, 1976: N. S. Rajbman (Editor): North-Holland Publishing Company, Amsterdam New York, 1978, 1900 pages, Dfl. 350.00, US$170.75. ISBN 0-444-85096-1. Mathematics and Computers in Simulation, 22 (4):380, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475480900828.

Anonymous:1980:BRId

[103] Anonymous. Book review: Identification of industrial processes. The application of computers in research and production control: N. S. Rajbman and V. M. Chadeev: North-Holland Publishing Company, Amsterdam, New York, 1978, Dfl. 95.00, US$46.25.

ISBN 0-444-85181-X. Mathematics and Computers in Simulation, 22(4): 380–381, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548090083X.

Anonymous:1980:BRPf

[104] Anonymous. Book review: A programming methodology in compiler construction, part 1: concepts.: J. Lewi, K. De Vlaminck, J. Huens and M. Huybrechts (Editors): North-Holland Publishing Company, Amsterdam, New York, 1979, 308 pages, Dfl. 85.00, US$41.50. ISBN 0-444-85288-3. Mathematics and Computers in Simulation, 22(4):381, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900841.

Anonymous:1980:BRBb

[105] Anonymous. Book review: Biological systems, modelling and control. IEE control engineering series 11: D. A. Linkens (Editors): Peter Peregrinus Ltd. G.B., 1979, 350 pages, £15.00, US$40.50. ISBN 0-906048-15. Mathematics and Computers in Simulation, 22(4):381, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900853.

Anonymous:1980:BRCs

[106] Anonymous. Book review: Computer aided tomography and ultra sound in medicine. Proceedings of the IFIP TC-4 working conference, August 8– 10, 1979, Haifa: J. Raviv, J. F. Greenleaf and G. T. Herman (Editors): North-Holland Publishing Company, Amsterdam, New York, 1979, 350 pages, Dfl. 100.00, US$48.75. ISBN 0-444-85299-9. Mathematics and Computers in Simulation, 22 (4):381, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900865.

Anonymous:1980:BRDb

[107] Anonymous. Book review: Distributed data bases. Proceedings of the international symposium on distributed data bases, Paris, March 12–14, 1980: C. Delobel and W. Litwin (Editors): North-Holland Publishing Company, Amsterdam, New York, 1980, Dfl. 100.00, US$48.75. ISBN 0-444-85471-1. Mathematics and Computers in Simulation, 22(4): 381–382, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900877.

Anonymous:1980:BRSe

[108] Anonymous. Book review: Similitude and modeling: E. Szucs: Elsevier Scientific Publishing Company, Amsterdam, 1980, Dfl. 135.00, US$65.75. ISBN 0-444-99780-6. Mathematics and Computers in Simulation, 22 (4):382, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900889.

Anonymous:1980:BRSf

[109] Anonymous. Book review: Simulation of distributed parameter and large-scale systems. Proceedings of the IMACS European simulation meeting, the University of Patras, Greece, 2–4 October 1979: S. G. Tzafestas (Editors): North-Holland Publishing Company, Amsterdam, 1980, Dfl. 120.00, US$58.50. ISBN 0-444-85447-9. Mathematics and Computers in Simulation, 22(4):382, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900890.

Anonymous:1980:B

[110] Anonymous. Bibliography. Mathematics and Computers in Simulation, 22 (4):383–398, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475480900907.

Anonymous:1980:AIV

[111] Anonymous. Author index: Volume XXII (1980). Mathematics and Computers in Simulation, 22(4):399– 400, December 1980. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475480900919.

Anonymous:1980:SIV

[112] Anonymous. Subject index: Volume XXII (1980). Mathematics and Computers in Simulation, 22 (4):401, December 1980. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475480900920.

Anonymous:1981:PM

[113] Anonymous. Pages 1–98 (March 1981). Mathematics and Computers in Simulation, 23(1):??, March 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1981:EB

[114] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 23(1):ii, March 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548190001X.

Jutila:1981:UCS

[115] Esa Jutila and Kauko Leiviskä. The use of computer simulation in the pulp and paper industry. Mathematics and Computers in Simulation, 23(1):1–11, March 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900021.

Ellison:1981:EAI

[116] D. Ellison. Efficient automatic integration of ordinary differential equations with discontinuities. Mathematics and Computers in Simulation, 23(1):12–20, March 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900033.

Taylor:1981:SPA

[117] James H. Taylor. Statistical performance analysis of nonlinear stochastic systems by the Monte Carlo method. Mathematics and Computers in Simulation, 23(1):21–33, March 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475481900045.

Gomm:1981:SAE

[118] W. Gomm. Stability analysis of explicit multirate methods. Mathematics and Computers in Simulation, 23(1):34–50, March 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900057.

Cooke:1981:AVN

[119] Charlie H. Cooke. On artificial vorticity in numerical simulation of incompressible flow. Mathematics and Computers in Simulation, 23(1):51– 55, March 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475481900069.

Wu:1981:PMC

[120] Chih Wu. Physical modeling and computer graphic simulation of the depletion of world energy reserve. Mathematics and Computers in Simulation, 23(1):56–62, March 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900070.

Klenke:1981:CFT

[121] M. Klenke and H. Thiem. A coupled flow and transport model for simulation of groundwater reservoirs. Mathematics and Computers in Simulation, 23(1):63–72, March 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900082.

Anonymous:1981:BRCb

[122] Anonymous. Book review: Computer performance. IFIP working group 7.3 on computer system modelling: M. K. Chandi and M. Reiser (Editors): 1978, 564 pages. North-Holland Publishing Company, Amsterdam, New York, Dfl. 120.00, US$58.50. ISBN 0-444-85038-4. Mathematics and Computers in Simulation, 23(1):73, March 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900100.

Anonymous:1981:BRCc

[123] Anonymous. Book review: Characteristics of software quality: B. W Boehm et al., 1978, 166 pages. North-Holland Publishing Company, Amsterdam, New York, Dfl. 65.00 and US$25.00. ISBN 0-444-85105-4. Mathematics and Computers in Simulation, 23(1):73, March 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900094.

Anonymous:1981:BRPa

[124] Anonymous. Book review: The performance of computer installations.

Evaluation and management. Proceedings of the international conference on the performance of computer installations: D. Ferrari (Editor): ICPC’78, 22–23 June 1978, Lake Garda, Italy. North-Holland Publishing Company, Amsterdam, New York, 1978, 332 pages, Dfl 100.00, US$44.50. ISBN 0-444-85186-0. Mathematics and Computers in Simulation, 23(1):74–75, March 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900124.

Anonymous:1981:BRCd

[125] Anonymous. Book review: Constructing quality software. Proceedings of the IFIP working conference on constructing quality software: P. G. Hibbard and S. A. Schuman: Novosibirsk, USSR, 23–28 May 1977. North-Holland Publishing Company, Amsterdam, New York, 1978, 520 pages, Dfl. 110.00, US$53.75. ISBN 0-444-85106-2. Mathematics and Computers in Simulation, 23(1):75, March 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900136.

Anonymous:1981:BRFa

[126] Anonymous. Book review: Formal descriptions of programming concepts. Proceedings of the working conference on August 1–5, 1977: E.J Neuhold: St. Andrews, Canada 1978, 648 pages, Dfl. 150.00, US$73.25. North-Holland Publishing Company, Amsterdam, 1978. ISBN 0-444-85107-0. Mathematics and Computers in Simulation, 23(1):

75–76, March 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475481900148.

Anonymous:1981:BRSa

[127] Anonymous. Book review: Software reliability. A study of large project reality: T. A. Thayer, M. Lipow and E. C. Nelson: 1978, 311 pages. North-Holland Publishing Company, Amsterdam, Dfl. 90.00. US$40.00. ISBN 0.-444-85217-4. Mathematics and Computers in Simulation, 23(1):76, March 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548190015X.

Anonymous:1981:BRCe

[128] Anonymous. Book review: Computers and instrumentation. A practical handbook of measurement, interfacing and control circuits: A. Carrick: Heyden and Son, London, Philadelphia and Rheine, 1979, £8.60, $19.50, DM 40.50.256 pages. ISBN 0-85501-452-0. Mathematics and Computers in Simulation, 23(1): 76–77, March 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475481900161.

Anonymous:1981:BRMa

[129] Anonymous. Book review: Methods for the numerical solution of partial differential equations: D.U von Rosenberg: G. L. Farrar and Ass., Printing Division, 5536E, 35th Street, Tulsa, Okla 74135, USA, 1980.

$12.50. ISBN 0-444-00049-6. Mathematics and Computers in Simulation, 23(1):77, March 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900173.

Anonymous:1981:Ba

[130] Anonymous. Bibliography. Mathematics and Computers in Simulation, 23(1):78–97, March 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900185.

Anonymous:1981:EMM

[131] Anonymous. Erratum: Mathematical modeling of dynamic decisions; priorities and hierarchies with time dependence: T. L. Saaty, Mathematics and Computers in Simulation XXI (1979) 352–358. Mathematics and Computers in Simulation, 23(1):98, March 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900197. See

[1].

Anonymous:1981:PJ

[132] Anonymous. Pages 99–234 (July 1981). Mathematics and Computers in Simulation, 23(2):??, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Jutila:1981:PCM

[133] P. Jutila, P. J. Orava, and B. Salmelin. A physico-chemical model and simulation of pH-process in continuous stirred tank reactors. Mathematics and Computers in Simulation, 23(2):99–106, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548190046X.

Murthy:1981:NSC

[134] D. N. P. Murthy, J. D. Day, and I. R. Webb. A note on the stability of CSDT methods for solving parabolic equations. Mathematics and Computers in Simulation, 23 (2):107–110, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900471.

Rideout:1981:MSE

[135] Vincent C. Rideout. The modeling of socio-economic-resource systems. Mathematics and Computers in Simulation, 23(2):111– 126, July 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475481900483.

Holnicki:1981:EDT

[136] Piotr Holnicki. On L ^{2}-estimates for discrete-time Galerkin approximation to second order hyperbolic equations. Mathematics and Computers in Simulation, 23(2):127– 132, July 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475481900495.

Baron:1981:CAN

[137] G. Baron and S. Wajc. Convergence acceleration of non-scalar sequences with non-linear transformations. Mathematics and Computers in Simulation, 23(2):133–141, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900501.

Neundorf:1981:IBM

[138] Werner Neundorf. Iterative block methods for the hybrid computer solution of the method of lines. Mathematics and Computers in Simulation, 23(2):142–148, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900513.

Bui:1981:SSD

[139] T. D. Bui. Solving stiff differential equations for simulation. Mathematics and Computers in Simulation, 23(2):149–156, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900525.

Murphy:1981:CSS

[140] C. P. Murphy and D. J. Evans. Chebyshev series solution of the two dimensional heat equation. Mathematics and Computers in Simulation, 23(2):157–162, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900537.

Tzafestas:1981:FDW

[141] S. Tzafestas and G. Frangakis. A fast digital Walsh-Hadamard signal processing system. Mathematics and Computers in Simulation, 23(2):163– 169, July 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475481900549.

Jarny:1981:ISV

[142] Yvon Jarny. Identification of spatially varying parameters for a distributed system — Application to a thermal process. Mathematics and Computers in Simulation, 23 (2):170–179, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900550.

Evans:1981:PSC

[143] D. J. Evans and A. Hadjidimos. Parallel solution to certain banded, symmetric and centro-symmetric systems by using the quadrant interlocking factorisation method. Mathematics and Computers in Simulation, 23(2):180–187, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900562.

Mathis:1981:EER

[144] F. H. Mathis. An L _{∞} error estimate for the Ritz–Trefftz approximation in optimal control. Mathematics and Computers in Simulation, 23(2):188–190, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900574.

Missirlis:1981:APS

[145] N. M. Missirlis and D. J. Evans. On the acceleration of the preconditioned simultaneous displacement method. Mathematics and Computers in Simulation, 23(2):191– 198, July 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475481900586.

Cooke:1981:PSK

[146] Charlie C. Cooke. Performance sensitivity of the Kalman filter to choice of state and observation vectors. Mathematics and Computers in Simulation, 23(2):199–205, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900598.

Tzafestas:1981:PSA

[147] S. G. Tzafestas and P. Stavroulakis. Partioned [sic] adaptive filtering and control of distributed systems with space-dependent unknown parameters. Mathematics and Computers in Simulation, 23(2):206– 212, July 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475481900604.

Anonymous:1981:BRN

[148] Anonymous. Book review: Numerical analysis for computer science: A. Dodes: North-Holland Publishing Company, Amsterdam, New York, 1978, 640 pages, US$20.75. ISBN 0-444-00238-3. Mathematics and Computers in Simulation, 23(2):213, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900616.

Anonymous:1981:BRMb

[149] Anonymous. Book review: Modeling the bipolar transistor: I. E. Getreu: Computer-aided design of electronic circuits, 1. North-Holland Publishing Company, Amsterdam, New York, 1978, 262 pages, US$43.50 Dfl. 100.00. ISBN 0-444-41722-2. Mathematics and Computers in Simulation, 23(2):213–214, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900628.

Anonymous:1981:BRIa

[150] Anonymous. Book review: An introduction to elementary computer and compiler design: D. R. Steel: North-Holland Publishing Company, Amsterdam, New York, 1978, 208 pages, US$16.50. ISBN 0-444-00243-X. Mathematics and Computers in Simulation, 23(2):214, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548190063X.

Anonymous:1981:BREa

[151] Anonymous. Book review: Error detecting codes, self-checking circuits and applications: J. F. Wakerly: North-Holland Publishing Company, Amsterdam New York, 1978, 224 pages, US$22.00. ISBN 0-444-00256-1. Mathematics and Computers in Simulation, 23(2):214, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900641.

Anonymous:1981:BRDa

[152] Anonymous. Book review: Data analysis and informatics: xxxxxx: Proceedings of the symposium organized by INRIA at Fontainebleau, 19–30 March 1979. Published by INRIA-service formation 1980, Service de Vente directe des Publications, B. P. 105 F-78150 Le Chesnay, France. ISBN 2-7261-0232-8, 310 pages, 90.00 FF + postage. Mathematics and Computers in Simulation, 23(2):214, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900653.

Anonymous:1981:BRCf

[153] Anonymous. Book review: Computing US systems reliability: T. Anderson and B. Randell: Cambridge University Press, New York, 1979, 482 pages, $29.50. ISBN 0-521-22767-4. Mathematics and Computers in Simulation, 23(2):214–215, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900665.

Anonymous:1981:BRIb

[154] Anonymous. Book review: Introduction to dynamic systems: theory, models and applications: D. G. Luenberger: Wiley, New York, 1979, 446 pages US$19.95. ISBN 0-471-0259-4. Mathematics and Computers in Simulation, 23(2):215, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900690.

Anonymous:1981:BRIc

[155] Anonymous. Book review: Introduction to automata theory, languages and computation: J. E. Hopcroft and J. D. Ullman: Addison-Wesley, Reading, MA, 1979, 418 pages, US$20.95. ISBN 0-201-02988-X. Mathematics and Computers in Simulation, 23 (2):215, July 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475481900689.

Anonymous:1981:BRId

[156] Anonymous. Book review: Inference and linear models: D. A. S. Fraser: McGraw-Hill, New York, 1979, 297 pages, US$34.50. ISBN 0-07-021910-9. Mathematics and Computers in Simulation, 23(2):215, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900677.

Anonymous:1981:BRDb

[157] Anonymous. Book review: Digital signal processing and control estimation theory: A. S. Willsky: M.I.T. Press, Cambridge, MA, 1979, 256 pages, US$22.50. ISBN 0-262-23091-7. Mathematics and Computers in Simulation, 23(2):215–216, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900719.

Anonymous:1981:BRDc

[158] Anonymous. Book review: Distributedprocessor communication architecture: K. J. Thurber and G. M. Masson: Lexington Press, 1979, 252 pages, US$23.00. ISBN 0-669-0214-9. Mathematics and Computers in Simulation, 23(2):215, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900707.

Anonymous:1981:BRIe

[159] Anonymous. Book review: Integer programming: S. Winograd: Heyden & Son Ltd., Spectrum House, Hillview Gardens, London NW4 2JQ, order no 91400 032 2, £5.5. Mathematics and Computers in Simulation, 23(2):216, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900720.

Anonymous:1981:BRAa

[160] Anonymous. Book review: Algorithmische sprache und programmentwicklung: F. L. Bauer and H. Wossner: Springer-Verlag, Berlin, Heidelberg, New York, 530 Seiten, 1980, DM 79.00, US$46.70. ISBN 3-540-09853-4.

Mathematics and Computers in Simulation, 23(2):216, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900732.

Anonymous:1981:BRCg

[161] Anonymous. Book review: Contentaddressable memories: T. Kohonen: Springer-Verlag, Berlin, Heidelberg, New York, 1980, 368 pages, DM 78,00, US$46.10. ISBN 3-540-09823-2. Mathematics and Computers in Simulation, 23(2):216, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900744.

Anonymous:1981:BRFb

[162] Anonymous. Book review: Fast Fourier Transform and convolution algorithms: H. J. Nussbaumer: Springer-Verlag, Berlin, Heidelberg, New York, 1980, 330 pages. ISBN 3-540-10159-4. Mathematics and Computers in Simulation, 23(2):216– 217, July 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475481900756.

Anonymous:1981:BRPb

[163] Anonymous. Book review: Petri net theory and the modeling of systems: J. L. Peterson: Prentice-Hall International, GB, 1980, 288 pages, £14.65, US$30.40. ISBN 13-661983-5. Mathematics and Computers in Simulation, 23(2):217, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900768.

Anonymous:1981:BRAb

[164] Anonymous. Book review: The art of digital design, an introduction to topdown design: D. Winkel: Prentice-Hall International, GB, 1980, 656 pages, £16.20, US$33.70. ISBN 13-056607-7. Mathematics and Computers in Simulation, 23(2):217, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548190077X.

Anonymous:1981:Bb

[165] Anonymous. Bibliography. Mathematics and Computers in Simulation, 23(2):218–233, July 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900781.

Anonymous:1981:PS

[166] Anonymous. Pages 235–332 (September 1981). Mathematics and Computers in Simulation, 23(3):??, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Webb:1981:HOR

[167] I. R. Webb and D. N. P. Murthy. Higher order Runge–Kutta type CSDT methods for solving parabolic equations. Mathematics and Computers in Simulation, 23(3):235–244, September 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475481900793.

Gustafsson:1981:SSE

[168] Leif Gustafsson. Split-simulation — an efficient method for studies of models with discrete choices. Mathematics and Computers in Simulation, 23 (3):245–252, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548190080X.

deFornel:1981:HSC

[169] B. de Fornel, H. C. Hapiot, J. M. Farines, and J. Hector. Hybrid simulation of a current fed asynchronous machine. Mathematics and Computers in Simulation, 23(3):253– 261, September 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475481900811.

Andersson:1981:NSH

[170] Torbjörn Andersson, Jan Johansson, and Hans Eklund. Numerical solution of the Hilbert transform for phase calculation from an amplitude spectrum. Mathematics and Computers in Simulation, 23(3): 262–266, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900823.

Spalding:1981:GPC

[171] D. B. Spalding. A general purpose computer program for multi-dimensional one- and two-phase flow. Mathematics and Computers in Simulation, 23 (3):267–276, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900835.

Kaplan:1981:ASG

[172] B. Z. Kaplan and D. Kottick. Another stabilized generator model for three phase sinewaves. Mathematics and Computers in Simulation, 23(3): 277–279, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900847.

Humeau:1981:CSP

[173] J. P. Humeau and M. Grolleau. Computer simulation of a parametric optimization problem for an industrial process. Mathematics and Computers in Simulation, 23(3):280– 284, September 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475481900859.

Lipitakis:1981:UGE

[174] Elias A. Lipitakis. The use of the generalized extended to the limit sparse factorization techniques for the solution of non-linear elliptic and parabolic difference equations. Mathematics and Computers in Simulation, 23(3): 285–292, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900860.

Foester:1981:EMS

[175] Hartmut Foester and Kristian Witsch. On efficient multigrid software for elliptic problems on rectangular domains. Mathematics and Computers in Simulation, 23(3):293–298, September 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475481900872.

Lawson:1981:CMT

[176] P. A. Lawson. Contraction mapping techniques applied to the hybrid computer solution of parabolic PDE’s. Mathematics and Computers in Simulation, 23(3):299–303, September 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475481900884.

El-Zorkany:1981:HCS

[177] H. I. El-Zorkany and R. Balasubramanian. Hybrid computer solution of PDE’s using Laplace-modified Galerkin approximation. Mathematics and Computers in Simulation, 23 (3):304–311, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900896.

Anonymous:1981:BRMc

[178] Anonymous. Book review: Mathematical models in water pollution control: A. James: Wiley, New York, 1978, 420 pages. Mathematics and Computers in Simulation, 23 (3):312, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900902.

Anonymous:1981:BRMd

[179] Anonymous. Book review: The modelling of systems with small observation sets: W. Maciejowski: Springer-Verlag, Berlin, 1978, 242 pages. Mathematics and Computers in Simulation, 23(3):312, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900914.

Anonymous:1981:BRSb

[180] Anonymous. Book review: System dynamics: K. Ogata: Prentice-Hall, Englewood Cliffs, 1978, 596 pages. Mathematics and Computers in Simulation, 23(3):312, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900926.

Anonymous:1981:BRMe

[181] Anonymous. Book review: Modeling, estimation and their applications for distributed parameter systems: Y. Sawaragi et al.: Springer-Verlag, Berlin, 1978. Mathematics and Computers in Simulation, 23 (3):312, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900938.

Anonymous:1981:BRW

[182] Anonymous. Book review: What computers can’t do: H. L. Dreyfus: Harper Colophon, New York, 1979, 354 pages. Mathematics and Computers in Simulation, 23(3):312, September 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847548190094X.

Anonymous:1981:BRPc

[183] Anonymous. Book review: Prozessrechnentechnik: G. Farber: Springer-Verlag, Berlin, 1979, 208 pages. Mathematics and Computers in Simulation, 23(3):312, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900951.

Anonymous:1981:BRMf

[184] Anonymous. Book review: Models and decision making in national economics: J. Janssen, L. Pau and A. Straszak: North-Holland Publishing Company, Amsterdam, 1979, 425 pages. Mathematics and Computers in Simulation, 23(3):312, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900963.

Anonymous:1981:BRPd

[185] Anonymous. Book review: Probabilistic models in engineering science. vol. I: Random variables and stochastic processes: H. J. Larson and B. O. Shubert: Wiley, New York, 1979, 544 pages, bound, US$24.95, ISBN 0-471-01751-5. Mathematics and Computers in Simulation, 23 (3):312, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900975.

Anonymous:1981:BRPe

[186] Anonymous. Book review: Probabilistic models in engineering sciences: H. J. Larson and B. O. Shubert: Vol. II: Random noise, signals and dynamic systems. Wiley, New York, 1979, 237 pages, bound US$26.95, ISBN 0-471-05179-9. Mathematics and Computers in Simulation, 23(3):312, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900987.

Anonymous:1981:BRIf

[187] Anonymous. Book review: Introduction to dynamic systems: Theory, models and applications: D. G. Luenberger: Wiley, New York, 1979, 446 pages. Mathematics and Computers in Simulation, 23(3):312, September 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475481900999.

Anonymous:1981:BRMg

[188] Anonymous. Book review: Modelling and optimization of complex systems: G. I. Marchuk: Springer-Verlag, Berlin, 1979, 293 pages. Mathematics and Computers in Simulation, 23 (3):312, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481901002.

Anonymous:1981:BRIg

[189] Anonymous. Book review: Introduction to simulation and SLAM: A. Pritsker and C. Pegden: Wiley, New York, 1979, 588 pages. Mathematics and Computers in Simulation, 23 (3):313, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481901014.

Anonymous:1981:BRMh

[190] Anonymous. Book review: Modeling and control of river quality: S. Rinaldi et al.: McGraw-Hill, New York, 1979, 177 pages. Mathematics and Computers in Simulation, 23 (3):313, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481901026.

Anonymous:1981:BRMi

[191] Anonymous. Book review: Modeling and simulation, volume 10. Proceedings of the 10th annual Pittsburg conference: W. G. Vogt and M. H. Mickle (Editors): Published by Instrument Society of America. Distributor: Wiley, New York, 1979. Mathematics and Computers in Simulation, 23(3):313, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481901038.

Anonymous:1981:BRO

[192] Anonymous. Book review: Optimal experiment design for dynamic system identification: M. B. Zarrop: Springer-Verlag, Berlin, 1979, 197 pages. Mathematics and Computers in Simulation, 23(3):313, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548190104X.

Anonymous:1981:BRIh

[193] Anonymous. Book review: Identification of control plants: Z. Bubnicki: Elsevier Scientific Publishing Company, Amsterdam, 1980, 312 pages, US$68.25, Dfl. 140.00, ISBN 0-444-99767-9. Mathematics and Computers in Simulation, 23 (3):313, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481901051.

Anonymous:1981:BRPf

[194] Anonymous. Book review: Practical applications of symbolic computation: J. C. Howard: IPC Science and Technology Press, UK, 1980, 394 pages £20.00 or US$52.00; ISBN 0-86103-036-2. Mathematics and Computers in Simulation, 23(3): 313–314, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481901063.

Anonymous:1981:BRT

[195] Anonymous. Book review: Theory and applications of some new classes of integral equations: A. Ramm: Springer-Verlag, Berlin, 1980, 350 pages, DM 31, US$18.80, ISBN 3-540-90540-5. Mathematics and Computers in Simulation, 23(3):314–315, September 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475481901075.

Anonymous:1981:BREb

[196] Anonymous. Book review: Experimental computer performance evaluation: D. Ferrari and M. Spandoni North-Holland Publishing Company, Amsterdam, 1981, 264 pages, US$36.50, ISBN 0-444-86129-7. Mathematics and Computers in Simulation, 23(3):315, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481901087.

Anonymous:1981:Bc

[197] Anonymous. Bibliography. Mathematics and Computers in Simulation, 23 (3):316–330, September 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481901099.

Anonymous:1981:AIW

[198] Anonymous. Announcement: 10th IMACS World Congress on System Simulation and Scientific Computation: August 8–13, Montreal, Canada. Mathematics and Computers in Simulation, 23(3):331–332, September 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475481901105.

Anonymous:1981:PD

[199] Anonymous. Pages 333–412 (December 1981). Mathematics and Computers in Simulation, 23(4):??, December 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Vichnevetsky:1981:EGV

[200] R. Vichnevetsky. Energy and group velocity in semi discretizations of hyperbolic equations. Mathematics and Computers in Simulation, 23(4): 333–343, December 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900203.

Vichnevetsky:1981:PTN

[201] R. Vichnevetsky. Propagation through numerical mesh refinement for hyperbolic equations. Mathematics and Computers in Simulation, 23(4):344– 353, December 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475481900215.

Markatos:1981:TDC

[202] N. C. Markatos and T. Mukerjee. Three-dimensional computer analysis of flow and combustion in automotive internal combustion engines. Mathematics and Computers in Simulation, 23(4):354–366, December 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475481900227.

Evans:1981:SEP

[203] D. J. Evans and G. Avdelas. The solution of elliptic partial differential equations in R–θ geometry by extrapolated A.D.I. methods. Mathematics and Computers in Simulation, 23(4): 367–372, December 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900239.

Morrison:1981:SMA

[204] K. A. Morrison and N. Therien. A simple model for aquatic plant management. Mathematics and Computers in Simulation, 23(4):373–383, December 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475481900240.

Busscher:1981:SID

[205] W. J. Busscher and C. Lin. Simulated infiltration from discontinuous clogged and unclogged sources. Mathematics and Computers in Simulation, 23(4): 384–388, December 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900252.

Anonymous:1981:BRCh

[206] Anonymous. Book review: Computer Science: 21st IEEE Symposium on Foundations of Computer Science, held in Syracuse, NY, October 13–15, 1980. IEEE Service Centre, Publication Sales Department, 445 Hoes Lane, Piscataway, NJ 08854. $30.00. 80CH1498-5. Mathematics and Computers in Simulation, 23 (4):389, December 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900264.

Anonymous:1981:BRDd

[207] Anonymous. Book review: Distributed Computing: Proceedings of the 21st IEEE (COMPCON Fall 80) Computer Society International Conference held September 23–25, 1980 in Washington, DC, 746 pages, $30.00. IEEE Service Centre, Publications Sales Department, 445 Hoes Lane, Piscataway, NJ 08845. 80CH1598-2. Mathematics and Computers in Simulation, 23 (4):389, December 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900276.

Anonymous:1981:BRCi

[208] Anonymous. Book review: Computer Software and Applications: IEEE Transactions on Software Engineering, Special issue on the 3rd International IEEE Conference on Computer Software and Applications, Vol. SE-6, No. 6, November, 1980, $10.00. Mathematics and Computers in Simulation, 23(4):389, December 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900288.

Anonymous:1981:BRCj

[209] Anonymous. Book review: Computer Software and Applications: Proceedings of the IEEE 4th International Conference on Computer Software and Applications (COMPSAC 80) held October 27–31, 1980 in Chicago. 874 pages, IEEE Service Centre, Publication Sales Department, 445 Hoes Lane, Piscataway, NJ 08854. $36.00. 80CH1607-1. Mathematics and Computers in Simulation, 23 (4):389, December 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548190029X.

Anonymous:1981:BRDe

[210] Anonymous. Book review: Distributed Data Acquisition, Computing and Control: Computer Conference Record of the Symposium held in Miami Beach, 1980. 212 pages, IEEE Service Centre, Publications Sales Department, 445 Hoes Lane, Piscataway, NJ 08854, $20.00, 80CH1571-9. Mathematics and Computers in Simulation, 23(4):389, December 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900306.

Anonymous:1981:BRPg

[211] Anonymous. Book review: Parallel Processing: Proceedings of the 9th International Conference on Parallel Processing, held in Harbor Springs, MI, 1980. 349 pages. IEEE Service Centre, Publications Sales Department, 445 Hoes Lane, Piscataway, NJ 08854, $25.00. 80CH1569-3. Mathematics and Computers in Simulation, 23(4):389, December 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900318.

Anonymous:1981:BREc

[212] Anonymous. Book review: EMC Computer Models: Special issue of IEEE Transactions on Electromagnetic Compatibility, Vol. EMC-22, 1980, No. 4, November, Part 1, $10.00. Mathematics and Computers in Simulation, 23(4):389, December 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548190032X.

Anonymous:1981:BRAc

[213] Anonymous. Book review: The Auerbach Annual 1979: Best Computer Papers: 1980, 352 pages. North-Holland Publishing Company, US$35.00, ISBN 0-444-00350-9. Mathematics and Computers in Simulation, 23(4):389– 390, December 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475481900331.

Anonymous:1981:BRFc

[214] Anonymous. Book review: Firmware, Microprogramming and Restructurable Hardware: Proceedings of the IFIP Working Conference on Firmware, Microprogramming and Restructurable Hardware, Linz, Austria, 28 April–1 May 1980, 310 pages. North-Holland Publishing Company, 1980, US$39.00, ISBN 0-444-86056-8. Mathematics and Computers in Simulation, 23 (4):390, December 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900343.

Anonymous:1981:BRSc

[215] Anonymous. Book review: Simulation of Systems ’79: Proceedings of the 9th IMACS Congress in Sorrento, Italy, September 24–28, 1979, 1172 pages. North-Holland Publishing Company, 1980. US$146.25 or Dfl. 300.00, ISBN 0-444-86123-8. Mathematics and Computers in Simulation, 23(4):390, December 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900355.

Anonymous:1981:BRPh

[216] Anonymous. Book review: Practice in Software Adaptation and Maintenance: Proceedings of the SAM Workshop, Berlin, April 1979. 456 pages. North-Holland Publishing Company, 1980, US$58.50, ISBN 0-444-85449-5. Mathematics and Computers in Simulation, 23(4):390–391, December 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475481900367.

Anonymous:1981:BRCk

[217] Anonymous. Book review: Computing Methods in Applied Sciences and Engineering: Proceedings of the Fourth International Symposium on Computing Methods in Applied Sciences and Engineering, Versailles, France, December 10–14, 1979, 724 pages. North-Holland Publishing Company, 1980, US$68.25, ISBN 0-444-86008-8. Mathematics and Computers in Simulation, 23(4):391, December 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900379.

Anonymous:1981:BRCl

[218] Anonymous. Book review: Computers: Applications in Industry and Management: Proceedings of the International Seminar held at the University of Patras, Greece, 29 July– 17 August 1979, 450 pages. North-Holland Publishing Company, 1980, US$48.75, ISBN 0-444-86053-3. Mathematics and Computers in Simulation, 23(4):391, December 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475481900380.

Anonymous:1981:BRQ

[219] Anonymous. Book review: Quadratic Form Theory and Differential Equations: Academic Press, 1980, Vol. 152 in Mathematics in Science and Engineering, $29.50, 237 pages. Mathematics and Computers in Simulation, 23 (4):391–392, December 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900392.

Anonymous:1981:BRCm

[220] Anonymous. Book review: The Computer in the School: Tutor, Tool, Tutee: Teachers College Press, Columbia University, 1980, $14.95, 274 pages. Mathematics and Computers in Simulation, 23(4):392, December 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900409.

Anonymous:1981:BRR

[221] Anonymous. Book review: Rapport d’activité du groupe de travail sur “La Simulation Numérique d’ensembles de Machines Tournantes – Convertisseurs Statiques”: ERA No. 070905, I.E.E.A. de l’E.N.S.E.M., 2 rue de la Citadelle, F-54011 Nancy, France, 1981. Mathematics and Computers in Simulation, 23(4):392, December 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900410.

Anonymous:1981:Bd

[222] Anonymous. Bibliography. Mathematics and Computers in Simulation, 23 (4):393–407, December 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475481900422.

Anonymous:1981:A

[223] Anonymous. Announcement. Mathematics and Computers in Simulation, 23(4):408–409, December 1981. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475481900434.

Anonymous:1981:AIV

[224] Anonymous. Author index to volume XXIII. Mathematics and Computers in Simulation, 23(4):410–411, December 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475481900446.

Anonymous:1981:SIV

[225] Anonymous. Subject index to volume XXIII. Mathematics and Computers in Simulation, 23(4):412, December 1981. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475481900458.

Anonymous:1982:PF

[226] Anonymous. Pages 1–112 (February 1982). Mathematics and Computers in Simulation, 24(1):??, February 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1982:EB

[227] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 24(1):ii, February 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900441.

Epstein:1982:UAF

[228] C. Epstein, W. L. Miranker, and T. J. Rivlin. Ultra-arithmetic I: Function data types. Mathematics and Computers in Simulation, 24 (1):1–18, February 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900453.

Epstein:1982:UAI

[229] C. Epstein, W. L. Miranker, and T. J. Rivlin. Ultra-arithmetic II: intervals of polynomials. Mathematics and Computers in Simulation, 24 (1):19–29, February 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900465.

Korn:1982:ED

[230] Granino A. Korn. Early desire. Mathematics and Computers in Simulation, 24(1):30–36, February 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900477.

Cheng:1982:ABB

[231] Sin-I. Cheng. Asymptotic behavior and best approximation in computational fluid dynamics. Mathematics and Computers in Simulation, 24 (1):37–48, February 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900489.

Evans:1982:FSI

[232] D. J. Evans and A. Benson. Fast solution of implicit methods for linear hyperbolic equations. Mathematics and Computers in Simulation, 24 (1):49–53, February 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900490.

McWilliams:1982:SHW

[233] B. McWilliams and D. Sprevak. The simulation of hourly wind speed and direction. Mathematics and Computers in Simulation, 24(1):54– 59, February 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482900507.

Metzner:1982:GAS

[234] John R. Metzner. Generating an autocorrelated sequence of random variates without distorting their distribution. Mathematics and Computers in Simulation, 24(1):60–64, February 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482900519.

Franke:1982:EAS

[235] O. L. Franke, G. F. Pinder, and E. P. Patten. An electric-analog simulation of elliptic partial differential equations using finite element theory. Mathematics and Computers in Simulation, 24(1):65–71, February 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475482900520.

Hadjidimos:1982:SAO

[236] A. Hadjidimos and A. Yeyios. Symmetric accelerated overrelaxation (SAOR) method. Mathematics and Computers in Simulation, 24(1):72–76, February 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482900532.

Bigman:1982:BSO

[237] David Bigman. Buffer stocks operation with different supply specification: a simulation analysis. Mathematics and Computers in Simulation, 24(1):77–87, February 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475482900544.

Kaufman:1982:CSC

[238] Arie Kaufman. Continuous simulation of a car-following equation. Mathematics and Computers in Simulation, 24(1):88–94, February 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475482900556.

Anonymous:1982:BRL

[239] Anonymous. Book review: The logical design of multiple-microprocessor systems: Prentice-Hall International, 1980, 272 pages, US$29.05. ISBN 13-539908-4. Mathematics and Computers in Simulation, 24(1):

95, February 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482900568.

Anonymous:1982:BRDa

[240] Anonymous. Book review: Distributed computer systems. Impact on management, design, analysis: North-Holland Publishing Company, 1980, 380 pages US$29.50. ISBN 0-444-86109-2. Mathematics and Computers in Simulation, 24(1):95–96, February 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548290057X.

Anonymous:1982:BRT

[241] Anonymous. Book review: Tutorial on software design techniques: 3rd Edition, 1980, 455 pages, IEEE Tutorials, LC 80-81351. EH0161-0, US$25.00. Mathematics and Computers in Simulation, 24(1): 96, February 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482900581.

Anonymous:1982:BRAa

[242] Anonymous. Book review: Algorithmes et programmes d’optimisation non linéaire avec contraintes: Masson, 120, Bd. Saint-Germain, 75280 Paris, 1980, 188 pages. Mathematics and Computers in Simulation, 24(1):96, February 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900593.

Anonymous:1982:BRB

[243] Anonymous. Book review: Biological systems, modeling and control. IEE control engineering series 11: Peter Peregrinus Ltd, London, 1979, 348 pages, US$40.50. ISBN 0-90648-15-X. Mathematics and Computers in Simulation, 24(1): 96, February 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548290060X.

Anonymous:1982:BRCa

[244] Anonymous. Book review: Computer applications in medical care. Proceedings of the 4th symposium, November 2–5, 1980 in Washington DC: 3 volumes, 1980, 1945 pages. IEEE Service Centre, Publication Sales Dept., 445 Hoes Lane, Piscataway, NJ 08854. US$68.00. ISBN 80CH170-1. Mathematics and Computers in Simulation, 24(1):96–97, February 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900611.

Anonymous:1982:BRCb

[245] Anonymous. Book review: Circuits and computers. Proceedings of the 1st IEEE International Conference on Circuits and Computers, in Port Chester, NY, 1980: 2 volumes, 1980, 1200 pages. IEEE Service Centre, Publication Sales Dept., 445 Hoes Lane, Piscataway, NJ 08854, US$50.00. ISBN 80CH1511-5. Mathematics and Computers in Simulation, 24(1):97, February 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900623.

Anonymous:1982:BRMa

[246] Anonymous. Book review: Multiprocessors. A comparative study: Prentice-Hall International, 1980, 208 pages, US$29.05. ISBN 13-605154-5. Mathematics and Computers in Simulation, 24(1):97, February 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900635.

Anonymous:1982:BRPa

[247] Anonymous. Book review: Protocols and techniques for data communication networks: Prentice-Hall International, 1981, 432 pages, US$40.45. ISBN 13-731729-8. Mathematics and Computers in Simulation, 24(1): 97, February 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482900647.

Anonymous:1982:BRIa

[248] Anonymous. Book review: Introduction to SIMULA 67: Friedr. Vieweg & Sohn Verlagsgesellschaft mbH., Postfach 5829, D-6200 Wiesbaden 1. 1981, 233 pages, DM 36,- . ISBN 3-528-03340-1. Mathematics and Computers in Simulation, 24 (1):97–98, February 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475482900659.

Anonymous:1982:BRPb

[249] Anonymous. Book review: Petri Net theory and the modeling of systems: Prentice-Hall International, 1981, 288 pages, US$30.40, ISBN 13-661983-5. Mathematics and Computers in Simulation, 24(1): 98, February 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482900660.

Anonymous:1982:BRCc

[250] Anonymous. Book review: Computer simulation in university teaching. Proceedings of the FEoLL Workshop Paderborn, 28–30 January 1980: North-Holland Publishing Company, 1981, 257 pages, US$48.75, Dfl. 100.00. ISBN 0-444-86142-4. Mathematics and Computers in Simulation, 24(1): 98, February 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482900672.

Anonymous:1982:Ba

[251] Anonymous. Bibliography. Mathematics and Computers in Simulation, 24 (1):99–112, February 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475482900684.

Anonymous:1982:PA

[252] Anonymous. Pages 113–255 (April 1982). Mathematics and Computers in Simulation, 24(2–3):??, April 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Milner:1982:EI

[253] C. Milner. Editorial introduction. Mathematics and Computers in Simulation, 24(2–3):113, April 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475482900921.

Nguyen:1982:NSE

[254] V. V. Nguyen and E. F. Wood. Nonlinear simulation of environmental resources systems. Mathematics and Computers in Simulation, 24 (2–3):114–121, April 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900933.

Demaree:1982:CTO

[255] G. Demarée. Comparison of techniques for the optimization of conceptual hydrological models. Mathematics and Computers in Simulation, 24 (2–3):122–130, April 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900945.

Cluckie:1982:RTS

[256] I. D. Cluckie and R. Harpin. A real-time simulator of the rainfallrunoff process. Mathematics and Computers in Simulation, 24(2–3): 131–139, April 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482900957.

Pimentel:1982:SGF

[257] K. D. Pimentel, J. V. Candy, S. G. Azevedo, and T. A. Doerr. Simplified groundwater flow modeling: an application of Kalman filter based identification. Mathematics and Computers in Simulation, 24(2–3):140– 151, April 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482900969.

Benedini:1982:PRM

[258] M. Benedini, G. Giuliano, B. Sirangelo, and S. Troisi. Preliminary results from a methodology study in the field of groundwater resources management. Mathematics and Computers in Simulation, 24(2–3):152– 160, April 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482900970.

Herraiz:1982:TSA

[259] Andrés Sahuquillo Herráiz, Jesús López García, and Bernardo López-Camacho. Transient simulation of an aquifer connected to a wetland ecosystem. Mathematics and Computers in Simulation, 24(2–3):161–172, April 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475482900982.

Knisel:1982:SEN

[260] W. G. Knisel. Systems for evaluating nonpoint source pollution: an overview. Mathematics and Computers in Simulation, 24(2–3):173– 184, April 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482900994.

Gagliardi:1982:HDE

[261] Francesco Gagliardi, Antonio Calabro, Federico Rossi, and Piero Rostirolla. Hydrothermal dispatching and estimate of pollution. Mathematics and Computers in Simulation, 24 (2–3):185–193, April 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482901008.

Kemeny:1982:RGE

[262] Leslie G. Kemeny. A review of the global environmental impact of fossil and nuclear fuels. Mathematics and Computers in Simulation, 24 (2–3):194–203, April 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548290101X.

Nihoul:1982:EFW

[263] Jacques C. J. Nihoul. The effect of fresh water discharges from artificial reservoirs on the water quality in a polluted river. Mathematics and Computers in Simulation, 24 (2–3):204–207, April 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475482901021.

Ichikawa:1982:AUS

[264] Arata Ichikawa and Yoshikazu Sekiguchi. Administrative usage of simulation techniques for environmental protection. Mathematics and Computers in Simulation, 24(2–3):208– 214, April 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482901033.

Carpinelli:1982:EDE

[265] Guido Carpinelli, Umberto de Martinis, Antonio Piccolo, and Federico Rossi. On the environmental disturbances of EHV and UHV transmission lines. Mathematics and Computers in Simulation, 24(2–3):215– 223, April 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482901045.

Caussade:1982:MTF

[266] B. Caussade, C. Chaussavoine, and L. Masbernat. Modelling of turbulent fluxes in water surface ecosystems studies. Mathematics and Computers in Simulation, 24(2–3):224– 229, April 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482901057.

Tartaglia:1982:MGE

[267] Angelo Tartaglia. A model of the growth and evolution of woods. Mathematics and Computers in Simulation, 24(2–3):230–232, April 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482901069.

Horst:1982:MMA

[268] Thomas J. Horst. A mathematical model to assess the effects of passage of zooplankton on their respective populations. Mathematics and Computers in Simulation, 24(2– 3):233–237, April 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482901070.

Anonymous:1982:BRFa

[269] Anonymous. Book review: Fundamentals of numerical computation (Computer-oriented numerical analysis): G. Alefeld (Editor) Computing, Supplementum 2, Springer-Verlag, Wien 1980, 229 pages, AS 690.- , ISBN 3-211-81566-X. Mathematics and Computers in Simulation, 24(2–3):238, April 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482901082.

Anonymous:1982:BRMb

[270] Anonymous. Book review: Magnetic bubble technology: A. H. Eschenfelder Springer Series in Solid-State Sciences 14. Springer-Verlag, New York, 1980, 317 pages, $49.80, ISBN 3-540-09822-4. Mathematics and Computers in Simulation, 24(2– 3):238, April 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482901094.

Anonymous:1982:BRMc

[271] Anonymous. Book review: Modelling of dynamical systems — vol. 1: N. Nicholson 1980, 256 pages, £23.40; Vol. 2, 1981, 288 pages, £27.00; IEE Control Engineering Series, Peter Peregrinus Ltd., P.O. Box 26, Hitchin Herts, SG5 1SA, England. Mathematics and Computers in Simulation, 24 (2–3):238–239, April 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482901100.

Anonymous:1982:BRIb

[272] Anonymous. Book review: Induction, recursion and programming: M. Wand, North-Holland Publishing Company, Amsterdam, 1980, 202 pages, Dfl. 70.00, ISBN 0-444-00322-3. Mathematics and Computers in Simulation, 24(2–3):239, April 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482901112.

Anonymous:1982:BRSa

[273] Anonymous. Book review: Software engineering environments: H. Hunke North-Holland Publishing Company, Amsterdam, 1981, 410 pages, Dfl. 120.00, ISBN 0-444-86133-5. Mathematics and Computers in Simulation, 24(2–3):239, April 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482901124.

Anonymous:1982:BRCd

[274] Anonymous. Book review: Continued fractions — analytic theory and applications: W. B. Jones and W. J. Thron Addison-Wesley Reading, MA, 1981, 460 pages, $45.00. Mathematics and Computers in Simulation, 24(2–3):240, April 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482901136.

Anonymous:1982:BRPc

[275] Anonymous. Book review: Parallele prozesse und damit zusammenhängende automaten: W. Knodel (Editor) Springer-Verlag, Wien, 1981, 203 pages, AS 630.00, ISBN 3-211-81606-2. Mathematics and Computers in Simulation, 24(2–3): 240, April 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482901148.

Anonymous:1982:BRSb

[276] Anonymous. Book review: Stochastic and deterministic averaging processors: P. Mars and T. Poppelbaum Peter Peregrinus Ltd., G.B., 1981, 176 pages, £21.00, ISBN 0-906048-44-3. Mathematics and Computers in Simulation, 24(2–3):240, April 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548290115X.

Anonymous:1982:BRIc

[277] Anonymous. Book review: An introduction to mini and micro computers: F. Monds and R. McLaughlin Peter Peregrinus Ltd., G.B., 1981, 144 pages, £11.00, ISBN 0-906048-48-6. Mathematics and Computers in Simulation, 24(2–3):240–241, April 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475482901161.

Anonymous:1982:BRNa

[278] Anonymous. Book review: Numerical optimization of computer models: H. P. Schwefel, Wiley, New York, 1981, 389 pages, £12.50, ISBN 0-471-09988-0. Mathematics and Computers in Simulation, 24(2–3): 241, April 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482901173.

Anonymous:1982:BRU

[279] Anonymous. Book review: UKSC ’81, proceedings of the 1981 UKSC conference on computer simulation: J. Stephenson (Editor) May 13–15, 1981 in Harrogate, England. Westbury House, IPC Science and Technology Press Ltd, P.O. Box 63, Guilford, Surrey GU2 58H, U.K., 1981, 452 pages, £30.00, ISBN 0-86103-051-6. Mathematics and Computers in Simulation, 24(2–3):241, April 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482901185.

Anonymous:1982:BRSc

[280] Anonymous. Book review: Simulation of energy systems: K. E. F. Watt (Editor) Parts 1 & 2. Society for Computer Simulation, P.O. Box 2228, La Jolla, CA 92038, USA, 100 and 120 pages, $20.00 each. Mathematics and Computers in Simulation, 24 (2–3):241–242, April 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482901197.

Anonymous:1982:Bb

[281] Anonymous. Bibliography. Mathematics and Computers in Simulation, 24 (2–3):243–255, April 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482901203.

Anonymous:1982:PJ

[282] Anonymous. Pages 257–371 (June 1982). Mathematics and Computers in Simulation, 24(4):??, June 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Rubinstein:1982:ERS

[283] Y. R. Rubinstein and G. Samorodnitsky. Efficiency of the random search method. Mathematics and Computers in Simulation, 24(4):257– 268, June 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482900696.

Mezencev:1982:HTT

[284] R. Mezencev, J. Descusse, and J. F. Lafay. Hybrid test of a triangular decoupling control algorithm applied to a large scale ship boiler. Mathematics and Computers in Simulation, 24(4):269–280, June 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900702.

Chrystall:1982:SSS

[285] M. S. Chrystall and P. Mars. Simulation study of switched circuit communication networks using learning automata routing. Mathematics and Computers in Simulation, 24 (4):281–287, June 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900714.

Tu:1982:VPM

[286] Yih-O Tu. Variational principle for magnetostatic potential. Mathematics and Computers in Simulation, 24 (4):288–291, June 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900726.

Vukobratovic:1982:MRC

[287] M. Vukobratović and V. Cvetković. Manipulation robots: Computer aided design and control synthesis. Mathematics and Computers in Simulation, 24(4):292–302, June 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900738.

Stavroulakis:1982:SPL

[288] P. Stavroulakis and S. G. Tzafestas. Singularity perturbed largescale distributed-parameter control systems: an application to nuclear reactor control. Mathematics and Computers in Simulation, 24(4):303– 313, June 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548290074X.

Deans:1982:IGT

[289] N. D. Deans and D. P. Mann. An improved generation technique for random number sequences. Mathematics and Computers in Simulation, 24 (4):314–325, June 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900751.

Danaee:1982:HPF

[290] A. Danaee and D. J. Evans. Hopscotch procedures for a fourth-order parabolic partial differential equation differential equation. Mathematics and Computers in Simulation, 24(4):326–329, June 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900763.

Elhennawey:1982:IBS

[291] M. S. Elhennawey, S. E. Aidarous, and H. I. H. Shahein. Identification of bilinear systems. Mathematics and Computers in Simulation, 24 (4):330–337, June 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475482900775.

Cooke:1982:APO

[292] C. H. Cooke. On acceleration parameters for the odd-even hopscotch method. Mathematics and Computers in Simulation, 24(4):338– 340, June 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482900787.

Pimenides:1982:FDC

[293] T. G. Pimenides and S. G. Tzafestas. Feedback decoupling-controller design of 3-D systems in state space. Mathematics and Computers in Simulation, 24(4):341–352, June 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475482900799.

Anonymous:1982:BRDb

[294] Anonymous. Book review: Data handling for science and technology: S. A. Rossmannler and D. G. Watson (Editors) North-Holland Publishing Company, 1980, 184 pages, Dfl. 60.00. ISBN 0-444-86012-6. Mathematics and Computers in Simulation, 24(4):353, June 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900805.

Anonymous:1982:BRSd

[295] Anonymous. Book review: Systems engineering models of human– machine interaction: W. B. Rouse North-Holland Publishing Company, 1980, 153 pages, Dfl 75.00. ISBN 0-444-00366-5. Mathematics and Computers in Simulation, 24(4): 353, June 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482900817.

Anonymous:1982:BRDc

[296] Anonymous. Book review: Data management system modelling: T. B. Steel and J. M. Gallitano, North-Holland Publishing Company, 1981, in preparation. Mathematics and Computers in Simulation, 24(4): 354, June 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482900829.

Anonymous:1982:BRCe

[297] Anonymous. Book review: Computer simulation in university teaching, Proceedings of the FEoLLworkshop, Paderborn, Germany, 28– 30 Jan. 1980: D. Wildenberg (Editor) North-Holland Publishing Company, 1981, 260 pages, Dfl. 100.000. ISBN 0-444-86142-4. Mathematics and Computers in Simulation, 24 (4):354, June 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482900830.

Anonymous:1982:BROa

[298] Anonymous. Book review: Optimization of computer ECG processing. Proceedings of the IFIP TC-4 working conference, Halifax, Nova Scotia, Canada, 5–7 June 1979: H. K. Wolf and P. W. Macfarlane (Editors) North-Holland Publishing Company, 1980, 346 pages, Dfl. 100.000. ISBN 0-444-85413-4. Mathematics and Computers in Simulation, 24(4):354, June 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475482900842.

Anonymous:1982:BRCf

[299] Anonymous. Book review: Case studies in mathematical modelling: R. Bradley, R. G. Gibson and M. Cross (Editors) Pentech Press, 1981, 176 pages, £16.00.ISBN 7273 0311 2. Mathematics and Computers in Simulation, 24(4):355, June 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475482900854.

Anonymous:1982:BRId

[300] Anonymous. Book review: Industrial model building: L. G. Lamit, Prentice-Hall International, 1981, 380 pages, £19.45, 13-461566-2. Mathematics and Computers in Simulation, 24(4):355–356, June 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900866.

Anonymous:1982:BRPd

[301] Anonymous. Book review: Petri net theory and the modeling of systems: J. L. Peterson, Prentice-Hall International, 1981, 288 pages £14.90, 13-661983-5. Mathematics and Computers in Simulation, 24 (4):356, June 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482900878.

Anonymous:1982:BRCg

[302] Anonymous. Book review: Computer Networks. Toward Distributed Processing Systems: A. Tanenbaum Prentice-Hall software series, 1981, 544 pages. £18.20, 13-165183-8. Mathematics and Computers in Simulation, 24(4):356, June 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548290088X.

Anonymous:1982:BRPe

[303] Anonymous. Book review: Principles of programming languages: R. D. Tennent Prentice-Hall International, 1981, 280 pages, £12.95, 13-709873-1. Mathematics and Computers in Simulation, 24(4):356, June 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900891.

Anonymous:1982:BRAb

[304] Anonymous. Book review: Advances in computer methods for partial differential equations IV: R. D. Vichnevetsky and R. S. Stepleman (Editors) Proceedings of the 4th IMACS symposium on computer methods for partial differential equations, Bethlehem, USA, June 30–July 2. 1981. Published by IMACS, 418 pages. Mathematics and Computers in Simulation, 24(4):356–357, June 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900908.

Anonymous:1982:Bc

[305] Anonymous. Bibliography. Mathematics and Computers in Simulation, 24(4):358–371, June 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548290091X.

Anonymous:1982:PO

[306] Anonymous. Pages 373–422 (October 1982). Mathematics and Computers in Simulation, 24(5):??, October 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Rubinstein:1982:SNP

[307] R. Y. Rubinstein. Solution of nonlinear programming with unknown distribution function. Mathematics and Computers in Simulation, 24(5): 373–384, October 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548290026X.

Gyergyek:1982:SOC

[308] L. Gyergyek and S. Pres̆ern. Simulation and optimal control of large water distribution systems. Mathematics and Computers in Simulation, 24 (5):385–392, October 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900271.

Morrison:1982:FSM

[309] K. A. Morrison and N. Thérien. Followup on a simple model for aquatic plant management. Mathematics and Computers in Simulation, 24(5):393– 395, October 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482900283.

Belyaev:1982:TBS

[310] V. I. Belyaev. Theoretical background for the simulation of marine system dynamics. Mathematics and Computers in Simulation, 24 (5):396–401, October 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900295.

Anonymous:1982:BRFb

[311] Anonymous. Book review: Fault tolerance: T. Anderson and P. A. Lee: Prentice-Hall International, 1981, 280 pages, $14.95, ISBN 13-308254-7. Mathematics and Computers in Simulation, 24(5): 402, October 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482900301.

Anonymous:1982:BROb

[312] Anonymous. Book review: Operational research ’81: J. P. Brans (Editor): North-Holland Publishing Company, 1981, 984 pages, Dfl. 250.00, ISBN 0-444-86223-4. Mathematics and Computers in Simulation, 24(5): 402, October 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482900313.

Anonymous:1982:BRPf

[313] Anonymous. Book review: Progress in boundary element methods: C. A. Brebbia (Editor): Pentech Press, 1981, 324 pages, $28.50. ISBN 7273-1610-9. Mathematics and Computers in Simulation, 24(5):402– 403, October 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482900325.

Anonymous:1982:BRCh

[314] Anonymous. Book review: Computational algorithms for closed queuing networks: S. C. Bruell and G. Balboo: North-Holland Publishing Company, 1980, 190 pages, Dfl. 60.00. ISBN 0-444-00421-1. Mathematics and Computers in Simulation, 24(5): 403, October 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482900337.

Anonymous:1982:BRH

[315] Anonymous. Book review: Hardware/software design of digital systems: R. E. H. Bywater: Prentice-Hall International, 1981, 496 pages, $15.95. ISBN 13-383950-8. Mathematics and Computers in Simulation, 24 (5):403–404, October 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900349.

Anonymous:1982:BRNb

[316] Anonymous. Book review: Numerical methods for engineering application: J. H. Ferziger: John Wiley & Sons Ltd., G.B., 1981, 270 pages, $16.60. ISBN 0-471-06336-3. Mathematics and Computers in Simulation, 24(5):404, October 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900350.

Anonymous:1982:BRAc

[317] Anonymous. Book review: Analog I/O design, acquisition, conversion, recovery: P. H. Garrett: Prentice-Hall International, 1981, 400 pages, $14.95. ISBN 8359-0208-0. Mathematics and Computers in Simulation, 24(5):404, October 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900362.

Anonymous:1982:BRSe

[318] Anonymous. Book review: Studies on graphs and discrete programming (Annals of discrete mathematics series No.

11 — Mathematics studies series no. 59): P. Hansen (Editor) North-Holland Publishing Company, 1981, 395 pages, Dfl. 195.00. ISBN 0-444-86216-1. Mathematics and Computers in Simulation, 24(5):404–405, October 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900374.

Anonymous:1982:BRCi

[319] Anonymous. Book review: Computer science & technologies 1982: T. Kitagawa (Editor): North-Holland Publishing Company, 1982, 350 pages, Dfl. 310.00/US$124.00. Mathematics and Computers in Simulation, 24 (5):405–406, October 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900386.

Anonymous:1982:BRMd

[320] Anonymous. Book review: Managing software reliability: M. G. Walker: North-Holland Publishing Company, 1981, 252 pages, Dfl. 70.00. ISBN 0-444-00381-9. Mathematics and Computers in Simulation, 24(5): 406, October 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482900398.

Anonymous:1982:BRR

[321] Anonymous. Book review: Recursive functions in computer theory. (Computers and their application series no. 9): R. Peter: John Wiley & Sons Ltd., 1981, 179 pages, $17.50. ISBN 0-85312-164-8. Mathematics and Computers in Simulation, 24(5):406–407, October 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900404.

Anonymous:1982:BRMe

[322] Anonymous. Book review: Manmachine communication in CAD/ CAM: T. Sata and E. Warman (Editors): North-Holland Publishing Company, 1981, 274 pages, Dfl. 75.00. ISBN 0-444-86224-2. Mathematics and Computers in Simulation, 24(5): 407, October 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482900416.

Anonymous:1982:BRIe

[323] Anonymous. Book review: Introduction to the theory of error-correcting codes: V. Pless: a volume in the Wiley Intersciences Series in Discrete Mathematics. John Wiley & Sons Ltd., 1982, 164 pages, $17.00. ISBN 0-471-08684-3. Mathematics and Computers in Simulation, 24(5): 407–408, October 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482900428.

Anonymous:1982:Bd

[324] Anonymous. Bibliography. Mathematics and Computers in Simulation, 24 (5):409–422, October 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548290043X.

Anonymous:1982:PD

[325] Anonymous. Pages 423–607 (December 1982). Mathematics and Computers in Simulation, 24(6):??, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Eisenfeld:1982:EI

[326] J. Eisenfeld and V. C. Rideout. Editorial introduction. Mathematics and Computers in Simulation, 24(6): 423–424, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475482906371.

Garfinkel:1982:CBM

[327] David Garfinkel. Complex biological models: Their construction and effective use. Mathematics and Computers in Simulation, 24(6):425– 429, December 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482906383.

Yates:1982:SPI

[328] F. Eugene Yates and Arthur S. Iberall. A skeleton of physical ideas for the dynamics of complex systems. Mathematics and Computers in Simulation, 24 (6):430–436, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475482906395.

Mikulecky:1982:NTS

[329] D. C. Mikulecky. Network thermodynamic simulation of biological systems: an overview. Mathematics and Computers in Simulation, 24(6): 437–441, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482906401.

Cobelli:1982:MME

[330] C. Cobelli, L. Finkelstein, and E. R. Carson. Mathematical modelling of endocrine and metabolic systems: Model formulation, identification and validation. Mathematics and Computers in Simulation, 24(6): 442–451, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482906413.

Jacquez:1982:IPC

[331] John A. Jacquez. The inverse problem for compartmental systems. Mathematics and Computers in Simulation, 24 (6):452–459, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482906425.

Bekey:1982:EPS

[332] G. A. Bekey and F. Y. Hadaegh. Equivalence problems in structural identifiability. Mathematics and Computers in Simulation, 24(6):460–465, December 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482906437.

Delforge:1982:APL

[333] Jacques Delforge. An approach to the problem of linear system identifiability through the investigation of structural properties of the connection, injection and observation matrices. Mathematics and Computers in Simulation, 24 (6):466–471, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475482906449.

Walter:1982:GAI

[334] Eric Walter and Yves Lecourtier. Global approaches to identifiability testing for linear and nonlinear state space models. Mathematics and Computers in Simulation, 24(6):472– 482, December 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482906450.

Brown:1982:RIV

[335] R. F. Brown. Role of identifiability in validation of compartmental models. Mathematics and Computers in Simulation, 24(6):483–486, December 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482906462.

Travis:1982:PIP

[336] C. C. Travis and L. W. White. Parameter identifiability for partial differential equations. Mathematics and Computers in Simulation, 24(6):487– 489, December 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482906474.

Pohjanpalo:1982:SSI

[337] H. Pohjanpalo and B. Wahlström. Software for solving identification and identifiability problems, e. g. in compartmental systems. Mathematics and Computers in Simulation, 24(6): 490–493, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482906486.

Eisenfeld:1982:NTS

[338] J. Eisenfeld. New techniques for structural identifiability for large linear and nonlinear compartmental systems. Mathematics and Computers in Simulation, 24(6):494–501, December 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482906498.

Bassingthwaighte:1982:OUS

[339] J. B. Bassingthwaighte, M. Chaloupka, and A. A. Goldstein. Optimization using sensitivity functions for fitting models to data. Mathematics and Computers in Simulation, 24(6): 502–506, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482906504.

Ladde:1982:SVD

[340] G. S. Ladde and M. Sambandham. Stochastic versus deterministic. Mathematics and Computers in Simulation, 24(6):507–514, December 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482906516.

Matis:1982:SMT

[341] J. H. Matis, D. R. Olson, and K. B. Gerald. On the statistical moments transformation in pharmacokinetic models: a study of the rate parameter and the mean residence time estimates. Mathematics and Computers in Simulation, 24(6): 515–524, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475482906528.

Landaw:1982:OMS

[342] Elliot M. Landaw. Optimal multicompartmental sampling designs for parameter estimation: Practical aspects of the identification problem. Mathematics and Computers in Simulation, 24(6):525–530, December 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847548290653X.

DiStefano:1982:ASS

[343] Joseph J. DiStefano. Algorithms, software and sequential optimal sampling schedule designs for pharmacokinetic and physiologic experiments. Mathematics and Computers in Simulation, 24(6):531–534, December 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482906541.

Norwich:1982:PAP

[344] Kenneth H. Norwich. Perception as an active process. Mathematics and Computers in Simulation, 24(6): 535–539, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482906553.

Capasso:1982:SPB

[345] V. Capasso and L. Maddalena. Saddle point behaviour for a reaction– diffusion system: Application to a class of epidemic models. Mathematics and Computers in Simulation, 24 (6):540–547, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482906565.

Duchting:1982:MST

[346] Werner Düchting and Thomas Vogelsaenger. Modeling and simulation as a tool for optimizing the treatment of tumor diseases. Mathematics and Computers in Simulation, 24 (6):548–557, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482906577.

Mori:1982:FEA

[347] Hideki Mori, Kunio Ohtake, Shigeaki Sawai, Takashi Mineyama, Kenichiro Ryu, Kouichi Mori, Akira Takekawa, Shinzo Kitamura, and Toshiro Ono. Feature extraction and auto-regressive simulation of hepatic disordered states based on time-seriesed clinical laboratory data. Mathematics and Computers in Simulation, 24(6):558– 563, December 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482906589.

Mori:1982:SPC

[348] H. Mori, S. Kitamura, T. Nagaoka, A. Takekawa, and N. Yamaguchi. Simulation and prediction of clinical states of leukemia by the autoregression model. Mathematics and Computers in Simulation, 24(6):564– 568, December 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482906590.

Kyner:1982:IAD

[349] W. T. Kyner and G. A. Rosenberg. Isotopic analysis of diffusion and convection in brain tissue. Mathematics and Computers in Simulation, 24 (6):569–571, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475482906607.

Witten:1982:MCA

[350] Matthew Witten. Modeling cellular agign and tumorigenic transformation. Mathematics and Computers in Simulation, 24(6):572–584, December 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475482906619.

Anonymous:1982:BROc

[351] Anonymous. Book review: Ordering in two dimensions: S. K. Sinha (Editor): Proceedings of the International Conference held at Lake Geneva, Wisconsin, USA, May 28– 30, 1980. North-Holland Publishing Company, 1980, 498 pages, $55.00, ISBN 0-444-00581-1. Mathematics and Computers in Simulation, 24 (6):585, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482906620.

Anonymous:1982:BRNc

[352] Anonymous. Book review: Numerical analysis of variational inequations: R. Glowinski, R. J. Lions and R. Tremolieres: North-Holland Publishing Company, 1981, 776 pages, $109.75, ISBN 0-444-86199-8. Mathematics and Computers in Simulation, 24(6):585, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482906632.

Anonymous:1982:BRIf

[353] Anonymous. Book review: An introduction to distributed systems: B. P. Lientz: Addison-Wesley, (U.S.A.), 1981, 224 pages, ISBN 0-201-04297-5. Mathematics and Computers in Simulation, 24(6):586, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/0378475482906644.

Anonymous:1982:BRCj

[354] Anonymous. Book review: Computer performance evaluation: Tools and techniques for effective analysis: F. Morris and P. F. Roth: Van Nostrand Reinhold, New York, 1981, 288 pages, $24.95, ISBN 0-442-80325-7. Mathematics and Computers in Simulation, 24(6):586, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475482906656.

Anonymous:1982:BRSf

[355] Anonymous. Book review: Statistical modeling techniques: S. S. Shapiro and A. J. Gross: Marcel Dekker AG, Elisabethenstrasse 19, Postfach 34, CH4010 Basel, Switzerland, 1981, 384 pages, SFr. 89.00, ISBN 0-8247-1387-7. Mathematics and Computers in Simulation, 24(6):586, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475482906668.

Anonymous:1982:BRFc

[356] Anonymous. Book review: Fundamental structures of computer science: W. Wulf, M. Shaw and P. Hilfinger: Addison-Wesley, U.S.A., 1981, 640 pages, ISBN 0-201-08725-1. Mathematics and Computers in Simulation, 24 (6):586–587, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548290667X.

Anonymous:1982:BRAd

[357] Anonymous. Book review: Advances in computers: M. C. Yovits (Editor): Volume 20. Academic Press, New York, London, 1981, 304 pages, $32.00, ISBN 0-12-012120-4. Mathematics and Computers in Simulation, 24(6): 587–588, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482906681.

Anonymous:1982:BRG

[358] Anonymous. Book review: Groping in the dark: the first decade of global modelling: D. Meadows, J. Richardson and G. Bruckmann: John Wiley and Sons Ltd., 1982, 311 pages, £11.00, ISBN 0-471-10027-7. Mathematics and Computers in Simulation, 24(6):588, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475482906693.

Anonymous:1982:BRCk

[359] Anonymous. Book review: Computer aided modelling and simulation: J. A. Spriet and G. C. Vansteenkiste: Academic Press, London, 1982, 489 pages, £21.60, ISBN 0-12-659050-8. Mathematics and Computers in Simulation, 24(6):588– 589, December 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847548290670X.

Anonymous:1982:Be

[360] Anonymous. Bibliography. Mathematics and Computers in Simulation, 24 (6):590–603, December 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475482906711.

Anonymous:1982:AIV

[361] Anonymous. Author index to volume XXIV. Mathematics and Computers in Simulation, 24(6):604–606, December 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482906723.

Anonymous:1982:SIV

[362] Anonymous. Subject index to volume XXIV. Mathematics and Computers in Simulation, 24(6):607, December 1982. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475482906735.

Anonymous:1983:PF

[363] Anonymous. Pages 1–104 (February 1983). Mathematics and Computers in Simulation, 25(1):??, February 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1983:EB

[364] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 25(1):ii, February 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900228.

Keramidas:1983:EEN

[365] G. A. Keramidas and T. S. Papatheodorou. On error estimates of the numerical solution of the convection– diffusion equation. Mathematics and Computers in Simulation, 25(1):1– 9, February 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548390023X.

Arons:1983:ESS

[366] H. de Swaan Arons. Expert systems in the simulation domain. Mathematics and Computers in Simulation, 25 (1):10–16, February 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900241.

Hasenjaeger:1983:MSN

[367] E. Hasenjaeger and M. Koehne. Modeling and simulation of nonlinear and adaptive observers. Mathematics and Computers in Simulation, 25 (1):17–26, February 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900253.

Kossidas:1983:CFS

[368] A. T. Kossidas and S. A. Pactitis. Comb filters — simulation and design. Mathematics and Computers in Simulation, 25(1):27–30, February 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475483900265.

Ohkita:1983:AHF

[369] Masaaki Ohkita, Yasuhiro Kobayashi, and Michio Inoue. Application of the Haar functions to solution of differential equations. Mathematics and Computers in Simulation, 25 (1):31–38, February 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483900277.

Tolla:1983:LNL

[370] Pierre Tolla. Linear and non-linear programming software validity. Mathematics and Computers in Simulation, 25(1):39–42, February 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483900289.

Standridge:1983:ACC

[371] Charles R. Standridge and A. Alan B. Pritsker. Advanced concepts and capabilities for the analysis of simulation output. Mathematics and Computers in Simulation, 25(1):43– 45, February 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475483900290.

Gorez:1983:SDC

[372] Raymond Gorez. Simulation of a digitally controlled system, using the SIMUL language. Mathematics and Computers in Simulation, 25(1):46– 50, February 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483900307.

Baccigalupi:1983:PPM

[373] A. Baccigalupi and C. Sanges. PMICE programming methodology. Mathematics and Computers in Simulation, 25 (1):51–54, February 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900319.

Wilson:1983:VRC

[374] James R. Wilson. Variance reduction: the current state. Mathematics and Computers in Simulation, 25 (1):55–59, February 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900320.

Rolland:1983:CAS

[375] Colette Rolland, François Ambendet, Dieudonné Isso-Ibeka, and O. Thiery. A conceptual approach for simulation software. Mathematics and Computers in Simulation, 25(1):60– 62, February 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483900332.

McCormick:1983:MMV

[376] S. F. McCormick. Multigrid methods for variation problems: the V-cycle. Mathematics and Computers in Simulation, 25(1):63–65, February 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483900344.

Javor:1983:DNE

[377] A. Jávor. Dual nature of events in discrete simulation. Mathematics and Computers in Simulation, 25 (1):66–69, February 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900356.

El-Zorkany:1983:HSMa

[378] H. I. El-Zorkany and R. Balasubramanian. Hybrid simulation of multidimensional P.D.E.’s via solution of many one dimensional problems I. Theoretical basis and hybrid implementation. Mathematics and Computers in Simulation, 25 (1):70–76, February 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900368.

El-Zorkany:1983:HSMb

[379] H. I. El-Zorkany and R. Balasubramanian. Hybrid simulation of multidimensional P.D.E.’s via solution of many one dimensional problems. II. Performance study. Mathematics and Computers in Simulation, 25 (1):77–84, February 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548390037X.

Anonymous:1983:BRCa

[380] Anonymous. Book review: The correctness problem in computer science: R. S. Boyer and J. S. Moore (Editors): Academic Press, New York, London, 1981, 296 pages, US$29.50. ISBN 0-12-122920-3. Mathematics and Computers in Simulation, 25(1): 85, February 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483900381.

Anonymous:1983:BRPa

[381] Anonymous. Book review: Progress in ecological engineering and management by mathematical modelling: D. M. Dubois (Editor): Proceedings of the Second International Conference held in Liège, Belgium from 18th to 24th April 1980, Editions Cebedoc, sprl. 2, rue Armand Stévard, B-4000 Liège (Belgium), 1981, 1024 pages, 3000 BFr. ISBN 2-87080-000-2. Mathematics and Computers in Simulation, 25(1):85–86, February 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900393.

Anonymous:1983:BREa

[382] Anonymous. Book review: Eurographics ’81: J. L. Encarnaçao (Editor): North-Holland Publishing Company, Amsterdam, 1981, 360 pages, Dfl.100.00, US$42.50. ISBN 0-444-86284-6. Mathematics and Computers in Simulation, 25(1):

86, February 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847548390040X.

Anonymous:1983:BRPb

[383] Anonymous. Book review: Performance ’81: F. J. Kylstra (Editor): North-Holland Publishing Company, Amsterdam, 1981, 545 pages, DFl.150.00, US$69.75, ISBN 0-444-86330-3. Mathematics and Computers in Simulation, 25(1): 86, February 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475483900411.

Anonymous:1983:BRNa

[384] Anonymous. Book review: Nonlinear programming 4: O. L. Mangasarian, R. R. Meyer and S. M. Robinson (Editors): Academic Press New York, London, 1981, 560 pages, US$40.00. ISBN 0-12-468662-1. Mathematics and Computers in Simulation, 25(1):86–87, February 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483900423.

Anonymous:1983:BREb

[385] Anonymous. Book review: Elliptic problem solvers: M. Schultz (Editor): Academic Press, New York, London, 1981, 456 pages, US$27.50. ISBN 0-12-632620-7. Mathematics and Computers in Simulation, 25 (1):87–88, February 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900435.

Anonymous:1983:BRSa

[386] Anonymous. Book review: Software reflected. The socially responsible programming of our computers: R. L. Baber: North-Holland Publishing Company, Amsterdam, 1982, 192 pages, Dfl.80.00. ISBN 0-444-86372-9. Mathematics and Computers in Simulation, 25(1): 88, February 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483900447.

Anonymous:1983:BRCb

[387] Anonymous. Book review: Circuit theory and design: R. Boite and P. Dewilde (Editors): North-Holland Publishing Company, Amsterdam, 1982, 1090 pages, Dfl.240.00. ISBN 0-444-86307-9. Mathematics and Computers in Simulation, 25(1): 89, February 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483900459.

Anonymous:1983:BRNb

[388] Anonymous. Book review: Numerical modelling in diffusion–convection: J. Caldwell and A. O. Moscardini (Editors): Pentech Press, London, 1982, 270 pages, £18.00. ISBN 7273-1404-1. Mathematics and Computers in Simulation, 25(1): 89, February 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475483900460.

Anonymous:1983:BRCc

[389] Anonymous. Book review: Computer science and technologies 1982. Japan annual review in electronics, computers & telecommunications: T. Kitagawa (Editor): North-Holland Publishing Company, Amsterdam, 1982, 365 pages. ISBN 0167-50036. Mathematics and Computers in Simulation, 25 (1):89–90, February 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483900472.

Anonymous:1983:BRNc

[390] Anonymous. Book review: Nondifferential and variational techniques in optimization: D. C. Sorensen and R. J. B. Wets: North-Holland Publishing Company, Amsterdam, 1982, 160 pages, Dfl.60.00. ISBN 0-444-86392-3. Mathematics and Computers in Simulation, 25(1):90, February 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483900484.

Anonymous:1983:BROa

[391] Anonymous. Book review: Optical devices and fibers 1982. Japan annual review in electronics computers & telecommunications: Y. Suematsu (Editor): North-Holland Publishing Company, Amsterdam, ISSN 0167-5036. Mathematics and Computers in Simulation, 25(1):90– 91, February 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483900496.

Anonymous:1983:Ba

[392] Anonymous. Bibliography. Mathematics and Computers in Simulation, 25 (1):92–104, February 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900502.

Anonymous:1983:PAa

[393] Anonymous. Pages 105–188 (April 1983). Mathematics and Computers in Simulation, 25(2):??, April 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Saaty:1983:IMS

[394] T. L. Saaty. Introduction to a modeling of social decision processes. Mathematics and Computers in Simulation, 25(2):105–107, April 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900721.

Alexander:1983:PPC

[395] Joyce M. Alexander. Priorities and preferences in conflict resolution. Mathematics and Computers in Simulation, 25(2):108–119, April 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900733.

Gerrish:1983:PCW

[396] Wes Gerrish. Pain in creative writing: Quick relief using the analytic hierarchy process. Mathematics and Computers in Simulation, 25 (2):120–124, April 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900745.

Gholamnezhad:1983:NME

[397] A. H. Gholamnezhad. A new model in the energy policy planning. Mathematics and Computers in Simulation, 25(2):125–134, April 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900757.

Mitchell:1983:MIS

[398] Kenneth H. Mitchell and Mary P. Soye. Measuring the intangibles in social decisions: Assessing benefits and costs of energy policy options. Mathematics and Computers in Simulation, 25(2):135–145, April 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900769.

Odynocki:1983:FBP

[399] Boris Odynocki. The forward and backward processes in health policy planning. Mathematics and Computers in Simulation, 25(2):146– 155, April 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483900770.

Vargas:1983:PAH

[400] Luis G. Vargas. Prediction and the analytic hierarchy process. Mathematics and Computers in Simulation, 25(2):156–167, April 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900782.

Anonymous:1983:BRSb

[401] Anonymous. Book review: Software engineering for telecommunications switching systems: Institution of Electrical Engineers, Book Publishing Department, Southgate House, P.O. Box 8, Stevenage, Herts SG1 1HQ, 1981, 39 papers, 219 pages, UK £23.00, USA $66.00, elsewhere £26.50. ISBN 0-85296-242-8, IEE Conference Publication No. 198. Mathematics and Computers in Simulation, 25(2):169, April 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900794.

Anonymous:1983:BRAa

[402] Anonymous. Book review: Adaptive control: the modal reference approach: Y. Landau: Marcel Dekker Verlag, AG Postfach 34, CH-4010 Basel, 1981, 432 pages, SFr. 165. ISBN 0-8247-6548-6. Mathematics and Computers in Simulation, 25 (2):169–170, April 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900800.

Anonymous:1983:BRSc

[403] Anonymous. Book review: Systems modelling and optimisation: P. Nash: Peter Peregrinus Ltd. Book Publishing Department, Southgate House, P.O. Box 8, Stevenage, Herts SG1 1HQ, 1981, 224 pages. UK £14.15, USA $43.00, elsewhere £17.00. ISBN 0-906048-63-X. Mathematics and Computers in Simulation, 25 (2):170, April 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483900812.

Anonymous:1983:BRCd

[404] Anonymous. Book review: Computeraided kinetics for machine design: D. L. Ryan: Marcel Dekker, Verlag, AG, Postfach 34, CH-4010 Basel, 1981, 288 pages, SFr. 74,00. ISBN 0-8247-1421-0. Mathematics and Computers in Simulation, 25(2):170– 171, April 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483900824.

Anonymous:1983:BRLa

[405] Anonymous. Book review: Local computer network technologies: C. Tropper: Academic Press, New York, London, 1981, 176 pages, $19.50. ISBN 0-12-700850-0. Mathematics and Computers in Simulation, 25 (2):171, April 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483900836.

Anonymous:1983:BRSd

[406] Anonymous. Book review: Stochastic control by functional analysis methods: A. Bensoussan: North-Holland Publishing Company, Amsterdam, 1982, 410 pages, $58.25. ISBN 0-444-864329-X. Mathematics and Computers in Simulation, 25 (2):171–172, April 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900848.

Anonymous:1983:BREc

[407] Anonymous. Book review: Evolutionary information systems: J. Hawgood (Editor): Proceedings of the IFIP TC 8 working conference, Budapest, Hungary, 1–3 September 1981. North-Holland Publishing Company, Amsterdam, 1982, 256 pages, $35.00. ISBN 0-444-86359-1. Mathematics and Computers in Simulation, 25(2):172, April 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548390085X.

Anonymous:1983:BRMa

[408] Anonymous. Book review: Modern applied mathematics — optimization and operations research: B. Korte (Editor): North-Holland Publishing Company, Amsterdam 1982, 675 pages, DFl. 280.00. ISBN 0-444-86134-3. Mathematics and Computers in Simulation, 25(2):172, April 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900861.

Anonymous:1983:BRRa

[409] Anonymous. Book review: The relationship between numerical computation and programming languages: J. K. Reid: Proceedings of the IFIP TC2 working conference Boulder, Colorado, August 1981. North-Holland Publishing Company, 1982, Amsterdam, 368 pages, $37.25. ISBN 0-444-86377-X. Mathematics and Computers in Simulation, 25(2):172, April 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483900873.

Anonymous:1983:BRAb

[410] Anonymous. Book review: Algorithms and theory in filtering and control: D. C. Sorensen and R. J. B. Wets (Editors): North-Holland Publishing Company, Amsterdam, 1982, 160 pages, Dfl. 60.00. ISBN 0-444-86399-0. Mathematics and Computers in Simulation, 25(2):173, April 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483900885.

Anonymous:1983:BRSe

[411] Anonymous. Book review: Synthetische programmierung auf dem HP-41C/CV: W. C. Wickes: Deutsch von H. Dalkowski. DM.30.00. Helderman Verlag, Berlin, 1982, 137 pages, ISBN 3-88538-800-6. Mathematics and Computers in Simulation, 25(2):173, April 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483900897.

Anonymous:1983:Bb

[412] Anonymous. Bibliography. Mathematics and Computers in Simulation, 25(2):174–187, April 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900903.

Anonymous:1983:PJ

[413] Anonymous. Pages 189–308 (June 1983). Mathematics and Computers in Simulation, 25(3):??, June 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Avdelas:1983:OSO

[414] G. Avdelas and A. Hadjidimos. Optimum second order stationary extrapolated iterative schemes. Mathematics and Computers in Simulation, 25(3):189–198, June 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900915.

Vichnevetsky:1983:NST

[415] R. Vichnevetsky. New stability theorems concerning one-step numerical methods for ordinary differential equations. Mathematics and Computers in Simulation, 25(3):199– 205, June 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483900927.

Cooke:1983:CDA

[416] Charlie H. Cooke. A correct derivation of acceleration parameter for hopscotch and checkerboard ( P, Q )- cyclic relaxation schemes. Mathematics and Computers in Simulation, 25(3):206–209, June 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900939.

Bekey:1983:RST

[417] George A. Bekey and Sami F. Masri. Random search techniques for optimization of nonlinear systems with many parameters. Mathematics and Computers in Simulation, 25(3):210– 213, June 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483900940.

Tzafestas:1983:WTT

[418] S. G. Tzafestas. Walsh transform theory and its application to systems analysis and control: an overview. Mathematics and Computers in Simulation, 25(3):214–225, June 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900952.

McNulty:1983:CSS

[419] G. J. McNulty, D. Allen-Booth, R. Gunson, and A. Tanchou. Computer simulation of sound in polygonal shaped enclosures. Mathematics and Computers in Simulation, 25 (3):226–228, June 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900964.

Allen-Booth:1983:CSA

[420] D. Allen-Booth and G. J. McNulty. Computer simulation of acoustic intensity in a rectangular enclosure. Mathematics and Computers in Simulation, 25(3):229–231, June 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900976.

Salas:1983:RDT

[421] Manuel D. Salas. Recent developments in transonic Euler flow over a circular cylinder. Mathematics and Computers in Simulation, 25 (3):232–236, June 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900988.

Looney:1983:MCS

[422] Carl G. Looney. Monte Carlo simulation with moment matching samples. Mathematics and Computers in Simulation, 25(3):237– 240, June 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548390099X.

Selim:1983:MSN

[423] H. M. Selim, M. Mehran, K. K. Tanji, and I. K. Iskandar. Mathematical simulation of nitrogen interactions in soils. Mathematics and Computers in Simulation, 25 (3):241–248, June 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483901003.

Ferguson:1983:MST

[424] R. S. Ferguson and D. Sprevak. A model for the simulation of a two MOS-transistor inverter. Mathematics and Computers in Simulation, 25 (3):249–258, June 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483901015.

Hooker:1983:HAV

[425] J. N. Hooker, A. B. Rose, and G. F. Roberts. A holistic approach to vehicle simulation. Mathematics and Computers in Simulation, 25 (3):259–267, June 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483901027.

Briggs:1983:FDS

[426] William L. Briggs and Talib Sarie. Finite difference solutions of dispersive partial differential equations. Mathematics and Computers in Simulation, 25(3):268–278, June 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483901039.

Breitenecker:1983:CST

[427] F. Breitenecker. The concept of supermacros in today’s and future simulation languages. Mathematics and Computers in Simulation, 25 (3):279–289, June 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483901040.

Anonymous:1983:BRCe

[428] Anonymous. Book review: Circuit theory and design: R. Boite and P. Dewilde (Editors): North-Holland Publishing Company, Amsterdam, 1981, 1090 pages, US $102.00/ Dfl. 240.00, ISBN 0-444-86307-9. Mathematics and Computers in Simulation, 25(3):290, June 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483901052.

Anonymous:1983:BRCf

[429] Anonymous. Book review: Computer hardware description languages (CHDL) and their applications: M. Brever and R. Hartenstein (Editors): North-Holland Publishing Company, Amsterdam, 1981, 250 pages, Dfl. 85.00, ISBN 0-444-86279-X. Mathematics and Computers in Simulation, 25(3):290, June 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483901064.

Anonymous:1983:BRPc

[430] Anonymous. Book review: Progress in modelling and simulation: F. E. Cellier (Editor): Academic Press, London, 1982, 466 pages, £30.00, ISBN 0-12-164780-3. Mathematics and Computers in Simulation, 25 (3):291, June 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475483901076.

Anonymous:1983:BRDa

[431] Anonymous. Book review: Discrete simulation and related fields: A. Javor (Editor): North-Holland Publishing Company, Amsterdam, 1982, 246 pages, Dfl. 75.00, ISBN 0-444-86429-6. Mathematics and Computers in Simulation, 25(3):291, June 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483901088.

Anonymous:1983:BRV

[432] Anonymous. Book review: Variational analysis and approximate solutions for transport phenomenas: G. A. Keramidas: Naval Research Laboratory Memorandum No. 4852, 1982, 82 pages, Washington, DC. Mathematics and Computers in Simulation, 25(3):291, June 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548390109X.

Anonymous:1983:BRFa

[433] Anonymous. Book review: Fifth generation computer systems: T. Moto-Oka (Editor): North-Holland Publishing Company, Amsterdam, 1982, 286 pages, Dfl. 149.00, ISBN 0-444-86440-7. Mathematics and Computers in Simulation, 25(3):291– 292, June 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475483901106.

Anonymous:1983:BRDb

[434] Anonymous. Book review: Distributed data bases: H. L. Schneider (Editor): North-Holland Publishing Company, Amsterdam, 1982, 364 pages, Dfl. 105.00, ISBN 0-444-86474-1. Mathematics and Computers in Simulation, 25(3):292, June 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483901118.

Anonymous:1983:BROb

[435] Anonymous. Book review: Optimization and control of dynamic operational research models: S. G. Tzafestas (Editor): North-Holland Publishing Company, Amsterdam, 1982, 436 pages, Dfl. 150.00, ISBN 0-444-86380-X. Mathematics and Computers in Simulation, 25(3):292–293, June 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548390112X.

Anonymous:1983:BRMb

[436] Anonymous. Book review: Mathematics of finite elements and applications IV: J. R. Whiteman: Academic Press, New York, 1982, 600 pages, US$51.50, ISBN 0.12.747254.1. Mathematics and Computers in Simulation, 25(3):293, June 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483901131.

Anonymous:1983:Bc

[437] Anonymous. Bibliography. Mathematics and Computers in Simulation, 25(3):294–308, June 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483901143.

Anonymous:1983:PAb

[438] Anonymous. Pages 309–404 (August 1983). Mathematics and Computers in Simulation, 25(4):??, August 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Mulpuru:1983:FSM

[439] S. R. Mulpuru. A flux splitting method for the numerical simulation of onedimensional compressible flow. Mathematics and Computers in Simulation, 25(4):309–320, August 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483900514.

Rubinstein:1983:AOM

[440] Y. R. Rubinstein and J. Kreimer. About one Monte Carlo method for solving linear equations. Mathematics and Computers in Simulation, 25 (4):321–334, August 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483900526.

Wysocki:1983:AOC

[441] Marian Wysocki. Application of orthogonal collocation to simulation and control of first order hyperbolic systems. Mathematics and Computers in Simulation, 25(4):335– 345, August 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483900538.

Eitelberg:1983:SSN

[442] Eduard Eitelberg. A simple A _{2}-stable numerical method for state space models with stiff oscillations. Mathematics and Computers in Simulation, 25 (4):346–355, August 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548390054X.

Sehitoglu:1983:AAS

[443] H. Sehitoglu. An alternative approach to state estimation in linear time-invariant systems. Mathematics and Computers in Simulation, 25 (4):356–365, August 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900551.

Grant:1983:VRT

[444] Floyd H. Grant and James J. Solberg. Variance reduction techniques in stochastic shortest route analysis: application procedures and results. Mathematics and Computers in Simulation, 25(4):366–375, August 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900563.

Carotenuto:1983:ATE

[445] Luciano Carotenuto, Alfredo Eisinberg, and Giancarlo Raiconi. On the approximation of the transport equation by the method of weighted residuals. Mathematics and Computers in Simulation, 25(4):376–383, August 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900575.

Anonymous:1983:BRAc

[446] Anonymous. Book review: Application and implementation of finite element methods: J. E. Akin Academic Press, New York, 1982, 272 pages, $34.50, ISBN 0.12.047650.9. Mathematics and Computers in Simulation, 25(4):384, August 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900587.

Anonymous:1983:BRLb

[447] Anonymous. Book review: Logic programming: K. L. Clark and S.A Tarnlund(Editor) Academic Press, New York, 1982, 366 pages, $34.50, ISBN 0.12.175520.7. Mathematics and Computers in Simulation, 25 (4):384–385, August 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900599.

Anonymous:1983:BRCg

[448] Anonymous. Book review: Computer network architectures and protocols: P. E. Green (Editor) Plenum Publishing Co., New York, 1982, 725 pages, $71.40, ISBN 0-306-40788-4. Mathematics and Computers in Simulation, 25(4):385, August 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900605.

Anonymous:1983:BRCh

[449] Anonymous. Book review: Control, identification and input optimization: R. Kalaba and K. Spingarn, Plenum Publishing Co., New York, 1982, 450 pages, $47.50, ISBN 0-306-40847-3. Mathematics and Computers in Simulation, 25(4): 385, August 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483900617.

Anonymous:1983:BRIa

[450] Anonymous. Book review: Involving micros in education: R. Lewis and E. D. Tagg, North-Holland Publishing Company, Amsterdam, 1982, 248 pages, Dfl. 85.00, US$39.50, ISBN 0-444-86459-8. Mathematics and Computers in Simulation, 25 (4):385–386, August 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900629.

Anonymous:1983:BRNd

[451] Anonymous. Book review: Numerical solutions of partial differential equations: B. J. Noye Proceedings of the 1981 Conf., Queens College, Melbourne University, Australia. North-Holland Publishing Company, Amsterdam, 1982, 648 pages, $93.00, ISBN 0-444-86356-7. Mathematics and Computers in Simulation, 25(4):386, August 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483900630.

Anonymous:1983:BRAd

[452] Anonymous. Book review: Applications of variational inequalities in stochastic control: A. Bensoussan and J. L. Lions: North-Holland Publishing Company, Amsterdam, 1982, 564 pages, $81.50 ISBN 0-444-86358-3. Mathematics and Computers in Simulation, 25 (4):386–387, August 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483900642.

Anonymous:1983:BRLc

[453] Anonymous. Book review: Local computer networks: P. Ravasio, G. Hopkins and N. Naffah (Editors): North-Holland Publishing Company, Amsterdam, 1982, 563 pages, Dfl. 140.00, US$65.00, ISBN 0.444.86386.9. Mathematics and Computers in Simulation, 25(4):387, August 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483900654.

Anonymous:1983:BRBa

[454] Anonymous. Book review: Biomedical statistics with computing: M. H. Regier, R. N. Mohapatra and S. N. Mohapatra: Wiley Ltd, England, 1982, 308 pages, £14.50, ISBN 0.471.10449.3. Mathematics and Computers in Simulation, 25(4): 387, August 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483900666.

Anonymous:1983:BREd

[455] Anonymous. Book review: Environmental systems analysis and management: S. Rinaldi (Editor): North-Holland Publishing Company, Amsterdam, 1982, 828 pages, Dfl. 200.00, US$93.00. ISBN 0-444-86406-7. Mathematics and Computers in Simulation, 25(4):387–388, August 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483900678.

Anonymous:1983:BRAe

[456] Anonymous. Book review: Automated tools for information systems design: H. J. Scheneider and A. I. Wasserman (Editors) Proceedings of the IFIP WG 8.1 Working Conf. New Orleans, USA, 26–28 Jan, 1982. North-Holland Publishing Company, Amsterdam 1982, 262 pages, Dfl. 85.00, US$39.50, ISBN 0-444-86338-9. Mathematics and Computers in Simulation, 25(4):388, August 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548390068X.

Anonymous:1983:BRDc

[457] Anonymous. Book review: Distributed parameter control systems: theory and applications: S. G. Tzafestas (Editor) International Series on Systems and Control, Volume 6. Pergamon Press, Oxford (U.K.), 1982, 497 pages, US$70.00, ISBN 0-08-027624-5. Mathematics and Computers in Simulation, 25(4):388–389, August 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483900691.

Anonymous:1983:BRDd

[458] Anonymous. Book review: Distributed data sharing systems: R. P. van de Riet and W. Litwin (Editors) Proceedings of the 2nd Int. Seminar, Amsterdam, 3– 5 June 1981. North-Holland Publishing Company, 1982, 320 pages, Dfl. 100.00, US$46.50, ISBN 0-444-86374-5. Mathematics and Computers in Simulation, 25(4):389, August 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483900708.

Anonymous:1983:Bd

[459] Anonymous. Bibliography. Mathematics and Computers in Simulation, 25 (4):390–403, August 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548390071X.

Anonymous:1983:PO

[460] Anonymous. Pages 405–480 (October 1983). Mathematics and Computers in Simulation, 25(5):??, October 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Lipitakis:1983:IIM

[461] Elias A. Lipitakis. Isomorphic iterative methods in solving singularly perturbed elliptic difference equations. Mathematics and Computers in Simulation, 25(5):405– 416, October 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483901398.

Findler:1983:MFE

[462] Nicholas V. Findler and Ernesto Morgado. Morph-fitting: an effective technique of approximation. Mathematics and Computers in Simulation, 25 (5):417–429, October 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483901404.

Lin:1983:SEF

[463] Chenfang Lin, Warren J. Busscher, and Lowell A. Douglas. Simulated effects of flow rate on phosphate retention in soils. Mathematics and Computers in Simulation, 25(5):430– 435, October 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483901416.

DAlessio:1983:NMS

[464] A. D’Alessio, P. L. De Angelis, A. Murli, and M. A. Pirozzi. Numerical models for the simulation of currents in Naples Gulf. Mathematics and Computers in Simulation, 25 (5):436–451, October 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483901428.

Spriet:1983:SSS

[465] J. A. Spriet and P. Herman. A simulation study of structure characterization methods. Mathematics and Computers in Simulation, 25(5): 452–459, October 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548390143X.

Anonymous:1983:BRCi

[466] Anonymous. Book review: Computational Hydraulics: C. A. Brebbia and A. J. Ferrante: Butterworths, London, 1982, 304 pages, £20.00. ISBN 0-408-01153-X. Mathematics and Computers in Simulation, 25(5):460, October 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483901441.

Anonymous:1983:BRSf

[467] Anonymous. Book review: The Specification of Complex Systems: B. Cohen: Addison-Wesley, Amsterdam, 1982, 356 pages. ISBN 0-201-14400-X. Mathematics and Computers in Simulation, 25(5):460–461, October 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483901453.

Anonymous:1983:BRFb

[468] Anonymous. Book review: File Structures and Data Bases for CAD: J. Encarnaçao and F. I. Krause: North-Holland, Amsterdam, 1982, 372 pages, Dfl. 110.00. ISBN 0-444-86462-8. Mathematics and Computers in Simulation, 25(5):461, October 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483901465.

Anonymous:1983:BRMc

[469] Anonymous. Book review: Mathematical Programming with Data Perturbations I: A. V. Fiacco (Editor): Lecture Notes in Pure and Applied Mathematics Series, Vol. 73, Marcel Dekker AG, Basel, Switzerland, 1982, 256 pages, SFr. 98.-. ISBN 0-8247-1543-8. Mathematics and Computers in Simulation, 25(5):461–462, October 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483901477.

Anonymous:1983:BRCj

[470] Anonymous. Book review: Computers in Cardiology: IEEE Publication: Proceedings of the 8th Conference held Sept. 23–25, 1981 in Florence, Italy, 1982, 576 pages, $36.00. ISSN 0276-6574. Mathematics and Computers in Simulation, 25(5):462, October 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483901489.

Anonymous:1983:BRCk

[471] Anonymous. Book review: Computer Architecture: IEEE Publication: Proceedings of the 9th Symposium on Computer Architecture, held April 26– 29, 1982 in Austin, TX, USA, 335 pages, $28.00. ISSN 0149-7111. Mathematics and Computers in Simulation, 25(5):462, October 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483901490.

Anonymous:1983:BRCl

[472] Anonymous. Book review: Computers and Communications: IEEE Publication: Proceedings of the 1st Phoenix Conference on Computers and Communications, held May 9– 12, 1982 in Phoenix, AZ, U.S.A., 444 pages. $32.00. Mathematics and Computers in Simulation, 25(5): 462, October 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475483901507.

Anonymous:1983:BRFc

[473] Anonymous. Book review: Fuzzy Techniques in Pattern Recognition: A. Kandel: Wiley, Chichester, England, 1982, 353 pages, 29.00. Mathematics and Computers in Simulation, 25 (5):462–463, October 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483901519.

Anonymous:1983:BRPd

[474] Anonymous. Book review: Parallel Algorithms on a Multiprocessor: P. N. Oleinick: Bowker Publishing Co., Epping, England, 1982, 126 pages. ISBN 0-8357-1327. Mathematics and Computers in Simulation, 25(5):463, October 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483901520.

Anonymous:1983:BRIb

[475] Anonymous. Book review: Introduction to Computer Simulation: the Systems Dynamics Approach: N. Roberts et al.: Addison-Wesley, Amsterdam, 1982. ISBN 0-201-06414-6. Mathematics and Computers in Simulation, 25(5):463, October 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483901532.

Anonymous:1983:BRRb

[476] Anonymous. Book review: On Reduction and Livelocks in Asynchronous Parallel Computation: Y.- S. Kwong: Bowker Publishing Co., Epping, England, 1983, 215 pages. ISBN 0-8357-1364. Mathematics and Computers in Simulation, 25(5): 463, October 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483901544.

Anonymous:1983:BRSg

[477] Anonymous. Book review: Systems and Models for Energy and Environmental Analysis: T. R. Lakshmanan and P. Nijkamp: Gower Publishing Company, Ltd., Alder-shot, England, 1983, 230 pages, £12.50/ $26.50. ISBN 0-566-00558-1. Mathematics and Computers in Simulation, 25(5):463–464, October 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483901556.

Anonymous:1983:BRCm

[478] Anonymous. Book review: Computational Methods in Elementary Numerical Analysis: J. L. I. Morris: Wiley, Chichester, England, 1983, 410 pages, £9.95. ISBN 0-471-10419-1. Mathematics and Computers in Simulation, 25 (5):464–465, October 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483901568.

Anonymous:1983:BRDe

[479] Anonymous. Book review: Disposition von Blutkonserven — Ein Modellesystem und dessen Anwendungsaspekte: H. Vollert: Paul Haupt Verlag, Bern, Stuttgart, 1983, 104 pages, SFR. 28.00/DM 34.00. Mathematics and Computers in Simulation, 25(5):465, October 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548390157X.

Anonymous:1983:Be

[480] Anonymous. Bibliography. Mathematics and Computers in Simulation, 25 (5):466–479, October 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475483901581.

Anonymous:1983:PD

[481] Anonymous. Pages 481–571 (December 1983). Mathematics and Computers in Simulation, 25(6):??, December 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Cooke:1983:ISS

[482] Charlie H. Cooke. Invariance of stiff stability under a small parameter variation. Mathematics and Computers in Simulation, 25(6):481– 484, December 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483901155.

Field:1983:DIT

[483] David A. Field. Determining invertible two and three dimensional quadratic isoparametric finite element transformations. Mathematics and Computers in Simulation, 25(6):485– 488, December 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483901167.

Jones:1983:MER

[484] John Jones. Matrix equations over rings. Mathematics and Computers in Simulation, 25(6):489–492, December 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483901179.

Austman:1983:MTD

[485] G. Austman, J. Szymanski, and W. I. Midvidy. Modcir: a threedimensional transient code used in thermal-hydraulic analyses of CANDU power reactors. Mathematics and Computers in Simulation, 25(6):493– 496, December 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483901180.

Maslowski:1983:SID

[486] Andrzej Maslowski. On state identification in distributed dynamics systems. Mathematics and Computers in Simulation, 25(6):497–501, December 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483901192.

Kalashnikov:1983:OSS

[487] V. V. Kalashnikov and B. V. Nemchinov. The organization of system studies by aggregative simulation system. Mathematics and Computers in Simulation, 25(6):502– 512, December 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483901209.

Vasilev:1983:PMM

[488] A. S. Vasil’ev, G. N. Deinega, and R. T. Fedotov. Prognostic model of the marine ecosystem state as applied to fisheries. Mathematics and Computers in Simulation, 25(6): 513–516, December 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483901210.

Faist:1983:DWL

[489] M. B. Faist, L. Gevirtzman, and C. E. Riese. Development of a wet lime/limestone flue gas desulfurization process model. Phase I: Species distribution model. Mathematics and Computers in Simulation, 25(6):517– 522, December 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483901222.

Chapyak:1983:VSE

[490] E. J. Chapyak, P. J. Blewett, and D. J. Cagliostro. Verification studies of entrained-flow gasification and combustion systems with the simmer-II code. Mathematics and Computers in Simulation, 25(6):523–526, December 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483901234.

Simons:1983:SPM

[491] Girard A. Simons and Paul F. Lewis. Single particle model for char oxidation. Mathematics and Computers in Simulation, 25(6):527–530, December 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483901246.

Delledonne:1983:CSC

[492] M. Delledonne, P. Nicoletti, and T. O’Brien. Computer simulation of a cold fluidized bed. Mathematics and Computers in Simulation, 25(6):

531–534, December 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483901258.

Klein:1983:TDR

[493] H. H. Klein and M. F. Scharff. A timedependent, reactive model of fluidizedbed chemical reactors. Mathematics and Computers in Simulation, 25(6): 535–537, December 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548390126X.

Ramos:1983:RSN

[494] J. I. Ramos. A review of some numerical methods for reaction–diffusion equations. Mathematics and Computers in Simulation, 25(6):538–548, December 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483901271.

Anonymous:1983:BRIc

[495] Anonymous. Book review: Identifiability of state space models with application to transformation systems: Springer-Verlag, Berlin, 1982, 202 pages, DM 32.50, ISBN 3-540-11590-0. Mathematics and Computers in Simulation, 25(6):549, December 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483901283.

Anonymous:1983:BRBb

[496] Anonymous. Book review: Basic computational techniques for engineers: Pentech Press, London, 1983, 208 pages, £12.50, ISBN 7273-0203-5. Mathematics and Computers in Simulation, 25(6):549, December 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483901295.

Anonymous:1983:BRSh

[497] Anonymous. Book review: Surfaces in computer aided geometric design: North-Holland, Amsterdam, 1983, 216 pages, Dfl. 110.00, ISBN 0-444-86550-0. Mathematics and Computers in Simulation, 25(6):549– 550, December 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483901301.

Anonymous:1983:BRPe

[498] Anonymous. Book review: Progress in boundary element methods — volume 2: Pentech Press, London, 1983, 226 pages, £28.50, ISBN 7273-1611-7. Mathematics and Computers in Simulation, 25 (6):550, December 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483901313.

Anonymous:1983:BRCn

[499] Anonymous. Book review: Computer programs for electronic analysis and design: Prentice-Hall International, England, 1983, 272 pages, £15.25, ISBN 8359-9874-7. Mathematics and Computers in Simulation, 25 (6):550–551, December 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483901325.

Anonymous:1983:BRMd

[500] Anonymous. Book review: Multirate digital signal processing: Prentice-Hall International, England, 1983, 336 pages, £24.60, ISBN 13-605162-

6. Mathematics and Computers in Simulation, 25(6):551, December 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475483901337.

Anonymous:1983:BREe

[501] Anonymous. Book review: Error control coding: Prentice-Hall International, England, 1983, 602 pages, £35.95, ISBN 13-283796-X. Mathematics and Computers in Simulation, 25(6):551, December 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483901349.

Anonymous:1983:BRNe

[502] Anonymous. Book review: Numerical methods in aeronautical fluid dynamics: Academic Press, London, 1983, 548 pages, £30.00, ISBN 0-12-592520-4. Mathematics and Computers in Simulation, 25(6):551, December 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483901350.

Anonymous:1983:Bf

[503] Anonymous. Bibliography. Mathematics and Computers in Simulation, 25 (6):552–565, December 1983. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475483901362.

Anonymous:1983:AIV

[504] Anonymous. Author index to volume XXV. Mathematics and Computers in Simulation, 25(6):566–568, December 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483901374.

Anonymous:1983:CVX

[505] Anonymous. Contents volume XXV. Mathematics and Computers in Simulation, 25(6):569–571, December 1983. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475483901386.

Anonymous:1984:PF

[506] Anonymous. Pages 1–85 (February 1984). Mathematics and Computers in Simulation, 26(1):??, February 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1984:EB

[507] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 26(1):iv, February 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900880.

LaVita:1984:SSR

[508] James A. LaVita. Simulation of a simple reacting gas flow with an upwind scheme. Mathematics and Computers in Simulation, 26(1):1– 10, February 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900892.

Roubicek:1984:HSW

[509] Tomás̆ Roubícĕk. Hybrid solution of weakly formulated boundaryvalue problems. Mathematics and Computers in Simulation, 26(1):11– 19, February 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900909.

Angus:1984:CBB

[510] J. E. Angus. The connection between the Barnard-Birnbaum Monte Carlo test and the two-sample Kolmogorov– Smirnov test. Mathematics and Computers in Simulation, 26(1):20– 22, February 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900910.

Dunn:1984:SLS

[511] Floyd E. Dunn and Frederick G. Prohammer. The SASSYS LMFBR systems analysis code. Mathematics and Computers in Simulation, 26 (1):23–26, February 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900922.

Tzafestas:1984:OCC

[512] S. G. Tzafestas. Optimal control of ε-coupled and singularly perturbed distributed-parameter systems. Mathematics and Computers in Simulation, 26(1):27–38, February 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900934.

Argintaru:1984:STP

[513] V. Argintaru and D. H. Willis. Simulation of the tidal propagation in the Bay of Fundy using a hybrid model. Mathematics and Computers in Simulation, 26(1):39–42, February 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900946.

Zamani:1984:RIF

[514] Nader Zamani. The role of implicitness in a finite element calculation of viscous flow. Mathematics and Computers in Simulation, 26 (1):43–53, February 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900958.

Basanez:1984:MSM

[515] L. Basañez, J. Riera, and J. Ayza. Modelling and simulation of multiarea power system load- frequency control. Mathematics and Computers in Simulation, 26(1):54–62, February 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548490096X.

Raffort:1984:CHS

[516] Lionel Raffort and Bernard Lang. Calculateur hybride en simulation de systemes continus complexes par methode de decomposition et coordination par relaxation. Mathematics and Computers in Simulation, 26 (1):63–71, February 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900971.

Anonymous:1984:BRNa

[517] Anonymous. Book review: Numerical models in geomechanics: R. Dungar, C. N. Pande and J. A. Studer (Editors): Proceedings of the international symposium of Zurich, 13– 17 September 1982. A. A. Balkema Publishers, Rotterdam Netherlands, 1982, 831 pages, Dfl.125.00, $48.00, ISBN 90-6191-225-3. Mathematics and Computers in Simulation, 26(1): 72, February 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900983.

Anonymous:1984:BRNb

[518] Anonymous. Book review: Numerical methods in geomechanics: Z. Eisenstein (Editor): Proceedings of the international conference of Edmonton, Canada, 31 May–4 June 1982. A. A. Balkema Publishers, Rotterdam, Netherlands, 1982, 1300 pages (3 Vols.) Dfl. 365.00, $145.00, ISBN 90-6191-246-6. Mathematics and Computers in Simulation, 26(1): 72, February 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900995.

Anonymous:1984:BRPa

[519] Anonymous. Book review: Performance control: T. A. Hoie: North-Holland, Amsterdam., 1982, 252 pages, Dfl.100.00, ISBN 0-444-86517-9. Mathematics and Computers in Simulation, 26(1):72–75, February 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484901009.

Anonymous:1984:BRPb

[520] Anonymous. Book review: Problem complexity and method efficiency in optimization: A. S. Nemirovsky and D. B. Yudin: Wiley, New York, 1983, 388 pages, £26.00, ISBN 0-471-10345-4. Mathematics and Computers in Simulation, 26(1):75, February 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484901010.

Anonymous:1984:BRSa

[521] Anonymous. Book review: Simulation with GPSS: K. Weber, R. Trzebiner and H. Temperlmeier: Haupt Verlag, Bern, Switzerland, 1983, ISBN 3-258-03076-6. Mathematics and Computers in Simulation, 26(1): 75, February 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475484901022.

VanRemoortere:1984:Ba

[522] P. Van Remoortere. Bibliography. Mathematics and Computers in Simulation, 26(1):76–85, February 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/0378475484901034.

Anonymous:1984:PAa

[523] Anonymous. Pages 87–179 (April 1984). Mathematics and Computers in Simulation, 26(2):??, April 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Kinnmark:1984:OSIa

[524] Ingemar P. E. Kinnmark and William G. Gray. One step integration methods with maximum stability regions. Mathematics and Computers in Simulation, 26(2):87–92, April 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475484900399.

Vichnevetsky:1984:EFE

[525] R. Vichnevetsky. The energy flow equation. Mathematics and Computers in Simulation, 26(2):93– 101, April 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900405.

Shonkwiler:1984:SOB

[526] R. Shonkwiler. Simulating the outcome of a birth and death process. Mathematics and Computers in Simulation, 26(2):102–112, April 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900417.

Sydow:1984:DCA

[527] Achim Sydow. Decentralized control approach to complex planning processes. Mathematics and Computers in Simulation, 26(2):113– 116, April 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900429.

Chauhan:1984:SIN

[528] N. S. Chauhan. Simulation of ionosphere: a numerical technique. Mathematics and Computers in Simulation, 26(2):117–121, April 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900430.

Lalwani:1984:MCS

[529] C. S. Lalwani and J. M. Brown. Models for crude steel production in the United Kingdom. Mathematics and Computers in Simulation, 26 (2):122–135, April 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475484900442.

Wasson:1984:CMM

[530] K. W. Wasson and T. Prasad. A computerized mathematical model for studying nonlinear functional characteristics of skeletal muscle activity. Mathematics and Computers in Simulation, 26(2):136– 147, April 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475484900454.

Furuta:1984:DRH

[531] Katsuhisa Furuta, Yasuhiro Ohyama, and Osamu Yamano. Dynamics of RC helicopter and control. Mathematics and Computers in Simulation, 26(2):148–159, April 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475484900466.

Tyran:1984:MLS

[532] Jean-Luc Tyran. MAXI-LILLIPUT: a simplified model of a French pressurized water reactor power plant. Mathematics and Computers in Simulation, 26(2):160–163, April 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475484900478.

Anonymous:1984:BREa

[533] Anonymous. Book review: Engineering software III: R. A. Adey (Editor): Proceedings of the Third International Conference & Exhibition, Imperial College, London, April 1983. CML Publications, Ashurst Lodge, Ashurst, New Forest, Hampshire, S04 2AA, England, 100 pages, £44.00, ISBN 0-905451-15-5. Mathematics and Computers in Simulation, 26(2):164, April 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548490048X.

Anonymous:1984:BRNc

[534] Anonymous. Book review: Nonlinear stochastic problems: R. S. Bucy and J. M. F. Moura (Editors): Reidel, Dordrecht, Holland, 1983, 623 pages. Dfl.183.00. ISBN 90-277-1590-4. Mathematics and Computers in Simulation, 26(2):164, April 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900491.

Anonymous:1984:BRQ

[535] Anonymous. Book review: qhypergeometric functions and applications: H. Exton: Wiley, New York. 1983, 347 pages, US$22.50. ISBN 0-85312-491-4. Mathematics and Computers in Simulation, 26 (2):164–167, April 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900508.

Anonymous:1984:BRAa

[536] Anonymous. Book review: Algorithmstructured computer arrays and networks. Architectures and processes for image, percepts, models information: L. Uhr: Academic Press, London, 1933, 363 pages, ISBN 0-12-706960-7. Mathematics and Computers in Simulation, 26(2):167, April 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548490051X.

Anonymous:1984:BRF

[537] Anonymous. Book review: Functional analysis, holomorphy and approximation theory: G. L. Zapata Marcel Dekker, Basel, 1983. 472 pages, SFr.156.00. ISBN 0-8247-1634-5. Mathematics and Computers in Simulation, 26(2):167, April 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475484900521.

Anonymous:1984:Ba

[538] Anonymous. Bibliography: 1981. Mathematics and Computers in Simulation, 26(2):168–169, April 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/0378475484900533.

Anonymous:1984:Bb

[539] Anonymous. Bibliography: 1982. Mathematics and Computers in Simulation, 26(2):169–172, April 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475484900545.

Anonymous:1984:Bc

[540] Anonymous. Bibliography: 1983. Mathematics and Computers in Simulation, 26(2):172–179, April 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475484900557.

Anonymous:1984:PJ

[541] Anonymous. Pages 181–318 (June 1984). Mathematics and Computers in Simulation, 26(3):??, June 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Kinnmark:1984:OSIb

[542] Ingemar P. E. Kinnmark and William G. Gray. One step integration methods of third-fourth order accuracy with large hyperbolic stability limits. Mathematics and Computers in Simulation, 26(3):181–188, June 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900569.

Nedoma:1984:OTF

[543] Jiři Nedoma. On one type of finite element simulation in dynamic elasticity. Mathematics and Computers in Simulation, 26(3):189– 209, June 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900570.

Babary:1984:MNS

[544] Jean-Pierre Babary and Jorge Lopes de Souza Leao. Modelling and numerical simulation of an epitaxial reactor. Mathematics and Computers in Simulation, 26(3):210– 217, June 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900582.

Mahrer:1984:NSG

[545] Y. Mahrer and R. Avissar. A numerical simulation of the greenhouse microclimate. Mathematics and Computers in Simulation, 26 (3):218–228, June 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900594.

Margeta:1984:COS

[546] Jure Margeta. Coupled optimizationsimulation water quality model for regional water quality management. Mathematics and Computers in Simulation, 26(3):229–242, June 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475484900600.

Vignes:1984:EIO

[547] Jean Vignes. An efficient implementation of optimization algorithms. Mathematics and Computers in Simulation, 26(3):243–256, June 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900612.

Missirlis:1984:MPS

[548] Nikolaos M. Missirlis and David J. Evans. The modified preconditioned simultaneous displacement (MPSD) method. Mathematics and Computers in Simulation, 26(3):257– 262, June 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900624.

Cvetkovic:1984:RLM

[549] V. Cvetković. Robot language for manipulation systems programming, teaching, and verification of control algorithms. Mathematics and Computers in Simulation, 26(3):263– 275, June 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900636.

Rajagopal:1984:PPF

[550] K. R. Rajagopal and E. Sciubba. Pulsating Poiseuille flow of a non-Newtonian fluid. Mathematics and Computers in Simulation, 26(3):276– 288, June 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900648.

Richard:1984:DST

[551] M. J. Richard and R. J. Anderson. Dynamic simulation of threedimensional rigid body systems using Vector-Network techniques. Mathematics and Computers in Simulation, 26(3):289–296, June 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548490065X.

Anonymous:1984:BRSb

[552] Anonymous. Book review: Software engineering problems: R. A. Adey (Editor): Computational mechanics, Ashurst Lodge, Ashurst, Southampton, SO4 2AA, England, 1983 150 pages, £15.00, ISBN 0-905451-16-3. Mathematics and Computers in Simulation, 26(3):297, June 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900661.

Anonymous:1984:BRBa

[553] Anonymous. Book review: Basic computational techniques for engineers: R. A. Adey and C. A. Brebbia: Pentech Press, Plymouth, 1983, 208 pages, ISBN 0-7273-0203-5. Mathematics and Computers in Simulation, 26(3):297, June 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900673.

Anonymous:1984:BRSc

[554] Anonymous. Book review: Stochastic systems: G. Adomian: Academic Press, London, 1983, 352 pages, ISBN 0-12-044370-8. Mathematics and Computers in Simulation, 26 (3):297–298, June 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900685.

Anonymous:1984:BRIa

[555] Anonymous. Book review: Introduction to computational complexity: A. V. Aho, J. E. Hopcroft and J. D. Ullman: Wesley, Reading, MA, 1983. ISBN 0-201-00018-0. Mathematics and Computers in Simulation, 26(3):298, June 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900697.

Anonymous:1984:BRCa

[556] Anonymous. Book review: Computer modelling of electrical power systems: J. Arrillaga, C. D. Arnold and A. J. Harker: Wiley, Chichester, England, 1983, 423 pages, £25.00. Mathematics and Computers in Simulation, 26(3):298, June 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900703.

Anonymous:1984:BREb

[557] Anonymous. Book review: Error control codes: R. E. Blahut: Addison-Wesley, Reading, MA, 1983, ISBN 0-201-10102-5. Mathematics and Computers in Simulation, 26 (3):298, June 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900715.

Anonymous:1984:BRCb

[558] Anonymous. Book review: Computational methods in subsurface flow: P. S. Huyalorn and G. F. Pinder:

Academic Press, 1983, 494 pages, ISBN 0-12-363480-6. Mathematics and Computers in Simulation, 26 (3):299, June 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900727.

Anonymous:1984:BRAb

[559] Anonymous. Book review: Applied abstract algebra: K. H. Kim and F. W. Roush: Wiley, Chichester, England, 1983, 265 pages, £25.00, ISBN 0-85312-563-5. Mathematics and Computers in Simulation, 26(3):299, June 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900739.

Anonymous:1984:BRAc

[560] Anonymous. Book review: Algorithms: R. Sedgewick: Addison-Wesley, Reading, MA, 1983, ISBN 0-201-06672-6. Mathematics and Computers in Simulation, 26(3):299– 300, June 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900740.

Anonymous:1984:BRBb

[561] Anonymous. Book review: Basic 2: W. K. Weber und C. W. Turschmann: (2. Auflage). Paul Haupt, Bern, Switzerland, 1982, 185 Seiten, DM 19.80 ISBN 3-258-03175-4. Mathematics and Computers in Simulation, 26(3):300, June 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900752.

Anonymous:1984:BRMa

[562] Anonymous. Book review: Microprocessors in signal processing, measurement and control: S. G. Tzafestas (Editor): Reidel, Dordrecht, Holland, 1983, 442 pages, Dfl, 135.00, US$59.00, ISBN 90-277-1497-5. Mathematics and Computers in Simulation, 26(3):300–301, June 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900764.

Anonymous:1984:Bd

[563] Anonymous. Bibliography. Mathematics and Computers in Simulation, 26(3):302–315, June 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900776.

Anonymous:1984:IIS

[564] Anonymous. IFIP 7th international symposium on computer hardware description languages and their applications August 29–31, 1985, Keidanren Kaikan, Tokyo, Japan. Mathematics and Computers in Simulation, 26(3):316–317, June 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900788.

Anonymous:1984:PAb

[565] Anonymous. Pages 319–400 (August 1984). Mathematics and Computers in Simulation, 26(4):??, August 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Bharucha-Reid:1984:I

[566] A. T. Bharucha-Reid and C. P. Tsokos. Introduction. Mathematics and Computers in Simulation, 26 (4):319–320, August 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475484900016.

Bharucha-Reid:1984:ASR

[567] A. T. Bharucha-Reid and M. J. Christensen. Approximate solution of random integral equations: general methods. Mathematics and Computers in Simulation, 26(4):321– 328, August 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475484900028.

Rao:1984:SCS

[568] A. N. V. Rao and W. J. Padgett. Stability of a class of stochastic differential systems. Mathematics and Computers in Simulation, 26(4):329– 333, August 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847548490003X.

Lakshmikantham:1984:SBV

[569] V. Lakshmikantham. Stochastic boundary value problems. Mathematics and Computers in Simulation, 26 (4):334–339, August 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900041.

Alves:1984:ASE

[570] João Bosco da Mota Alves. Alternative state estimation from nongaussian observation. Mathematics and Computers in Simulation, 26 (4):340–343, August 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900053.

Maksimey:1984:TSE

[571] I. V. Maksimey. On technology of simulation experiment with dynamic stochastic queuing networks models. Mathematics and Computers in Simulation, 26(4):344– 348, August 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900065.

Palamara:1984:MSM

[572] A. Orsi Palamara and G. Rizzi. A mathematical stochastic model for the particles diffusion in atmosphere. Mathematics and Computers in Simulation, 26(4):349– 354, August 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900077.

Tien:1984:SMP

[573] Chi Tien. Stochastic modelling of particle deposition on collectors. Mathematics and Computers in Simulation, 26(4):355–356, August 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475484900089.

Pinsky:1984:BME

[574] Mark A. Pinsky. Brownian motion, exit times and stochastic Riemannian geometry. Mathematics and Computers in Simulation, 26(4):357– 360, August 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475484900090.

Bellomo:1984:MMP

[575] N. Bellomo, E. Cafaro, and G. Rizzi. On the mathematical modelling of physical systems by ordinary differential stochastic equations. Mathematics and Computers in Simulation, 26 (4):361–367, August 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475484900107.

dAlessandro:1984:RBM

[576] P. d’Alessandro, A. Germani, and M. Piccioni. Relationships between measures induced by Itô and white noise linear equations. Mathematics and Computers in Simulation, 26 (4):368–372, August 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475484900119.

delGrosso:1984:PME

[577] Gabriella del Grosso, Anna Gerardi, and Federico Marchetti. Probabilistic methods for eigenvalue estimates. Mathematics and Computers in Simulation, 26(4):373– 376, August 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900120.

Snider:1984:LRD

[578] Arthur David Snider. Linear regression with delta-modulated integrated signals. Mathematics and Computers in Simulation, 26(4):377– 382, August 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900132.

Anonymous:1984:BR

[579] Anonymous. Book Review. Mathematics and Computers in Simulation, 26 (4):383–387, August 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900144.

Anonymous:1984:Be

[580] Anonymous. Bibliography. Mathematics and Computers in Simulation, 26 (4):388–400, August 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900156.

Anonymous:1984:PO

[581] Anonymous. Pages 401–473 (October 1984). Mathematics and Computers in Simulation, 26(5):??, October 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Faro:1984:CNA

[582] Alberto Faro, Orazio Mirabella, and Corrado Nigro. Computer network analysis by using a generalized Petri network simulator. Mathematics and Computers in Simulation, 26(5):401– 411, October 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484901046.

Pronzato:1984:GPG

[583] Luc Pronzato, Eric Walter, Alain Venot, and Jean-Francois Lebruchec. A general-purpose global optimizer: Implementation and applications. Mathematics and Computers in Simulation, 26(5):412–422, October 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484901058.

Kleptsyn:1984:DLS

[584] A. F. Kleptsyn, M. A. Krasnoselśkii, N. A. Kuznetsov, and V. S. Kozjakin. Desynchronization of linear systems. Mathematics and Computers in Simulation, 26(5):423– 431, October 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548490106X.

Tzafestas:1984:MSS

[585] S. G. Tzafestas and N. J. Theodorou. Multidimensional state-space models: a comparative overview. Mathematics and Computers in Simulation, 26 (5):432–442, October 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484901071.

Michell:1984:MCE

[586] J. A. Michell, J. M. Drake, and S. Bracho. A method for the calculation of eigenfunction expansions. Mathematics and Computers in Simulation, 26 (5):443–447, October 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484901083.

Surma:1984:CSH

[587] T. Surma and A. Maslowski. Computer solution of hyperbolic equation in head-on collision problem. Mathematics and Computers in Simulation, 26(5):448–453, October 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484901095.

Anonymous:1984:BRSd

[588] Anonymous. Book review: Simulation in engineering sciences: J. Burger and Y. Jarny (Editors): North-Holland, Amsterdam, 1983, 438 pages, Dfl. 135.00, ISBN 0-444-86795-3. Mathematics and Computers in Simulation, 26(5):454, October 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484901101.

Anonymous:1984:BRSe

[589] Anonymous. Book review: Systems analysis and design: a structured approach: W. S. Davis: Addison-Wesley, Reading, MA, U.S.A., 1983, ISBN 0-201-10271-4. Mathematics and Computers in Simulation, 26(5): 454–455, October 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484901113.

Anonymous:1984:BRIb

[590] Anonymous. Book review: An introduction to operating systems: H. M. Deitel: Addison-Wesley, U.S.A., 1983, 700 pages, ISBN 0-201-14473-5. Mathematics and Computers in Simulation, 26(5):455, October 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484901125.

Anonymous:1984:BRSf

[591] Anonymous. Book review: Security, IFIP / SEC’83: V. Fak (Editor): North-Holland, Amsterdam, 1983, 300 pages, Dfl. 95.00, ISBN 0-444-86669-8. Mathematics and Computers in Simulation, 26(5):455– 456, October 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484901137.

Anonymous:1984:BRMb

[592] Anonymous. Book review: Matrix computations: G. H. Golub and C. F. van Loan: Oxford Academic Press, England, 1984, Hard ISBN 0-946536-00-7, £35.00; Soft ISBN 0-946536-05-8, £19.50. Mathematics and Computers in Simulation, 26(5): 456, October 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484901149.

Anonymous:1984:BRAd

[593] Anonymous. Book review: Applied abstract algebra: K. H. Kim and F. W. Roush: Wiley, England, 265 pages, £25.00, ISBN 0-85312-563-5. Mathematics and Computers in Simulation, 26(5):456, October 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484901150.

Anonymous:1984:BRLa

[594] Anonymous. Book review: Logic design and computer organization: M. H. Lewin: Addison-Wesley, Reading, MA, U.S.A., 1983, ISBN 0-201-04144-8. Mathematics and Computers in Simulation, 26(5):456– 457, October 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484901162.

Anonymous:1984:BRLb

[595] Anonymous. Book review: Linear programming: K. G. Murty: Wiley, England, 482 pages, 1983, £31.00, ISBN 0-471-09725-X. Mathematics and Computers in Simulation, 26(5): 457–458, October 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484901174.

Anonymous:1984:BRIc

[596] Anonymous. Book review: Information, uncertainty, complexity: J. F. Traub, G. Wasilowski and H. Wozniakowski: Addison-Wesley, Reading, MA, U.S.A., 1983 199 pages, ISBN 0-201-07890-2. Mathematics and Computers in Simulation, 26(5): 458, October 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484901186.

Anonymous:1984:BRBc

[597] Anonymous. Book review: BASIC-Software: K. Weber: HAUPT Akademische Buchhandlung, Bern, Switzerland. Mathematics and Computers in Simulation, 26(5): 458, October 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484901198.

Anonymous:1984:Bf

[598] Anonymous. Bibliography. Mathematics and Computers in Simulation, 26 (5):459–473, October 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484901204.

Anonymous:1984:PD

[599] Anonymous. Pages 475–571 (December 1984). Mathematics and Computers in Simulation, 26(6):??, December 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Carver:1984:TDM

[600] M. B. Carver and M. A. Salcudean. Two dimensional modelling of the homogenizing effects of flow obstructions in two fluid flow. Mathematics and Computers in Simulation, 26 (6):475–484, December 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900272.

Cooke:1984:SCF

[601] Charlie H. Cooke. On shock capturing by first and second order accurate Godunov methods. Mathematics and Computers in Simulation, 26 (6):485–496, December 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900284.

Oguztoreli:1984:ORS

[602] M. N. Oğuztöreli, W. J. Gibb, and B. Özüm. Optimal reactor size distribution with fixed terminal constraints. Mathematics and Computers in Simulation, 26(6):497–501, December 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900296.

Gal:1984:OEC

[603] S. Gal, R. Y. Rubinstein, and A. Ziv. On the optimality and efficiency of Common Random Numbers. Mathematics and Computers in Simulation, 26(6):502–512, December 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900302.

Dolique:1984:PMC

[604] J.-M. Dolique, J.-R. Roche, and F. Chatelin. Pegase: a multifluid code for computational study of high power relativistic electron beam propagation in atmosphere. Mathematics and Computers in Simulation, 26(6): 513–522, December 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900314.

Jutila:1984:PPT

[605] P. Jutila and A. Visala. Pilot plant testing of an adaptive pHcontrol algorithm based on physicochemical modelling. Mathematics and Computers in Simulation, 26(6): 523–533, December 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900326.

Stavroulakis:1984:RAS

[606] Peter Stavroulakis. Recent advances in the study of network and system sensitivity. Mathematics and Computers in Simulation, 26(6):534– 545, December 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475484900338.

VanRemoortere:1984:BRC

[607] P. Van Remoortere. Book review: Computational Methods for transient analysis. Volume 1 in: Computational methods in mechanics: T. Belytschko and T. J. R. Hughes (Editors): North-Holland, Amsterdam, 1983, 523 pages, US$88.00, Dfl.225.00. ISBN 0-444-86479-2. Mathematics and Computers in Simulation, 26(6): 546–548, December 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548490034X.

VanRemoortere:1984:BRP

[608] P. Van Remoortere. Book review: Physical systems theory in terms of bond graphs: P. C. Breedveld: Vakgroep Besturingsystemen en Computertechniek, THT, Afdeling Electrotechniek, Postbus 217, 7500 AE Enschede, The Netherlands. 1984, 200 pages, ISBN 90-90-00599-4. Mathematics and Computers in Simulation, 26(6): 548–551, December 1984. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475484900351.

VanRemoortere:1984:Bb

[609] P. Van Remoortere. Bibliography. Mathematics and Computers in Simulation, 26(6):552–556, December 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475484900363.

Anonymous:1984:AIV

[610] Anonymous. Author index to volume XXVI. Mathematics and Computers in Simulation, 26(6):567–568, December 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475484900375.

Anonymous:1984:CVX

[611] Anonymous. Contents volume XXVI. Mathematics and Computers in Simulation, 26(6):569–571, December 1984. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475484900387.

Anonymous:1985:PF

[612] Anonymous. Pages 1–91 (February 1985). Mathematics and Computers in Simulation, 27(1):??, February 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1985:EB

[613] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 27(1):ii, February 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475485900096.

Humeau:1985:CCD

[614] J. P. Humeau and M. Grolleau. Coefficient control for a distributed parabolic system. Mathematics and Computers in Simulation, 27(1):1– 17, February 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900102.

Vancraeynest:1985:ATD

[615] Jan Prosper Vancraeynest. An application of the three-dimensional Hadamard transform in particle accelerator simulations. Mathematics and Computers in Simulation, 27 (1):19–25, February 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900114.

Mathis:1985:DAM

[616] Frank H. Mathis. A direct approximation method for parameter estimation in partial differential equations. Mathematics and Computers in Simulation, 27(1):27–35, February 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900126.

Grotjahn:1985:GVE

[617] Richard Grotjahn. Group velocity errors caused by finite differencing: a meteorological perspective. Mathematics and Computers in Simulation, 27 (1):37–46, February 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475485900138.

Charette:1985:ESB

[618] André Charette, Rung T. Bui, and Guy Simard. The effect of stringent boundary conditions on the solution of a transient heat conduction problem. Mathematics and Computers in Simulation, 27(1):47–60, February 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548590014X.

Zamani:1985:SRU

[619] Nader Zamani. Some remarks on the use of the boundary element method in transient heat conduction problems. Mathematics and Computers in Simulation, 27(1):61– 64, February 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475485900151.

Anonymous:1985:BRCa

[620] Anonymous. Book review: The computer package STATCAT. Source programs and user manual: H. David: North-Holland, Amsterdam, 1982, 780 pages, US$102.00, Dfl. 240.00, ISBN 0-444-86453-9. Mathematics and Computers in Simulation, 27(1): 65, February 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475485900163.

Anonymous:1985:BRIa

[621] Anonymous. Book review: Information processing ’83: North-Holland, Amsterdam, 1983, 976 pages, US$98.00 (USA & CND), Dfl. 260.00 (Rest of world), ISBN 0-444-86729-5. Mathematics and Computers in Simulation, 27(1):65–66, February 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900175.

Anonymous:1985:BRO

[622] Anonymous. Book review: Optical communications: North-Holland, Amsterdam, 1983, 572 pages, US$85.00 (USA & CND), Dfl. 200.00 (Rest of world), ISBN 0-444-86797-X. Mathematics and Computers in Simulation, 27(1):66, February 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900187.

Anonymous:1985:BRIb

[623] Anonymous. Book review: Information technology and the computer network: Springer, Berlin, 1984, 271 pages, DM 92.00, US$34.40, ISBN 3-540-12883-2. Mathematics and Computers in Simulation, 27(1):66– 67, February 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900199.

Anonymous:1985:BRP

[624] Anonymous. Book review: Prüfungsauggaben informatik: Springer, Berlin, 1984, 460 pages, DM 39.80, ISBN 3-540-12772-0. Mathematics and Computers in Simulation, 27(1):67–68, February 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900205.

Anonymous:1985:BRCb

[625] Anonymous. Book review: On conceptual modelling: Springer, Berlin, 1984, 510 pages, DM 52.00, US$20.00, ISBN 3-540-90482-0 [bad ISBN]. Mathematics and Computers in Simulation, 27 (1):68–69, February 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900217.

Anonymous:1985:BRCc

[626] Anonymous. Book review: Computation mathematics: an introduction to numerical approximation: Wiley, New York, 1984, 431 pages, £25.00, ISBN 0-85312-565-1. Mathematics and Computers in Simulation, 27(1):69, February 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900229.

Anonymous:1985:BRCd

[627] Anonymous. Book review: Computer methods for the range of functions: Wiley, New York, 1984, 168 pages, £16.95, ISBN 0-85312-703-4. Mathematics and Computers in Simulation, 27(1):69–70, February 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900230.

VanRemoortere:1985:B

[628] P. Van Remoortere. Bibliography. Mathematics and Computers in Simulation, 27(1):71–88, February 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475485900242.

Anonymous:1985:SCC

[629] Anonymous. Simulation computer code abstract. Mathematics and Computers in Simulation, 27(1):89– 91, February 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900254.

Anonymous:1985:PAa

[630] Anonymous. Pages 93–294 (April 1985). Mathematics and Computers in Simulation, 27(2–3):??, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Phillips:1985:I

[631] F. G. Phillips and J. R. Coleby. Introduction. Mathematics and Computers in Simulation, 27(2–3):93– 94, April 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900266.

Gipps:1985:MSM

[632] P. G. Gipps and B. Marksjö. A micro-simulation model for pedestrian flows. Mathematics and Computers in Simulation, 27(2–3):95– 105, April 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900278.

Warren:1985:AAS

[633] H. J. Warren, R. A. Low, P. D. Hughes, and G. J. Gray. AU-TOSIM: an automatic simulation program generator. Mathematics and Computers in Simulation, 27(2–3): 107–114, April 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548590028X.

Baunach:1985:SCT

[634] G. R. Baunach, E. S. Wibberley, and B. R. Wood. Simulation of a coal transhipment terminal: Batam Island, Indonesia. Mathematics and Computers in Simulation, 27(2–3): 115–120, April 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900291.

Benyon:1985:EVC

[635] Peter R. Benyon. Exploiting vector computers for simulation. Mathematics and Computers in Simulation, 27 (2–3):121–127, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900308.

Hulskamp:1985:CSS

[636] John P. Hulskamp. Continuous system simulation using digital signal processing elements. Mathematics and Computers in Simulation, 27(2– 3):129–136, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548590031X.

Freeman:1985:SBM

[637] T. Graham Freeman. Selecting the best model to fit data. Mathematics and Computers in Simulation, 27 (2–3):137–140, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900321.

Cumming:1985:STD

[638] Malcolm J. Cumming. Simplification of a two-dimensional heat and mass transfer model by a transformation of coordinates. Mathematics and Computers in Simulation, 27 (2–3):141–145, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900333.

Rudenno:1985:SPE

[639] V. Rudenno and M. Richardson. The simulation of the proportional effect for ore reserve analysis. Mathematics and Computers in Simulation, 27 (2–3):147–153, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900345.

Brooker:1985:SSC

[640] P. I. Brooker. Simulation of spatially correlated data in two dimensions. Mathematics and Computers in Simulation, 27(2–3):155– 157, April 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475485900357.

Williams:1985:RRM

[641] B. J. Williams and W. G. Field. Rainfall runoff models in flood forecasting applications. Mathematics and Computers in Simulation, 27(2– 3):159–165, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475485900369.

Jakeman:1985:CDS

[642] A. J. Jakeman, R. W. Simpson, and J. A. Taylor. Combining deterministic and statistical models for ill-defined systems: Advantages for air quality assessment. Mathematics and Computers in Simulation, 27 (2–3):167–178, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475485900370.

Jakeman:1985:MUU

[643] A. J. Jakeman and G. A. Thomas. Modelling under uncertainty: the scientific method revisited. Mathematics and Computers in Simulation, 27 (2–3):179–189, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900382.

Green:1985:SSP

[644] D. G. Green, A. P. N. House, and S. M. House. Simulating spatial patterns in forest ecosystems. Mathematics and Computers in Simulation, 27 (2–3):191–198, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900394.

Leech:1985:MPM

[645] J. W. Leech. A management planning model of a large plantation forest. Mathematics and Computers in Simulation, 27(2–3):199– 206, April 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900400.

Mayer:1985:SSS

[646] David G. Mayer. Simulation of small systems using microcomputers. Mathematics and Computers in Simulation, 27(2–3):207–210, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900412.

Hansen:1985:RFB

[647] B. R. Hansen and G. T. Oborne. Redevelopment of fruit blocks in the riverland region: an intertemporal programming approach. Mathematics and Computers in Simulation, 27 (2–3):211–222, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475485900424.

Pearson:1985:EMS

[648] K. R. Pearson and Russell J. Rimmer. An efficient method for the solution of large computable general equilibrium models. Mathematics and Computers in Simulation, 27 (2–3):223–229, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475485900436.

Haritos:1985:MRT

[649] N. Haritos. Modelling the response of tension-leg platforms to the effects of wind using simulated traces. Mathematics and Computers in Simulation, 27(2–3):231–240, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475485900448.

Anderson:1985:SMJ

[650] S. J. Anderson. Simulation and modelling for the Jindalee over-the-horizon radar. Mathematics and Computers in Simulation, 27(2–3): 241–248, April 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847548590045X.

Henderson:1985:MWS

[651] R. G. Henderson, R. N. Johnson, and J. A. Wills. Modelling of weapon system effects. Mathematics and Computers in Simulation, 27 (2–3):249–257, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900461.

Tyler:1985:CSP

[652] D. W. Tyler and A. K. Quibell. Computer simulation in process control. Mathematics and Computers in Simulation, 27(2–3):259– 266, April 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900473.

Anonymous:1985:BRE

[653] Anonymous. Book review: EURO-GRAPHICS 83: P. W. J. ten Hagen (Editor): North-Holland, Amsterdam, 1983, 396 pages, US$46.25, Dfl. 120.00, ISBN 0-444-86770-8. Mathematics and Computers in Simulation, 27(2–3):267, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900485.

Anonymous:1985:BRF

[654] Anonymous. Book review: Finite elements — computational aspects — volume — III: C. F. Carey and J. T. Oden: Prentice-Hall, 1984, 350 pages, ISBN 0-13-317107-8. Mathematics and Computers in Simulation, 27 (2–3):267–268, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900497.

Anonymous:1985:BRR

[655] Anonymous. Book review: Ring technology local area networks: I. N. Dallas and E. B. Spratt (Editors): North-Holland, Amsterdam, 1984, 170 pages, US$36.50, Dfl. 85.00, ISBN 0-444-86852-6. Mathematics and Computers in Simulation, 27(2– 3):268, April 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900503.

Anonymous:1985:BRN

[656] Anonymous. Book review: Numerical solution of singular integral equations: A. Gerasoulis and R. Vichnevetsky (Editors): 1984, 141 pages, To be ordered at: IMACS Symposium, Rutgers University, Dept. of Computer Sciences, New Brunswick, NJ 08903, USA. Mathematics and Computers in Simulation, 27(2–3): 268, April 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900515.

Anonymous:1985:BRM

[657] Anonymous. Book review: Mathematical computer performance and reliability: G. Iazelooa, P. J. Courtois and A. Hordijk: North-Holland, Amsterdam, 1984, 430 pages, US$50.00, Dfl. 130.00, ISBN 0-444-86892-5. Mathematics and Computers in Simulation, 27(2–3):268–269, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900527.

Anonymous:1985:BRIc

[658] Anonymous. Book review: Introduction to numerical methods for parallel computers: U. Schendel: Wiley, England, 1984, 151 pages, £15.00, ISBN 0-85312-597-X. Mathematics and Computers in Simulation, 27(2– 3):269, April 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900539.

Anonymous:1985:BRAa

[659] Anonymous. Book review: Analysis and simulation of semiconductor devices: S. Selberherr: Springer, Wien, New York, 1984, 294 pages, DM 138.00, OS 966.00, ISBN 3-211-81800-6. Mathematics and Computers in Simulation, 27(2–3):269–270, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900540.

Anonymous:1985:BRAb

[660] Anonymous. Book review: Advances in computer methods for partial differential equations V: R. Vichnevetsky and R. S. Stepleman (Editors): 1984, 558 pages, To be ordered at: IMACS Symposium, Rutgers University, Dept. of Computer Science, New Brunswick, NJ 08903, USA. Mathematics and Computers in Simulation, 27(2–3):270, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475485900552.

Anonymous:1985:B

[661] Anonymous. Bibliography. Mathematics and Computers in Simulation, 27 (2–3):271–290, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475485900564.

Anonymous:1985:NIa

[662] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 27(2–3):291–293, April 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475485900576.

Anonymous:1985:PAb

[663] Anonymous. Pages 295–386 (August 1985). Mathematics and Computers in Simulation, 27(4):??, August 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Cendes:1985:CEB

[664] Z. J. Cendes and D. N. Shenton. Complementary error bounds for foolproof finite element mesh generation. Mathematics and Computers in Simulation, 27(4):295–305, August 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475485900011.

Kleinstreuer:1985:ASN

[665] Clement Kleinstreuer and Ali Eghlima. Analysis and simulation of new approximation equations for boundary-layer flow on curved surfaces. Mathematics and Computers in Simulation, 27 (4):307–325, August 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900023.

Vichnevetsky:1985:GVG

[666] Robert Vichnevetsky. Group velocity in general waveforms. Mathematics and Computers in Simulation, 27 (4):327–331, August 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900035.

Ahmad:1985:SFA

[667] A. I. Sheikh Ahmad. Simulation of a feedback and adaptive systems. Mathematics and Computers in Simulation, 27(4):333–338, August 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900047.

Haryono:1985:AMR

[668] Haryono and B. D. Sivazlian. Analysis of the machine repair problem: a diffusion process approach. Mathematics and Computers in Simulation, 27 (4):339–364, August 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900059.

Demetriou:1985:PVM

[669] I. C. Demetriou and D. J. Evans. The preconditioned variational methods for solving large linear systems. Mathematics and Computers in Simulation, 27(4):365–372, August 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900060.

Missirlis:1985:CRC

[670] N. M. Missirlis and D. J. Evans. A comparison of rates of convergence for the Modified Alternating Direction Preconditioning (MADP) method. Mathematics and Computers in Simulation, 27 (4):373–382, August 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900072.

Anonymous:1985:NIb

[671] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 27(4):383–386, August 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900084.

Anonymous:1985:PO

[672] Anonymous. Pages 387–647 (October 1985). Mathematics and Computers in Simulation, 27(5–6):??, October 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Tzafestas:1985:GE

[673] Spyros G. Tzafestas. Guest editorial. Mathematics and Computers in Simulation, 27(5–6):387, October 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900588.

Rao:1985:DPC

[674] Ganti Prasada Rao and Spyros G. Tzafestas. A decade of piecewise constant orthogonal functions in systems and control. Mathematics and Computers in Simulation, 27(5–6): 389–407, October 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548590059X.

Lai:1985:SPO

[675] David C. Lai. Signal processing with orthonormalized exponentials. Mathematics and Computers in Simulation, 27(5–6):409– 420, October 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900606.

Rao:1985:DT

[676] K. R. Rao. Discrete transforms. Mathematics and Computers in Simulation, 27(5–6):421–430, October 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475485900618.

Moraga:1985:SAC

[677] Claudio Moraga. On some applications of the Chrestenson functions in logic design and data processing. Mathematics and Computers in Simulation, 27 (5–6):431–439, October 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548590062X.

Corinthios:1985:DDA

[678] Michael J. Corinthios and Youssef Geadah. 1-D and 2-D algorithms for orthogonal transformations. Mathematics and Computers in Simulation, 27 (5–6):441–452, October 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900631.

Paraskevopoulos:1985:SAS

[679] P. N. Paraskevopoulos. System analysis and synthesis via orthogonal polynomial series and Fourier series. Mathematics and Computers in Simulation, 27(5–6):453– 469, October 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900643.

Chen:1985:ACS

[680] C. F. Chen and W. K. Leung. Algorithms for converting Sequency-, Dyadic-, and Hadamard-ordered Walsh functions. Mathematics and Computers in Simulation, 27(5–6):471– 478, October 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900655.

Kawaji:1985:MRW

[681] S. Kawaji and T. Shiotsuki. Model reduction by Walsh function techniques. Mathematics and Computers in Simulation, 27(5–6):479– 484, October 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900667.

Bohn:1985:OWF

[682] E. V. Bohn. Optimal Walsh function input signals for parameter identification in identification systems. Mathematics and Computers in Simulation, 27(5–6):485– 490, October 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900679.

Frangakis:1985:FWT

[683] G. P. Frangakis, G. Papakonstantinou, and S. G. Tzafestas. A fast Walsh transform-based data compression multi-microprocessor system: application to ECG signals. Mathematics and Computers in Simulation, 27 (5–6):491–502, October 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900680.

Chen:1985:AMS

[684] W. L. Chen and S. G. Wu. Analysis of multirate sampled-data systems by block-pulse functions. Mathematics and Computers in Simulation, 27 (5–6):503–510, October 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900692.

Ohkita:1985:EAF

[685] Masaaki Ohkita. Evaluation of analytic functions by generalized digital integration. Mathematics and Computers in Simulation, 27(5–6): 511–517, October 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475485900709.

Harmuth:1985:OEC

[686] Henning F. Harmuth. Orthogonal expansions and the causality law. Mathematics and Computers in Simulation, 27(5–6):519– 529, October 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475485900710.

Fikioris:1985:COE

[687] John G. Fikioris. On the convergence of orthogonal eigenfunction series. Mathematics and Computers in Simulation, 27(5–6):531– 539, October 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475485900722.

Clement:1985:AGL

[688] P. R. Clement. Applications of generalized Laguerre functions. Mathematics and Computers in Simulation, 27 (5–6):541–550, October 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475485900734.

Campanella:1985:SPI

[689] M. Campanella, U. Lo Faso, and G. Mamola. On the solution of a particular integral equation through Gram–Schmidt orthogonalization of sinc functions. Mathematics and Computers in Simulation, 27(5–6): 551–557, October 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475485900746.

Pitas:1985:TDT

[690] I. Pitas and A. N. Venetsanopoulos. Transform decomposition in two-dimensional digital filter realization. Mathematics and Computers in Simulation, 27(5–6):559– 583, October 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900758.

Rajan:1985:SDT

[691] P. Karivaratha Rajan and Harnatha C. Reddy. Symmetrical decomposition and transformation. Mathematics and Computers in Simulation, 27(5– 6):585–598, October 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548590076X.

Cappellini:1985:IDC

[692] Vito Cappellini and Enrico Del Re. Image data compression by the discrete cosine transform. Mathematics and Computers in Simulation, 27(5– 6):599–608, October 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475485900771.

Alexandridis:1985:DFM

[693] N. A. Alexandridis, P. D. Tsanakas, and S. V. Konstantinidou. A dataflow multiprocessor architecture for the transform coding of regularly decomposed images. Mathematics and Computers in Simulation, 27(5–6): 609–622, October 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475485900783.

Goutis:1985:UDO

[694] C. E. Goutis. Unified dual optimisation and algorithms for image reconstruction and spectral estimation. Mathematics and Computers in Simulation, 27(5–6):623– 634, October 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475485900795.

Anonymous:1985:NIc

[695] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 27(5–6):635–640, October 1985. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/0378475485900801.

Anonymous:1985:AIV

[696] Anonymous. Author index to volume 27. Mathematics and Computers in Simulation, 27(5–6):641– 644, October 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900813.

Anonymous:1985:CV

[697] Anonymous. Contents volume 27. Mathematics and Computers in Simulation, 27(5–6):645–647, October 1985. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475485900825.

Anonymous:1986:PF

[698] Anonymous. Pages 1–90 (February 1986). Mathematics and Computers in Simulation, 28(1):??, February 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1986:EB

[699] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 28(1):ii, February 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900820.

Miller:1986:IRT

[700] John J. H. Miller. On the inclusion of the recombination term in discretizations of the semiconductor device equations. Mathematics and Computers in Simulation, 28(1):1– 11, February 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475486900832.

Greenspan:1986:PSC

[701] Donald Greenspan. Particle simulation of compression waves. Mathematics and Computers in Simulation, 28 (1):13–24, February 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486900844.

Birta:1986:SRA

[702] Louis G. Birta and Usha Deo. A sequential refinement approach for parameter optimization in continuous dynamic models. Mathematics and Computers in Simulation, 28 (1):25–39, February 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486900856.

Dube:1986:ECD

[703] S. K. Dube, P. C. Sinha, and G. D. Roy. The effect of a continuously deforming coastline on the numerical simulation of storm surges in Bangladesh. Mathematics and Computers in Simulation, 28(1):41–56, February 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486900868.

Amyot:1986:SCH

[704] J. R. Amyot and G. A. Camiré. Stability of a class of hybrid computer models of dynamical systems. Mathematics and Computers in Simulation, 28(1):57–64, February 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548690087X.

Diou:1986:RDS

[705] Alain Diou. Resolution directe des systemes algebriques par approximations. Mathematics and Computers in Simulation, 28(1):65– 75, February 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475486900881.

Anonymous:1986:BRDa

[706] Anonymous. Book review: Decision making, models and algorithms: Wiley, Chichester, United Kingdom, 1985. 412 pages, £38.00, ISBN 0-471-80963-2. Mathematics and Computers in Simulation, 28(1): 77, February 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475486900893.

Anonymous:1986:BRAa

[707] Anonymous. Book review: Applied statistics algorithms: Royal Statistical Society/Ellis Horwood, Chichester, United Kingdom, 1985, 307 pages, £30.00, ISBN 0-85312-772-7. Mathematics and Computers in Simulation, 28(1):77, February 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548690090X.

Anonymous:1986:BRFa

[708] Anonymous. Book review: Fundamentals of queueing: 2nd Ed.. Wiley, New York, 1985, 587 pages, £43.95, ISBN 0-471-89067-7. Mathematics and Computers in Simulation, 28(1):77–78, February 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900911.

Anonymous:1986:BRL

[709] Anonymous. Book review: Linear and nonlinear programming: Wiley, Chichester, United Kingdom, 1985, 221 pages, £9.95, ISBN 0-85312-848-0. Mathematics and Computers in Simulation, 28(1): 78, February 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475486900923.

Anonymous:1986:BRDb

[710] Anonymous. Book review: Designing for reliability and safety control: Prentice-Hall, Englewood Cliffs, 1985, 527 pages, ISBN 0-13-201229-4. Mathematics and Computers in Simulation, 28(1):78, February 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900935.

Anonymous:1986:BRNa

[711] Anonymous. Book review: Numerical methods for scientific and engineering computation: Wiley Eastern, New Delhi, 1985, 406 pages, (paper) £4.95, ISBN 0-85226-434-8. Mathematics and Computers in Simulation, 28(1):78, February 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900947.

Anonymous:1986:BRT

[712] Anonymous. Book review: Theory and applications of linear differential and difference equations: Ellis Horwood, Chichester, United Kingdom, 1984, 183 pages, £18.50, ISBN 0-85312-670-4. Mathematics and Computers in Simulation, 28(1): 79, February 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475486900959.

Anonymous:1986:BRFb

[713] Anonymous. Book review: Fundamentals of the average case analysis of particular algorithms: Wiley, Chichester. United Kingdom, 233 pages, £18.50, ISBN 0-471-90322-1. Mathematics and Computers in Simulation, 28(1):79, February 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900960.

Anonymous:1986:BRFc

[714] Anonymous. Book review: The foundations of program verification: Wiley, Chichester, United Kingdom, 1985, 230 pages, £18.50, ISBN 0-519-02101-3 [bad ISBN]. Mathematics and Computers in Simulation, 28(1):79, February 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900972.

Anonymous:1986:BREa

[715] Anonymous. Book review: The electronic oracle: Wiley, New York, 1985, 445 pages, £24.95, ISBN 0-471-90558-5. Mathematics and Computers in Simulation, 28(1):79– 80, February 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475486900984.

Anonymous:1986:BRCa

[716] Anonymous. Book review: Computational functional analysis: Ellis Horwood, Chichester, United Kingdom, 1985. 156 pages, (cloth) £15.00, (paper) £7.95, ISBN 0-85312-807-3. Mathematics and Computers in Simulation, 28(1):80, February 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486900996.

Anonymous:1986:BRCb

[717] Anonymous. Book review: Combinatorial optimization: Annotated bibliographies.: Wiley, Chichester, United Kingdom, 1985, 204 pages, (paper) £14.95, ISBN 0-471-90490-2. Mathematics and Computers in Simulation, 28(1):80, February 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548690100X.

Anonymous:1986:BRCc

[718] Anonymous. Book review: Computation and automata: Cambridge University Press, Cambridge, 1985, 282 pages, £39.50. ISBN 0-521-30245-5. Mathematics and Computers in Simulation, 28(1):80–81, February 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486901011.

Anonymous:1986:BRCd

[719] Anonymous. Book review: Cluster dissection and analysis: Ellis Horwood, Chichester, United Kingdom, 1984, 183 pages £25.00. ISBN 0-85312-736-0. Mathematics and Computers in Simulation, 28(1): 81, February 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475486901023.

Anonymous:1986:NIa

[720] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 28(1):83–86, February 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486901035.

Anonymous:1986:ICEa

[721] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 28(1):86–90, February 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475486901047.

Anonymous:1986:PAa

[722] Anonymous. Pages 91–168 (April 1986). Mathematics and Computers in Simulation, 28(2):??, April 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Vergis:1986:CAC

[723] Anastasios Vergis, Kenneth Steiglitz, and Bradley Dickinson. The complexity of analog computation. Mathematics and Computers in Simulation, 28(2):91–113, April 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486901059.

Galanis:1986:DCM

[724] S. Galanis, A. Hadjidimos, and D. Noutsos. On different classes of monoparametric stationary iterative methods for the solution of linear systems. Mathematics and Computers in Simulation, 28(2):115– 128, April 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475486901060.

Korn:1986:BDP

[725] Granino A. Korn and Hamid Vakilzadian. Beyond desire: Pascal implementations of direct-executing simulation languages. Mathematics and Computers in Simulation, 28 (2):129–139, April 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486901072.

OGrady:1986:PPP

[726] E. Pearse O’Grady. Problem partitioning in parallel processor simulation systems. Mathematics and Computers in Simulation, 28(2):141– 150, April 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475486901084.

Adomian:1986:IM

[727] G. Adomian. Inversion of matrices. Mathematics and Computers in Simulation, 28(2):151– 153, April 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475486901096.

Adomian:1986:SAE

[728] G. Adomian. Solution of algebraic equations. Mathematics and Computers in Simulation, 28(2):155– 157, April 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475486901102.

Anonymous:1986:BREb

[729] Anonymous. Book review: Error-free polynomial matrix computations: E. V. Krishnamurthy: Springer, New York, 1985, 154 pages, ISBN 0-387-96146-1. Mathematics and Computers in Simulation, 28(2):159, April 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486901114.

Anonymous:1986:BRDc

[730] Anonymous. Book review: Delay equations, approximation and application: G. Meinardus and G. Nürnberger (Editors): ISNM 74, Birkhäuser, Basel, 1985, 351 pages, ISBN 3-7643-1733-7. Mathematics and Computers in Simulation, 28 (2):159–160, April 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486901126.

Anonymous:1986:BRNb

[731] Anonymous. Book review: Numerical analysis of parametrized nonlinear equations: W. C. Rheinboldt: Wiley, New York, 1986, 299 pages, £35.75, ISBN 0-471-88814-1. Mathematics and Computers in Simulation, 28(2):160, April 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486901138.

Anonymous:1986:NIb

[732] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 28(2):161–167, April 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548690114X.

Anonymous:1986:PJ

[733] Anonymous. Pages 169–270 (June 1986). Mathematics and Computers in Simulation, 28(3):??, June 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Rubinstein:1986:IMU

[734] Y. R. Rubinstein and J. Kreimer. Inventory models under uncertainty: an adaptive approach. Mathematics and Computers in Simulation, 28 (3):169–188, June 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548690025X.

Mohapatra:1986:EAE

[735] Pratap K. J. Mohapatra and Sushil K. Sharma. An eigenvalue analysis for the evaluation of policy parameters in a system dynamics model. Mathematics and Computers in Simulation, 28(3):189–208, June 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900261.

Mooney:1986:SRD

[736] J. R. Mooney. Simulation of a reaction– diffusion system on large dimpled surfaces using a vector computer. Mathematics and Computers in Simulation, 28(3):209–225, June 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900273.

Lang:1986:ARA

[737] B. Lang, J. C. Miellou, and P. Spiteri. Asynchronous relaxation algorithms for optimal control problems. Mathematics and Computers in Simulation, 28(3):227–242, June 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900285.

Sciubba:1986:TDS

[738] Enrico Sciubba, Steven C. Gonzales, and Richard L. Peskin. A two-dimensional, second-order model for turbulent flow in the upper atmosphere. Mathematics and Computers in Simulation, 28(3):243– 260, June 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475486900297.

Anonymous:1986:BRMa

[739] Anonymous. Book review: Mathematics in the social and life sciences: theories, models and methods: M. A. Ball, Ellis Horwood, Chichester, 1985, £12.50, ISBN 0-85312-887-1. Mathematics and Computers in Simulation, 28(3):261, June 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486900303.

Anonymous:1986:BRSa

[740] Anonymous. Book review: Spectral approximation of linear operators: F. Chatelin, Academic Press, New York, 1983, 458 pages, $69.50, ISBN 0-12-170620-6. Mathematics and Computers in Simulation, 28 (3):261, June 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475486900315.

Anonymous:1986:BRMb

[741] Anonymous. Book review: Multi-grid methods and applications: W. Hackbusch, Springer, Heidelberg, 1985, 377 pages, DM 178, ISBN 3-540-12761-5. Mathematics and Computers in Simulation, 28(3):262, June 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900327.

CB:1986:BRP

[742] CB. Book review: Primality and cryptography: E. Kranakis, Wiley-Teubner, 1986, 235 pages, £22.30, ISBN 0-471-90934-3. Mathematics and Computers in Simulation, 28 (3):262, June 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475486900339.

Anonymous:1986:BRIa

[743] Anonymous. Book review: Integral methods in science and engineering: F. R. Payne, C. C. Corduneanu, A. Haji-Sheikh and Tseng Huang Proceedings of a Conference held at the University of Texas at Arlington, March 18–21, 1985. Hemisphere/Springer, 1986, 653 pages, $95.50, ISBN 0-89116-461-8˘29f83-540-16308-5. Mathematics and Computers in Simulation, 28(3):262, June 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900340.

Anonymous:1986:BRIb

[744] Anonymous. Book review: An introduction to the numerical solution of differential equations: D. Quinney, Research Studies Press/Wiley, 1985, 283 pages, £15, ISBN 0-86380-033-5/0-471-90849-5. Mathematics and Computers in Simulation, 28(3):263, June 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486900352.

Anonymous:1986:BRIc

[745] Anonymous. Book review: Introduction to applied mathematics: G. Strang, Wellesley-Cambridge Press, 1986, 758 pages, $36, ISBN 0-9614088-0-4. Mathematics and Computers in Simulation, 28(3):263, June 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486900364.

Anonymous:1986:NIc

[746] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 28(3):265–270, June 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486900376.

Anonymous:1986:PAb

[747] Anonymous. Pages 271–350 (August 1986). Mathematics and Computers in Simulation, 28(4):??, August 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Wu:1986:SBH

[748] Qiming Wu and François E. Cellier. Simulation of bipolar highvoltage devices in the neighborhood of breakdown. Mathematics and Computers in Simulation, 28(4):271– 284, August 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475486900480.

Gipps:1986:RCG

[749] P. G. Gipps. The role of computer graphics in validating simulation models. Mathematics and Computers in Simulation, 28(4):285– 289, August 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475486900492.

Gipps:1986:MMS

[750] P. G. Gipps. Multsim: a model for simulating vehicular traffic on multi-lane arterial roads. Mathematics and Computers in Simulation, 28 (4):291–295, August 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900509.

Cremer:1986:FSM

[751] M. Cremer and J. Ludwig. A fast simulation model for traffic flow on the basis of boolean operations. Mathematics and Computers in Simulation, 28(4):297–303, August 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900510.

Young:1986:MVM

[752] William Young. A model of vehicles movements in parking facilities. Mathematics and Computers in Simulation, 28(4):305–309, August 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486900522.

Joshi:1986:SME

[753] Dhananjay S. Joshi and Lawrence J. Bain. Strategies for minimizing the effects of late truck arrivals on newspaper distribution systems. Mathematics and Computers in Simulation, 28 (4):311–315, August 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486900534.

Friedman:1986:PSD

[754] Moshe F. Friedman. On the properties of the sum of durations of perceived conflicts in air traffic control. Mathematics and Computers in Simulation, 28(4):317–324, August 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486900546.

Kiss:1986:CSH

[755] Lajos Kiss. Computer simulation of holography for seismic purposes. Mathematics and Computers in Simulation, 28(4):325–330, August 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486900558.

Dovi:1986:IRA

[756] V. G. Doví, O. Paladino, E. P. Arato, L. Maga, and A. Passerone. An improved regression analysis for automatic surface tension measurements. Mathematics and Computers in Simulation, 28(4):331– 335, August 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548690056X.

Anonymous:1986:BRSb

[757] Anonymous. Book review: Surfaces in CAGD ’84: R. E. Barnhill and W. Boehm (Ed.), North-Holland, Amsterdam, 1985, 236 pages, US $55.50/ Dfl.150.00, ISBN 0-444-87798-3. Mathematics and Computers in Simulation, 28(4):337, August 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900571.

Anonymous:1986:BRAb

[758] Anonymous. Book review: Artificial intelligence (methodology, systems, applications): W. Biebel and B. Petkoff, North-Holland, Amsterdam, 1985, 248 pages, US $44.50/ Dfl.120.00, ISBN 0-444-87743-6. Mathematics and Computers in Simulation, 28(4):337, August 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900583.

Anonymous:1986:BRAc

[759] Anonymous. Book review: Ablaufsteuerung von montagefertigungen: P. Biendl, Verlag Paul Haupt, Bern, 1984, SFr. 48.00/DM 58.00, ISBN 3-258-03417-6. Mathematics and Computers in Simulation, 28(4):

337, August 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475486900595.

Anonymous:1986:BRP

[760] Anonymous. Book review: Protocol specification, testing, and verification: M. Diaz (Ed.), V. North-Holland, Amsterdam, 1986, 544 pages, US $65/ Dfl.210.00, ISBN 0-444-87881-5. Mathematics and Computers in Simulation, 28(4):337–338, August 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486900601.

Anonymous:1986:BRCe

[761] Anonymous. Book review: Computer hardware description, languages and their applications: C. J. Koomen and T. Moto-Oka (Eds.), North-Holland, Amsterdam, 1985, 506 pages, US $64.75/Dfl.175.00, ISBN 0-444-87826-2. Mathematics and Computers in Simulation, 28(4): 338, August 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475486900613.

Anonymous:1986:BRSc

[762] Anonymous. Book review: The shallow water wave equations: formulation, analysis and application: I. Kinnmark, Lecture Notes in Engineering Vol. 15, Springer, Berlin, 1986, 187 pages, DM39.00 (soft cover), ISBN 3-540-16031-0. Mathematics and Computers in Simulation, 28 (4):338–339, August 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900625.

Anonymous:1986:BRZ

[763] Anonymous. Book review: Zwiechenauswertungen und vorzeitiger abbruch von therapiestudien: W. Kopcke, Springer, Berlin, 1984, 197 pages, DM39.50, ISBN 3-540-13373-9. Mathematics and Computers in Simulation, 28(4): 339, August 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475486900637.

Anonymous:1986:BRSd

[764] Anonymous. Book review: Systemanalyse biologischer prozesse: D. P. F. Moller (Ed.), Springer, Berlin, 1984, 226 pages, DM39.50, ISBN 3-540-13371-2. Mathematics and Computers in Simulation, 28(4): 339, August 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475486900649.

Anonymous:1986:BRMc

[765] Anonymous. Book review: Modelling techniques and tools for performance analysis: D. Potier (Ed.), North-Holland, Amsterdam, 1985, 640 pages, US $74.00/Dfl.200.00, ISBN 0-444-87696-0. Mathematics and Computers in Simulation, 28(4): 339, August 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475486900650.

Anonymous:1986:BRSe

[766] Anonymous. Book review: Solving elliptic problems using ELLPACK: J. R. Rice and R. F. Boisvert, Springer, New York, 1985, 497 pages, DM138.00, ISBN 3-540-90910-9. Mathematics and Computers in Simulation, 28 (4):339–340, August 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486900662.

Anonymous:1986:BRH

[767] Anonymous. Book review: Human– computer interaction — Interact 84: B. Shackel (Ed.), North-Holland, Amsterdam, 1985, US $79.50/Dfl.275.00, ISBN 0-444-87773-8. Mathematics and Computers in Simulation, 28 (4):340–341, August 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486900674.

Anonymous:1986:BRMd

[768] Anonymous. Book review: Metalevel inference: B. Silver, North-Holland, Amsterdam, 1986, 222 pages, US$35.00/Dfl.120.00, ISBN 0-444-87900-5. Mathematics and Computers in Simulation, 28(4): 341, August 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475486900686.

Anonymous:1986:BRSf

[769] Anonymous. Book review: Sortimentsplanung einer unternehmung: K. Spalti, Verlag Paul Haupt, Bern, Switzerland, 1985, 109 pages, SFr.28.00/DM34.00, ISBN 3-258-03423-0. Mathematics and Computers in Simulation, 28(4):341, August 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475486900698.

Anonymous:1986:BRDd

[770] Anonymous. Book review: Digital techniques in simulation, communication and control: S. G. Tzafestas (Ed.), North-Holland, Amsterdam, 1985, 556 pages, US$63.00/ Dfl.170.00, ISBN 0-444-87695-2. Mathematics and Computers in Simulation, 28(4):341–342, August 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900704.

Anonymous:1986:NId

[771] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 28(4):343–350, August 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900716.

Anonymous:1986:PO

[772] Anonymous. Pages 351–424 (October 1986). Mathematics and Computers in Simulation, 28(5):??, October 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Rubinstein:1986:SFA

[773] Reuven Y. Rubinstein. The score function approach for sensitivity analysis of computer simulation models. Mathematics and Computers in Simulation, 28(5):351–379, October 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486900728.

Kladas:1986:AMC

[774] A. Kladas and A. Razek. Analytic macroelements combined with finite elements: Application in modelling of electromagnetic systems. Mathematics and Computers in Simulation, 28 (5):381–390, October 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548690073X.

Jones:1986:AMC

[775] D. I. Jones and M. H. Lorenz. An application of a Markov chain noise model to wind generator simulation. Mathematics and Computers in Simulation, 28(5):391–402, October 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486900741.

Vajravelu:1986:NSE

[776] K. Vajravelu and L. Debnath. Numerical study of the effects of variable fluid property and internal heat generation on flat-plate thermometer. Mathematics and Computers in Simulation, 28 (5):403–406, October 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900753.

Anonymous:1986:BRMe

[777] Anonymous. Book review: Methods for Approximation: R. E. Bellman and R. S. Roth, Reidel, Dordrecht, 1986, xv + 224 pages, US $49.00/ Dfl.120.00, ISBN 90-277-2188-2. Mathematics and Computers in Simulation, 28(5):407, October 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900765.

Anonymous:1986:BRFd

[778] Anonymous. Book review: Foundations of Optimum Experimental Design: A. Pezman, Reidel, Dordrecht, 1986, xv + 225 pages, US $39.00/ Dfl.110.00, ISBN 90-277-1865-2. Mathematics and Computers in Simulation, 28(5):407, October 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900777.

Anonymous:1986:BRB

[779] Anonymous. Book review: The Beauty of Fractals: H.-O. Peitgen and P. H. Richter, Springer, Berlin, 1986, 199 pages, DM 78.00, ISBN 3-540-15851-0. Mathematics and Computers in Simulation, 28(5): 407, October 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475486900789.

Anonymous:1986:BRNc

[780] Anonymous. Book review: Nonlinear Functional Analysis and its Applications: S. P. Singh(Ed.), Reidel, Dordrecht, 1986, 418 pages, US$69.95, ISBN 90-277-2211-0. Mathematics and Computers in Simulation, 28(5): 408, October 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475486900790.

Anonymous:1986:NIe

[781] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 28(5):409–415, October 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486900807.

Anonymous:1986:ICEb

[782] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 28(5):415– 424, October 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475486900819.

Anonymous:1986:PD

[783] Anonymous. Pages 425–516 (December 1986). Mathematics and Computers in Simulation, 28(6):??, December 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Kinnmark:1986:MNW

[784] Ingemar P. E. Kinnmark and William G. Gray. The minimum nonaliased wavelength in an orthogonal multidimensional grid. Mathematics and Computers in Simulation, 28(6):425– 431, December 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475486900388.

Cao:1986:OAP

[785] Xi-Ren Cao and Yves Dallery. An operational approach to perturbation analysis of closed queuing networks. Mathematics and Computers in Simulation, 28(6):433–451, December 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548690039X.

Yousif:1986:EGR

[786] W. S. Yousif and D. J. Evans. Explicit group over-relaxation methods for solving elliptic partial differential equations. Mathematics and Computers in Simulation, 28(6):453– 466, December 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475486900406.

Swiniarski:1986:MCS

[787] R. Swiniarski. Modelling and computer simulation of nonlinear biotechnological processes. Mathematics and Computers in Simulation, 28(6): 467–471, December 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475486900418.

Moussiopoulos:1986:SSL

[788] Nicolas Moussiopoulos. Simulation of the sea- and land-breezes in Athens. Mathematics and Computers in Simulation, 28(6):473–478, December 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847548690042X.

Walter:1986:ENU

[789] E. Walter, H. Piet-Lahanier, and J. Happel. Estimation of nonuniquely identifiable parameters via exhaustive modeling and membership set theory. Mathematics and Computers in Simulation, 28(6):479– 490, December 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475486900431.

Elperin:1986:PMC

[790] T. Elperin. Proposal for a Monte Carlo simulation of turbulent mixing in reactive flows. Mathematics and Computers in Simulation, 28(6): 491–499, December 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475486900443.

Anonymous:1986:NIf

[791] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 28(6):501–510, December 1986. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/0378475486900455.

Anonymous:1986:AIV

[792] Anonymous. Author index to volume 28. Mathematics and Computers in Simulation, 28(6):511–513, December 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475486900467.

Anonymous:1986:CV

[793] Anonymous. Contents volume 28. Mathematics and Computers in Simulation, 28(6):515–516, December 1986. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475486900479.

Anonymous:1987:PF

[794] Anonymous. Pages 1–74 (February 1987). Mathematics and Computers in Simulation, 29(1):??, February 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1987:EB

[795] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 29(1):ii, February 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487900619.

Lipitakis:1987:ESD

[796] Elias A. Lipitakis and David J. Evans. Explicit semi-direct methods based on approximate inverse matrix techniques for solving boundary-value problems on parallel processors. Mathematics and Computers in Simulation, 29 (1):1–17, February 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487900620.

Witten:1987:SMS

[797] M. Witten, B. L. Bodnar, and A. C. Liu. Simulation and modeling of a single bus tightly coupled multiprocessor system. Mathematics and Computers in Simulation, 29(1):19– 31, February 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475487900632.

Holodniok:1987:DIT

[798] M. Holodniok and M. Kubíček. Determination of invariant tori bifurcation points. Mathematics and Computers in Simulation, 29(1):33– 39, February 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475487900644.

Findler:1987:PSO

[799] Nicholas V. Findler, Cher Lo, and Ron Lo. Pattern search for optimization. Mathematics and Computers in Simulation, 29(1):41–50, February 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475487900656.

Quinn:1987:FEM

[800] Dennis W. Quinn. A finite element method for computing sound propagation in ducts. Mathematics and Computers in Simulation, 29 (1):51–64, February 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487900668.

Anonymous:1987:NIa

[801] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 29(1):65–68, February 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548790067X.

Anonymous:1987:ICE

[802] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 29(1):68–74, February 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475487900681.

Anonymous:1987:PA

[803] Anonymous. Pages 75–166 (April 1987). Mathematics and Computers in Simulation, 29(2):??, April 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Enting:1987:MSC

[804] I. G. Enting. A modelling spectrum for carbon cycle studies. Mathematics and Computers in Simulation, 29(2):75–85, April 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487900991.

Kinnmark:1987:PCO

[805] Ingemar P. E. Kinnmark. A principle for construction of one-step integration methods with maximum imaginary stability limits. Mathematics and Computers in Simulation, 29 (2):87–106, April 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487901005.

Quinn:1987:ML

[806] Dennis W. Quinn. Modeling of lightning. Mathematics and Computers in Simulation, 29(2):107– 118, April 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475487901017.

Adomian:1987:VOSa

[807] G. Adomian. Vibration in offshore structures: an analysis for the general nonlinear stochastic case — part I. Mathematics and Computers in Simulation, 29(2):119– 122, April 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475487901029.

Kreimer:1987:MDD

[808] Joseph Kreimer. Minimization of the delay in a data-communication network with centralized control under uncertainty. Mathematics and Computers in Simulation, 29(2):123– 127, April 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475487901030.

Korn:1987:TTN

[809] Granino A. Korn. Tricks and treats: Nonlinear operations in digital simulation. Mathematics and Computers in Simulation, 29(2):129– 143, April 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475487901042.

Anonymous:1987:BRNa

[810] Anonymous. Book review: Nonlinear approximation theory: D Braess: Springer, Berlin, 1986, xiv + 290 pages, DM 148.00, ISBN 3-540-13625-8. Mathematics and Computers in Simulation, 29(2):145, April 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487901054.

Anonymous:1987:BRNb

[811] Anonymous. Book review: The numerical analysis of ordinary differential equations: J. C. Butcher: Wiley, Chichester, 1987, xv + 512 pages, £38.50, ISBN 0-471-91046-5. Mathematics and Computers in Simulation, 29(2):145, April 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487901066.

Anonymous:1987:BRMa

[812] Anonymous. Book review: M. C. Escher: Art and science: H. S. M.

Coxeter et al. (Eds.): North-Holland, Amsterdam, 1986, xiv + 402 pages, $50.00, ISBN 0-444-70011-0. Mathematics and Computers in Simulation, 29(2):146, April 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487901078.

Anonymous:1987:BRMb

[813] Anonymous. Book review: Multivariate approximation theory: Selected topics: E. W. Cheney: SIAM, Philadelphia, 1986, 68 pages, ISBN 0-89871-207-6. Mathematics and Computers in Simulation, 29(2): 146, April 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548790108X.

Anonymous:1987:BRL

[814] Anonymous. Book review: Large scale eigenvalue problems: J. Cullum and R. A. Willoughby (Eds.): North-Holland, Amsterdam, 1986, viii + 330 pages, $51.75, ISBN 0-444-70074-9. Mathematics and Computers in Simulation, 29(2):146, April 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487901091.

Anonymous:1987:BRIa

[815] Anonymous. Book review: Industrial numerical analysis: C. M. Elliot and S. McKee (Eds.), Clarendon Press, Oxford, 1986, xvii + 256 pages, £22.50, ISBN 0-19-853190-7. Mathematics and Computers in Simulation, 29(2):146, April 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487901108.

Anonymous:1987:BRV

[816] Anonymous. Book review: The VNR concise encyclopedia of mathematics: W. Gellert et al.: Van Nostrand Reinhold, New York, 1977, 816 pages, £21.95, ISBN 0-442-22646-2. Mathematics and Computers in Simulation, 29(2):147, April 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548790111X.

Anonymous:1987:BRDa

[817] Anonymous. Book review: Dynamic analysis of enzyme systems: K. Hayashi and N. Sakamoto: Springer, Berlin, 1986, xiii + 370 pages, DM 98.00, ISBN 3-540-15485-X. Mathematics and Computers in Simulation, 29(2):147, April 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487901121.

Anonymous:1987:BRCa

[818] Anonymous. Book review: Combinatorics: Theory and applications: V. Krishnamurthy: Ellis Horwood, Chichester, 1986, xxxii + 483 pages, £39.50, ISBN 0-7458-0100-5. Mathematics and Computers in Simulation, 29(2):147, April 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487901133.

Anonymous:1987:BRMc

[819] Anonymous. Book review: Mathematical programming: Theory and applications: M. Minoux: Wiley, Chichester, 1986, xxviii + 489 pages, £34.95, ISBN 0-471-90170-9. Mathematics and Computers in Simulation, 29(2):147, April 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487901145.

Anonymous:1987:BRCb

[820] Anonymous. Book review: Computer simulation and modelling: F. Neelamkavil: Wiley, Chichester, 1987, xvi + 307 pages, £17.50, $29.95, ISBN 0-471-91130-5. Mathematics and Computers in Simulation, 29 (2):147, April 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475487901157.

Anonymous:1987:BRNc

[821] Anonymous. Book review: Numerical methods in engineering and science: C. E. Pearson: Van Nostrand Reinhold, New York, 1986, x + 214 pages, £32.45, ISBN 0-442-27344-4. Mathematics and Computers in Simulation, 29(2):148, April 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487901169.

Anonymous:1987:BRCc

[822] Anonymous. Book review: Computational numerical methods: C. Phillips and B. Cornelius: Ellis Horwood, Chichester, 1986, 375 pages, £39.50, ISBN 0-85312-495-7. Mathematics and Computers in Simulation, 29(2):148, April 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487901170.

Anonymous:1987:BRSa

[823] Anonymous. Book review: System modelling and optimization: A. Prekopa, J. Szelezsan and B. Strazicky (Eds.): Springer, Berlin, 1986, xii + 1046 pages, DM 228.00, ISBN 3-540-16854-0. Mathematics and Computers in Simulation, 29 (2):148, April 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475487901182.

Anonymous:1987:BRDb

[824] Anonymous. Book review: Discrete iterations: a metric study: F. Robert: Springer, Berlin, 1986, xi + 195 pages, DM 138.00, ISBN 3-540-13623-1. Mathematics and Computers in Simulation, 29(2):148, April 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487901194.

Anonymous:1987:BRMd

[825] Anonymous. Book review: Monte Carlo, optimization, simulation and sensitivity of queuing networks: R. V. Rubinstein: Wiley, New York, 1986, xi + 260 pages, £33.75, ISBN 0-471-81929-8. Mathematics and Computers in Simulation, 29 (2):148, April 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475487901200.

Anonymous:1987:BRT

[826] Anonymous. Book review: Theory of linear and integer programming: A. Schrijver: Wiley, Chichester, 1986, xi + 471 pages, £37.50, ISBN 0-471-90854-1. Mathematics and Computers in Simulation, 29(2):149, April 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487901212.

Anonymous:1987:NIb

[827] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 29(2):151, April 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487901224.

Anonymous:1987:AFI

[828] Anonymous. Announcement: First IFAC symposium on modelling and control in biomedical systems: April 6–8, 1988, Venice, Italy. Mathematics and Computers in Simulation, 29(2):152, April 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487901236.

Anonymous:1987:CPI

[829] Anonymous. Call for papers: IMACS conference on dynamical systems modelling: October 1987, Kishinev, U.S.S.R. Mathematics and Computers in Simulation, 29(2):152– 153, April 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475487901248.

Anonymous:1987:CE

[830] Anonymous. Calendar of events. Mathematics and Computers in Simulation, 29(2):154–157, April 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548790125X.

Anonymous:1987:ITC

[831] Anonymous. IMACS technical committees. Mathematics and Computers in Simulation, 29(2):158– 159, April 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475487901261.

Anonymous:1987:SCP

[832] Anonymous. Second call for papers: 12th IMACS world congress: Paris, France, July 18–22, 1988. Mathematics and Computers in Simulation, 29(2):160–166, April 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487901273.

Riganti:1987:I

[833] R. Riganti. Introduction. Mathematics and Computers in Simulation, 29 (3–4):167–168, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487901285.

West:1987:MIU

[834] Bruce J. West. Measurement, information and uncertainty. Mathematics and Computers in Simulation, 29 (3–4):169–189, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487901297.

vomScheidt:1987:SSR

[835] J. vom Scheidt and B. Fellenberg. Some simulation results as to weakly correlated processes. Mathematics and Computers in Simulation, 29 (3–4):191–208, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487901303.

Kellogg:1987:DPQ

[836] Stuart D. Kellogg and J. Wesley Barnes. The distribution of products, quotients, and powers of two dependent H-function variates. Mathematics and Computers in Simulation, 29 (3–4):209–221, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487901315.

Ladde:1987:SSR

[837] G. S. Ladde and S. Seikkala. On sample solutions of random initial value and Nicoletti boundary value problems. Mathematics and Computers in Simulation, 29(3–4):223– 231, July 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475487901327.

Bonzani:1987:CAP

[838] I. Bonzani, M. G. Zavattaro, and N. Bellomo. On the continuous approximation of the probability density and of the entropy functions for nonlinear stochastic dynamical systems. Mathematics and Computers in Simulation, 29(3–4):233–241, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487901339.

Spigler:1987:MCT

[839] Renato Spigler. Monte Carlo-type simulation for solving stochastic ordinary differential equations. Mathematics and Computers in Simulation, 29(3–4):243–251, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487901340.

Liske:1987:SST

[840] Horst Liske and Eckhard Platen. Simulation studies on time discrete diffusion approximations. Mathematics and Computers in Simulation, 29 (3–4):253–260, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487901352.

Serebrovski:1987:SOC

[841] A. P. Serebrovski. On the synthesis of optimal control for the continuous-time linear stochastic systems with singular mean-square performance index. Mathematics and Computers in Simulation, 29(3–4):261–273, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487901364.

Adomian:1987:NOP

[842] G. Adomian. Nonlinear oscillations in physical systems. Mathematics and Computers in Simulation, 29 (3–4):275–284, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487901376.

Gabetta:1987:WPU

[843] E. Gabetta. On the wave propagation in unbounded random media. Mathematics and Computers in Simulation, 29(3–4):285–290, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487901388.

Tsokos:1987:SMF

[844] Chris P. Tsokos and Albert D. Liles. Stochastic models for forecasting. Mathematics and Computers in Simulation, 29(3–4):291– 311, July 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548790139X.

Christakos:1987:STS

[845] George Christakos. The space transformation in the simulation of multidimensional random fields. Mathematics and Computers in Simulation, 29(3–4):313–319, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487901406.

Anonymous:1987:BRSb

[846] Anonymous. Book review: A short course on functional equations: J. Aczél: D. Reidel, Dordrecht, 1987, 169 pages, US$49.50, ISBN 90-277-2376-1. Mathematics and Computers in Simulation, 29(3–4):321, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487901418.

Anonymous:1987:BRCd

[847] Anonymous. Book review: Computational methods in the mechanics of fracture: S. N. Atluri (Ed.): (Volume 2 in Computational Methods in Mechanics, A Series of Handbooks). North-Holland, Amsterdam, 1986, 414 pages, US$75.00/Dfl. 250.00, ISBN 0-444-87955-2. Mathematics and Computers in Simulation, 29(3– 4):321, July 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548790142X.

Anonymous:1987:BRQa

[848] Anonymous. Book review: Quality circles und kreativität (in German) (Quality circles and creativity): P. Berigir: Volume 53 of a series from the Institute of Economic Management of the University of Zurich. Paul Haupt, Bern, 1986, 244 pages, SFr. 49.00/ DM 59.00, ISBN 3-258-03596-2. Mathematics and Computers in Simulation, 29(3–4):322, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487901431.

Anonymous:1987:BRH

[849] Anonymous. Book review: The higher calculus: a history of real and complex analysis from Euler to Weierstrass: U. Bottazzini: Springer, Heidelberg, 1986, vii + 332 pages, DM 96.00, ISBN 3-540-96302-2. Mathematics and Computers in Simulation, 29(3–4):322, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487901443.

Anonymous:1987:BRIb

[850] Anonymous. Book review: Inverse problems in vibration: G. M. L. Gladwell: Martinus Nijhoff, Dordrecht, 1986, x + 263 pages, ISBN 90-247-3408-8. Mathematics and Computers in Simulation, 29(3– 4):322, July 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475487901455.

Anonymous:1987:BRFa

[851] Anonymous. Book review: Foundation for human-computer communication: K. Hopper and I. A. Newman (Eds.): (Proceedings of the IFIP WG 2.6 Conference on the Future of Command Languages, Rome, 1985) North-Holland, Amsterdam, 1986, 524 pages, US$68.50/Dfl. 185.00, ISBN 0-444-70028-5. Mathematics and Computers in Simulation, 29 (3–4):322–323, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487901467.

Anonymous:1987:BRMe

[852] Anonymous. Book review: Macht in kooperativen führungsstil (in German) (Authority in Cooperative Leadership): R. Kiechl: Volume 48 of a series from the Institute of Economic Management of the University of Zürich. Paul Haupt, Bern, 1985, 424 pages, SFr. 55.00/DM 66.00, ISBN 3-258-03537-7. Mathematics and Computers in Simulation, 29(3–4):323, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487901479.

Anonymous:1987:BRSc

[853] Anonymous. Book review: Singular perturbation methods in control: analysis and design: P. Kokotovic, H. K. Khalil and J. O’Reilly: Academic Press, London, 1986, 371 pages, £43.00, ISBN 0-12-417635-6. Mathematics and Computers in Simulation, 29(3–4):323–324, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487901480.

Anonymous:1987:BRAa

[854] Anonymous. Book review: Applied analysis: A. M. Krall: D. Reidel, Dordrecht, 1986, xi + 561 pages, US$89.50, ISBN 90-277-2328-1. Mathematics and Computers in Simulation, 29(3–4):324, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487901492.

Anonymous:1987:BRAb

[855] Anonymous. Book review: Algorithms: their complexity and efficiency: L. Kronsjö: Wiley, New York, 2nd ed., 1987, xii + 363 pages, £24.95, ISBN 0-471-91201-8. Mathematics and Computers in Simulation, 29(3–4):324, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487901509.

Anonymous:1987:BRPa

[856] Anonymous. Book review: The predator–prey model: M. Peschel and W. Mende: (Do we live in a Volterra world?) Springer, Wien, 1986, 251 pages, DM 52.00… ÖS 364.00, ISBN 3-211-81848-0. Mathematics and Computers in Simulation, 29(3–4):324–325, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487901510.

Anonymous:1987:BREa

[857] Anonymous. Book review: Eurographics ’86: A. Requicha (Ed.): (Proceedings of the European computer graphics conference and exhibition, Lisbon, 25–29 August 1986). North-Holland, Amsterdam, 1986, 392 pages, US$72.25/Dfl. 195.00, ISBN 0-444-70065-X. Mathematics and Computers in Simulation, 29(3–4): 325, July 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475487901522.

Anonymous:1987:BRPb

[858] Anonymous. Book review: Protocol specification, testing and verification VI: B. Sarikaya and G. V. Bockmann (Eds.): North-Holland, Amsterdam, 1986, 428 pages, US$76.00/ Dfl. 190.00 ISBN 0-444-70126-5. Mathematics and Computers in Simulation, 29(3–4):325, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487901534.

Anonymous:1987:BRAc

[859] Anonymous. Book review: Analysis and mathematical physics: H. Triebel: D. Reidel, Dordrecht, 1986, xxviii + 456 pages, US$89.00, ISBN 90-277-2077-0. Mathematics and Computers in Simulation, 29(3–4): 325–326, July 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475487901546.

Anonymous:1987:BRQb

[860] Anonymous. Book review: A quarter century of IFIP: H. Zemanek (Ed.): (Proceedings of the 25th anniversary celebration of IFIP, Munich, 1985). North-Holland, Amsterdam, 1986, 542 pages (plus 17 pages of index), US$45.00/Dfl. 150.00, ISBN 0-444-70003-X. Mathematics and Computers in Simulation, 29(3– 4):326, July 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475487901558.

Anonymous:1987:NIc

[861] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 29(3–4):327–342, July 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548790156X.

Anonymous:1987:PO

[862] Anonymous. Pages 343–460 (October 1987). Mathematics and Computers in Simulation, 29(5):??, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Goby:1987:NCE

[863] F. Goby and A. Razek. Numerical calculation of electromagnetic forces. Mathematics and Computers in Simulation, 29(5):343– 350, October 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475487900693.

Adomian:1987:VOSb

[864] G. Adomian. Vibration in offshore structures — part II. Mathematics and Computers in Simulation, 29 (5):351–356, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548790070X.

Rattay:1987:MSE

[865] Frank Rattay. Modelling and simulation of electrically stimulated nerve and muscle fibers: a review. Mathematics and Computers in Simulation, 29(5):357–366, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487900711.

Miller:1987:PGM

[866] John J. H. Miller. On a Petrov– Galerkin method for the electrical and thermal behaviour of semiconductor devices in two dimensions. Mathematics and Computers in Simulation, 29 (5):367–372, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487900723.

Day:1987:MFM

[867] B. Day, P. Fitt, P. Kimber, and D. Richardson. A mathematical framework for modular simulation. Mathematics and Computers in Simulation, 29(5):373–383, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487900735.

Neghabat:1987:MCR

[868] M. Neghabat, J. L. Ringuest, and T. R. Gulledge. A model of the controller responses of the human temperature regulating system to changes in water temperature. Mathematics and Computers in Simulation, 29 (5):385–397, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487900747.

Carver:1987:TFM

[869] M. B. Carver and M. Salcudean. Twofluid modelling of phase redistribution by obstructions. Mathematics and Computers in Simulation, 29(5): 399–412, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487900759.

Drenovac:1987:MNI

[870] Vojin Drenovac. A method for the numerical integration of mechanical systems with unilateral constraints: study of impact in multibody systems. Mathematics and Computers in Simulation, 29(5):413–420, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487900760.

Eijsenga:1987:HSE

[871] Willem J. Eijsenga. Hybrid simulation of electrical machines in combination with power electronics. Mathematics and Computers in Simulation, 29(5):421–443, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487900772.

Ames:1987:BRE

[872] W. F. Ames and C. Brezinski. Book review: Essential LISP: John R. Anderson, A. T. Corbett and B. J. Reiser: Addison-Wesley, Reading, MA, 1987, 352 pages, US$24.95, ISBN 0-201-11148-9. Mathematics and Computers in Simulation, 29(5): 445, October 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475487900784.

Anonymous:1987:BREb

[873] Anonymous. Book review: Expert systems and knowledge engineering: T. Bernold (Ed.): (Proceedings of the 2nd international conference, Rüschlikon, Switzerland, 1985). North-Holland, Amsterdam, 1986, 322 pages, Dfl. 140.00/US$51.75, ISBN 0-444-70057-9. Mathematics and Computers in Simulation, 29(5):445– 446, October 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475487900796.

Anonymous:1987:BRFb

[874] Anonymous. Book review: Finite element analysis — from concepts to applications: D. S. Burnett: Addison-Wesley, Reading, MA, 1987, 844 pages, ISBN 0-201-10806-2. Mathematics and Computers in Simulation, 29(5):446, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487900802.

Anonymous:1987:BRIc

[875] Anonymous. Book review: Industrial vibration modelling: J. Caldwell and B. Bradley (Eds.): (Proceedings of the 9th annual conference of the Northeast Polytechnics Mathematical Modelling and Computer Simulation Group, Newcastle upon Tyne, 1986). Kluwer Academic Publishers Group, Dordrecht, Netherlands, 1987, 251 pages, Dfl. 150.00/US$69.00/UK £49.95, ISBN 90-247-3423-1. Mathematics and Computers in Simulation, 29(5):446, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487900814.

Anonymous:1987:BRU

[876] Anonymous. Book review: Understanding statistics: B. J. Chalmer: Marcel Dekker, New York, 1987, 448 pages, US and Canada $39.75, others US$47.75, ISBN 0-8247-7322-5. Mathematics and Computers in Simulation, 29(5):446–447, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487900826.

Anonymous:1987:BRSd

[877] Anonymous. Book review: Solving ordinary differential equations I. nonstiff problems: E. Hairer, S. P. Norsett and G. Wanner: (Springer series in computational mathematics, vol. 8).

Springer, Berlin, 1987, 480 pages, US$69.00, ISBN 3-540-17145-2. Mathematics and Computers in Simulation, 29(5):447, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487900838.

Anonymous:1987:BRNd

[878] Anonymous. Book review: Numerical methods in engineering and applied science: B. M. Irons and N. G. Shrive: Ellis Horwood, Chichester, United Kingdom, 1987, 248 pages, UK £10.95 (paper), UK £29.95/U.S. $54.20 (Cloth), ISBN 0-7458-0101-3. Mathematics and Computers in Simulation, 29(5): 448, October 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548790084X.

Anonymous:1987:BRSe

[879] Anonymous. Book review: The state of the art in numerical analysis: A. Iserles, M. J. D. Powell (Eds.): Oxford University Press, Oxford, 1987, xiv + 719 pages, UK £55.00, ISBN 0-19-853614-3. Mathematics and Computers in Simulation, 29(5):448, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487900851.

Anonymous:1987:BRCe

[880] Anonymous. Book review: Computer arithmetic — scientific computation and programming languages: E. Kaucher, U. Kulisch and Ch. Ullrich (Eds): B. G. Teubner, Stuttgart, 1987, 456 pages, DM 28.00, ISBN 3-519-02448-9. Mathematics and Computers in Simulation, 29(5):448– 449, October 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475487900863.

Anonymous:1987:BRPc

[881] Anonymous. Book review: Pacific computer communications ’85: K. H. Kim, K. Chon and C. V. Ramamoorthy (Eds.) (Proceedings of a Seoul Symposium 1985). North-Holland, Amsterdam, 1986, 658 pages, Dfl. 250.00/ US$80.00, ISBN 0-444-70022-6. Mathematics and Computers in Simulation, 29(5):449, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487900875.

Anonymous:1987:BRAd

[882] Anonymous. Book review: Algorithms for approximation: J. C. Mason and M. G. Cox (Eds.): Clarendon Press, Oxford, 1987, xvi + 694 pages, UK £60.00, ISBN 0-19-853612-7. Mathematics and Computers in Simulation, 29(5):449–450, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487900887.

Anonymous:1987:BRMf

[883] Anonymous. Book review: Methods of functional analysis in approximation theory (ISNM 76): C. A. Micchelli, D. V. Pai and B. V. Limaye (Eds.):

Birkhäuser, Basel, 1986, 410 pages, 88.00 SFr. ISBN 3-7643-1761-2. Mathematics and Computers in Simulation, 29(5):450, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487900899.

Anonymous:1987:BRNe

[884] Anonymous. Book review: Numerical algorithms: J. L. Mohammed and J. Walsh (Eds.): Clarendon Press, Oxford, 1986, xii + 356 pages, UK £30.00, ISBN 0-19-853364-0. Mathematics and Computers in Simulation, 29(5):450, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487900905.

Anonymous:1987:BRNf

[885] Anonymous. Book review: Numerical approximation of partial differential equations: E. L. Ortez: (Papers presented at the International Symposium of Numerical Analysis, Madrid, 1985). North-Holland, Amsterdam, 1987, 433 pages, Dfl. 175.00, US$77.75, ISBN 0-444-70140-0. Mathematics and Computers in Simulation, 29(5):450–451, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487900917.

Anonymous:1987:BRB

[886] Anonymous. Book review: The Bellman continuum: R. S. Roth: World Scientific, Singapore, 1986, distributed by Wiley, xxii + 868 pages, UK £67.60, ISBN 9971-5-0090-6. Mathematics and Computers in Simulation, 29(5):451, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487900929.

Anonymous:1987:BRCf

[887] Anonymous. Book review: Comparative worldwide national computer policies: H. Sackman (Ed.): (Proceedings of the 3rd IFIP-TC 9 Conference, Stockholm, 1985). North-Holland, Amsterdam, 1986, 486 pages, Dfl. 160.00/ US$59.25, ISBN 0-444-70056-0. Mathematics and Computers in Simulation, 29(5):451, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487900930.

Anonymous:1987:BRAe

[888] Anonymous. Book review: Akaike information criterion statistics: Y. Sakamoto, M. Ishiguro and G. Kitagawa: Kluwer Academic Publishers Group, Dordrecht, Netherlands, 1986, 290 pages, Dfl. 215.00/US $89.00/UK £62.00, ISBN 90-277-2253-6. Mathematics and Computers in Simulation, 29(5):452, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475487900942.

Anonymous:1987:BRCg

[889] Anonymous. Book review: Computer networks and simulation III: S. Schoemaker (Ed.): North-Holland, Amsterdam, 1986, 412 pages, Dfl 160.00/ US$59.25, ISBN 0-444-70053-6. Mathematics and Computers in Simulation, 29(5):452, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475487900954.

Anonymous:1987:BRW

[890] Anonymous. Book review: Workload characterization of computer systems and computer networks: G. Serazzi (Ed.): (An International Workshop held at Parma, Italy, 1985). North-Holland, Amsterdam 1986, 178 pages, Dfl. 140.00/US$51.75, ISBN 0-444-70051-X. Mathematics and Computers in Simulation, 29(5):452– 453, October 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475487900966.

Anonymous:1987:BRNg

[891] Anonymous. Book review: Numerical mathematics and applications: R. Vichnevetsky and J. Vignes (Eds.): (IMACS Transactions on Scientific Computation ’85; A selection of papers from the 11th IMACS World Congress, Oslo, 1985). North-Holland, Amsterdam, 1986, 429 pages, Dfl. 185.00US$68.50, ISBN 0-444-70067-6. Mathematics and Computers in Simulation, 29(5): 453, October 1987. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475487900978.

Anonymous:1987:NId

[892] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 29(5):455–460, October 1987. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548790098X.

Anonymous:1988:BRLa

[893] Anonymous. Book review: Laurent series and their Padé approximations: A. Bultheel: Birkhäuser, Basel, 1987, xi + 274 pages, SFr. 88.00, ISBN 3-7643-1940-2. Mathematics and Computers in Simulation, 30(1– 2):195–196, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488901267.

Anonymous:1988:EB

[894] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 30(1–2):ii, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548890095X.

Anonymous:1988:I

[895] Anonymous. Introduction. Mathematics and Computers in Simulation, 30 (1–2):1–2, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488900961.

Jun:1988:SDM

[896] Bai Jun, A. J. Jakeman, and J. A. Taylor. Statistical distribution modelling: Function, methods and application to air quality management. Mathematics and Computers in Simulation, 30(1–2):3–9, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900973.

Brooker:1988:CDV

[897] Peter I. Brooker. Changes in dispersion variance consequent upon inaccurately modelled semi-variograms. Mathematics and Computers in Simulation, 30 (1–2):11–16, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900985.

Benyon:1988:PMP

[898] P. R. Benyon. Presenting models in plain English. Mathematics and Computers in Simulation, 30(1– 2):17–25, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900997.

McCall:1988:UCD

[899] D. G. McCall and R. J. Townsley. A use of calibration in the development of simulation models. Mathematics and Computers in Simulation, 30 (1–2):27–32, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488901000.

Green:1988:SAM

[900] D. G. Green, R. E. Reichelt, and R. G. Buck. Self-adaptive modelling algorithms. Mathematics and Computers in Simulation, 30(1–2):33– 38, February 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475488901012.

Green:1988:EFL

[901] D. G. Green, R. H. Bradbury, and S. J. Bainbridge. Embodiment of formal languages. Mathematics and Computers in Simulation, 30(1–2):39– 44, February 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475488901024.

Doukas:1988:IEC

[902] Louis Doukas. Integrated environmental control model for coal-to-electricity power plants. Mathematics and Computers in Simulation, 30 (1–2):45–53, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488901036.

Vint:1988:STB

[903] M. K. Vint and D. B. Gilmore. Simulation of transit bus regenerative braking systems. Mathematics and Computers in Simulation, 30(1– 2):55–61, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488901048.

Vucetic:1988:PSC

[904] Branka Vucetic, Julien Nicolas, and David Skellern. Performance study of coding on satellite channels by simulation. Mathematics and Computers in Simulation, 30(1–2):63– 68, February 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548890105X.

Vucetic:1988:MSM

[905] Branka Vucetic, David Skellern, Michael Miller, and Liren Zhang. Modelling and simulation of M-QAM digital radio systems. Mathematics and Computers in Simulation, 30(1– 2):69–73, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488901061.

Korzeniowski:1988:SMC

[906] Krzysztof Korzeniowski. Simulation methods in current transformer investigation and design. Mathematics and Computers in Simulation, 30 (1–2):75–80, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488901073.

Haritos:1988:ECS

[907] Nicholas Haritos. The excitation of cable-stayed masts by turbulent wind. Mathematics and Computers in Simulation, 30(1–2):81– 86, February 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475488901085.

Haritos:1988:MCS

[908] Nicholas Haritos. Monte Carlo simulation of ocean beacon response to environmental loading. Mathematics and Computers in Simulation, 30(1– 2):87–92, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488901097.

Feteris:1988:SPO

[909] S. M. Feteris and O. Sitnai. Simulation of pyrolysis in oil shale particles. Mathematics and Computers in Simulation, 30(1–2):93– 98, February 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475488901103.

Simmonds:1988:GCP

[910] Ian Simmonds and Glenn Trigg. Global circulation and precipitation changes induced by sea surface temperature anomalies to the North of Australia in a general circulation model. Mathematics and Computers in Simulation, 30 (1–2):99–104, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488901115.

Frederiksen:1988:SMR

[911] C. S. Frederiksen and J. S. Frederiksen. Simulation and models of the role of topographic instability in the formation of atmospheric teleconnection patterns. Mathematics and Computers in Simulation, 30(1–2): 105–110, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488901127.

Fisher:1988:SRL

[912] I. H. Fisher and P. J. Ring. Structuring rainfall-landuse-runoff models for a large catchment in N.S.W. Mathematics and Computers in Simulation, 30(1–2):111–117, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475488901139.

Retnam:1988:IER

[913] M. T. P. Retnam and B. J. Williams. Input errors in rainfall-runoff modelling. Mathematics and Computers in Simulation, 30(1–2):119– 131, February 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475488901140.

Beer:1988:AES

[914] Tom Beer. Applied environmetrics: Simulation applied to the physical environment. Mathematics and Computers in Simulation, 30(1–2):133– 138, February 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475488901152.

Dolman:1988:SSD

[915] G. S. Dolman. Simulating sediment deposition to establish a chronology for an urban lake. Mathematics and Computers in Simulation, 30(1–2): 139–144, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488901164.

Reichelt:1988:SSC

[916] R. E. Reichelt, S. J. Bainbridge, and D. G. Green. A simulation study of crown of thorns starfish outbreaks on the Great Barrier Reef. Mathematics and Computers in Simulation, 30(1– 2):145–150, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488901176.

Barrett:1988:MCR

[917] J. E. Barrett and F. G. Phillips. A model of the circadian rhythm of deep body temperature. Mathematics and Computers in Simulation, 30(1– 2):151–158, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488901188.

McCaskill:1988:MTC

[918] Malcolm R. McCaskill and Graeme J. Blair. Medium-term climatic variation of the northern tablelands of N.S.W. Mathematics and Computers in Simulation, 30(1–2):159– 164, February 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548890119X.

Bywater:1988:MAG

[919] A. C. Bywater, J. W. Oltjen, R. L. Baldwin, and N. R. St.-Pierre. Modelling animal growth. Mathematics and Computers in Simulation, 30(1– 2):165–174, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488901206.

Diggle:1988:RRG

[920] A. J. Diggle. Rootmap: a root growth model. Mathematics and Computers in Simulation, 30(1–2): 175–180, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488901218.

Wilson:1988:STM

[921] S. G. Wilson. Simulation of thermal and moisture boundary-layers during aeration of cereal grain. Mathematics and Computers in Simulation, 30(1– 2):181–188, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548890122X.

Rickert:1988:CMF

[922] K. G. Rickert and G. M. McKeon. Computer models of forage management on beef cattle farms. Mathematics and Computers in Simulation, 30(1– 2):189–194, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488901231.

Anonymous:1988:BRLb

[923] Anonymous. Book review: Littlewood’s miscellany: B. Bollobás (Ed.): Cambridge University Press, Cambridge, 1986, 200 pages, US$11.95, ISBN 0-521-33058-0. Mathematics and Computers in Simulation, 30 (1–2):195, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488901255.

Anonymous:1988:BRSa

[924] Anonymous. Book review: Stochastic analysis of computer storage: O. I. Aven, E. G. Coffman, Jr. and Y. A. Kogan: Mathematics and its applications, Reidel, Dordrecht, Netherlands, 1987, 254 pages, US$89.00/Dfl. 185.00/ £63.00, ISBN 90-277-2515-2. Mathematics and Computers in Simulation, 30(1–2):195, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488901243.

Anonymous:1988:BRI

[925] Anonymous. Book review: Iterative aggregation theory: L. M. Dudkin, I. Rabinovich and I. Vakhutinsky: Marcel Dekker, New York, 1987, xiv + 273 pages, ISBN 0-8247-7570-8. Mathematics and Computers in Simulation, 30(1–2):196, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488901280.

Anonymous:1988:BRNa

[926] Anonymous. Book review: Numerical methods of approximation theory, vol. 8: L. Collatz, G. Meinardus and G. Nü (Eds.): ISNM 81, Birkhäuser, Basel, 1987, 261 pages, SFr. 68.00, ISBN 3-7643-1855-4. Mathematics and Computers in Simulation, 30(1–2):196, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488901279.

Anonymous:1988:BRPa

[927] Anonymous. Book review: Practical methods of optimization: R. Fletcher: Wiley, Chichester, 1987, xiv + 436 pages, £24.50, ISBN 0-471-91547-5. Mathematics and Computers in Simulation, 30(1–2):196– 197, February 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475488901292.

Anonymous:1988:BRCa

[928] Anonymous. Book review: Calculus: R. M. Johnson: Wiley, New York, 1987, 333 pages, £14.95, ISBN 0-7458-0288-5. Mathematics and Computers in Simulation, 30(1–2): 197, February 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475488901322.

Anonymous:1988:BRMa

[929] Anonymous. Book review: The mathematics of Plato’s academy: D. H. Fowler: Clarendon Press, Oxford, 1987, xxi + 401 pages, £50.00, ISBN 0-19-853912-6. Mathematics and Computers in Simulation, 30 (1–2):197, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488901309.

Anonymous:1988:BRNb

[930] Anonymous. Book review: Numerical analysis: I. Jacques and C. Judd: Chapman and Hall, London, 1987, vii + 326 pages, ISBN 0-412-27960-6. Mathematics and Computers in Simulation, 30(1–2): 197, February 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475488901310.

Anonymous:1988:BRNc

[931] Anonymous. Book review: Numerical integration: P. Keast and G. Fairweather (Eds.): Reidel, Dordrecht, Netherlands, 1987, xiii + 394 pages, US$89.00, ISBN 90-277-2514-4. Mathematics and Computers in Simulation, 30(1–2):197, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488901334.

Anonymous:1988:BRFa

[932] Anonymous. Book review: Finitedimensional spaces, algebra, geometry and analysis, vol. I: W. Noll:

Martinus Nijhoff, Dordrecht, Netherlands, 1987, 393 pages, US$97.50/ Dfl. 235.00/£64.50, ISBN 90-247-3581-5 (Hardback). Also available in paperback (ISBN 90-247-3582-3). Mathematics and Computers in Simulation, 30(1–2):198, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488901358.

Anonymous:1988:BRNd

[933] Anonymous. Book review: Nonlinear elliptic and parabolic equations of the second order: N. N. Krylov: Mathematics and its applications — Soviet series, Reidel, Dordrecht, Netherlands, 1987, 462 pages, US $99.00/Dfl. 245.00/£87.00, ISBN 90-277-2289-7. Mathematics and Computers in Simulation, 30 (1–2):198, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488901346.

Anonymous:1988:BRNe

[934] Anonymous. Book review: Numerical recipes: W. H. Press, B. P. Flannery, S. A. Teukolsky and W. T. Vetterling: Cambridge University Press, Cambridge, 1986, xx + 818 pages, ISBN 0-521-30811-9. Mathematics and Computers in Simulation, 30(1– 2):198–199, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548890136X.

Anonymous:1988:BRCb

[935] Anonymous. Book review: Current trends in matrix theory: F. Uhlig and R. Grone (Eds.): North-Holland, Amsterdam, 1987, 448 pages, US $65.00/ Dfl. 225.00, ISBN 0-444-01181-1. Mathematics and Computers in Simulation, 30(1–2):199, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488901383.

Anonymous:1988:BRNf

[936] Anonymous. Book review: Nonstandard asymptotic analysis: I. van den Berg: LNM 1249, Springer, Heidelberg, 1987, ix + 187 pages, DM 28.50, ISBN 3-540-17767-1. Mathematics and Computers in Simulation, 30(1–2):199, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488901395.

Anonymous:1988:BRPb

[937] Anonymous. Book review: Periodic solutions of Hamiltonian systems and related topics: P. H. Rabinowitz, A. Ambrosetti, I. Ekeland and E. J. Zehnder (Eds.): NATO ASI series C: Mathematical and physical sciences Vol. 209, Reidel Dordrecht, Netherlands, 1987, 284 pages US $59.00/Dfl. 145.00/ £39.95, ISBN 90-277-2553-5. Mathematics and Computers in Simulation, 30(1–2):199, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488901371.

Anonymous:1988:BRAa

[938] Anonymous. Book review: Autowave processes in kinetic systems: V. A. Vasiliev, Yu. M. Romanovskii, D. S. Chernavskii and V. G. Yakhmo: Mathematics and its applications — Soviet series, Reidel, Dordrecht, Netherlands, 1987, 262 pages, US $69.00/Dfl. 160.00/ £56.00, ISBN 90-277-2379-6. Mathematics and Computers in Simulation, 30(1–2):200, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488901413.

Anonymous:1988:BRSb

[939] Anonymous. Book review: Simulated annealing: Theory and applications: P. J. M. Van Laarhoven and E. H. L. Aarts: Reidel, Dordrecht, Netherlands, 1987, xi + 186 pages, ISBN 90-277-2513-6. Mathematics and Computers in Simulation, 30(1–2):200, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488901401.

Anonymous:1988:BRTa

[940] Anonymous. Book review: Z transform theory and applications: R. Vich: Mathematics and its applications — East European series, Reidel, Dordrecht, Netherlands, 1987, 246 pages, Dfl. 150.00/US$69.00/£54.00, ISBN 90-277-1917-9. Mathematics and Computers in Simulation, 30 (1–2):200, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488901425.

Anonymous:1988:BRNg

[941] Anonymous. Book review: Numerical fluid dynamics: R. Vichnevetsky (Ed.): North-Holland, Amsterdam, 1987, 213 pages, Dfl. 200.00, ISBN 0-444-70225-3. Mathematics and Computers in Simulation, 30 (1–2):201, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488901449.

Anonymous:1988:BRTb

[942] Anonymous. Book review: Z transform theory and applications: R. Vich: Mathematics and its applications — East European series, Reidel, Dordrecht, Netherlands, 1987, xii + 246 pages, US$69.00, ISBN 90-277-1917-9. Mathematics and Computers in Simulation, 30(1–2): 201, February 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475488901437.

Anonymous:1988:NIa

[943] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 30(1–2):203–206, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/0378475488901450.

Anonymous:1988:CE

[944] Anonymous. Calendar of events. Mathematics and Computers in Simulation, 30(1–2):206–212, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475488901462.

Anonymous:1988:PF

[945] Anonymous. Pages 1–212 (February 1988). Mathematics and Computers in Simulation, 30(1–2):??, February 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1988:PJ

[946] Anonymous. Pages 213–298 (June 1988). Mathematics and Computers in Simulation, 30(3):??, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Galanis:1988:ESI

[947] S. Galanis, A. Hadjidimos, and D. Noutsos. On the equivalence of the k-step iterative Euler methods and successive overrelaxation (SOR) methods for k-cyclic matrices. Mathematics and Computers in Simulation, 30(3):213–230, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900018.

Riganti:1988:ETP

[948] Riccardo Riganti. Evolution of the n th probability density and entropy function in stochastic systems. Mathematics and Computers in Simulation, 30(3):231–242, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548890002X.

Zamani:1988:CFE

[949] Nader G. Zamani and Weiwei Sun. Collocation finite element solution of a compressible flow. Mathematics and Computers in Simulation, 30 (3):243–251, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900031.

Lau:1988:SSN

[950] H. T. Lau. On solving systems of nonlinear equations by simulation. Mathematics and Computers in Simulation, 30(3):253–256, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900043.

Clement:1988:RPI

[951] Thierry Clement and Sylviane Gentil. Reformulation of parameter identification with unknown-but-bounded errors. Mathematics and Computers in Simulation, 30(3):257– 270, June 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475488900055.

Adomian:1988:DMA

[952] G. Adomian, R. Rach, and M. Elrod. The decomposition method applied to stiff systems. Mathematics and Computers in Simulation, 30 (3):271–276, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900067.

Anonymous:1988:BRMb

[953] Anonymous. Book review: Mathematical statistics and probability theory, vol. B: (Proc. 6th Pannonian Symp., Bad Tatzmannsdorf, Austria, 1986). Reidel, Dordrecht, The Netherlands, 1987, 262 pages, Dfl. 148.00/US$49.00/ £40.75, ISBN 90-277-2581-0. Mathematics and Computers in Simulation, 30(3):277, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900079.

Anonymous:1988:BRMc

[954] Anonymous. Book review: Multivariate statistical modeling and data analysis: (Proc. 1986 Symp. at James Madison University). Reidel, Dordrecht, The Netherlands, 1987, 189 pages, Dfl. 110.00/US$54.00/£33.95, ISBN 90-277-2592-6. Mathematics and Computers in Simulation, 30 (3):277, June 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475488900080.

Anonymous:1988:BRCc

[955] Anonymous. Book review: Computational and applied mathematics for engineering analysis: Springer, Heidelberg, Fed. Rep. Germany, 1987, vii + 380 pages, DM 118.00, ISBN 3-540-17505-9. Mathematics and Computers in Simulation, 30(3):277– 278, June 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475488900092.

Anonymous:1988:BRRa

[956] Anonymous. Book review: Robot dynamics algorithms: Kluwer Academic Publishers, (Reidel), Dordrecht, The Netherlands, 1987, 211 pages, Dfl. 105.00/US$42.50/£28.50, ISBN 0-89838-230-0. Mathematics and Computers in Simulation, 30 (3):278, June 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475488900109.

Anonymous:1988:BRCd

[957] Anonymous. Book review: Computational acoustics: North-Holland, Amsterdam, The Netherlands, 1988, 2 volumes, viii + 422 pages resp. viii + 434 pages, Dfl. 273.60/ US$144.00 and ISBN 0-444-70351-9 (for the set), Dfl. 185.00/US$97.25 and ISBN 0-444-70349-7 (for Vol. 1), Dfl. 185.00/US$97.25 and ISBN 0-444-70350-0 (for Vol. II). Mathematics and Computers in Simulation, 30(3):278, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488900110.

Anonymous:1988:BRNh

[958] Anonymous. Book review: Numerical modelling: applications to marine systems: North-Holland, Amsterdam, The Netherlands, 1987, 296 pages, Dfl. 150.00, ISBN 0-444-70305-5. Mathematics and Computers in Simulation, 30(3):278–279, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900122.

Anonymous:1988:BRMd

[959] Anonymous. Book review: Mathematical statistics and probability theory, vol. A: (Proc. 6th Pannonian Symp., Bad Tatzmannsdorf, Austria, 1986). Reidel, Dordrecht, The Netherlands, 1987, 326 pages, Dfl. 168.00/US$68.00/ £46.25, ISBN 90-277-2580-2. Mathematics and Computers in Simulation, 30(3):279, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900134.

Anonymous:1988:BRCe

[960] Anonymous. Book review: Classical principles and optimization problems: a volume in Mathematics and Its Application (Soviet Series). Reidel, Dordrecht, The Netherlands, 1987, 513 pages, Dfl. 270.00/US$129.00/ £82.25, ISBN 90-277-2605-1. Mathematics and Computers in Simulation, 30(3):279–280, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900146.

Anonymous:1988:BRFb

[961] Anonymous. Book review: Functional analysis and control theory: a volume in Mathematics and Its Applications (East European Series). Reidel, Dordrecht, The Netherlands, 1987, 524 pages, Dfl. 295.00/US$118.00/ £80.95, ISBN 90-277-2186-6. Mathematics and Computers in Simulation, 30(3):280, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488900158.

Anonymous:1988:BRSc

[962] Anonymous. Book review: Scientific computing on vector computers: (Special Topics in Supercomputing Vol. 2), North-Holland (Elsevier), Amsterdam, The Netherlands, 1987, 488 pages, ISBN 0-444-70288-1. Mathematics and Computers in Simulation, 30(3):280–281, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548890016X.

Anonymous:1988:BRPc

[963] Anonymous. Book review: Power series from a computational point of view: Springer, Heidelberg, Fed. Rep. Germany, 1987, viii + 132 pages, DM 45.00, ISBN 3-540-96516-5. Mathematics and Computers in Simulation, 30(3):281, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488900171.

Anonymous:1988:BREa

[964] Anonymous. Book review: Expert systems and artificial intelligence in decision support systems: (Proc. 2nd Mini-Euroconf., Lunteren, The Netherlands, 1985), 1987, 338 pages, Reidel, Dordrecht, The Netherlands, Dfl. 135.00/£47.25, ISBN 90-277-2437-7. Mathematics and Computers in Simulation, 30(3):281, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900183.

Anonymous:1988:BRH

[965] Anonymous. Book review: A Haar memorial conference: 2 volumes. North-Holland, Amsterdam, The Netherlands, 1987, 1018 pages, Dfi. 295.45/US$155.50, ISBN 0-444-80095-1 [bad ISBN]. Mathematics and Computers in Simulation, 30 (3):281–282, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900195.

Anonymous:1988:BRSd

[966] Anonymous. Book review: System fault diagnostics, reliability and related knowledge-based approaches, vol. I: fault diagnostics and reliability: (Proc. 1st European Workshop, Rhodes, Greece, 1986). Reidel, Dordrecht, The Netherlands, 1987, 431 pages, Dfl. 195.00/US$78.00/£54.00, ISBN 90-277-2550-0. Mathematics and Computers in Simulation, 30 (3):282, June 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475488900201.

Anonymous:1988:BRSe

[967] Anonymous. Book review: System fault diagnostics, reliability and related knowledge-based approaches, vol. II, knowledge-based and faulttolerant techniques: (Proc. 1st European Workshop, Rhodes, Greece, 1986). Reidel, Dordrecht, The Netherlands, 1987, 438 pages, Dfl. 195.00/ US$78.00/£54.00, ISBN 90-277-2551-9. Mathematics and Computers in Simulation, 30(3):282, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900213.

Anonymous:1988:BRAb

[968] Anonymous. Book review: Asymptotics for orthogonal polynomials: Vol. 1265 in the Lecture Notes in Mathematics series, Springer, Heidelberg, Fed. Rep. Germany, 1987, vi + 201 pages, DM 35.00, ISBN 3-540-18023-0. Mathematics and Computers in Simulation, 30(3):282– 283, June 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475488900225.

Anonymous:1988:BRSf

[969] Anonymous. Book review: Simulated annealing: theory and applications.: Reidel, Dordrecht, The Netherlands, 1987, 198 pages, Dfl. 120.00/ £42.50, ISBN 90-277-2513-6. Mathematics and Computers in Simulation, 30(3):283, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900237.

Anonymous:1988:BREb

[970] Anonymous. Book review: Encyclopaedia of mathematics: (updated and annotated translation of the Soviet “Mathematical Encyclopaedia”). Reidel (Kluwer Academic Publishers), Dordrecht, The Netherlands, 1988, ISBN- 1-55608-010-7 (for the set), ISBN 1-55608-000-X (for Vol. I), Vol. I, 488 pages, Dfl. 365.00/ US$149.00/£112.50. Mathematics and Computers in Simulation, 30 (3):283–284, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900249.

Anonymous:1988:BRCf

[971] Anonymous. Book review: Chinese mathematics, a concise history: Clarendon Press, Oxford, 1987, xiii + 290 pages, £25.00, ISBN 0-19-858181-5. Mathematics and Computers in Simulation, 30(3):284, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900250.

Anonymous:1988:NIb

[972] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 30(3):285–298, June 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900262.

Anonymous:1988:PS

[973] Anonymous. Pages 299–388 (September 1988). Mathematics and Computers in Simulation, 30(4):??, September 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Papatheodorou:1988:TCC

[974] Theodore S. Papatheodorou. Tridiagonal C ^{1}-collocation. Mathematics and Computers in Simulation, 30(4): 299–309, September 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498900015.

Papatheodorou:1988:CSF

[975] Theodore S. Papatheodorou. C^{1}- collocation semidiscretization of u_{t} + cu_{x} = 0: its Fourier analysis and equivalence to the Galerkin method with linear splines. Mathematics and Computers in Simulation, 30(4): 311–323, September 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498900027.

Adomian:1988:ADM

[976] G. Adomian. An adaptation of the decomposition method for asymptotic solutions. Mathematics and Computers in Simulation, 30(4):325– 329, September 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475498900039.

Azadivar:1988:ODV

[977] Farhad Azadivar and Young-Hae Lee. Optimization of discrete variable stochastic systems by computer simulation. Mathematics and Computers in Simulation, 30(4):331–345, September 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498900040.

Ho:1988:EGD

[978] Yu-Chi Ho, Shu Li, and Pirooz Vakili. On the efficient generation of discrete event sample paths under different system parameter values. Mathematics and Computers in Simulation, 30(4): 347–370, September 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498900052.

Ames:1988:BRMa

[979] W. F. Ames. Book review: Mathematics for dynamic modeling: E. Beltrami, Academic Press, Boston, 1987, 288 pages, US$27.50, ISBN 0-12-085555-0. Mathematics and Computers in Simulation, 30(4):371, September 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498900064.

Ames:1988:BRMb

[980] W. F. Ames. Book review: Means and their inequalities (mathematics and its applications — East European series): P. S. Bullen, D. S. Matrinović and P. M. Vasić, D. Reidel Publishing Co., 1988, 459 pages, Dfl. 190.00/US$89.00/£58.75, ISBN 90-277-2629-9. Mathematics and Computers in Simulation, 30(4): 371–372, September 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498900076.

Ames:1988:BRW

[981] W. F. Ames. Book review: Writing for your peers: S. P. Carter, Prager Publishers, New York, 1987, 129 pages, US$29.95, ISBN 0-275-92630-3 (hardbound); US$9.95, ISBN 0-275-92229-4 (paperback). Mathematics and Computers in Simulation, 30(4):372, September 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498900088.

Ames:1988:BRI

[982] W. F. Ames. Book review: Intuition in science and mathematics: E. Fischbein, D. Reidel Publishing Co., 1987, 226 pages, Dfl. 140.00/US$56.00/ £38.50, ISBN 90-277-2506-3. Mathematics and Computers in Simulation, 30(4):372, September 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847549890009X.

Ames:1988:BRD

[983] W. F. Ames. Book review: Designing games and simulations: C. S. Greenblat, Sage Publishing, Newbury Park, CA, 159 pages, US$19.95, ISBN 0-8039-2956-0. Mathematics and Computers in Simulation, 30 (4):372, September 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498900106.

Ames:1988:BRN

[984] W. F. Ames. Book review: Numerical analysis for applied mathematics, science and engineering: D. Greenspan and V. Casulli, Addison-Wesley Publishing Co., Redwood City, CA, 1988, 341 pages, ISBN 0-201-09286-7. Mathematics and Computers in Simulation, 30(4):373, September 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498900118.

Ames:1988:BRS

[985] W. F. Ames. Book review: Special classes of linear operators and other topics (proceedings of the 11th international conference on operator theory, Bucharest, 1986): H. Helson, C. M. Pearcy, F. H. Vasilescu and D. Voiculescu(Editors), Birkhäuser Basel, 1988, 322 pages, US$68.00, ISBN 3-7643-1970-4. Mathematics and Computers in Simulation, 30 (4):373, September 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847549890012X.

Ames:1988:BRV

[986] W. F. Ames. Book review: Variational principles of continuum mechanics with engineering applications, vol. 2: Introduction to optimal design theory: V. Komkov, D. Reidel Publishing Co., 1988, 276 pages, Dfl. 140.00/US$69.00/£43.00, ISBN 90-277-2639-6. Mathematics and Computers in Simulation, 30(4):373– 374, September 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475498900131.

Ames:1988:BRT

[987] W. F. Ames. Book review: Theory of difference equations: Numerical methods and applications(vol. 181, mathematics in science and engineering): V. Lakshmikantham and D. Trigiante, Academic Press, Boston, 1988, 242 pages, US$49.95, ISBN 0-12-434100-4. Mathematics and Computers in Simulation, 30 (4):374, September 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498900143.

Ames:1988:BRC

[988] W. F. Ames. Book review: Computational acoustics — vol. 1: Wave propagation, vol 2: Algorithms and applications (proc. 1st IMACS symp. on computational acoustics, Yale Univ., August 1986): D. Lee, R. L. Sternberg and M. H. Schultz (Editors), North-Holland, Amsterdam, 1988, 422 pages (Vol. 1), 434 pages (Vol. 2), ISBN 0-444-70351-9 (for the set), ISBN 0-444-70349-7 (Vol. 1), ISBN 0-444-70350-0 (Vol. 2). Mathematics and Computers in Simulation, 30(4): 374–375, September 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498900155.

Brezinski:1988:BRN

[989] C. Brezinski. Book review: Nonnegative matrice: H. Minc, Wiley, New York, 1988, xiii + 206 pages, ISBN 0-471-83966-3. Mathematics and Computers in Simulation, 30 (4):375, September 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498900167.

Ames:1988:BRA

[990] W. F. Ames. Book review: Algorithmic logic: G. Mirkowska and A. Salwicki, D. Reidel Publishing Co. and PWN — Polish Scientific Publishers, 1987, 372 pages, Dfl. 185.00/US$79.00/ £55.25, ISBN 90-277-1928-4. Mathematics and Computers in Simulation, 30(4):375, September 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498900179.

Ames:1988:BRP

[991] W. F. Ames. Book review: Partial differential equations for scientists and engineers(3rd edition): T. Myint-U and Debnath, Elsevier Publishing Co., Amsterdam, 1987, 554 pages, ISBN 0-444-01173-0. Mathematics and Computers in Simulation, 30 (4):375–376, September 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498900180.

Brezinski:1988:BRS

[992] C. Brezinski. Book review: Special functions of mathematical physics: A. F. Nikiforov and V. B. Uvarov, Birkhäuser, Basel, 1988, xviii + 427 pages, Sfr 168.00, ISBN 3-7643-3183-6. Mathematics and Computers in Simulation, 30(4):376, September 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475498900192.

Brezinski:1988:BRR

[993] C. Brezinski. Book review: Rational approximation of real functions: P. P. Petrushev and V. A. Popov, Cambridge University Press, Cambridge, 1987, xi + 371 pages, US$69.50, ISBN 0-521-33107-2. Mathematics and Computers in Simulation, 30(4):376, September 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498900209.

Ames:1988:BRR

[994] W. F. Ames. Book review: The Riemann boundary problem on Riemann surfaces (mathematics and its applications — Soviet series): Yu. L. Rodin, D. Reidel Publishing Co., 1988, 199 pages, Dfl. 130.00/US$64.00/ £39.50, ISBN 90-277-2653-1. Mathematics and Computers in Simulation, 30(4):377, September 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498900210.

Ames:1988:BRE

[995] W. F. Ames. Book review: From equilibrium to chaos (practical bifurcation and stability analysis): R. Seydel, Elsevier, Amsterdam, 1988, 367 pages, US$55.00, ISBN 0-444-01250-8. Mathematics and Computers in Simulation, 30(4):377, September 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498900222.

Anonymous:1988:NIc

[996] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 30(4):379–388, September 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475498900234.

Anonymous:1988:PN

[997] Anonymous. Pages 389–478 (November 1988). Mathematics and Computers in Simulation, 30(5):??, November 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Carotenuto:1988:OSP

[998] L. Carotenuto, P. Muraca, and G. Raiconi. Observation strategy for a parallel connection of discretetime linear systems. Mathematics and Computers in Simulation, 30(5): 389–403, November 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900535.

Marcos:1988:PEA

[999] Bernard Marcos and Guy Payre. Parameters estimation of an aquatic biological system by the adjoint method. Mathematics and Computers in Simulation, 30(5):405–418, November 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475488900547.

Ohkita:1988:ARH

[1000] Masaaki Ohkita and Yasuhiro Kobayashi. An application of rationalized Haar functions to solution of linear partial differential equations. Mathematics and Computers in Simulation, 30(5): 419–428, November 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488900559.

Alaylioglu:1988:FEC

[1001] Ayse Alaylioglu. A finite element code for thin plate dynamics. Mathematics and Computers in Simulation, 30 (5):429–440, November 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488900560.

Ong:1988:ANQ

[1002] Kim L. Ong and Michael R. Taaffe. Approximating nonstationary ph(t)/ph(t)/1/c queueing systems. Mathematics and Computers in Simulation, 30(5):441– 452, November 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475488900572.

Murthy:1988:MAS

[1003] C. Siva Ram Murthy and V. Rajaraman. A multiprocessor architecture for solving nonlinear partial differential equations. Mathematics and Computers in Simulation, 30(5):

453–464, November 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900584.

Anonymous:1988:BRPd

[1004] Anonymous. Book review: Perturbation methods in optimal control: A. Bensoussan:, Bordas, Paris, France, 1988, xiv + 573 pages, ISBN 2-04-015748-4. Mathematics and Computers in Simulation, 30 (5):465, November 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900596.

Anonymous:1988:BRMe

[1005] Anonymous. Book review: Means and their inequalities: P. S. Bullen, D. S. Mitrinović and P. M. Vasić:, Reidel, Dordrecht, The Netherlands, 1988, xiv + 459 pages, US$89.00, ISBN 90-277-2629-9. Mathematics and Computers in Simulation, 30 (5):465, November 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900602.

Anonymous:1988:BRK

[1006] Anonymous. Book review: Kalman filtering: C. K. Chui and G. Chen:, Springer-Verlag, Heidelberg, Fed. Rep. Germany, 1987, xv + 191 pages, DM 67.50, ISBN 3-540-18395-7. Mathematics and Computers in Simulation, 30 (5):465–466, November 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488900614.

Anonymous:1988:BRO

[1007] Anonymous. Book review: Orthogonal matrix-valued polynomials and applications: I. Gohberg (Editor):, OT 34, Birkhäuser-Verlag, Basel, Switzerland, 1988, ix + 214 pages, SFr 78.00, ISBN 3-7643-2242-X. Mathematics and Computers in Simulation, 30(5):466, November 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488900626.

Anonymous:1988:BRRb

[1008] Anonymous. Book review: The rapid evaluation of potential fields in particle systems: L. F. Greengard:, The MIT Press, Cambridge, MA, U.S.A., 1988, 91 pages, US$33.75, ISBN 0-262-7110-X. Mathematics and Computers in Simulation, 30 (5):466, November 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488900638.

Anonymous:1988:BRNi

[1009] Anonymous. Book review: Numerical analysis for applied mathematics, science, and engineering: D. Greenspan and V. Casulli: Addison-Wesley, Reading, MA, U.S.A., 1988, ix + 341 pages, ISBN 0-201-09286-7. Mathematics and Computers in Simulation, 30(5):467, November 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548890064X.

Anonymous:1988:BRMf

[1010] Anonymous. Book review: Moments in mathematics: H. J. Landau (Editor):, American Mathematical Society, Providence, U.S.A., 1987, ix + 154 pages, US$25.00, ISBN 0-8218-0114-7. Mathematics and Computers in Simulation, 30(5):467, November 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900651.

Anonymous:1988:BRSg

[1011] Anonymous. Book review: Scientific discovery: P. Langley et al.: the MIT Press, Cambridge, MA, U.S.A., 1987, 357 pages, US$13.50, ISBN 0-262-62052-9. Mathematics and Computers in Simulation, 30 (5):467, November 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900663.

Anonymous:1988:BRRc

[1012] Anonymous. Book review: Reduce — software for algebraic computation: G. Rayna: Springer-Verlag, Heidelberg, Fed. Rep. Germany, 1987, ix + 329 pages, DM 59.00, ISBN 3-540-96598-X. Mathematics and Computers in Simulation, 30(5):467, November 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900675.

Anonymous:1988:NId

[1013] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 30(5):469–478, November 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/0378475488900687.

Anonymous:1988:PD

[1014] Anonymous. Pages 479–587 (December 1988). Mathematics and Computers in Simulation, 30(6):??, December 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Vignes:1988:F

[1015] J. Vignes. Foreword. Mathematics and Computers in Simulation, 30 (6):479, December 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488900699.

Vignes:1988:RSA

[1016] J. Vignes. Review on stochastic approach to round-off error analysis and its applications. Mathematics and Computers in Simulation, 30(6): 481–491, December 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488900705.

Toutounian:1988:PME

[1017] Faezeh Toutounian. Practical methods for evaluating the accuracy of the eigenelements of a symmetric matrix. Mathematics and Computers in Simulation, 30(6):493–504, December 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475488900717.

Alt:1988:FPE

[1018] René Alt. Floating-point error propagation in iterative methods. stochastic methods in round-off error analysis. Mathematics and Computers in Simulation, 30(6):505–517, December 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475488900729.

Abadie:1988:VCM

[1019] J. Abadie and F. Dekhli. A variant of the CESTAC method and its application to constrained optimization. Mathematics and Computers in Simulation, 30(6):519–529, December 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475488900730.

Alliot:1988:DEA

[1020] Nicole Alliot. Data error analysis in unconstrained optimization problems with the CESTAC method. Mathematics and Computers in Simulation, 30 (6):531–539, December 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900742.

Pichat:1988:APC

[1021] M. A. Pichat. All possible computed results in correct floating-point summation. Stochastic methods in round-off error analysis. Mathematics and Computers in Simulation, 30 (6):541–552, December 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900754.

Ton-That:1988:NAC

[1022] Long Ton-That. Numerical accuracy control in fixed-point arithmetic. Mathematics and Computers in Simulation, 30(6):553–561, December 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475488900766.

Anonymous:1988:BRMg

[1023] Anonymous. Book review: Mathematical modeling in combustion and related topics: C. M. Brauner and C. Schmidt-Laine (editors) (NATO ASI series E: applied sciences, vol. 140). Martinus Nijhoff Publishers, Dordrecht, The Netherlands, 1988, 592 pp., Dfl195.00/U.S.$99.00/£59.00, ISBN 90-247-3656-0. Mathematics and Computers in Simulation, 30 (6):563, December 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900778.

Anonymous:1988:BRV

[1024] Anonymous. Book review: Variational and Hamiltonian control systems: P. E.

Crouch and A. J. van der Schaft: (lecture notes in control and information sciences, vol. 101). Springer, Berlin, Fed. Rep. Germany, 1987, 221 pp., DM36.00, ISBN 3-540-18372-8. Mathematics and Computers in Simulation, 30(6):563, December 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548890078X.

Anonymous:1988:BRTc

[1025] Anonymous. Book review: Theory of multicodimensional (n + 1)-webs: V. V. Goldberg: (mathematics and its applications). Kluwer Academic Publishers, Dodrecht, The Netherlands, 1988, 466 pp., Dfl260.00/U.S.$139.00 /£78.00, ISBN 90-277-2756-2. Mathematics and Computers in Simulation, 30(6):564, December 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900791.

Anonymous:1988:BRLc

[1026] Anonymous. Book review: Large scale systems stability under structural and singular perturbations: L. T. Grujic, A. A. Martynyuk and M. Ribbens-Pavella: (lecture notes in control and information sciences, vol. 92). Springer, Berlin, Fed. Rep. Germany, 1987, 368 pp., DM84.00, ISBN 3-540-18300-0. Mathematics and Computers in Simulation, 30(6): 564–565, December 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900808.

Anonymous:1988:BRFc

[1027] Anonymous. Book review: First European symposium on mathematics in industry: M. Hazewinkel, R. M. M. Mattheij and E. W. C. Van Groesen (editors): D. Reidel Publishing Co., Dordrecht, The Netherlands, 1988, 238 pp., Dfl.130.00/U.S.$69.00/ £ 39.50, ISBN 90-277-2730-9. Mathematics and Computers in Simulation, 30(6):565, December 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548890081X.

Anonymous:1988:BRD

[1028] Anonymous. Book review: Distributed parameter systems: F. Kappel, K. Kunisch and W. Schappacher (editors): (lecture notes in control and information sciences, vol. 102). Springer, Berlin, Fed. Rep. Germany, 1987, 343 pp., DM84.00, ISBN 3-540-18468-6. Mathematics and Computers in Simulation, 30 (6):565, December 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488900821.

Anonymous:1988:BRAc

[1029] Anonymous. Book review: Algebraic analysis: D. Przeworska-Rolewicz: (mathematics and its applications — East European series). D. Reidel Publishing Co, Dordrecht, The Netherlands, 1988, 624 pp. Dfl320.00/U.S.$149.00/£98.00, ISBN 90-277-2443-1. Mathematics and Computers in Simulation, 30(6):565– 566, December 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475488900833.

Anonymous:1988:BRB

[1030] Anonymous. Book review: Biomathematics and related computational problems: L. M. Ricciardi (editor): Kluwer Academic Publishers, Dordrecht, The Netherlands, 1988, 733 pp., Dfl.295.00/U.S.$195.00/£ 89.00, ISBN 90-277-2726-0. Mathematics and Computers in Simulation, 30 (6):566, December 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900845.

Anonymous:1988:BRMh

[1031] Anonymous. Book review: Multiple criteria decision analysis in regional planning — concepts, methods and applications: F. Seo and M. Sakawa: D. Reidel Publishing Co., Dordrecht, The Netherlands, 1988, 535 pp., Dfl120.00/U.S.$59.00/£ 37.00, ISBN 90-277-2641-8. Mathematics and Computers in Simulation, 30(6): 566–567, December 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900857.

Anonymous:1988:BRMi

[1032] Anonymous. Book review: Mathematical problems of statistical hydromechanics: M. J. Vishik and A. V. Furskov: (mathematics and its applications — Soviet series). Kluwer Academic Publishers, Dordrecht, The Netherlands, 1988, 576 pp., Dfl.295.00/U.S.$149.00/ £90.00, ISBN 90-277-2336-2. Mathematics and Computers in Simulation, 30(6):567, December 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488900869.

Anonymous:1988:BRTd

[1033] Anonymous. Book review: Tauberian theorems for generalized functions: V. S. Vladimorov, Yu. N. Drozzinov and B. I. Zavialov: (mathematics and its applications — Soviet series). Kluwer Academic Publishers, Dordrecht, The Netherlands, 1988, 293 pp. Dfl185.00/U.S.$89.00/ £57.25, ISBN 90-277-2383-4. Mathematics and Computers in Simulation, 30(6):568, December 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475488900870.

Anonymous:1988:IVP

[1034] Anonymous. IMACS-(Vichnevetsky) prize. Mathematics and Computers in Simulation, 30(6):569–570, December 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475488900882.

Anonymous:1988:IIS

[1035] Anonymous. IMACS international symposium on approximation, optimization and computing. Mathematics and Computers in Simulation, 30 (6):570, December 1988. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475488900894.

Anonymous:1988:NVC

[1036] Anonymous. NASECODE VI conference — 6th international conference on the numerical analysis of semiconductor devices and integrated circuits. Mathematics and Computers in Simulation, 30(6):571– 572, December 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475488900900.

Anonymous:1988:IGI

[1037] Anonymous. IMACS-GAMM international symposium on computer arithmetic and self-validating numerical methods. Mathematics and Computers in Simulation, 30(6):572, December 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475488900912.

Anonymous:1988:CIE

[1038] Anonymous. Calendar of IMACS Events. Mathematics and Computers in Simulation, 30(6):573–580, December 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475488900924.

Anonymous:1988:AIV

[1039] Anonymous. Author index volume 30 (1988). Mathematics and Computers in Simulation, 30(6):581– 584, December 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475488900936.

Anonymous:1988:CV

[1040] Anonymous. Contents volume 30 (1988). Mathematics and Computers in Simulation, 30(6):585–587, December 1988. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475488900948.

Anonymous:1989:PF

[1041] Anonymous. Pages 1–164 (February 1989). Mathematics and Computers in Simulation, 31(1–2):??, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1989:EB

[1042] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 31(1–2):ii, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475489900475.

Riganti:1989:F

[1043] R. Riganti. Foreword. Mathematics and Computers in Simulation, 31 (1–2):1–2, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475489900487.

Bellomo:1989:SPD

[1044] N. Bellomo and F. Flandoli. Stochastic partial differential equations in continuum physics — on the foundations of the stochastic interpolation method for ITO’s type equations. Mathematics and Computers in Simulation, 31 (1–2):3–17, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489900499.

Kubrusly:1989:SMD

[1045] C. S. Kubrusly. On stochastic modelling for discrete bilinear systems in Hilbert space. Mathematics and Computers in Simulation, 31(1–2):19– 30, February 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489900505.

Ladde:1989:LSS

[1046] G. S. Ladde and O. Sirisaengtaksin.

Large-scale stochastic singularly perturbed systems. Mathematics and Computers in Simulation, 31(1–2):31– 40, February 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489900517.

Grasman:1989:AAN

[1047] Johan Grasman. Asymptotic analysis of nonlinear systems with small stochastic perturbations. Mathematics and Computers in Simulation, 31 (1–2):41–54, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475489900529.

Wedig:1989:BSS

[1048] Walter V. Wedig. Bifurcations in stochastic systems — models, analysis and simulation. Mathematics and Computers in Simulation, 31(1– 2):55–69, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475489900530.

Devroye:1989:RVG

[1049] Luc Devroye. On random variate generation when only moments or Fourier coefficients are known. Mathematics and Computers in Simulation, 31 (1–2):71–89, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475489900542.

Kellogg:1989:BFD

[1050] S. D. Kellogg and J. W. Barnes. The bivariate H-function distribution. Mathematics and Computers in Simulation, 31(1–2):91–111, February 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475489900554.

Cao:1989:SRP

[1051] Xi-Ren Cao. System representations and performance sensitivity estimates of discrete event systems. Mathematics and Computers in Simulation, 31(1– 2):113–122, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489900566.

Kreimer:1989:GSA

[1052] Joseph Kreimer. Generalized sensitivity analysis of ergodic stochastic systems. Mathematics and Computers in Simulation, 31(1–2):123– 136, February 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489900578.

Anonymous:1989:BRSa

[1053] Anonymous. Book review: Sequence transformations: J. P. Delahaye: (Volume 11, Springer Series in Computational Mathematics). Springer-Verlag, Berlin, 1988, 252 pp., DM 128, ISBN 3-540-15283-0. Mathematics and Computers in Simulation, 31 (1–2):137, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548990058X.

Anonymous:1989:BRSb

[1054] Anonymous. Book review: Sequence transformations: J. P. Delahaye: Springer Series in Computational Mathematics, vol. 11. Springer-Verlag, New-York, ISBN 0-387-15283-0, xxi + 252 pages, $79.50. Mathematics and Computers in Simulation, 31(1–2):137, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489900591.

Anonymous:1989:BRNa

[1055] Anonymous. Book review: Numerical methods for initial value problems in ordinary differential equations: S. O. Fatunla: (Volume in Computer Science and Scientific Computing). Academic Press, San Diego, 1988, 295 pp., US$44.50, ISBN 0-12-249930-1. Mathematics and Computers in Simulation, 31(1–2):137– 138, February 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489900608.

Anonymous:1989:BRFa

[1056] Anonymous. Book review: Frequency and domain properties of scalar and multivariable feedback systems: J. S. Freudenberg and D. P. Looze: (Volume 104, Lecture Notes in Control and Information Sciences). Springer-Verlag, Berlin, 1988, 281 pp., DM 64, ISBN 3-540-18869-X. Mathematics and Computers in Simulation, 31 (1–2):138, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548990061X.

Anonymous:1989:BRCa

[1057] Anonymous. Book review: Contributions to operator theory and its applications: I. Gohberg, J. W. Helton and L. Rodman (Eds.): (Vol. 35 of Operator Theory: Advances and Applications). Birkhäuser-Verlag, Basel, 1988, US$99.50, ISBN 3-7643-2221-7. Mathematics and Computers in Simulation, 31(1–2):138, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489900621.

Anonymous:1989:BRPa

[1058] Anonymous. Book review: Probability, random processes, and ergodic properties: R. M. Gray. Mathematics and Computers in Simulation, 31 (1–2):139, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489900633.

Anonymous:1989:BRNb

[1059] Anonymous. Book review: Numerical methods: D. Greenspan and P. Rózsa (Eds.): Colloquia Mathematica Societatis János Bolyai, vol. 50, North-Holland, Amsterdam, 1988, ISBN 0-444-70209-1, 688 pages, $223.75. Mathematics and Computers in Simulation, 31(1–2):139, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489900645.

Anonymous:1989:BRNc

[1060] Anonymous. Book review: Nonlinear problems in abstract cones: D. Guo and V. Lakshmikantham: (Vol. 5 of Notes and Reports in Mathematics in Science and Engineering). Academic Press, San Diego, 1988, 275 pp., US$34.95, ISBN 0-12-293475-X. Mathematics and Computers in Simulation, 31(1–2):139, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489900657.

Anonymous:1989:BRNd

[1061] Anonymous. Book review: Nonlinear time series and signal processing: R. R. Mohler (Ed.): (Volume 106, in Lecture Notes on Control and Information Sciences). Springer-Verlag, Berlin, 1988, 145 pp., (soft cover), DM 36, ISBN 3-540-18861-4. Mathematics and Computers in Simulation, 31(1– 2):139–140, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475489900669.

Anonymous:1989:BRCb

[1062] Anonymous. Book review: Computational techniques and applications: CTAC-87: J. Noye and C. Fletcher (Eds).: North-Holland, Amsterdam, 1988, ISBN 0-444-70400-0, xix + 685 pages, $184.25. Mathematics and Computers in Simulation, 31 (1–2):140, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475489900670.

Anonymous:1989:BRCc

[1063] Anonymous. Book review: Computational fluid dynamics: G. de Vahl Davis and C. Fletcher (Eds.): North-Holland, Amsterdam, 1988, ISBN 0-444-70430-2, xiii + 792 pages, $131.50. Mathematics and Computers in Simulation, 31(1–2):140, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475489900682.

Anonymous:1989:BRDa

[1064] Anonymous. Book review: Discrete optimization: R. G. Parker and R. L. Rardin, Academic Press, San Diego, 1988, 472 pp., US$69.95, ISBN 0-12-545075-3. Mathematics and Computers in Simulation, 31 (1–2):140, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489900694.

Anonymous:1989:BRNe

[1065] Anonymous. Book review: The numerical solution of ordinary and partial differential equations: G. Sewell, Academic Press, San Diego, 1988, 271 pp., US$39.95, ISBN 0-12-637475-9. Mathematics and Computers in Simulation, 31(1–2):140– 141, February 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489900700.

Anonymous:1989:BRSc

[1066] Anonymous. Book review: Systems and control encyclopedia: M. G. Singh (Editor-in-Chief): (Theory, Technology, Applications). Pergamon Press, Oxford, 1988, 5611 pp. Mathematics and Computers in Simulation, 31(1– 2):141–142, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489900712.

Anonymous:1989:BRIa

[1067] Anonymous. Book review: Informationbased complexity: J. F. Traub, G.

W. Wasilkowski and H. Wozniakowski: (Volume in Computer Science and Scientific Computing). Academic Press, San Diego, 1988, 523 pp., US$64.50, ISBN 0-12-697545-0. Mathematics and Computers in Simulation, 31 (1–2):142, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489900724.

Anonymous:1989:BRSd

[1068] Anonymous. Book review: Structural analysis and design of multivariable control systems: Y. T. Tsay, L.-S. Shieh and S. Barnett: (Volume 107, Lecture Notes on Control and Information Sciences). Springer-Verlag, Berlin, 1988, 208 pp., DM 54, ISBN 3-540-18916-5. Mathematics and Computers in Simulation, 31(1– 2):142–143, February 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489900736.

Anonymous:1989:BRCd

[1069] Anonymous. Book review: Common Lisp drill: T. Yuasa: Academic Press, San Diego, 1988, 232 pages (paperback), US$19.95, ISBN 0-12-774861-X. Mathematics and Computers in Simulation, 31(1–2): 143, February 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489900748.

Anonymous:1989:BRBa

[1070] Anonymous. Book review: Boundary control and boundary variations: J. P. Zolésio (Ed.): (Volume 100, Lecture Notes on Control and Information Sciences). Springer-Verlag, Berlin, 1988, 398 pp., DM 96, ISBN 3-540-18546-1. Mathematics and Computers in Simulation, 31(1–2): 143, February 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548990075X.

Anonymous:1989:IVP

[1071] Anonymous. IMACS-(Vichnevetsky)

prize. Mathematics and Computers in Simulation, 31(1–2):145– 164, February 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489900761.

Anonymous:1989:PJ

[1072] Anonymous. Pages 165–314 (July 1989). Mathematics and Computers in Simulation, 31(3):??, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Abou-Rabia:1989:CVS

[1073] Osman Abou-Rabia. A class of variable stepsize formulas for the parallel solution of ODE’s. Mathematics and Computers in Simulation, 31 (3):165–169, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901559.

Kampke:1989:MUR

[1074] Thomas Kämpke. Multiple use of random numbers in discrete-event simulation. Mathematics and Computers in Simulation, 31(3):171– 176, July 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475489901560.

Barron:1989:CIP

[1075] Ronald M. Barron. Computation of incompressible potential flow using von Mises coordinates. Mathematics and Computers in Simulation, 31(3):177–188, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475489901572.

Piriou:1989:SES

[1076] F. Piriou and A. Razek. Simulation of electromagnetic systems by coupling of magnetic and electric equations. Mathematics and Computers in Simulation, 31(3):189–194, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475489901584.

Franksen:1989:ECL

[1077] Ole Immanuel Franksen and Ivor Grattan-Guinness. The earliest contribution to location theory? Spatioeconomic equilibrium with Lamé and Clapeyron, 1829. Mathematics and Computers in Simulation, 31(3):195– 220, July 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489901596.

Zamani:1989:BEM

[1078] Nader G. Zamani. The boundary element method in cascade flow computation. Mathematics and Computers in Simulation, 31(3):221– 226, July 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489901602.

Jiang:1989:FEC

[1079] Song Jiang. A C ^{1} finite element collocation method for the equations of one-dimensional nonlinear thermoviscoelasticity. Mathematics and Computers in Simulation, 31 (3):227–239, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901614.

Spall:1989:NOO

[1080] Michael A. Spall and Allan R. Robinson. A new open ocean, hybrid coordinate primitive equation model. Mathematics and Computers in Simulation, 31(3):241–269, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901626.

Rotella:1989:ESS

[1081] F. Rotella and P. Borne. Explicit solution of Sylvester and Lyapunov equations. Mathematics and Computers in Simulation, 31(3):271– 281, July 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489901638.

Anonymous:1989:BRLa

[1082] Anonymous. Book review: Linear algebra and its applications: D. H. Griffel, Ellis Horwood Ltd., Chichester, 1989. Vol. 1: ISBN 0-85312-946-0, XIII + 289 pages, £29.95. Vol. 2: ISBN 0-7458-0581-7, X + 233 pages, £29.95. Mathematics and Computers in Simulation, 31(3):283, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548990164X.

Anonymous:1989:BRNf

[1083] Anonymous. Book review: Numerical mathematics: R. P. Agarwal, Y. M. Chow and S. J. Wilson (Eds), Singapore, 1988. ISNM Vol. 86, Birkhäuser, Basel, 1989. ISBN 3-7643-2255-1, XIII + 526 pages, SFr. 118. Mathematics and Computers in Simulation, 31(3):283, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901651.

Anonymous:1989:BRFb

[1084] Anonymous. Book review: Fundamentals of pattern recognition: M. Pavel, Marcel Dekker, New York, 1989. ISBN 0-8247-8025-6, IX + 183 pages. Mathematics and Computers in Simulation, 31(3):283–284, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901663.

Anonymous:1989:BRFc

[1085] Anonymous. Book review: Finite element approximation for optimal shape design, theory and applications: J. Haslinger and P. Neittaanmäki, J. Wiley, Chichester, 1988. ISBN 0-471-92079-7, XII + 335 pages, £35. Mathematics and Computers in Simulation, 31(3):284, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901675.

Anonymous:1989:BRAa

[1086] Anonymous. Book review: The averaged moduli of smoothness: B. Sendov and V. A. Popov, J. Wiley, Chichester, 1988. ISBN 0-471-91952-7, X + 181 pages, £24.95. Mathematics and Computers in Simulation, 31(3):284, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901687.

Anonymous:1989:BRSe

[1087] Anonymous. Book review: The science of fractal images: H. O. Peitgen and D. Saupe (Eds), Springer-Verlag, Berlin, 1988. ISBN 3-540-96608-0, XIV + 312 pages, DM 69. Mathematics and Computers in Simulation, 31 (3):284–285, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901699.

Anonymous:1989:BRH

[1088] Anonymous. Book review: Handbook of differential equations: D. Zwillinger, Academic Press, Boston, 1989. 673 pp., $49.95, ISBN 0-12-784390-6. Mathematics and Computers in Simulation, 31(3):285, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475489901705.

Anonymous:1989:BRDb

[1089] Anonymous. Book review: Differential equations with discontinuous right hand sides: A. F. Filippov, (mathematics and its applications — Soviet series), Kluwer Academic Publishers, Dordrecht, 1988. 304 pp., Dfl. 198, US$99, U.K. £59, ISBN 90-277-2699-X. Mathematics and Computers in Simulation, 31(3):285–286, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475489901717.

Anonymous:1989:BRSf

[1090] Anonymous. Book review: Séminaire de théorie des nombres Paris 1986– 1987: C. Goldstein, (Vol. 75 progress in mathematics), Birkhäuser, Basel, Boston, 1988. 475 pp., $45, ISBN 3-7643-3414-2 (Basel), 0-8176-3414-2 (Boston). Mathematics and Computers in Simulation, 31(3): 286, July 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475489901729.

Anonymous:1989:BRPb

[1091] Anonymous. Book review: Principles of optimal design-modeling and computation: P. Y. Papalambros and D. J. Wilde, Cambridge University Press, Cambridge, New York, 1988. 416 pp., $49.50, ISBN 0-521-30674-4. Mathematics and Computers in Simulation, 31(3):286, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901730.

Anonymous:1989:BRAb

[1092] Anonymous. Book review: Andreotti-Grauert theory by integral formulas: G. M. Henkin and J. Leiterer, (Vol. 74 of progress in mathematics), Birkhäuser, Boston, Basel, 1988. 270 pp., $44.90. ISBN 3-7643-3413-4 (Basel), 0-8176-3413-4 (Boston). Mathematics and Computers in Simulation, 31(3):286, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901742.

Anonymous:1989:BRG

[1093] Anonymous. Book review: Geothermal reservoir engineering: E. Okandan (Ed.), (Vol. 150 in NATO Advanced Science Institute Series), Kluwer Academic Publishers, Dordrecht, 1988. 332 pp. Dfl 175, US$89, U.K. £52, ISBN 90-247-3751-6. Mathematics and Computers in Simulation, 31 (3):287, July 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489901754.

Anonymous:1989:BRCe

[1094] Anonymous. Book review: Continued means and boundary value problems in function spaces: E. M. Polishchuk, (Vol. 31 in operator theory: Advances and applications), Birkhäuser, Basel, Boston, 1988. 160 pp. $42.00, ISBN 3-7643-2217 (Basel), 008176-2217-9 (Boston). Mathematics and Computers in Simulation, 31 (3):287, July 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475489901766.

Anonymous:1989:BRSg

[1095] Anonymous. Book review: Statistical reliability theory: I. B. Gertsbakh, (Vol. 4 in probability: Pure and applied), Marcel Dekker, New York, 1989. 352 pp. $94.75 (USA and Canada), $113.50 (elsewhere), ISBN 0-8247-8019-1. Mathematics and Computers in Simulation, 31 (3):287, July 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475489901778.

Anonymous:1989:BRBb

[1096] Anonymous. Book review: Boundary elements IX — Vol. 1 — mathematical and computational aspects. Vol. II — stress analysis applications. Vol. III — fluid flow and potential applications: C. A. Brebbra, W. F. Wendland and G. Kuhn (Eds), computational mechanics publications, Southampton UK; Springer-Verlag Berlin 1987. I — 607 pp, £78, ISBN 0-905451-67-8 (UK), ISBN 0-387-18140-7 (Ger). II — 641 pp, £78, ISBN 0-905451-94-5 (UK), ISBN 0-387-18141-5 (Ger.). III — 642 pp, £78, ISBN 0-905451-95-3 (UK), ISBN 0-387-18143-1 (Ger.). The set £195. Mathematics and Computers in Simulation, 31(3):287–288, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548990178X.

Anonymous:1989:BRBc

[1097] Anonymous. Book review: Boundary element fundamentals — basic concepts and recent developments in the Poisson equation: G. S. Gipson, (Vol. 2 in topics in engineering), Computational Mechanics Publications, Southampton, UK, 1987. 288 pp., UK £39, ISBN 0-905451-80-5 (US 0-931215-70-6). Mathematics and Computers in Simulation, 31(3):288, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901791.

Anonymous:1989:BRSh

[1098] Anonymous. Book review: Special purpose computers: B. J. Alder (Ed.), (Volume 5 in computational techniques), Academic Press, Boston, 1988. 282 pp. $49.95, ISBN 0-12-049260-1. Mathematics and Computers in Simulation, 31(3):288, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901808.

Anonymous:1989:BRCf

[1099] Anonymous. Book review: A computational logic handbook: R. S. Boyer and J. S. Moore, (perspectives on computing, V. 23), Academic Press, Boston, 1988. 408 pp., US$39.95, ISBN 0-12-122952-1. Mathematics and Computers in Simulation, 31(3):289, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847548990181X.

Anonymous:1989:BRBd

[1100] Anonymous. Book review: The boundary element reference book: J. Mackerle and C. A. Brebbia (Eds), computational mechanics publications, Southaampton; Springer-Verlag, Berlin, 1988. 382 pp., UK £51, ISBN 0-905451-74-0 (Comp. Mech., South.), 3-540-18584-4 (Springer). Mathematics and Computers in Simulation, 31 (3):289, July 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475489901821.

Anonymous:1989:BRMa

[1101] Anonymous. Book review: Mathematical modelling: J. N. Kapur, John Wiley, New York, 1988. 259 pp., US$21.50, ISBN 0-470-20088-X. Mathematics and Computers in Simulation, 31(3):289–290, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475489901833.

Anonymous:1989:BRNg

[1102] Anonymous. Book review: New computer methods for global optimization: H. Ratschek and J. Rokne, Ellis Horwood Ltd., Chichester, 1988. 229 pp., US$29.95, ISBN 0-7458-0139-0. Mathematics and Computers in Simulation, 31(3):290, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901845.

Anonymous:1989:BRDc

[1103] Anonymous. Book review: Dynamics reported: U. Kirchgraber and H. O. Walther (Eds), (Vol. I), B. G. Teubner, Stuttgart and John Wiley and Sons, New York, 1988. 306 pp., US$32.50, ISBN 3-519-02150-1 (Teubner), 0-471-91661-7 (Wiley). Mathematics and Computers in Simulation, 31(3):290–291, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901857.

Anonymous:1989:BRAc

[1104] Anonymous. Book review: Advances in computers — Vol. 27: M. C. Yovits (Ed.), Academic Press, Boston, 1988. 480 pp., US$69, ISBN 0-12-012127-1. Mathematics and Computers in Simulation, 31(3):291, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901869.

Anonymous:1989:BRPc

[1105] Anonymous. Book review: Physicallybased modelling and simulation of climate and climate change — Part 1: M. E. Schlesinger, (NATO ASI series), Kluwer Academic Press Publishers, Dordrecht, 1988. 624 pp., Dfl. 280, US$149, UK £79, ISBN 90-277-2788-0. Mathematics and Computers in Simulation, 31(3):291, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901870.

Anonymous:1989:BRMb

[1106] Anonymous. Book review: Matrix norms and their applications: G. R. Belitsky and Yu. I. Lyubich, OT 36, Birkhäuser-Verlag, Basel, Switzerland, 1988. VIII + 209 pages, SFr. 118, ISBN 3-7643-2220-9. Mathematics and Computers in Simulation, 31(3):291, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901882.

Anonymous:1989:BRCg

[1107] Anonymous. Book review: Collected papers: M. Riesz, Springer-Verlag, Berlin, F. R. Germany, 1988. VI + 897 pages, ISBN 3-540-18115-6. Mathematics and Computers in Simulation, 31(3):292, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901894.

Anonymous:1989:BRLb

[1108] Anonymous. Book review: The lost notebook and other unpublished papers: S. R. Ramanujan, Springer-Verlag, Berlin, F. R. Germany, 1988.

XXVII + 419 pages, DM 118, ISBN 3-540-18726-X. Mathematics and Computers in Simulation, 31 (3):292, July 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489901900.

Anonymous:1989:BRSi

[1109] Anonymous. Book review: Strong asymptotics for extremal polynomials associated with weights on r: D. S. Lubinsky and E. B. Saff, lecture notes in mathematics vol. 1305, Springer-Verlag, Berlin, F. R. Germany, 1988. VII + 153 pages, ISBN 3-540-18958-0. Mathematics and Computers in Simulation, 31(3):292, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901912.

Anonymous:1989:BRDd

[1110] Anonymous. Book review: Dynamics reported — a series in dynamical systems and their applications: U. Kirchgraber and H. O. Walther (Eds), Vol. 2, John Wiley, New York and B. G. Teubner, Stuttgart, 1989. 269 pp., US$34.95, ISBN 0-471-91958-6. Mathematics and Computers in Simulation, 31(3):292–293, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901924.

Anonymous:1989:BRDe

[1111] Anonymous. Book review: Deformations of mathematical structures — complex analysis with physical applications: J. Lawrynowicz (Ed.), Kluwer Academic Publishers, Dordrecht, 1989. 352 pp., Dfl. 220, US$119, UK £71, ISBN 0-7923-0023-8. Mathematics and Computers in Simulation, 31(3):293, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901936.

Anonymous:1989:BRCh

[1112] Anonymous. Book review: Computer aided engineering systems handbook: J. Puig-Pey and C. A. Brebbia, (2 Volumes), computational mechanics publications, Southampton, UK; Springer Verlag Berlin, 1987. Vol. I, 300 pp., £55; Vol. II, 280 pp., £47, ISBN 0-905451-49-X (UK), ISBN 3-540-17936-4 (Springer). Mathematics and Computers in Simulation, 31(3):293, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901948.

Anonymous:1989:BRR

[1113] Anonymous. Book review: Random fluctuations and pattern growth — experiments and models: H. E. Stanley and N. Ostrowsky (Eds), Kluwer Academic Publishers, Dordrecht, 1988. 355 pp., Dfl. 65, US$29, UK £19.95, ISBN 0-7923-0073-4 (paper), 0-7923-0072-6 (hardback). Mathematics and Computers in Simulation, 31(3):293–294, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847548990195X.

Anonymous:1989:BRSj

[1114] Anonymous. Book review: Sparse matrices — numerical aspects with applications for scientists and engineers: U. Schendel, translated from the German sparse matrizen. Ellis Horwood series in mathematics and its applications; John Wiley-distributors, 1989. 115 pp., US$21.95, Horwood ISBN 0-7458-0635-X; Wiley ISBN 0-470-21406-6. Mathematics and Computers in Simulation, 31(3):294, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901961.

Anonymous:1989:BRSk

[1115] Anonymous. Book review: Structured programming using Turbo Basic: W. Ellis, Jr. and E. Lodi, Academic Press, Boston, 1988. 337 pp. US$23.95, ISBN 0-12-237460-6 (paper); 0-12-237459-2 (hard cover). Mathematics and Computers in Simulation, 31(3):294, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901973.

Anonymous:1989:BRCi

[1116] Anonymous. Book review: Computers and mathematics — the use of computers in undergraduate instruction: D. A. Smith, G. J. Porter, L. C. Leinbach and R. H. Wenger (Eds), MAA Notes #9, Mathematical Association of America, Publisher; Marketed and distributed by John Wiley & Sons, 1988. 147 pp., US$7.65, ISBN 0-88385-059-1. Mathematics and Computers in Simulation, 31(3):294–295, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475489901985.

Anonymous:1989:BRDf

[1117] Anonymous. Book review: Digital signal processing — applications to communications and algebraic coding theories: Academic Press, Boston, 1989. 233 pp., US$37.95, ISBN-0-12-506995-2. Mathematics and Computers in Simulation, 31(3):295, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475489901997.

Anonymous:1989:NI

[1118] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 31(3):297–314, July 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475489902000.

Houstis:1989:ESN

[1119] E. N. Houstis, J. R. Rice, and R. Vichnevetsky. Expert systems for numerical computing. Mathematics and Computers in Simulation, 31(4–5):315– 517, October 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). Special double issue.

Barnett:1989:EES

[1120] D. Barnett and D. Kahaner. Experiences with an expert system for ODEs. Mathematics and Computers in Simulation, 31(4–5):315– 323, October 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489901262.

Moore:1989:TAN

[1121] Peter K. Moore, Can Ozturan, and Joseph E. Flaherty. Towards the automatic numerical solution of partial differential equations. Mathematics and Computers in Simulation, 31(4– 5):325–332, October 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901274.

Dyksen:1989:EEE

[1122] Wayne R. Dyksen and Carl R. Gritter. Ellipic expert: an expert system for elliptic partial differential equations. Mathematics and Computers in Simulation, 31(4–5):333– 342, October 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489901286.

Kowalski:1989:AAA

[1123] A. Daniel Kowalski, Richard L. Peskin, and Mark F. Russo. Anatomy of AGNES: an automatic generator of numerical equation solutions. Mathematics and Computers in Simulation, 31 (4–5):343–352, October 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901298.

Konno:1989:IVD

[1124] Chisato Konno, Yukio Umetani, Mitsuyoshi Igai, and Tadashi Ohta. Interactive/visual DEQSOL: interactive creation, debugging, diagnosis and visualization of numerical simulation. Mathematics and Computers in Simulation, 31(4–5):353– 369, October 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489901304.

Peskin:1989:SNG

[1125] Richard L. Peskin, Sandra S. Walther, and Andy M. Froncioni. Smalltalk — the next generation scientific computing interface? Mathematics and Computers in Simulation, 31(4–5): 371–381, October 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901316.

Balaban:1989:KRA

[1126] David Balaban, Joseph Garbarini, William Greiman, and Mark Durst. Knowledge representations for the automatic generation of numerical simulators for PDEs. Mathematics and Computers in Simulation, 31(4–5): 383–393, October 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901328.

Ford:1989:NKB

[1127] B. Ford, S. J. Hague, and R. M. J. Iles. Numerical knowledge-based systems. Mathematics and Computers in Simulation, 31(4–5):395– 400, October 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548990133X.

Hake:1989:PSD

[1128] J.-Fr. Hake and W. Homberg. PDE software for distributed computing. Mathematics and Computers in Simulation, 31(4–5):401–408, October 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489901341.

Schonauer:1989:CDU

[1129] W. Schönauer, A. Schreiner, and J. Ecke-Schüth. A critical discussion of the user interface of a PDE solver in the context of numerical expert systems. Mathematics and Computers in Simulation, 31(4–5): 409–417, October 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901353.

Tong:1989:CSM

[1130] Siu Shing Tong. Coupling symbolic manipulation and numerical simulation for complex engineering designs. Mathematics and Computers in Simulation, 31(4–5):419– 430, October 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489901365.

Macedo:1989:KBS

[1131] C. G. Macedo, J. C. Diaz, and R. E. Ewing. A knowledge-based system for the determination of activity indicators for self-adaptive grid methods. Mathematics and Computers in Simulation, 31(4–5):431– 439, October 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489901377.

King:1989:KBF

[1132] David King, Tracy Mullen, Bart Rice, Phillip Topping, and Nicholas Weyland. Knowledge-based FFT and convolution synthesis for any number of points. Mathematics and Computers in Simulation, 31(4–5): 441–451, October 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901389.

Boisvert:1989:GAM

[1133] Ronald F. Boisvert. The guide to available mathematical software advisory system. Mathematics and Computers in Simulation, 31(4–5): 453–463, October 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901390.

Berkbigler:1989:GAP

[1134] K. P. Berkbigler and P. A. Max. The graphics advisor: a prototypical advisory expert system. Mathematics and Computers in Simulation, 31(4– 5):465–473, October 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901407.

Hiebert-Dodd:1989:AET

[1135] Kathie L. Hiebert-Dodd. Assisting experts in their data analysis tasks

— an AI approach. Mathematics and Computers in Simulation, 31(4– 5):475–484, October 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901419.

Fox:1989:UNN

[1136] G. Fox, W. Furmanski, and J. Koller.

The use of neural networks in parallel software systems. Mathematics and Computers in Simulation, 31(4– 5):485–495, October 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901420.

Houstis:1989:PEE

[1137] E. N. Houstis, J. R. Rice, and T. S.

Papatheodorou. Parallel (∥) ELL-PACK: an expert system for parallel processing of partial differential equations. Mathematics and Computers in Simulation, 31(4–5):497– 507, October 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489901432. Reprinted

in [1149].

Anand:1989:TEC

[1138] Tejwansh S. Anand and Rajiv Gupta.

A tool for evaluating compiler-based parallelization strategies. Mathematics and Computers in Simulation, 31 (4–5):509–516, October 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901444.

Alekseev:1989:ESI

[1139] A. S. Alekseev, G. N. Erokhin, and N. L. Podkolodny. Expert systems for inverse problems of wave propagation theory. Mathematics and Computers in Simulation, 31(4–5): 517–522, October 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901456.

Anonymous:1989:BRAd

[1140] Anonymous. Book review: Advanced turbo C programming: Academic Press, Boston, 1988. 558 pp., US$45.95. ISBN 0-12-742689-2, 0-12-742690-6 (paper). Mathematics and Computers in Simulation, 31(4–5): 523, October 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489901468.

Anonymous:1989:BRE

[1141] Anonymous. Book review: Enriques surfaces I: Birkhäuser, Basel, Boston, 1989. 397 pp., US$42.50. ISBN 0-8176-3417-7 (Boston); 3-7643-3417-7 (Basel). Mathematics and Computers in Simulation, 31(4–5): 523, October 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548990147X.

Anonymous:1989:BRAe

[1142] Anonymous. Book review: Advanced topics in the theory of dynamical systems: Academic Press, Boston, 1989. 266 pp., US$39.95. ISBN 0-12-269990-4. Mathematics and Computers in Simulation, 31 (4–5):523, October 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901481.

Anonymous:1989:BRIb

[1143] Anonymous. Book review: Introduction to hyperfunctions: D. Reidel Publishing Co., Dordrecht, 1988. 458 pp., Dfl. 320, US$149, UK£ 99. ISBN 90-277-2837-2. Mathematics and Computers in Simulation, 31 (4–5):524, October 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475489901493.

Anonymous:1989:BRV

[1144] Anonymous. Book review: Variational methods in nonconservative phenomena: Academic Press, Boston, 1989. 371 pp., US$59.95. ISBN 0-12-728450-8. Mathematics and Computers in Simulation, 31(4–5): 524, October 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847548990150X.

Anonymous:1989:BRDg

[1145] Anonymous. Book review: Dynamics of curved fronts: Academic Press, San Diego, 1988. 515 pp., US$69.50. ISBN 0-12-550355-5. Mathematics and Computers in Simulation, 31 (4–5):525, October 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475489901511.

Anonymous:1989:BRCj

[1146] Anonymous. Book review: Curves and surfaces for computer aided geometric design — a practical guide: Academic Press, Boston, 1988. 334 pp., US$39.95. ISBN 0-12-249050-9. Mathematics and Computers in Simulation, 31(4–5):525, October 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475489901523.

Anonymous:1989:BRCk

[1147] Anonymous. Book review: Charles Conley memorial volume: Cambridge University Press, 1988. 409 pp., US$69.50. ISBN 0-521-36929-0. Mathematics and Computers in Simulation, 31(4–5):525, October 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475489901535.

Anonymous:1989:ICE

[1148] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 31(4–5):527– 540, October 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475489901547.

Houstis:1990:PEE

[1149] E. N. Houstis, J. R. Rice, and T. S.

Papatheodorou. Parallel ELLPACK: An expert system for parallel processing of partial differential equations. In Houstis et al. [1150], pages 63–73. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988. Reprint of [1137].

Houstis:1990:IMS

[1150] E. N. Houstis, John Rischard Rice, and Robert Vichnevetsky, editors. Intelligent mathematical software systems: proceedings of the First IMACS/ IFAC International Conference on Expert Systems for Numerical Computing, Purdue University, USA., 5–7 December, 1988. North-Holland, Amsterdam, The Netherlands, 1990. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.